A Complete Bibliography of Publications in Fisheries Research (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

12 June 2023 Version 1.02

Title word cross-reference

$162 [War95]. $163.25 [McI91a]. $178 [Bro97d]. $24 [Kes97a]. $24.95 [Mci97c]. $27.50 [Mci95c]. $30.00 [All96]. $35 [Pea93]. $39.95 [Loc93]. $40.00 [Whi91]. $48.50/£ [Mil91]. $50.00 [Mci95c]. $52.15 [All90]. $54.95 [Bru90]. $69.00 [Gat90]. $69.50 [How93]. $89.50 [Wil95]. < [TKH94]. F_{0.1} [MG97]. M [CA99].

0 [All90, All91, All93, Ano92c, Bea95, Bla90, Bla92, Bro97a, Bro97b, Bru90, Cri93, Ear93, Fra96, Gol95, Hea96, Hob91, How96b, Lai93, Lai95a, Lai95b, Luc90, Luc92, Luc95, Luc96, Luc97b, Mac93, McA93, McC90a, McI91a, McI91b, McI91d, McI91c, McI96d, McI97a, McI97b, McV93, Mil97, Nak93, Pat94, Paw90, Paw96, Plu91, Pri91, Pri95, Pri96a, Pri96b, Pur93, Ral96, She93, Sto90, Wil90, Bro97e, Gat90, Loc93, McI90, McI93, Mci92, Ste95, War95, Wil95]. 0-12-634345-4 [Loc93]. 0-13-957994-X [War95]. 0-216-93324-2 [Ear93]. 0-295-96840-0 [Knu91]. 0-295-97192-4 [Han94]. 0-295-97193-2 [Han94]. 0-295-97245-9 [Ste95]. 0-295-97475-3 [Mil97].

0-306-44532-8 [Wil95]. 0-309-05181-9 [Pri96a]. 0-309-05225-4 [McI96d]. 0-387-18769-3 [Luc90]. 0-412-06941-5 [Bro97a]. 0-412-32140-8 [Paw90]. 0-412-32200-5 [Ano92c]. 0-412-33040-7 [McA93]. 0-412-33060-1 [Nak93]. 0-412-33760-6 [McI91c]. 0-412-34920-5 [Luc92]. 0-412-35140-4 [Mac93]. 0-412-37380-7 [Fra96]. 0-412-42960-8 [Luc96]. 0-412-49370-5 [Gol95]. 0-412-55030-X [How96b]. 0-412-55050-4 [Luc97b]. 0-412-57350-4 [Paw96]. 0-412-58080-2 [Pri95]. 0-412-59540-0 [McI96c]. 0-412-59550-8 [Bro97b]. 0-412-62950-X [Bea95]. 0-442-23346-9 [Hob91]. 0-442-31428-0 [Bla92]. 0-442-31836-7 [McV93]. 0-444-43031-8 [Sto90]. 0-444-87308-2 [McI91a]. 0-471-54735-2 [Pat94]. 0-471-57020-6 [Cri93]. 0-471-57856-8 [Lai95b]. 0-471-83237-5 [Wil90]. 0-471-91151-8 [All90]. 0-521-25727-1 [McI91d]. 0-521-32211-1 [Ral96]. 0-521-34150-7 [Gat90]. 0-521-39032-x [McI95d]. 0-521-48448-0 [Bro97e]. 0-632-02974-9 [Lai93]. 0-7020-1314-5 [Plu91]. 0-7514-0022-X [Pri96b]. 0-7514-0243-5 [Pri96b]. 0-7923-1051-9 [All91]. 0-8047 [How93]. 0-8047-1790-7 [Mci92]. 0-8138-0147-8 [Bru90]. 0-8448-1635-3 [McI91b]. 0-85238-156-5 [Bla90]. 0-85238-177-8 [All93]. 0-85238-187-5 [She93]. 0-85238-188-3 [Pur93]. 0-85238-195-6 [Luc95]. 0-85238-207-3 [Hea96]. 0-85238-218-9 [Lai95a]. 0-85238-223-5 [Bro97c]. 0-85238-230-8 [McI97a]. 0-85238-231-6 [McI97b]. 0-88192-105 [McC90a]. 0-913235-58-X [McI93]. 0-919666-79-5 [Mci97c]. 0-943676-5 [Pea93]. 0-9603078-2-6 [McI90]. 0-group [TG95]. 0115 [Whi91]. 0115-4435 [Whi91].

1 [Kes96b, McI91d, McI94, McI95f, Nak93, Ral96, Tur97, All96, Bro97d, Kes97a, Mci95c]. 1-55963-255-0 [Mci95c]. 1-55963-256-9 [Mci95c]. 1-56612-021-7 [All96]. 1-56612-034-9 [Tur97]. 1-872795-63-3 [McI94]. 1-881570-38-X [McI95f]. 1-886106-10-X [Kes96b]. 1-886106-48-7 [Bro97d]. 1-888691-01-8 [Kes97a]. 105 [McC90a]. 1322 [Kes96c]. 1322-655X [Kes96c]. 14^{o} [Urb98]. 1870 [Bro97e]. 1980s [SM96]. 1988 [JM91]. 1990 [Mci97c]. 1994 [Kes96c]. 1996 [STM^{+}97].

2 [Ear93, McI91a, Pat94, Pri95]. 2162 [How93]. 2162-9 [How93]. 21of [Ano95e]. 24 [EE96a, IH95b]. 28 [DK97, Gar97, McI97d]. 2nd [All90, Pri96b].

3 [Hea96, Luc90, Luc97a, McI91b, McI94, McI95e, Mil97, Pri91, Pur93, McI95a]. 3-540-51927-0 [Pri91]. 3-7643-5321-X [Luc97a]. 3Ps [CM99b].

4 [Luc97b, Mac93, McI96d, Paw96, Loc93, War91]. 4-page [Pur93]. 41 [HCF^{+}99a]. 42 [Bas00]. 4435 [Whi91]. 4th [McI90].

5 [Ano92c, Bla90, Bro97a, Gol95, Luc92, Plu91, Pri96b, She93, Wil90, Ear91].

6 [Cri93, Luc95, McI91c, McI97b, McI90, Pea93]. 6-8 [Pea93]. 655X [Kes96c].

7 [Fra96, McA93, McV93, All96, Bro97d, Gat90, Mci92].

8 [All90, All93, Bro97b, Bru90, Lai95b, Luc96, McI97a, Paw90, Sto90, Hop92, Kes97a, Pea93, Wil95]. 89-16422 [Knu91].

9 [All91, Hob91, How93, Lai93, Lai95a, Loc97b, Luc93, Mil91, Pri96a, Tur97, Mci95c, Ste95]. 90 [Loc97b, Luc93, Mil91, Ear91]. 90-5410-637-9 [Loc97b]. 90-5485-001-9 [Luc93]. 90-6191-415-9 [Mil91]. 90-6191-479-5 [Ear91]. 91 [HH93]. 91/67/EEC [HH93]. 93-02 [All96]. 95-04 [Tur97]. 95.00/ [Ear91]. 95.00/US [Mil91]. 971 [Hop92, McI95a]. 971-10-2247-8 [Hop92]. 971-10-2264-8 [Whi91]. 971-8709-05-3 [McI95a].

Âncora [FP99].

A$20 [Kes96c]. Aaron [Pea93]. abalone [HGC97, HG97, McS94, McS96, PdP93, SB96, SB97a, SB98b, SJ97, WA97, WAB98]. Aberdeenshire [Smi90]. Abramis [MKLvD94, PB97]. Absorption [Tor91]. Abundance [HGC97, HG97, SG97, AMIA98, AN97, BBF94, EE95, EE96a, EBG96, FFGG94, FBF90, GS92, Han93, KPH^{+}96, Loc97a, MH97, McS94, PL92, PK99, PG95, PBHS94b, SRMdC95, San96, TGdB98, WMB92]. abundance/biomass [PK99]. abundances [KGM^{+}93]. abyssal [PM98]. Academic [All91, Ear93, Loc93]. Academy [McI96d, Pri96a]. acanthias [RSB^{+}98, Vin91]. Acanthuridae [Pla93]. Acanthurus [Pla93]. acarne [PS96, SME95]. Acceptability [SAV92]. acceptance [JM90]. acclimation [Bha94]. according [FGS92]. Accounting [KPH^{+}96]. Accuracy [PO98]. acidic [UYD^{+}94]. Acknowledgement [Ano96a, Ano97v]. Acknowledgements [Ano94a]. Acoustic [AMIA98, KD98a, Orł98, SI98, TR94, EBG96, FAC97, FGS92, GF92, JHB96, KW98c, LHW98, MP96a, Sve95, Tor91, TGdB98]. acoustic-midwater [FAC97]. Acoustics [Arg92, Tho92a, BSA92, Ros92, Tho92b, Tho98, Ven92, WTM92, Nak93]. Act [MH90]. activities [MG92]. activity [GM91, KK95b, LBBB95, Mil96]. aculeatus [JLMP96]. adapted [HG98]. adaptive [GS93]. addendum [Mil93b]. additive [Das99]. adjacent [Mil91, Nak98]. adjoining [Fur94]. adjustment [Acq92]. Administration [Kes97a]. Admiral [McI94]. Adriatic [DK97, Dul97, AGA98, DK95, DK96, Dul98, KDCP96, KD97, MP96b, PCD^{+}98]. adult [BHF99, GM97, RJS98, TR94]. advance [Ste90]. advice [BRCK90, LAG93, SBR98]. Aegean [SP94, AGA98, KBKO99, LTT^{+}97, SP93, SPP97, TLT^{+}98, VMT97]. affecting [MRB96, RD95, RR93]. affinis [CLGK97]. affinities [RTC99]. Africa [BB97, BP98, Cav93, CMGB99, ES98, FVF94, Gov94, Gov95, KK95a, vdWB97]. African [Luc97b, Bux92, Gri97, JB93c]. aft [IVLK92]. after [Dah94, EE93, McI96e, MB95, RB93, SLB96]. agboyiensis [EI98]. Age [And98, Ark93, AGA98, CRLG95, CDE93, CMGB99, DK95, GM97, JT92, KD90, KRS97, KBKO99, KDCP96, KD97, LSP99, LP99, MP97, MNSC96, MS95, NE96, RH90, SS95, VT96, YMSU97, vdWB97, dFB90, BPvD^{+}92, Dul95, GS92, Gov94, GP99, HPM94, Hor98, LS99, LP95a, MG97, PMA95, SÍW91, She90a, ZM92]. age-structured [BPvD^{+}92, GP99, HPM94, PMA95]. Ageing [LP97a, PG96, Ark93, Gra99, MBH99]. agent [FL93]. aggregating [IASS96]. aggregation [KMNM96]. aggregations [DCPM93]. Agriculture [Hea96]. Agulhas [BB97, BP98]. AHP [LMNP98]. aid [Cor98, ELCM92, HF90]. Airmail [Whi91]. Aiyetoro [KO90]. al [Kes96c]. Alabama [TMD97]. alalunga [Hsu99]. Alan [Mci97c]. Alaska [All96, IH95a, IH95b, Tur97, Coo93, Egg92, GS93, Kyl94, Kyl96, LFT91, Meg91, TM93, VLAM93, WZ99]. Albacore [Hsu99, LLL98]. Albanian [PVM^{+}96]. albula [AJ94]. albus [GM95]. Aldrichetta [ELCM92]. Algarve [SM98a, SME95, SMEL98]. Alheit [Paw96]. alizarin [VM97]. Allison [Knu91]. allocating [Mil93a]. Allocation [Eck98, HAS98, MK97, Mil95]. Allozyme [BP96, LMC98]. along [HJSNB98, Mon90, SD90, TG95]. alternate [PCS96]. alternate-haul [PCS96]. alternative [CW93, EC98, Fra95, IH95a, IH95b, LM98, Løk91, WA97]. amberjack [MP97]. Amblygaster [KW91]. American [Bro97e, How93, McI90, McI93, BMB97, BMB98, MR96, MW93, PL92, PKH^{+}95, WMCH92]. Ames [Bru90]. Ammodytes [DVY95, MMV98, PLW99]. among [Cio91, CW93, DRE97, GR94a, LMC98, WN92]. Amphipoda [Bro97d]. amphipods [Bro97d]. Amsterdam [McI91a, Sto90]. anadromous [EMJ90, Lan99b]. analogue [Fra95]. analyse [Bas99, Bas00, GGBC92]. analyses [DPB91, KGM^{+}93, Mci97c]. Analysis [MW92, Mil93b, fN94, Par98, TF94, Aze95, dFB90, CA99, CF99, Car93, CCHW95, Cor98, DE90, Erz90, FDD99, GR91, GAA99, HGC97, HBB97, HPM94, HMP96, IH95a, IH95b, JAC^{+}99, LAG93, LP95a, LC99, LFT91, MP96a, MNA97, MM98, MS95, NLT99, Pea99, PRMH99, PLL97, PvDM^{+}97b, PBHS94b, PdP93, SB97a, SZ98, SP94, SB90, SD91, Sum98, VMT97, Wal99, WN92]. analytic [LMNP98]. analytical [Cha98]. Anambra [EI98]. Anarhichadidae [Mil91]. Anarhichadidae-CLXXV [Mil91]. anchor [OKS98]. Anchorage [Tur97]. anchovy [CfyCaL^{+}96, Dul97, GNN^{+}97, IZJT93, KD90, LLW96, LL96, Mil99, PLL97, Ste91, Tud99]. Anderson [How96b]. Andrew [Kes97a]. angle [CT92]. Angling [TMD97]. Anguilla [Wes98]. animals [McI91c, WH93, War95]. Anisakis [Asp95]. Announcement [Ano90a, Ano90b, Ano91a, Ano93a, Ano94b, Ano96b, Ano97a]. Annual [Kes96a, TS92, BBC^{+}96, Cru97, Mil95, NLT99, PWMM95, RKCB90, SC96]. annularis [GM97, MMH93, PS95a]. anomala [WC95]. anomalies [Buc93]. Antarctic [MS91]. antarcticus [WTHC98]. antenna [LMA^{+}99]. antennatus [CCD^{+}99, RB96, TS92]. anthropogenic [SRSR90]. Antillean [Gob98]. Aphanius [LS98, LS99]. Aphanopus [MNSC96]. Apostolos [Mci96b]. apparent [WTHC98]. appearance [WCMG91]. Appendix [Ano94c, Ano94d]. Applicability [PK99]. Application [Bas99, Bas00, BMB98, GAA99, LPH^{+}98, LP95a, PB92, PLL97, SX91, Yam98b, All93, AKB93, Bux92, CP94, Eml91, Gri97, HAS98, JP90, Mac97, MW92, PT95, Ros92, SRG96, TMW98, WN92]. applications [Beg98]. Applied [SB97a, SB98b, BKH99, Orł98, PAL97]. applying [AN97]. Appraisal [VSJ^{+}99]. approach [BW99, BBBF91, BAR99, Cad97, DS99, Gar97, Kes96d, Kes96e, McI97d, SB98a, Bla92]. Approaches [ABB^{+}98, Xia98a]. appropriate [DC97, WH93]. approximate [PBN96]. April [Ano91q, Ano92n, Ano93o, Ano94m, Ano96o, Ano98z, Ano99n, JM91]. aquaculture [McI95a, WH90b, Bro97a, Cri93, How96b, Lai95b, Lai95a]. aquarium [MW93, SJ95]. Aquatic [Hop92, McI95a, Gol95, Pea93, McI97a]. Arabian [AJ90, Hus92, WH97]. Arafura [EW96]. archaeological [KK93]. archipelago [Por94]. Arctic [DJ93]. area [AOH90, Bow90, Caz96, Cru97, HAN97b, HAS98, JM91, RF97]. areas [Aco94, Mil91, Nak98, SM98a]. argentea [PBN96]. Argentina [Cio91, Caz96]. Argentine [CS97, MMPS96]. Argentinian [DLB91, ST96]. argentinus [Ark93, BBC^{+}96, SH97]. Argopecten [MJ93, SG97]. argus [EE95, EE96a, Med98]. Argyrosomus [BK94, Gri97]. Aristeidae [RB96]. Aristeus [CCD^{+}99, RB96, TS92]. Arizona [SM99]. armatus [PM98]. Arnason [McI95e]. arrangement [Hea96]. array [LMA^{+}99]. art [SBR98]. Artificial [BD98, PW97, HML97, Løk90, RJR91, SM97, SM98a, WK93, Loc93]. artificially [HLES93]. artisanal [AAP96, AL95, DC98, GH95, Nor95, VS97]. ascaridoid [HLES93]. Ashgate [Mci96b]. ASI [All91]. Asian [DMA^{+}91]. Aspects [Dun99, SME95, CBB97, IO98, VSJ^{+}99]. aspera [AMON95]. assemblage [ES99, GSM99]. assemblages [SM97, WN92]. assess [Mac97, PLL97, RMR^{+}99, SLB96]. assessed [HMP96]. Assessing [CG97, FAC97, SM98b, ABB^{+}98]. Assessment [AB92, Bae91, BBC^{+}96, BEMW96, Kyl94, NLVF98, ABD94, Arg92, BRCK90, BFP99, Bla90, BREB98, Bri92, Bro97c, Cor98, DMA^{+}91, DLB91, EL97, EE96b, FD97, FF96, FGS92, GW92, Gov95, JLMP96, KW98c, LAG95, Lyo98, MP96a, MPVB98, PG94, Pri96a, RAT91, RKCB90, Sha94, TCA^{+}97, UYD^{+}94, Wil90, Xia97, Xia98a, Kes96b]. Assessments [BKI96, PLW99, Mci97c, Ric91]. associated [FVF94, HGC94, Kai94, KCS^{+}99, MYMP91, Pea99]. association [JB93a, JB93b]. associations [TMK92]. Atherina [LP99]. Atlantic [CM99b, CDB^{+}97, ET91, HCF^{+}99a, Har95, HGHLC98, Paw90, Pri96a, Pur93, BM93, BBC^{+}96, Bow90, BMB98, BP94b, CNW92, CK94, CMK97, Dah96, DVY95, ES99, EMJ90, Fur94, GLGB94, GRG94, GSSM99, GSM99, HCF^{+}99b, Han90, He91, HJH^{+}93, HLES93, Hur98, JB93a, JB93b, JAC^{+}99, Jen90, JAG^{+}99, JHB96, KCS^{+}99, KK93, Ken96, Kly98, LPP96, LSP99, Loc97a, Lon99b, LP99, MR93, MFCG98, MPDW97, MCC90b, Mil93a, MMV98, MBT^{+}99, OKS98, PL96, PBHS94a, PM93, Pow96, PM98, RSB^{+}98, RF99, RMR^{+}99, RTC99, SÍW91, She90a, She90b, SJ95, STM^{+}97, SMS99, SMT95, SBM^{+}98, WH90a, ZM92, Zhe96]. Atmospheric [Kes97a, Dah94]. attachment [HM99]. attenuation [Tor91]. attractants [Løk90]. attracted [PM98]. attraction [HW99, PW97]. attractive [KMNM96]. August [Ano90n, Ano93p, Ano94o, Ano95r, Ano96m, Ano97r, Ano98s, Ano99i]. aurata [KD97]. auratus [FW95, IO98, OCU96, QK96, WMCW93]. aureti [BKH99]. Aurivillius [ELK96]. AusTED [MMM99, MBB95, RTBD95, RM99]. Australia [Kes96c, SB96, SB97a, SB98b, AKB93, Ano95r, BK94, BKI96, BK96, Col98, Fle92, KLB98, Kes96a, LK96, Mon90, NW95, Pea99, SK98, TMW98, WTHC98, WBB^{+}97, WBCG98, WT93, WMCW93, WA97, WAB98]. Australian [AJTP95, Bae91, BREB98, Gri99, KMNP97, MK97, MYMP91, MBB95, NW98, RTBD95]. Author [Ano90c, Ano90d, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano93b, Ano93c, Ano93d, Ano93e, Ano94e, Ano94f, Ano95a, Ano96c, Ano97b]. Automated [PW94, Kai94]. AutoRegressions [Ste91]. autumn [DE90]. Availability [YA99]. Available [Bro97c, Pur93, McI97a]. Avebury [Mci96b]. Avoidance [Mis93, SLW97, GF92]. axillary [SME95]. axisymmetrical [LH90]. Ayres [CS98a]. azevia [And98]. Azores [PM94, Por94, SIE99]. Azov [BSVC99].

B [Luc97b, McI93, Mil91, Plu91, Pri91, Pri96b, Sto90]. Back [GL90, Gra99, Lai93, Lon98, RvLV90]. Back-calculation [GL90, Gra99, RvLV90]. bacterial [FL93]. Bacteriological [AKKD92, RKK90]. Bagridae [IO98]. Bahía [Caz96, Urb98]. Bahrain [AN95]. Bailliere [Plu91]. Bait [He96, CSM95, FJM91, HW99, JM90, KMT^{+}95, Løk90, Løk91, LJ92, LB95, NT96]. baited [FL94, PM98]. baitfish [BMR^{+}90]. baitfishing [BMR^{+}90]. Baja [ACGMGS^{+}99, HJSNB98]. Bakun [Kes97a]. balances [Cad97]. Balearic [MNM97, QAC^{+}98, CCD^{+}99]. Balfour [Gat90]. Balkema [Ear91, Loc97b, Mil91]. Baltic [AW90, CM99a, EE93, HAN97a, HAN97b, HLK^{+}95, Hor98, KK97, LR98, MMPL98, Orł98, Rub94, Wes98]. Bandar [Whi91]. banded [KBKO99]. bandfish [SP93]. Bank [BB97, BP98, CDB^{+}97, FVF94, VCMCA97]. Banks [PL92]. Barbara [McI91a]. barbatus [LAG93, PS96, VP93]. Barents [Asp95, GB98b, HAS98, Nak98, SHH98, TB98, TGdB98, YA99]. barren [Kyl96]. Barrier [NW95, WBCG98]. bartrami [YMSU97]. based [ACC96, Ark93, DMA^{+}91, DLT92, GK91, Gov95, HBB97, HG98, JB93c, KK93, LAG93, LBK97, LCU93, Loc97a, MP96a, MJ93, MNA97, NLT99, PLW99, PLL97, PSMSvD99, PvDM^{+}97b, PBN96, PG96, SAV92, SRG96, TJE^{+}99b, TJ99, Pun97]. Basel [Luc97a]. bases [SG97]. Basic [Sto90]. basis [MM98]. bass [CM97, Jen91, MR93, RP92, SM99, TC92]. bastard [And98]. bathymetric [VO98]. Bay [BD98, Col98, Egg92, RT94, HAN97a, FP97, Mil99, RR93, TC92]. Bayesian [AP96, HPM94, Sch98]. BCS [Kes97a]. be [RMR^{+}99]. Beach [LT92, VCMCA97]. beam [CT92, FM92, GG98, Gui98, KD98b, RJS98, RHLE99, SI98]. beaming [KW98c]. Bear [Kyl94]. Beaumont [Gol95]. bed [LMA^{+}99]. beds [SG97].

Begawan [Whi91]. Behavior [Pri91, Bub96, SMT95]. Behaviour [Ano92c, BBF94, CSR99, GW95, HW99, HF90, HG98, KD98b, LMA^{+}99, MSP97, Mis92, Mis93, Nor95, PB93, SGK98, SJ95, SLW97, Fra96]. Behavioural [LLW96, ZS97]. behaviourally [KS93]. Belgian [FM92]. belly [IVLK92]. benchmark [MG97]. benefit [KGM^{+}93]. Benefits [HAN97a, Egg92, GP99]. Benin [IO98]. benthic [Lon99a, VWL91]. Benthosema [Hus92]. Bering [WZ99]. Berlin [Luc90, Pri91]. Bermuda [EE96a, EE95]. Bernard [Bro97b]. Bertalanffy [LLS97, LR90]. Bethesda [McI93]. between [Azu95, BP96, Cor96, FBB98, GG98, GSM99, HM92, HJH^{+}93, KK95a, LLL98, PM98, SB98c, She90a, She98, SMS99, Smi90, TJA92, VMT97, WH90b, WA97, YB96]. Beverton [Hol95]. Beyond [Tow98]. bias [FBAM99, LBK97]. big [LC99]. bight [MH90, ET91, Har95]. bilinearis [HC99]. billfish [McI90]. Bio [SB96]. Bio-economic [SB96]. biochemical [USS93]. biodiversity [McI96d]. Bioeconomic [LFT91, SBR98, DS99, Sum98]. Bioeconomics [MSP97, Fla98]. Biological [Azu95, LP95b, McI96d, PCD^{+}98, TM93, Gov95, HLM91, MA98, MH97, Mci95c, MM98, VSJ^{+}99]. Biologicas [Kes97a]. Biology [AOH90, CCA97, Luc92, Luc96, McC90a, PM94, QAC^{+}98, VCMCA97, Bla90, Boy90, BP94b, Bro97c, CCD^{+}99, Cor96, EI98, Hus92, IZJT93, MG99, MFCG98, MMN97, MdCP94, PL98, Pri95, Ral96, SME95, WC95, WH97, WMCW93, McI93, Pri96b, Whi91]. biomass [BBBF91, Das99, DC97, FK95, HPM94, PK99]. Bionomics [AN95, KO90]. biotechnology [AA94]. Birgus [FBF90]. Birkhäuser [Luc97a]. bit [Lon98]. Bivalvia [AGA98, GFM99]. Black [Avs95, Aze95, GWGR95, GW95, MNSC96, SIE99, TMD97, Avs94, BSVC99, Das99, GNN^{+}97, Ism95a, Ism95b, IB99, KD90, OB97]. Blackie [Ear93, Pri96b]. blacklip [HGC97, HG97, SB96, SB97a, SB98b, WA97, WAB98]. Blackwell [All93, Bro97c, Hea96, Lai93, Luc95, McI95e, McI97a, McI97b, Pur93, She93]. blainvillei [CRLG95]. Blanca [Caz96]. Blaxter [Pri96b]. Bleeker [KBKO99]. Blennioidei [Mil91]. Blennioidei-13 [Mil91]. Bloch [dS94]. Blue [KMNM96, Tro98, ACGMGS^{+}99, DKH99, Fri95, HLW96, SMS99, Tro97]. bluefin [Pri96a, SZ98]. bluemouth [SIE99]. Board [Ano90g, Ano90h, Ano91i, Ano91j, Ano92f, Ano92g, Ano92h, Ano93i, Ano93j, Ano93k, Ano94h, Ano94i, Ano94j, Ano95f, Ano95g, Ano95h, Ano96h, Ano96i, Ano96j, Ano96k, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j]. Body [SD94, LB95, MKS99]. bogaraveo [SIE99]. bogue [TF91]. Bohai [CLZS97]. Bolinus [MSR95]. Bone [Pri96b, ESNA95]. bonga [KO90]. Book [All90, All91, All93, All96, Ano92c, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano99a, Ano99b, Ano99c, Ano99d, Bea95, Bla90, Bla92, Bro97a, Bro97b, Bro97c, Bro97d, Bro97e, Bru90, Cri93, Ear91, Ear93, Fra96, Gat90, Gol95, Han94, Hea96, Hob91, Hop92, How93, How96b, Kes96b, Kes96c, Kes97a, Knu91, Lai93, Lai95a, Lai95b, Loc93, Loc97b, Luc90, Luc92, Luc93, Luc95, Luc96, Luc97a, Luc97b, Mac93, McA93, McC90a, McI90, McI91a, McI91b, McI91d, McI91c, McI93, McI94, McI95a, McI95e, McI95f, McI95d, McI96c, McI96d, McI97a, McI97b, McV93, Mci92, Mci95c, Mci96b, Mci97c, Mil91, Mil97, Nak93, Pat94, Paw90, Paw96, Pea93, Plu91, Pri91, Pri95, Pri96a, Pri96b, Pur93, Ral96, She93, Ste95, Sto90, Tur97]. Book [War91, War95, Whi91, Wil95, Wil90]. Books [Bla90, Bro97c, Hea96, Lai95a, McI95e, McI97a, McI97b, Pur93, She93, All93]. Books/Blackwell [McI97b]. Boops [TF91]. borealis [Tho92c]. boscii [CDE93, LP97a, Lan99a]. both [LHS95, She90a]. bottle [FF96]. bottom [AS91, ET91, FAC97, GW92, He96, LTT^{+}97, MKLvD94, TLT^{+}98, Tor91, YA99, Bro97a]. bottom-set [He96, MKLvD94]. bottomfishing [KK95b]. Boulenger [Iko96]. Boundary [Tre98a]. Bowdich [KO90]. Box [McI90]. Boyd [Bro97a]. brackish [Dol98, LV98]. brama [MKLvD94, PB97]. brandaris [MSR95]. Brazil [dCB95, Beg98, MH97, PAL97, SH97]. Brazilian [LMC98, PAL97]. bream [CAI96, KDCP96, KD97, MKLvD94, PCD^{+}98, PB97, SIE99]. breams [SMEL98]. Breeding [WH97, CP98a, McA93]. Brienomyrus [Iko96]. brightness [KW98a]. Brisbane [Ano95r]. Bristol [Egg92, RR93]. British [GW99, SB90, Vin91]. broken [GL90, PG96]. brook [AF99]. Brosme [FL94]. Brown [KDD^{+}94, Mil97, Ell93, HH90, HJ92, MPDW97, PMA95, Rob92, TKH94, VA96]. Bruce [Mil91, Mil97]. Brunei [Whi91]. brush [SU93]. Buccinum [Kid94, VSJ^{+}99]. Buckland [She93]. Buijse [Luc93]. building [Mci95c]. bullhead [JAG^{+}99]. Bureau [Kes96c]. burnt [PG96]. Burrowing [CSR99]. bursts [MB95]. bus [KMNP97, MK97]. bus-route [KMNP97, MK97]. Bush [CK94, CMK97]. butterfish [WC95]. By-catch [AJTP95, LK96, BK94, BKI96, BK96, IVLK92, KLB98, Rob92]. Bycatch [BREB98, HC99, KT98, MS98, RD95, RT94, RRdS^{+}97, WBCG98]. bypass [JSE92].

C [All96, Bro97a, Hob91, Knu91]. Côte [ES98]. CA [How93, McI90, Mci92]. Cabo [ACGMGS^{+}99]. Caddy [Wil90]. Cadiz [Mil99]. caeruleus [NMCCMLB99]. caged [HLES93, JHB96]. Calabar [NE96]. calculation [GL90, Gra99, RvLV90]. calculations [Har95]. California [ACGMGS^{+}99, HJSNB98, Kes97a, McI90, NMCCMLB99]. Caligidae [TW93]. Callao [Hop92, SQM95]. Callinectes [CD90]. Cambridge [Bro97e, Gat90, McI91d, McI95d, Ral96, War91]. cameras [Har95, PM98]. Cameroon [GH95]. Campeche [VCMCA97]. Can [CP98b, Mci97c, RMR^{+}99]. Canada [HC99, Hur98, SB90]. Canadian [BMB98, Cha98, Mil93a, Ros92, UYD^{+}94]. canal [PO98, GHHEE98]. canaliculatus [WH97]. Canary [HGHLC98, LP95b, LP99, PL96, PL98]. Cancer [AB92, SB90]. canoe [Loc97a]. canoe-based [Loc97a]. Cantabrian [VO98]. canvas [SX91]. capacity [Coo93]. Cape [Cru97, GM91]. capelanus [PP94]. capelin [GB98b, JS98, TGdB98]. Capello [And98]. capensis [GM91]. Capiz [BAG98]. captive [QK96]. captivity [SGK98]. capture [ABD94, Bub96, CD90, Dic93a, Dic93b, HW99, Jen91, KW98a, MA97, Mis92, Mis93, MMH93, PJ96]. captured [CAI96]. Caranx [HLW96]. carbo [MNSC96]. Carcharhinus [LSP99]. careful [KT98]. Caribbean [Bon97, FBAM99, GMHSR98, RU98]. Carl [McI93]. Carlo [Che96]. Carolina [VWL91]. carp [AKKD92, Bas99, Bas00, JT92]. carried [Cav93]. carrying [Coo93]. Carvalho [Bea95]. case [EL97, GAA99, Med98, San93, SB96, SBSW99, YB96]. Castellón [GAA99]. Catalan [MSR95, MS95, PLL97, SM98b]. Catch [CLZS97, CSM95, PvDM^{+}97a, YF92, Aco94, AMON95, AJTP95, Aze95, Bae91, dFB90, BSVH99, BK94, BKI96, BK96, Che96, CIA96, CfyCaL^{+}96, DMA^{+}91, ET91, EPCPW96, FW95, FK95, GAA99, GM91, Har95, HG97, HM92, IVLK92, Jen96, KLB98, LK96, LP95a, LP97b, LMF92b, Mac97, NW95, NW98, OK96, Otu90, PR96, PMA95, PSMSvD99, RLR92, RP92, RHO99, Rob92, SEJ97, SB98c, SM98a, SM98b, Tro97, WBCG98, WAB98, YDM95, All93]. catch-at-age [dFB90, LP95a]. catch-curve [Mac97]. catch-effort [AMON95, HG97]. catch-per-unit-effort [YDM95]. catchability [KR96, MSP97, PG95]. catches [CTCN93, CCJ^{+}99, FVF94, Gob94, GPV90, Gri99, Kub92, LBK97, MBT^{+}99, RSB^{+}98, Smi90, SP94, SC96, SCP97]. Catching [AA95, LMMK92, NP99, HLK^{+}95, MYMP91, RN98, Sha94]. catfish [AMON95, ELK99, IO98, McV93]. Catla [Bas99, Bas00, JT92, TJA92]. caught [HLES93, PM94, QAC^{+}98, VLAM93]. causative [FL93]. Causes [TH95]. cavillone [CCA97]. Cell [VN98]. Celtic [Du 95, Du 96]. Center [Hop92]. Central [CDB^{+}97, HGHLC98, WMB92, BAG98, HH90, KK95a, LAG95, Lev96, NW95, PT95, PL96, PPS93, Pur93, SG97, CCA97, LP99]. Central-East [CDB^{+}97, HGHLC98, LP99]. Centre [McI95a]. Centro [Kes97a]. Centroberyx [CLGK97]. century [KK93]. Cephalopod [Boy90, PG94]. Cephalopoda [GLGB94, CDB^{+}97, CGG99b, CGG99a, DC98, GRG94, GR94b, PBHK94, PM94, SH97, dSD97]. cephalopods [Lip99]. Cepola [SP93]. Cerastoderma [BP96, CS98b, CSR99]. Chaceon [DLB91]. chain [FS94]. Chairman [Pri96a]. chalcogramma [EPCPW96, Meg91, Par98]. challenges [Ull98]. chameleons [SF99]. change [Bux92, HG97, Ken96, Kly98, Knu91, MKKO95, SRSR90, WTHC98]. change-in-ratio [HG97]. Changes [Bha94, BMB97, CM99a, FGS92, HB97, JT96, LR98, AKKD92, Cad96, FS94, KD98b, KPH^{+}96, LVIG90, Luc97b, PR96, PG95, Pol91, RN98, SAB^{+}93, GSM99]. changing [Car93, DP97]. chanius [TJA92]. Channa [dS94]. Channel [BTBP98, Dun99, JHF93, LAG93, STM^{+}97, McV93]. Channidae [dS94]. Chaoborus [Eck98]. Chapman [Ano92c, Bea95, Bla92, Bro97a, Bro97b, Fra96, Gol95, Hob91, How96b, Luc92, Luc96, Luc97b, Mac93, McA93, McI91c, McI96c, Nak93, Paw90, Paw96, Pri95]. Characterisation [BN92, FJM91]. Characteristics [PVM^{+}96, PWMM95, DLT92, FB97, Iko96, Ken90, MS97, Mil99, Nor95, WZ99]. characterization [AL95]. characters [ME99, VMT97]. Checks [Cad97]. Cheilodactylidae [SQM95]. Cheilodactylus [SQM95]. Chela [TJA92]. chemical [LJ92]. chemoreception [Mac93]. Chenderoh [AL95]. Chesapeake [TC92]. Chichester [All90]. Chile [DC98, GM95, LF93, SG97]. Chilean [CAB99]. China [IZJT93, YB96, CCHW95, fN94]. Chinese [DLT92]. Chinook [BHF99]. Chlamys [Cio91]. Choosing [Med98]. Choptank [SRSR90]. Christopher [Mci97c]. Chromis [DK95]. chronic [Hea96]. Chrysichthys [IO98]. chub [PPS93, YAM98a]. chum [Azu95]. Cichlid [Ano92c]. Cienaga [RU98, MM96, RSM99]. ciliata [BKB97]. circulation [War91]. circumference [BK96]. cirrhosa [WMB92]. clam [RU98, Urb98]. Clarence [LK96]. Clarias [EI98]. Clariidae [EI98]. class [Mac97, PL92]. classes [CM99b, Dol98]. Classification [VA96]. climate [Kly98]. Close [LMC98]. closing [ESW97]. closure [AAP96, CP98a]. Cloth [McI93, Mci95c, Han94]. Cloud [Wil95]. clouds [Mil97]. club [GHR99]. Clupea [DE90, GB98b, HLK^{+}95, Hol94, Hol96, JM91, Mis93, TGdB98, TG95]. Clupeidae [KW91, VCMCA97]. clupeoids [MRB96]. Cluster [HBB97]. CLXXV [Mil91]. Clyde [McI95f, TCA^{+}97]. cm [TKH94]. Co [Sto90]. coarse [TSS97]. Coast [KO90, SD90, ACGMGS^{+}99, AB92, And98, Avs94, Avs95, BF97, CDE93, CCV^{+}98, CDB^{+}97, ET91, ELK96, GBL^{+}97, GLGB94, Gov94, HJSNB98, Ism95a, Ism95b, KW91, LMC98, MSR95, Mon90, MdCP94, NE96, PP94, SMEL98, SÍW91, She98, SK98, Sve99, TG95, TW93, vdWB97, BSA92, IB99, PCEP98, Ros92]. coastal [CfyCaL^{+}96, FM92, Gri99, LL96, MR93, MNJ91, NGH92, PLL97, PSMSvD99, WC95, McI91a]. coasts [KBKO99]. Cochin [SS93]. cockle [BP96, CS98b]. cockles [CSR99]. coconut [FBF90]. Cod [Lon98, PCEP98, TLT^{+}98, Asp95, CM99a, CM99b, DJ93, Dah96, Du 95, EMS^{+}95, FJM91, FL94, GFM94, GL90, GS92, HCF^{+}99a, HCF^{+}99b, He91, HLM91, HLES93, Hor98, HR92, JS98, JM90, KT98, KK93, LVIG90, LHML97, LTT^{+}97, LR96, MMPL98, Mci97c, NLVF98, NP99, OK96, OKS98, PC91, PP94, RJR91, RTC99, Sal91, SU92, Sal95, SB98a, SLB96, SHH98, She98, SPP97, Sve95, VN98, YA99]. Cod-end [PCEP98, TLT^{+}98, GFM94, LTT^{+}97, LR96, OK96]. cod-ends [SLB96, SPP97]. coded [LPH^{+}98]. Codend [WMCH92, EPCPW96]. codends [BK94, BKI96, BK96, CNW92, ESW97, FBB98, FM92, GWGR95, GW95, HCF^{+}99a, HCF^{+}99b, LTS98, MMPL98, SEO96, Tho92c]. Coho [Kyl94, VLAM93]. cohort [CA99, Car93, HMP96, PLL97]. coindetii [GRG94]. Colin [McA93]. collapse [MSP97, PPS93]. collapsing [BMB97]. collections [LT92]. College [All96, Kes97a, Pea93, Tur97]. Collette [Mil91]. Colombia [MM96, RSM99]. Colombian [RU98]. colour [Pur93, WCMG91]. coloured [CWG^{+}91]. Columbia [SB90, SI98]. comb [BSVC99]. combination [KMT^{+}95]. combined [EBG96]. Commercial [BTBP98, SLB96, Bae91, CA99, CTCN93, CCJ^{+}99, FVF94, FL94, Hor98, KK93, Ken96, Kly98, Knu91, KPH^{+}96, MP96b, MCC90b, PL92, Pea99, RP92, Sid91, SP94, TJE99a, TJE^{+}99b, TJ99, TC92, WBCG98, RM99]. commercial-length [WBCG98]. commercially [CCHW95, GHHEE98]. commerson [DAB92, Gov94, Gov95]. Commission [McI90]. Committee [McI96d, Pri96a]. Common [DLB91, ST96, KD98a, MNM97, PL98, SME95]. communication [Pri91]. communities [ADNT94, ST96, UYD^{+}94, VWL91, VS97, WBB^{+}97]. community [ACC96, ES99, GSSM99, GSM99, HB97, JAC^{+}99, KCS^{+}99, Kyl96, MS91, Piv92, RMR^{+}99, SAV92]. community-based [ACC96, SAV92]. Company [Mci96b, AAP96]. Comparative [BM93, CMK97, MMH93, Njo91, Was91, BPSW90, CDB^{+}97, DMA^{+}91, HG97, He96, KR96, Wal99]. compared [TG95]. comparing [WBCG98]. Comparison [AJ94, BACH99, BI97, Hor98, HR92, MKKO95, PMA95, Ped97, PPSvD95, RHLE99, SM98a, CAI96, Col90, CT92, FAC97, FBB98, KK95b, MBH99, PH91, PK99, SB98b, SB98c, She98, SP93, SC96, SCP97, TJE99a, TCA^{+}97, WF93]. comparisons [Dol98, RMR^{+}99]. competition [SPP96]. Complementary [USS93]. complex [Har95, JS98, Ste91]. complexity [Ste99]. component [KW98c]. composition [CLZS97, ET91, GPV90, HGC94, JT96, NE96, PM94, SD94, SPP96, Thr99, WBCG98]. compromise [WA97]. computed [Cri94]. Computers [Bro97b]. concentration [LBBB95, TJA92, TJA94]. concentrations [AJ90]. concept [Boo99, Sec99]. conceptual [Per92]. condition [CCV^{+}98, CA95, KT98, Mil99, SD94]. conditioning [FJM91]. conditions [Bru93, CM99a, Dah94, JP90, MN92, SRMdC95]. Conference [Hop92]. Congress [Ano95r, Kes96b]. conical [CT92]. Conrad [Sto90]. consequence [CB98]. consequences [AH93, BSVC99, BAR99, PWMM95]. conservation [Beg98, Bri92, EGR^{+}97, HLW96, Mci95c, Wil95, Luc97a]. Conservative [DCS99]. conserving [USS93]. consideration [Bux92]. considerations [Cav93, GK91, MDB92]. constrained [LP95a]. consumed [Fur94]. consumption [AA94, BPV95, JHB96, SG98]. Contents [Ano90e, Ano90f, Ano91e, Ano91f, Ano91g, Ano92d, Ano92e, Ano93f, Ano93g, Ano93h, Ano94g, Ano95e, Ano95b, Ano95c, Ano95d, Ano96d, Ano96e, Ano96f, Ano96g, Ano97c, Ano97d, Ano97e, ATJ91, HM95]. context [Boy90, She90b]. continent [RF99]. continental [CDE93, DBJ98, FFGG94, MMPS96, NLT99, VT96]. contingent [Sec99]. continuous [SL98, WS96]. contradictory [Ric91]. contrast [KS93, KW98a]. Contrasting [SK98, Ell93]. Contributions [MG99]. control [CW95, GWGR95, GW95, She98, YNEK93]. controlling [GP99]. conventional [MB95]. convergence [MH97]. convert [BI97]. converted [PMA95]. cooperatives [BAG98]. Copepoda [TW93]. coral [Aco94, AA95, ACC96, RHO99]. Corbiculidae [RU98]. Coregonus [AJ94, LPH^{+}98]. correlation [TJA94]. Correspondence [Ano98k]. Corrigendum [Ano91h, DK97]. Coryphaena [MMN97]. Coryphaenoides [PM98]. Cost [BPV95, KGM^{+}93]. cost-benefit [KGM^{+}93]. Cost-efficient [BPV95]. costs [Egg92]. Cottus [JAG^{+}99]. Counting [Col98]. countries [DMA^{+}91, Ven92, McI95a]. country [McI90]. covariation [WA97]. cove [DC98]. cover [All93, Kes96b, Kes96c, Kes97a, McI95e, McI95f, OK96, Pea93, Pri91]. Cowx [All93, Luc95]. CPUE [PRMH99, VSJ^{+}99]. crab [AB92, CSM95, DLB91, FBF90, PW94, SB90, Wen92, ZS97]. crabs [CD90, HW99, KMT^{+}95, PW94, SBK95]. Craig [Luc96, McV93]. Crangon [PR99]. crappie [MMH93]. crappies [TMD97]. Crassostrea [MM96, PKH^{+}95]. crayfish [GW99, HML97, McC90a]. Crean [McI97b]. creel [KMNP97, MK97]. Creighton [Bro97e]. Cretan [TMK92, VT96]. crisis [McI97b, Mci92]. critical [KS93]. critique [PJ96]. croaker [LMC98, Nor95]. Croatian [DK97, DK96]. Croix [GGBC92]. Croom [McC90a]. Cross [ELK99]. crucial [KW98c]. Crustacea [AM93, ES98, FF96, RB96]. Crustacea-Aristeidae [RB96]. Crustacean [Lai93]. ctenophores [MMPS96]. cultured [CGG99b, HJH^{+}93]. Current [Tho92b, FL93]. currents [BC93]. curve [Bas99, Bas00, Cri94, Mac97, TOSM96]. curves [EC98, Jen96, PMA95]. Cushing [McI91d]. cutlassfish [MH97]. cuttlefish [BD98, Dun99, GHHEE98, dSD97]. Cuvier [BKB97, Caz96, LP99, QAC^{+}98]. CXLV [Mil91]. cycle [Ell93, GFM99, TS92]. cycles [Gui98]. Cyclopterus [Hol93]. cynoglossus [Bow90, NGNH91]. Cynoscion [Caz96, HGC94]. cyprinid [Ama90, Luc92]. cyprinids [PPSvD95]. Cyprinodontidae [LS98, LS99]. Czechoslovakia [Kub92, Piv92].

D [All96, GM97, Hob91, Hop92, Luc93, Luc97a, McC90a, McI91d, McI95d, Nak93]. D. [GM97, VM97]. dactylopterus [SIE99, SIE99]. Daily [Gui98, Dul95, GB98a]. Dam [JSE92]. Damage [FBB98, SBK95, UYD^{+}94]. dams [SI98]. damselfish [DK95]. Dana [SB90]. Daniel [Lai93, McI96c]. Danish [DKH99]. dart [QK96]. Darussalam [Whi91]. data [AMON95, Bae91, BLQ91, Bas99, Bas00, BI97, BMB98, CD96, DMA^{+}91, Erz90, FK95, GNN^{+}97, Gov94, GGBC92, HPM94, Hor98, KK93, Kes97b, KPH^{+}96, LAG93, LLS97, LP95a, MP96a, MG97, MM98, PBHS94b, Pla93, RP92, Ric91, S.94, SBSW99, Sid91, SJ97, TF91, USS93, VS97, WS96]. data-limited [S.94]. David [McI97b]. Day [CB98, McI94, NW98]. Day-to-day [CB98]. daytime [TW92]. DC [Mci95c, Pri96a, Ste95]. debate [PW97]. decalcified [ZM92]. Decapoda [FF96, FF98]. December [Ano91o, Ano92l, Ano93n, Ano96v, Ano97t, Ano97o, Ano98q, Ano98w, Ano99l, JM91]. Decision [IH95a, IH95b, HPM94, Mci95c, Med98]. decline [MDMG^{+}93, MSWP93]. Declining [PKH^{+}95]. Dee [Smi90]. deep [ADNT94, Bae91, BTBP98, DBJ98, DLB91, LM98, RB96, TS92, dS91]. deep-sea [DLB91]. deep-water [Bae91, RB96, TS92]. defined [EC98]. del [Hop92, Kes97a]. delay [BHF99]. delineation [ME99]. delineator [Thr99]. Delta [Cav93, IO98, AM93]. DeLury [JP90]. demand [SM96]. Demersal [GFM94, Gob94, GBL^{+}97, KGM^{+}93, MS91, PM98, RMXP93, RMR^{+}99, SIE99]. Demographic [Cor98]. demography [MFCG98]. Dendo [Hop92]. dense [Tor91]. densities [LCU93, RHLE99]. Density [Gob98, HMP96, Lan99a, AH93, AJ94, HJ92, Hol96, PM98, SI98]. Density-dependent [Gob98, Lan99a]. Dentex [JDDK95, MNM97, JDDK95, MNM97]. Department [Kes96a, Whi91]. depend [AP96]. dependent [CLGK97, Gob98, Lan99a, OCU96]. depletion [HMP96, JP90]. deposition [VM97]. depression [Eml91]. Depth [CW95, Bow90, NW98, SEJ97]. Derek [Paw90]. derivation [SAB^{+}93]. derived [Hor98]. dermal [Rob93]. Describing [Ste91, CP94]. Description [Urb98, FGS92, Sha94]. desert [SM99]. design [BPSW90, BK96, FvMH96, HF90, KMNP97, LTT^{+}97, MBB95, PW97, MMM99]. designed [AS91, CW98, MKS99]. designs [BPV95, Col90]. details [AP96]. detecting [PK99]. Detection [ESNA95, LBK97, PW94, Tre98a, Tre98b]. determinants [RL90]. Determination [RP92, AGA98, CRLG95, RH90, VT96]. determine [BKH99, Cru97, KGM^{+}93, LCU93]. determined [CM97, He96, IKA90, KS93]. determining [TOSM96]. Deutsche [Hop92, McI95a]. developed [BMB98]. Developing [Ano95r, JSE92, McI95a, Ven92]. Development [CIA96, Dic93a, MBB95, RTBD95, Whi91, CY91, JDDK95, Lai95a, MNJ91, Rub94, Sal95]. developmental [Bas99, Bas00]. developments [AA94]. device [HM95, MMM99, MBB95, RD95, RT94, RTBD95, RM99]. Devices [BREB98, IASS96, KMNM96, RAT91, RRdS^{+}97]. Dfl [Ear91, Mil91, Sto90]. diamond [FBB98, FM92, HCF^{+}99a, HCF^{+}99b, SLB96, SPP97, WMCH92]. diamond-mesh [FBB98]. Dicentrarchus [Jen91, RP92]. Didymozoid [WMCW93]. Diego [Loc93]. Diel [FW95, NW95, TS92]. Diet [ACGMGS^{+}99, Du 96, HAN97b, RF97, BPD90, CDB^{+}97, DBJ98, LMTE97, RJR91, dSD97]. Diets [BMR^{+}90, GLGB94, PBHS94a]. Differences [Cio91, WN92, GNN^{+}97, GM97, PHG^{+}94, RN98]. different [Beg98, CM99b, HLK^{+}95, LTT^{+}97, Wen92]. differential [DP97]. differentiated [Pla93]. differentiation [Avs94]. digestion [LHS95]. dinoflagellates [MPMC97]. Diplodus [GM97, PS95a, SMEL98, VM97]. direct [SP93]. directed [Gui98]. direction [LP97b]. directions [Ros92]. Directive [HH93]. directly [Ped97]. discarded [KLB98, LBK97, SBK95, WH93]. discarding [CG97]. discards [LF99, TJE99a, TJE^{+}99b, TJ99]. discriminating [DRE97]. Discrimination [DVY95, MMV98, VMT97, MA98]. disease [FB93, FL93, HH93, dC93]. disease-induced [dC93]. Diseases [MSWP93, Buc93]. Dispersal [BF97, HLW96, Pea93]. dispersion [CSR99, HM92]. distance [HW99]. distinguish [CW93]. distributed [YA99]. Distribution [FVF94, MH97, PS99, PBHS94b, SRMdC95, TMK92, WMB92, AMIA98, BBBF91, Cav93, CLGK97, CCA97, ES99, FFGG94, Hol93, IZJT93, JHF93, JAG^{+}99, KPH^{+}96, TJA94, Yam98b]. distributions [EC98, Pow96, RHLE99]. disturbance [CSR99, KCS^{+}99]. Diurnal [GM91, KD98b]. diver [HGC97, HG97]. diver-survey [HGC97, HG97]. divergent [Sec99]. diverse [Sch98]. diversified [JAG^{+}99]. Diversity [McI96d, ES99, GSSM99, GSM99, JAC^{+}99, JT96, KCS^{+}99, Mci95c, RMR^{+}99, SFM91, SB97b]. diving [Loc97a]. Division [Aze95]. divisions [PLW99]. d’Ivoire [ES98]. DM [Luc90, Luc97a, Pri91]. DNA [CCV^{+}98, Dah96, GR91, GR94a, MMV98]. Do [NT96, SM98b]. dobsoni [SD90]. Does [CCV^{+}98, Lon99a, WA97]. dogfish [RSB^{+}98]. dollfusi [GHHEE98]. dolomieu [CM97]. dolphin [KK95b, MMN97]. dolphin-fish [MMN97]. dominance [FS94]. dominant [JT96, fN94]. Donald [Pat94]. Donax [GFM99]. d’Orb. [Cio91]. d’Orbigny [Iri91]. Dordrecht [All91]. Double [Lon99b]. down [JLMP96]. down-looking [JLMP96]. Downstream [SMT95, BHF99, Mor93]. drag [BD93, Jen91, WF93]. Drammenselv [Han90]. dredge [MYMP91]. dredged [HML97]. Dredging [DKH99]. drifting [ADNT94, GMHSR98, SEJ97]. driftnet [MS98, YDM95]. drum [GR91]. dual [KD98b]. dual-beam [KD98b]. dubius [MMV98]. due [RN98, SFM91]. duration [Cav93, CT92, GPV90, KS93, KGM^{+}93, WH93, WBCG98]. duration-in-beam [CT92]. Durham [McI94]. during [AKKD92, Bas99, Bas00, Bha94, Bub96, CM99a, GF92, GW92, GS92, Kub92, LR98, LVIG90, MSP97, MKKO95, SM96]. dwarf [FB97]. dynamic [DKH99, PS98]. Dynamics [Hop92, AM93, All90, CP98b, CCA97, Dun99, ELK96, ELK99, HQ91, HF91, MM96, McS96, Meg91, PS92, PvDM^{+}97b, PKH^{+}95, RSB^{+}98, RU98, RSM99, SD90, WZ99, Kes97a, Luc93].

early [Bas99, Bas00, Loc97b, Bla92]. East [BSA92, CDB^{+}97, HGHLC98, Ros92, AB92, Col98, ET91, Mon90, PL96, She98, SB97b, SK98, vdWB97, IZJT93, PvDM^{+}97a, PvDM^{+}97b, YB96, LP99]. eastern [DK95, DK96, DK97, EW96, EGR^{+}97, Fur94, Ism95b, LMC98, Mon90, MB96, PCD^{+}98, PPS93, PP94, Ste90, Ste91, WZ99, DPB91, SG98]. eblanae [GRG94]. Ebro [AM93]. EC [HH93]. Echinodermata [GM95]. Echinoidea [GM95]. echo [Eck98, GG98, GGBC92, JLMP96, LL95, RHLE99, Tor91]. echo-integration [GGBC92]. echo-signal [LL95]. echo-sounding [JLMP96]. echograms [LL95]. echosounder [KD98b]. echosounding [HMP96, JRM99]. Ecological [CR99, SB98a, USS93]. Ecology [Ano92c, HG98, dCB95, CS98a, EE95, EE96a, Han94, HU94, How93, IO98, Paw96, PP94, Paw90]. Economic [Mci97c, DC98, KK95b, SB96, Sha94]. economics [Njo91]. ecosystem [BSVC99, Hop92, JS98, MH97]. ecosystems [Pea93]. ectoparasite [BBF94]. Ecuador [PS92]. edible [AB92, AJ90]. Edited [Bro97b, Luc97b, McI97a]. Edition [Plu91, Pri96b]. Editor [All90, All91, Bru90, Ear93, Gol95, Luc90, McI95f, Mil91, Plu91, Wil90, All93, Ano92c, How93, Kes96b, Luc95, Luc96, Mci95c]. Editorial [Ano98l, Ano99e, McI95b, Ano90g, Ano90h, Ano91i, Ano91j, Ano92f, Ano92g, Ano92h, Ano93i, Ano93j, Ano93k, Ano94h, Ano94i, Ano94j, Ano95f, Ano95g, Ano95h, Ano96h, Ano96i, Ano96j, Ano96k, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j]. Editors [All96, Bea95, Hea96, Hob91, Hop92, Loc93, Loc97b, Luc97a, McC90a, McI91a, McI97b, Paw96, Pea93, War95, Whi91, Kes96c, Luc92, McI93, Wil95]. edn [All90, McI90]. education [Tho92b]. edule [BP96, CS98b, CSR99]. edulis [Dol98, DKH99]. edwardsii [TMW98]. Edwin [McV93]. EEC [HH93]. eel [Sve99, Wes98]. eelpout [KK97]. eels [Gra99]. Effect [GPV90, Kid94, MKLvD94, PR99, YDM95, GWGR95, GW95, GS92, GL91, LMF92a, LR96, NGH92, OK96, PG95, SHH98, WCMG91, WK90]. effecting [SM96]. Effective [IKA90, CT92, RX95]. Effectiveness [KMT^{+}95, MB96, WBB^{+}97]. Effects [BK96, BKB97, CD90, CfyCaL^{+}96, EPCPW96, FF98, GB98b, Gri99, HLK^{+}95, HJ92, LP97b, MG97, NW98, Otu90, QK96, RHO99, SEJ97, SÍW91, Aco94, AF99, BMR^{+}90, Car93, CR99, DP97, DKH99, ES99, GSSM99, GSM99, Hea96, Hol96, JAC^{+}99, JRM99, KCS^{+}99, Kyl96, LBBB95, LTT^{+}97, LB95, McS96, RMR^{+}99, VWL91]. efficiencies [MMH93]. Efficiency [FBAM99, AA95, Dic93a, Dic93b, FF96, FTHM96, JP90, LMMK92, LMF92b, MKLvD94, MMM99, MYMP91, MBB95, Njo91, PR96, RTBD95, RM99, RN98]. efficient [BPV95]. efficiently [EBG96]. effort [AMON95, All93, Bae91, Car93, CfyCaL^{+}96, DMA^{+}91, ES99, FDD99, FK95, GSSM99, GSM99, HAS98, Har95, HG97, JAC^{+}99, KCS^{+}99, LV98, Mil95, PT95, PSMSvD99, PvDM^{+}97a, Pol91, RLS92, RMR^{+}99, SAB^{+}93, YDM95]. egg [NLVF98]. Eggers [All96]. eggs [AF99, BD98, GNN^{+}97, Rub94, VN98]. eighteenth [KK93]. elasmobranch [Luc90]. electric [BC93, Fra96]. Electro [Pri91]. Electro-communication [Pri91]. electrofishing [FBAM99, LCU93]. electrophoresis [DVY95]. electrophoretic [Pla93]. Eledone [WMB92]. Elemental [Thr99]. elements [ELCM92, GF92]. Elliott [Mci95c]. Ellis [War95]. Elsevier [McI91a, Sto90]. Embryonic [JDDK95]. emerging [Ste99]. emigration [AMON95]. Empirical [CAB99, CP94, FvMH96, FDD99, PI93]. employed [Jen91]. employing [ME99]. enabling [Azu95]. encountering [SLW97]. Encrasicholina [MG99]. encrasicolus [Dul97, GNN^{+}97, KD90, Mil99, Tud99]. end [GFM94, LTT^{+}97, LR96, OK96, PCEP98, TLT^{+}98]. endangered [Luc97a]. Endocrine [Pic93]. Endocrine-induced [Pic93]. ends [SLB96, SPP97]. endurance [He91]. energetics [Bla92]. energy [Tor91]. England [Bla90, AB92, Lyo98, TJE^{+}99b, TJ99]. English [Dun99, JHF93]. Engraulidae [MG99]. Engraulis [Dul97, GNN^{+}97, KD90, LL96, Mil99, Tud99]. enhanced [TM93]. enhancing [fN94]. ensis [CTCN93]. entangling [Aco94, LMMK92, LMF92a, LMF92b]. entitlements [San93]. environment [Das99, PPS93, McI95a]. environmental [CM99a, LL96, Orł98, SRSR90, WTM92, WK93, Pri95]. environmentally [BEMW96]. environments [Wen92, McI91b]. eperlanus [PRMH99]. epidemics [BBF94]. epidemiological [NYK93]. epidemiology [Whi91]. Epinephelus [LVIG90]. epistemological [Ull98]. EPR [ESNA95]. equation [LLS97]. equatorial [LSP99]. equilibrium [Cad96]. equipped [RRdS^{+}97]. Erlend [Bro97b]. Erratum [Bas00, EE96a, HCF^{+}99a, IH95b]. error [CP98b, CM97, SL98]. erythrinus [PL98, PS96, SME95]. escape [ABR98, BKB97, GWGR95, GW95, LTS98, MMPHW99, SLB96, SLW97, TMW98, YF92]. escape-gaps [TMW98]. Escapement [GK91, GW92, AP96, Cri94, EPCPW96, RJS98]. escaping [CA95, FBB98, SEO96]. Eschborn [Hop92, McI95a]. escolar [LP95b]. Esk [She90a]. Esox [HM99, SD94]. estimate [BSVH99, FBF90, Fur94, KK95b, LS93]. estimated [BBBF91, HW99, KDD^{+}94]. estimates [CB98, CAB99, EE95, EE96a, Fle92, JB93c, KW91, LBK97, Loc97a, MG97, MYMP91, MJ93, PO98, PR99, SJ97]. Estimating [HGC94, Hol96, JRM99, LMF92a, McS94, S.94, AP96, BPV95, CL95, EBG96, FK95, KPH^{+}96, RKK90, RP99, WS96]. Estimation [Dic93a, Dic93b, Dul95, HH90, PRMH99, Sid91, AMIA98, AMON95, AJ94, Car93, Che96, CP98b, HG93, HGC97, HG97, HPM94, Jen96, PPSvD95, PWMM95, PBN96, SPC93, SP93, SD91, TGdB98, TF91, TF94]. estimator [SL98]. estimators [CT92]. estuaries [Pea99]. estuarine [BKI96, ELK96]. estuary [CCHW95, Gui98, KD97, LK96, NE96, CD90, Rob92]. Étang [GG98]. Ethmalosa [KO90]. Etolikon [LS98, LS99]. Euboikos [SPP96]. Eugerres [RSM99]. Eurasian [Kub92]. Europe [Luc97a]. European [Mci96b, Cor96, KD90, MdCP94, Tud99, VO98, Wes98]. euxinus [Ism95a, IB99, PG96]. Evaluating [LMNP98, BPvD^{+}92, CW98]. Evaluation [HC99, KMNP97, KK95b, MR93, MK97, MW93, Mts96, OKS98, RT94, RRdS^{+}97, WZ99, ABD94, Bub96, CR99, Coo93, CAB99, GHR99, HG97, MNJ91]. evaluations [RTBD95]. Evidence [BP98, Pla93, BPD90, DCPM93, HQ91, RTC99]. Evolution [McI95e, Gol95, Ano92c]. evolutionary [SB98a]. example [AP96, AMON95, BBC^{+}96, Cru97, HGC94]. examples [McS94]. excluder [RT94]. exclusion [Ken90]. Expandability [Li99]. Expansion [AAP96, RB93]. expectations [Mat92]. Experience [MH90, Bro97e, Cha98, HJH^{+}93]. Experiences [GHL98, HAN97a, Tho98]. experiment [AH93, MW93]. Experimental [GM95, BAR99, EKP93, NLVF98, Sal91, VN98, JB93b]. experiments [AS91, BM93, BPSW90, FBF90, FM92, JP90, Løk91, LMF92b, MMPL98, MMPHW99, NP99, OKS98, PCS96, Pri91, SLB96, SBK95, WH93]. Explanations [Ull98]. Exploitation [Han90, BMB97, CK94, Dun99, LS93, Luc93, MSR95, McS96, Nak98, PW97, Pur93, SU92, SG97, Luc92, Luc96, McC90a]. Exploited [All96, ELK99, HB97, HH90, RB93, SD90, SK98, Sve99]. exploiting [DP97]. explorations [CD96]. exponential [CD96]. extension [LR90]. external [FGS92, HM99, MW93]. eye [ELCM92, LC99].

F [Bro97d, Luc96, Wil90]. face [MG97]. facilities [PW94]. factor [MDMG^{+}93, SD94, TS92]. Factors [RD95, RR93, SM96, FF98, LL96, MRB96, fN94]. FAD [IKA90]. FADs [IASS96]. Fahe [ES98]. Failures [Tho92a, BSA92, Ven92, WTM92]. Fairbanks [All96, Tur97]. Falkland [ABB^{+}98, BRCK90]. Fam [Mil91]. familyprinting [LK99]. Fangliao [LL96]. far [WBCG98]. farmed [Bri93, PM93, TW93]. farming [McV93, Mun90, Sve91, Lai93]. Farnham [Bla90]. Faro [GFM99, SM98a]. Faroe [BTBP98, STM^{+}97]. Faroese [SÍW91]. fasciatus [LS98, LS99]. Fauna [CS97, GL91, KCS^{+}99]. faunas [RMR^{+}99]. February [Ano91l, Ano92o, Ano93l, Ano94k, Ano95i, Ano96p, Ano97p, Ano98p, Ano99f]. Fecundity [Ism95a, CQB^{+}94, GHHEE98, YAM98a]. fed [HLES93]. federal [Tho92a, SBM^{+}98]. feed [PC91]. Feeding [CS98a, LMTE97, PP94, RJR91, SGK98, VP93, Du 91, Du 95, FJM91, GM91, Iko96, JM90, Løk90, SH97, VO98]. females [LR98]. Fennoscandia [JAG^{+}99]. Ferguson [Bru90]. ferruginea [PL92]. fibulatum [Hus92]. field [AH93, EE96b, SBK95, JB93a, RTBD95]. filamentosus [MR97]. filleted [LVIG90]. fillets [Asp95]. filtered [EMS^{+}95]. fimbriata [KO90]. fin [BRCK90]. Fine [AF99, CA99]. finfish [SS93]. fingerprint [Dah96]. fingerprinting [DRE97]. Finlayson [Mci97c]. finned [Ark93, SH97]. Finnish [RL90, SM96, Sip98, VA96]. firms [MG92]. First [LAG95]. Firth [TCA^{+}97]. Fish [All96, Ano92c, Bas00, BBBF91, EE96a, FP99, Fra96, HCF^{+}99a, HF90, IH95b, Luc92, McI90, MMPS96, RLR92, RP99, SM97, ADNT94, AS91, AJ90, ATJ91, BW99, Bla92, BPV95, BBF94, BSVH99, BEMW96, BK94, BKI96, Cad97, CL95, CW98, Che96, CP98b, CA95, Col90, CP94, CB98, Das99, DK96, DK97, Eck98, EL97, ES99, ET91, ESNA95, EKP93, EGR^{+}97, EBG96, FAC97, FBB98, FBAM99, GB98a, GF92, GWGR95, GW95, Gob98, GW92, GBL^{+}97, GG98, GL91, Gri97, GGBC92, Gui98, HB97, Han93, HAN97a, HM95, HGC94, HJ92, HG98, Hus92, IKA90, IASS96, IVLK92, JT96, KS93, KMT^{+}95, KMNM96, KW98a, Kly98, KD98a, KD98b, KW98c, LF93, LF99, LTS98, LT92, LLS97, LM98, Løk91, LL95, Luc97a, LMA^{+}99, Lyo98, LHS95, MA98, Mac97, MSP97]. fish [MP96a, MMN97, MA97, McA93, McI93, McI96c, MS91, MSWP93, MR97, MG92, Mil96, ME99, MNSC96, MKS99, NW95, NGH92, Orł98, Otu90, Pau96, PL92, PO98, PK99, PPSvD95, PS95b, PS96, PS99, Pic93, Piv92, PM93, Pow96, PB93, PM98, RL90, RLS92, RAT91, RHO99, Rob92, RMR^{+}99, RHLE99, Sch91, TJA92, TJA94, TLT^{+}98, Tre98a, TF91, TSS97, UYD^{+}94, VS97, WN92, War91, WBB^{+}97, Xia97, Xia98a, YNEK93, Bru90, Hea96, She93, All90, Ear93, How96b, Mac93, Plu91]. fish-processing [MG92]. fish-releasing [IKA90]. fished [MMPL98, MHWMP99, PBHS94b, PB97]. Fisherian [Sch98]. Fisheries [Ano92c, Ano95e, Ano95r, Bis92, DK97, EE95, EE96a, FSSU98, Fra96, FP97, FB97, Kes96a, Kes96b, Kes99, Luc92, McA93, McI93, McI96e, Mci92, Pri95, Sym97, Tho92a, Whi91, Acq92, AAP96, All93, ACC96, Ano90e, Ano90f, Ano91f, Ano91g, Ano93g, Ano93h, Ano97e, BFP99, Beg98, Boo99, Boy90, BKI96, BSA92, Bro97b, CY91, Cha98, CLZS97, Chu98, Cor96, DMA^{+}91, DC97, DP97, FD97, FDD99, GH95, Gar97, GHR99, GMHSR98, GBL^{+}97, GAA99, Hoe98, HU94, HLW96, HJSNB98, Hug98, KT98, KK95a, Kes96d, LPSS98, LMNP98, LP95a, LTT^{+}97, Mat92, McI95e, McI97d, Mil95, MBT^{+}99, MBB95, MNJ91, MM98, Pat94, Pau96, PH91, Paw96, PJ96, Pea99, PVM^{+}96, PSMSvD99, PvDM^{+}97a, PvDM^{+}97b, PW97, PG94, Pol91, RTBD95, Ros92, SÍW91, SHH98, SBSW99, SAV92, Sid91]. fisheries [Ste96, Ste99, SP94, SC96, SPP96, SCP97, SB90, SBM^{+}98, Sum98, Tho92b, TS92, Ull98, Ven92, VS97, YB96, You98, Ano94g, Gar97, McI97d, Loc93, Bea95, Luc95, McI95d, Wil90, Ano91e, Ano92d, Ano92e, Ano93f, Ano95b, Ano95c, Ano95d, Ano96d, Ano96e, Ano96f, Ano96g, Ano97c, Ano97d, Bro97c, McI97b, Mci96b, Nak93]. fishermen [Ama90, SM98b]. fishers [MSP97]. Fishery [AM93, AA94, BP94b, IH95b, Kes97b, MFCG98, MH90, San93, She90b, AL95, AKB93, AJTP95, BBC^{+}96, BAG98, BRCK90, BMR^{+}90, BREB98, Bri92, BAR99, BK94, BK96, BPvD^{+}92, CCD^{+}99, Col98, DCS99, DC98, DKH99, ET91, EW96, Fle92, FM92, FS94, Fri95, Gob94, Gov95, GSR94, Gul98, HC99, HBB97, HGHLC98, Hur98, IH95a, KW91, KMNP97, LF99, LL96, LBK97, Lon99b, MMPL98, MMPHW99, MK97, McI95f, MS98, Med98, MCC90b, Mil93a, MNM97, Mts96, NMCCMLB99, fN94, Nor95, Okp91, PT95, PB92, PLL97, PSMSvD99, PWMM95, PBHS94b, PBN96, Por94, PdP93, QAC^{+}98, RMXP93, RD95, Rob92, RU98, SBR98, SB96, SB97a, SB98b, San96, SÍW91, SRG96, SD90, Ste90, Sto90, Tho92c, TMW98, Urb98, VSJ^{+}99, VCMCA97, WMCW93, WAB98, YDM95]. fishery [YB96, dCM94, dS91, dS94, Bla90, Kes96c]. fishery-effects [DKH99]. Fishes [Ano92c, Luc92, Arg92, BMR^{+}90, Fra96, GR91, GR94a, HG93, How96a, LHW98, LMA^{+}99, RKK90, RDG96, Thr99, VM97, Gat90, Wil95, How93, Luc90, Pri91, Pri95, Pri96b, Mil91]. Fishing [BAG98, Bla90, Bro97c, DLB91, ES99, FL94, GSSM99, GSM99, Hea96, JAC^{+}99, KCS^{+}99, Lai95a, Løk91, Mac92, McI97a, McI97b, Pur93, RMR^{+}99, ST96, She93, SBK95, All93, BPSW90, BC93, CM96, Car93, CA95, CfyCaL^{+}96, CS98b, CSR99, Fla98, GHR99, He96, HBB97, HM92, JRM99, Ken96, KW98a, KW98b, KPH^{+}96, LJ92, LB92, McI95e, MYMP91, MG92, Mil96, MR96, MDB92, MNJ91, PT95, RLS92, RHO99, SQM95, SEJ97, SLB96, SM97, SM98a, She98, SM96, Sip98, SFM91, SU93, SAB^{+}93, TJE99a, TJE^{+}99b, TJ99, dCB95, dSD97, Knu91, All91, Mci97c]. fishways [CSHW96]. fit [DCS99]. fitness [SB98a]. Fitting [CD96, EC98, LLS97]. five [McI96e, MRB96]. fixed [RJS98, She98]. fixed-location [RJS98]. fjord [Sal95, SB98a]. flat [LMA^{+}99]. flat-bed [LMA^{+}99]. Flatfish [Tur97, Bub96, NM99, PCEP98, SD91]. fleet [AAP96, GAA99, MNJ91, SAB^{+}93]. flexible [LPSS98]. flexuosa [PS96]. floating [CM96]. floodplain [HU94]. floodriver [EI98]. Florida [TR94]. flounder [Bow90, NGNH91, PL92]. flow [BP96, BD93, Smi90]. fluctuating [PPS93]. Fluctuations [Kub92]. fluvial [EMJ90]. fluviatilis [Kub92, KR96]. flying [YMSU97]. following [CSR99, KT98, Kyl94]. Food [Du 91, Du 95, Hea96, SH97, SG98, BPV95, Caz96, Hob91, IASS96, JHB96, KBKO99, MMPS96]. forage [Arg92]. forbesi [GR94b, MdCP94, PBHS94b, PBHS94a, PBHK94, PHG^{+}94, PM94, Por94]. forecasting [Ste91, SC96, SCP97, TC92]. form [SL98, War91]. forms [LPH^{+}98]. Forsskål [Was91]. forsteri [ELCM92]. found [WN92]. Foundation [She93]. four [CT92, LP97a, Lan99a, PS96, SLB96, SBSW99]. Fowler [MG99]. Fox [SL98]. Fox-form [SL98]. fractions [Bha94]. fragment [MMV98]. frame [MMH93]. France [BD98, GG98]. Francis [McI91b]. French [PMA95, Pla93]. frequencies [CA99]. frequency [BI97, EMS^{+}95, Erz90, KK93, MNA97, MS95, SJ97, TF91, WS96, YAM98a]. frequency-filtered [EMS^{+}95]. fresh [AKKD92, ES98, RU98]. fresh-water [ES98, RU98]. freshwater [All93, ATJ91, Bas99, Bas00, Bri93, FBAM99, Gra99, Hea96, How93, KS93, KW98c, LT92, Loc93, Luc95, Luc97a, Mts96, She93, McC90a]. Fridays [SM98b]. friendly [BEMW96]. fuegensis [SRMdC95]. full [WF93]. full-scale [WF93]. function [Bow90, Xia98b]. fungus [Hol94]. furnieri [LMC98, Nor95]. Furunculosis [MDMG^{+}93, NYK93]. future [Mat92, Nak98, Pur93, San93, VSJ^{+}99]. fuzzy [CW95]. fyke [Kub92]. fyke-net [Kub92].

G [All93, All96, Bea95, Han94, Luc95, McI94, Mil91, War95, Wil95]. gadoid [CP94]. Gadus [Asp95, CM99a, CM99b, DJ93, Dah96, Du 95, FJM91, FL94, He91, HLM91, HLES93, JM90, KK93, LHML97, MMPL98, NLVF98, OKS98, PC91, RJR91, RTC99, Sal91, SU92, Sal95, Sve95, YA99]. gahi [ABB^{+}98]. Gale [ES99]. Galician [FFGG94, GR94b]. gap [Cor96]. gaps [TMW98]. Garhwal [SS95]. Gari [Urb98]. Gary [Kes96b]. Gasterosteus [JLMP96]. gastric [HM95]. gastropod [EGR^{+}97]. Gastropoda [FD97, MSR95]. gayi [CAB99]. Gear [OCU96, SB98c, Bub96, CW95, Dic93a, Dic93b, FTHM96, Hol96, KW98a, KW98b, Mac92, SPP96]. Gear-dependent [OCU96]. gears [CA95, SBK95]. Gempylidae [LP95b]. gender [McI96c]. gene [BP96]. general [BLQ91]. generalized [Cru97, Das99, GAA99]. Genetic [GR91, SB97b, DE90, GS93, LMC98, RTC99, SBSW99, SFM91, USS93, Wil95]. genetically [Tud99]. genetics [Bea95, Gol95, McA93]. genotype [LK99]. genus [SMEL98, SK98]. Geoffrey [War91, IO98]. geometry [WF93]. George [Ear93, McI90, Ral96]. Georgia [MS91]. geostatistics [GGBC92]. Germany [Hop92, McI95a]. Gesellschaft [Hop92, McI95a]. Ghana [Acq92]. Gill [PS95a, PS96, BPvD^{+}92, HH90, KW91, KDD^{+}94, KR96, MHWMP99, PB97, RP92, SMEL98]. gill-net [RP92, SMEL98]. gillnet [CWG^{+}91, HGC94, PRMH99, PPSvD95, Sal91, SME95]. gillnets [AA95, ADNT94, HGN99, LMF92b, MKLvD94, Njo91, RP99, SEJ97]. gilt [KD97]. gilt-head [KD97]. Glasgow [Ear93]. Global [Mci95c]. Glubokoe [PB93]. Glyptocephalus [Bow90, NGNH91]. GmbH [Hop92]. goal [MNJ91]. goals [AP96, GK91]. goatfish [KBKO99]. gobio [JAG^{+}99]. Gobioidei [Mil91]. Gobioidei/ [Mil91]. golden [KBKO99]. gonadal [KK97]. Gonatus [DBJ98]. good [PI93, Sym97]. Gould [NW98]. gouldi [NW98]. governance [Sym97]. Grand [PL92]. Grande [MM96, RSM99, dCB95]. grandisquamis [NE96]. Grant [All96, Han94, Kes97a, Pea93, Tur97]. granulomatous [PM93]. Grate [RD95]. grates [HC99]. Gray [SS95]. Great [Arg92, NW95, WBCG98]. greater [MP97, SHH98]. Greece [KRS97, LS98, LS99, SP93, SPP96, MKKO95, PP94, TMK92, VP93, VT96]. Greek [FDD99, PS95a, PS95b, PS96, SP94]. green [AN95, KMNM96].

Greenland [DBJ98, HR92, HGN99, Jen90, RF99]. Greenstreet [McI97a]. grenadiers [PM98]. grey [MKKO95]. grid [IVLK92, TOSM96]. groenlandica [NGH92]. grooved [Sid91]. ground [CP98a]. groundfish [GSM99, Ken96, KPH^{+}96, Mil93a, RX95, RN98]. groundfishery [Cha98]. grounds [CA99, Ken96]. group [TG95]. grouper [Har95]. grouping [Erz90]. groups [DLT92]. Growth [CM99b, CGG99b, CGG99a, Dul97, ES98, GH95, GFM99, Gov94, GW99, Ism95b, MNA97, RB96, RB93, SJ97, dS91, AAP96, And98, Ark93, AGA98, BLQ91, Bas99, Bas00, CRLG95, CDE93, CL95, CMGB99, Cio91, Cru97, Dol98, DK95, Dul95, Dul98, Fra95, GM95, GL90, GM97, HQ91, HAN97b, HG93, HJ92, Iko96, JT92, KD90, KRS97, KBKO99, Kid94, KDCP96, KD97, LPP96, Lan99a, LPH^{+}98, LSP99, LLS97, LR90, LP99, MBH99, MP97, MG97, MJ93, MKKO95, MNSC96, MS95, NE96, NLT99, OM98, OB97, PH91, QK96, RvLV90, RH90, RR93, SS95, SG97, TF91, TF94, VT96, WTHC98, WS96, Was91, WT93, WA97, YMSU97, dS94, vdWB97]. grunt [AOH90, TR94]. GTZ [Hop92]. Guadalquivir [GL91]. Guadeloupe [FBAM99]. Guaíra [AMON95]. Guayas [CD90]. guide [MA98]. Guimarães [All91]. Gulf [Cio91, EL97, IH95b, SPP96, VP93, Bon97, GR94a, IH95a, KRS97, LFT91, MFCG98, Meg91, MNA97, NMCCMLB99, RD95, RTC99, WH97]. Gulland [All90]. gummy [WTHC98]. Gunderson [Pat94]. Gustaaf [Whi91]. guttatus [EE95, EE96a]. Gymnammodytes [DVY95, MMV98]. Gyrodactylus [BM93, JB93a, JB93b, MM93].

H [Ano92c, Bru90, McI91d, McI95a, McV93, Pri96b, Sto90, War91, Wil95]. Haan [AN95, Sid91]. Habitat [CP98a, EMJ90, GK91, HMP96]. habitats [JAG^{+}99, Loc93, WN92]. habits [Iko96, LMTE97, VP93]. Hackney [Hob91]. haddock [GS92, HCF^{+}99a, HCF^{+}99b, LR96, SLB96]. Haemoglobin [DJ93]. Haemulidae [AOH90]. Haemulon [TR94]. hairtail [YB96]. hake [CA99, CAB99, Du 96, GM91, HC99, NLT99, RF97, VO98, Paw96]. half [CT92]. half-angle [CT92]. halfbeak [LLS97]. Halibut [Kai94, DBJ98, HQ91, HGN99, KT98]. Haliotid [PdP93]. Haliotis [HGC97, HG97, McS94, McS96, SB96, SB97a, SB98b, SJ97, WA97, WAB98]. Haliporoides [OM98]. Hall [Ano92c, Bea95, Bla92, Bro97a, Bro97b, Ear93, Fra96, Hob91, How96b, Luc92, Luc96, Luc97b, Mac93, McI91c, McI96c, Nak93, Paw90, Paw96, Pri95, Gol95, McA93]. Hallegraeff [Whi91]. Ham [Bas00, Bas99]. Hamilton [JT92]. Hanabe [Ear91]. hand [SRG96]. hand-jig [SRG96]. handbook [McV93]. handline [San93]. Handling [BHF99]. Handling-induced [BHF99]. hanging [MKLvD94, SEJ97]. Hara [Mac93]. Hard [Pri91, All93, ET91, Kes96b, Lai93, McI95e, McI95f, Pea93]. Hardback [Bro97a, Bro97b, Bro97d, Cri93, Lai95a, Loc93, McI96d, McI97a, Paw96, Pri96b, Pur93, Tur97, McI95d, Mci96b]. hardcover [All90, Paw90]. harengus [DE90, GB98b, HLK^{+}95, Hol94, Hol96, JM91, Mis93, TGdB98, TG95]. harness [HM99]. harp [NGH92]. Harris [McI91c]. Hart [Luc97b]. harvest [CW98, FP97, IH95a, IH95b, Iri91, TC92]. harvested [WTHC98]. Harvesting [DC98, JS98, Li99, RB93]. hastatus [ELK96]. hatchery [Bra93, EE93, HJ92, Qui93, SMT95, WK93]. hatchery-produced [Qui93]. hatchery-reared [HJ92, SMT95]. haul [PCS96]. haumela [YB96]. Hawaii [HBB97, LMNP98]. Hawaii-based [HBB97]. Hawaiian [KK95b, FP97]. Hawkesbury [BK94]. hbk [McI97b]. head [KD97]. healing [WK90]. heavy [PO98]. Hefti [Luc97a]. Helicolenus [SIE99]. Helm [McC90a]. hematopoietic [YNEK93]. Hemirhamphus [LLS97]. Hepher [Gat90]. Hepsetia [LP99]. herded [LLW96]. Herding [RX95]. Herklots [ES98]. Herring [JM91, Zhe96, DE90, EMS^{+}95, FB97, GB98b, HLK^{+}95, Hol94, Hol96, LR98, Mis93, MB95, SHH98, SEO96, SLW97, Tor91, TGdB98, TG95, Tre98b, VCMCA97, ZM92]. herrings [PS92]. Hess [Gri99]. heterogeneity [CR99, GB98a, MS97]. Heterosis [Eml91]. hexamitosis [PM93]. Hickey [McI91a]. hierarchy [LMNP98]. high [AMON95, Bha94, Chu98, JP90, MS98]. high-resolution [JP90]. highly [PvDM^{+}97a]. Hilaire [IO98]. Hills [SS95]. Hilsa [DRE97]. Himalayan [SS95]. Hippoglossoides [BMB97, BMB98, MW93, PL92, DBJ98, HGN99]. hippurus [MMN97]. historical [How93, Ull98]. History [Fra96, BHS99, Bla92, Boo99, FB97, FL93, GR94b, MNM97, PL96, PBHK94, Pur93, RF99]. hoferi [Hol94, Hol96]. Holdich [McC90a]. holistic [BW99]. hololepidotus [BK94]. Holt [Hol95]. home [She90a]. homing [Qui93]. homogeneous [Tud99]. Hong [CTCN93]. hook [CAI96, HF90]. hooks [Kai94, LB95]. hoop [MMH93]. Horizontal [KW98c, CB98, GB98a, Gui98, Hug98, KDD^{+}94, SX91]. horizontal-directed [Gui98]. horse [KRS97, TF91]. Horwood [War95]. host [BBF94]. Houttuyn [CS97, PPS93]. hubbsi [DCPM93, NLT99, RF97, ST96]. human [PK99, RMR^{+}99]. hundred [Pau96]. hybrid [TMD97]. Hydroacoustic [JSE92, BACH99, Han93, Hug98, Lyo98, MKS99, PB92]. hydroacoustical [BBBF91]. Hydroacoustics [Mat92, PRMH99, RJS98]. hydrodynamic [RD95]. hydrographer [McI94]. hydrographic [SRMdC95]. hydrology [KK95a]. Hyperiid [Bro97d]. Hyperiidea [Bro97d]. hypotheses [CW93]. hypothesis [Sec99].

IA [Bru90]. Ian [Luc92]. Ibacus [SK98]. Iberian [dFB90, SRG96]. ice [JRM99, PB93, TSS97]. iced [AKKD92]. Icelandic [McI95e, EMJ90, JS98, SÍW91, Tho92c]. ICES [Aze95]. ichthyofauna [FVF94]. Ichthyophonus [Hol94, Hol96]. Ichthyophthirius [BBF94]. ICLARM [Hop92]. ideal [MN92]. identification [BW99, BFP99, CF99, ELCM92, JM90, LPSS98, LK99, LAA^{+}94, LL95, Mon90, PJ96, PvDM^{+}97b, RF99, SF99]. identifying [MR93, USS93]. identity [Vin91]. if [Lon98]. IHNV [YNEK93]. II [All96, Dic93b, GW95, GSSM99, HG97, JB93b, MMM99, MK97, RTBD95, RM99]. III [JAC^{+}99]. ilisha [DRE97]. ill [McI97b]. Illex [Ark93, BBC^{+}96, GRG94, SH97]. illicia [MBH99]. illumination [LBBB95]. illustrations [Pur93]. image [CF99, LL95, SD91]. IMARPE [Hop92]. immigration [AMON95]. immobilizing [How96a]. Impact [Gob92, Luc97b, Sip98, dC93]. impacts [Che96, CS98b, KK95b, PK99, RMR^{+}99, WTM92, WK93]. implementation [AP96, Mil93b]. Implications [Acq92, RSB^{+}98, All90, Bon97, GW92, PSMSvD99, WMCW93]. import [AA94]. Importance [IASS96, BAG98, GS93, LJ92, Sch91, Wal99]. important [CCHW95, Gri97]. improve [CP98b]. Improved [Gra99, WAB98, HM95, HF90, LTS98, MMM99, RM99]. Improving [AN97]. Incidental [MBT^{+}99]. including [LAG93, Whi91]. income [Ama90]. incorporated [PC91]. increase [BAR99, HML97]. increased [RHO99]. increases [RSB^{+}98]. increasing [LB95]. increment [VM97]. increments [Dul95, RH90]. incurred [SBK95]. Independencia [Urb98]. independent [Kes97b]. Index [Ano98m, Ano98n, Ano90c, Ano90d, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano93b, Ano93c, Ano93d, Ano93e, Ano94e, Ano94f, Ano94q, Ano95a, Ano95s, Ano96c, Ano96x, Ano97b, Ano97w]. India [Bas00, SS93, SS95, Bas99, JT92, SD90]. Indian [Hsu99, San93]. indicator [VN98]. indicators [BHS99, DC97]. indices [AN97, GS92, PG95, WAB98]. indirect [MYMP91, SP93]. individual [Cio91, Dol98, HG98, SB96, SB97a]. individual-based [HG98]. Indo [DAB92]. Indonesia [EW96, EGR^{+}97, HU94, MG99, PvDM^{+}97a, PvDM^{+}97b]. Indonesian [PSMSvD99]. induced [BHF99, HQ91, Pic93, dC93]. Indus [ATJ91]. industrial [Kes96d, Mil96]. industry [Sha94, McI95e]. inedible [LB95]. infection [Asp95, IB99, WMCW93]. infections [LHS95]. infectious [YNEK93]. infestation [TW93]. Influence [Nor95, FGS92, GB98a, LMF92b, PW97, Sal95, SRSR90]. information [LS93, PR96, Sch98, ZM92]. informative [GK91]. infrastructure [GS93]. initial [Fle92]. injected [PW94]. Injury [KT98, Kai94]. inland [DMA^{+}91]. inodorus [Gri97]. insect [Eck98]. inserted [Ken90]. inshore [RRdS^{+}97, RDG96, San96, YB96]. Institute [Mci97c]. Institutional [You98]. Instituto [Hop92]. integrated [CSHW96, LMA^{+}99]. integrating [Sch98]. integration [GGBC92]. integrator [Eck98]. integrity [JHF93]. intensity [ESNA95]. Inter [McI90, Cru97]. Inter-American [McI90]. inter-annual [Cru97]. inter-area [Cru97]. Interaction [PM93, CD90]. Interactions [HJH^{+}93, Bri93, CB98, KK95b, MH97, TM93, WK93, WH90b, YB96, Hop92]. intermidius [ELK99]. internal [FDD99]. internally [PW94]. International [All96, DN98, Hop92, McI95a, McI95f, Tur97, WH90a, Chu98, DPB91, Hoe98, Tho92a, You98, JAC^{+}99]. Interscience [All90]. interspecific [TMK92]. Intertidal [Iri91]. Interval [HG93]. interventions [CW98]. Intra [NLT99]. Intra-annual [NLT99]. introduced [GW99, OB97]. Introduction [QH91, Cri93]. invading [NGH92]. invasion [BSVC99]. inverse [FBF90]. invertebrate [Wil90]. Investigaciones [Kes97a]. Investigation [LLS97, MA97, FX96, NW95, SU93]. investigations [Ark93, JM90]. Iowa [Bru90]. Ireland [CK94, CMK97, MPDW97, MDMG^{+}93, SAB^{+}93, TW93]. iris [McS96]. Irish [SG98, ABR98, Bri92, Bri95, BAR99, TJE^{+}99b, TH95]. irradiated [ESNA95]. irrigation [dS94]. ISBN [All90, All91, All93, All96, Ano92c, Bea95, Bla90, Bla92, Bro97a, Bro97b, Bro97c, Bro97d, Bro97e, Bru90, Cri93, Ear91, Ear93, Fra96, Gat90, Gol95, Han94, Hea96, Hob91, Hop92, How93, How96b, Kes96b, Kes97a, Knu91, Lai93, Lai95a, Lai95b, Loc93, Loc97b, Luc90, Luc92, Luc93, Luc95, Luc96, Luc97a, Luc97b, Mac93, McA93, McC90a, McI90, McI91a, McI91b, McI91d, McI91c, McI93, McI94, McI95a, McI95e, McI95f, McI95d, McI96c, McI96d, McI97a, McI97b, McV93, Mci92, Mci95c, Mci96b, Mci97c, Mil91, Mil97, Nak93, Pat94, Paw90, Paw96, Pea93, Plu91, Pri91, Pri95, Pri96a, Pri96b, Pur93, Ral96, She93, Ste95, Sto90, Tur97, War91, War95, Whi91, Wil95, Wil90]. Island [ACC96, Mci95c, RU98, MMN97, FBAM99, KK95b, MS91]. Islands [CCD^{+}99, HGHLC98, MNM97, YAM98a, EE96b, ABB^{+}98, BRCK90, BMR^{+}90, FD97, Fri95, FB97, LP95b, LP99, PL96, PL98]. Isles [Vin91]. ISSN [Kes96c, Whi91]. issues [Ste99]. Istiophoridae [Fri95]. Italian [LAG93]. iterative [LLS97]. ITQs [Tow98]. iuy [OB97]. IV [GR91, RMR^{+}99]. ix [All93]. IXa [Aze95]. Izu [YAM98a].

J [All90, All96, Bea95, Bla90, Hop92, Knu91, Luc90, Luc92, Luc97b, Mac93, McI95a, Nak93, Paw96, Plu91, Pri96a, Pri96b, Wil95]. Jack [OM98]. Jack-knife [OM98]. jacobaeus [MP96b]. James [McI90, Mci92]. January [Ano90k, Ano91m, Ano92k, Ano93r, Ano95p, Ano96l, Ano97k, Ano99q]. Japan [Mil91, OM98, YAM98a, YNEK93]. Japanese [CSM95, MS98, NYK93, WC95, YAM98a, YDM95]. japonicus [CS97, FTHM96, LL96, PPS93, YAM98a]. Jasus [JB93c, MB96, TMW98]. Java [PvDM^{+}97a, PvDM^{+}97b]. Jay [Whi91]. jelly [BSVC99]. Jenkinsia [FB97]. jig [SRG96]. jigging [NW98, PM94]. Jizan [AOH90]. Jobling [Pri95]. John [Cri93, Lai93, Luc96, Mci97c, Pri96a]. Jolla [McI90]. José [Cio91]. Joseph [Luc92, McI90]. Jr [Loc93]. July [Ano90i, Ano95k, Ano95r, Ano96r, Ano97m, Ano99p]. June [Ano92j, Ano94l, Ano95n, Ano96n, Ano97s, Ano98t, Ano98-27, Ano99m, STM^{+}97]. Juvenile [MPDW97, BF97, BK94, BKB97, Coo93, Eck98, GM97, Jen91, Mon90, MNA97, MW93, NM99, Rob92, SZ98, Sve95, TW92]. juveniles [PC91, SRMdC95, VM97].

Kainji [Otu90]. Kamler [Bla92]. Karagiannakos [Mci96b]. Kariba [KK95a, Mts96, San96]. Kattegat [DE90]. Keenleyside [Ano92c]. Kenchington [McI91b]. Kesteven [Gar97, McI97d]. Keven [McI97b]. Kevin [Kes96c]. keys [Hor98]. kidney [FB93, FL93]. King [Bro97c, Gov94, Gov95, Mon90, PW94, ZS97]. Kingdom [BP96, Hug98]. Kirchhofer [Luc97a]. kisutch [Kyl94]. kitt [JHF93]. Klawe [McI90]. Klícava [Piv92]. Kluwer [All91]. knife [OM98]. knowledge [AN97]. Knox [Ral96]. Kong [CTCN93]. Korea [Par98]. Kousa [Ear91]. Kramer [Pri91]. Krasyukova [Mil91]. Kribi [GH95]. Krøyer [TW93]. Kruse [All96]. kuruma [FTHM96]. Kuwait [Sid91]. KwaZulu [CMGB99]. KwaZulu-Natal [CMGB99]. Kyushu [OM98].

L [How93, Loc93, McI95a, McI97a, Whi91]. L. [AB92, AW90, AGA98, Asp95, BM93, BD98, dFB90, Bow90, CNW92, CS98b, CSR99, CK94, CMK97, CWG^{+}91, DJ93, Dah96, Dol98, DKH99, Du 91, Du 95, DK95, Dul97, EMJ90, Ell93, FBF90, FJM91, GFM99, GNN^{+}97, Han90, HAN97a, HAN97b, HLK^{+}95, He91, HLM91, HLES93, HH90, HJ92, Hol93, Hol94, Hol96, Ism95a, Ism95b, JB93a, JB93b, Jen91, JM90, JAG^{+}99, KD90, KRS97, KK97, KDCP96, KD97, KR96, Lan99a, LV98, Lan99b, LHML97, LPH^{+}98, LCU93, MKLvD94, MMN97, MPDW97, Med98, Mil99, NGNH91, OKS98, PC91, PRMH99, PB97, RF99, RP92, RvLV90, RJR91, Rub94, SD94, Sal91, SU92, SGK98, She90a, SJ95, STM^{+}97, Smi90, SP93, Sve95, TF91, TF94, TW93, TH95, TKH94, VA96, VO98, VMT97, Vin91, Wes98, YA99, dSD97]. Labeo [AKKD92]. Laboratory [BM93, JM90, SBK95, AF99, Kid94, GWGR95]. Labrador [Bow90, BMB97, She98]. labrax [Jen91, RP92]. Lacepede [PP94, CM97, Was91]. lacustrine [DMA^{+}91, EMJ90, MPDW97]. lacustris [VA96]. Lagoon [PB97, BMR^{+}90, GG98, Gri99, SU93]. lagoons [LS98, LS99, PVM^{+}96]. Lagos [SU93]. Lake [Kyl94, Njo91, TJA92, Ama90, AJ94, SM99, Bis92, CS98a, FAC97, GGBC92, KK95a, MP96a, Mts96, fN94, Otu90, PS99, PBN96, PB93, San96]. lakes [ADNT94, CB98, HJ92, Kyl96, LT92, RL90, RLR92, RLS92, SM96, Sip98, UYD^{+}94, VA96, Arg92, Luc97b]. lalandii [JB93c]. Lamarck [CDB^{+}97, SG97]. Lampam [HU94]. lamprotaenia [FB97]. Landau [Cri93]. landing [AB92, GSSM99]. landings [MA97, Par98]. Landry [McI91a]. laniarius [BB97]. Lanka [Ama90, KW91, MNJ91, dS91, dS94]. Lankan [PPSvD95, PWMM95, SEJ97]. lantern [Hus92]. large [CM99a, FvMH96, HAS98, HGC97, HM99, KMNP97, LLW96, MK97, MDB92, NE96]. large-mesh [LLW96]. large-meshed [MDB92]. large-scale [HGC97]. large-scaled [NE96]. largemouth [SM99]. larvae [CCV^{+}98, Dul95, Dul97, Dul98, JM91, RJR91, SRMdC95]. Larval [Dul98, BBF94, CfyCaL^{+}96, Eck98, HLES93, JDDK95, LLW96, LL96, LT92, NLVF98, RDG96, RTC99]. Latreille [EE96a, EE95]. latro [FBF90]. lavage [HM95]. lavaretus [LPH^{+}98]. Lawrence [RTC99]. LC [Knu91]. LCA [CA99]. lead [MDB92]. learn [CW93]. Lee [Lai93]. Legal [Chu98]. legislation [Kes99]. lemon [JHF93]. Length [MJ93, PvDM^{+}97b, Aco94, Bas99, Bas00, BI97, CA99, Car93, CG97, CLGK97, DK96, DK97, Erz90, GPV90, GBL^{+}97, Gov94, Hor98, Hsu99, KK93, MKLvD94, MG97, MKKO95, MNA97, MS95, PMA95, PLL97, PS95b, SQM95, SJ97, TJA92, TF91, VO98, WS96, WBCG98, Xia98b, dSD97]. length-at-age [MG97]. Length-based [MJ93, PvDM^{+}97b]. length-cohort [PLL97]. length-converted [PMA95]. length-dependent [CLGK97]. length-frequency [BI97, Erz90, KK93, MNA97, MS95, SJ97, TF91, WS96]. length-structured [CG97]. leniusculus [GW99]. lentic [JAG^{+}99]. lentjan [Was91]. Lepeophtheirus [TW93]. Lepidorhombus [CDE93, LPP96, LP97a]. Lepidotrigla [CCA97]. Lepomis [KS93]. lepturus [MH97]. Leslie [FBF90, JP90]. lesser [DVY95, MMV98, PLW99]. Lethrinidae [Was91]. Lethrinus [Was91]. level [BSVH99, CWG^{+}91, Kyl94, Pol91]. levels [KGM^{+}93]. leydei [BSVC99]. Life [MNM97, PL96, BHS99, Bla92, Ell93, FB97, GR94b, Loc97b, PBHK94]. lift [Mts96, RDG96]. lift-net [Mts96, RDG96]. Light [CWG^{+}91, SD91, TS92]. likelihood [SPC93]. Limanda [PL92]. limit [Cad97, CCV^{+}98, LFT91, WA97]. limitations [Tre98a]. limited [S.94]. Limnothrissa [Mts96]. Lindberg [Mil91]. line [CAI96, SÍW91]. linear [Cru97, GAA99, PH91]. Linnaeus [Dun99, JDDK95, LMTE97, SME95, Wen92]. Linné [LBBB95, TGdB98]. Lions [MNA97]. list [Ano96w]. Lithognathus [BKH99, KDCP96]. Little [VLAM93]. littoral [PAL97, TG95]. live [HM95, How96a]. lived [MRB96]. liver [BN92, Bha94]. Living [Cha98, Hop92, McI95a, EKP93, Pea93]. Liza [MKKO95, NE96]. Lloyd [Hea96, She93]. lobster [Car93, Cru97, EE96b, JB93c, Med98, MR96, RB93, TMW98, TCA^{+}97]. lobsters [EE95, EE96a, MB96, SK98]. local [NGH92]. location [RJS98]. lochs [KDD^{+}94]. Lockwood [Bla90]. locus [Eml91]. logic [CW95]. logistic [CD96]. logs [CM96]. Loire [Rob92]. Loliginidae [GR94b, PBHK94]. Loligo [ABB^{+}98, CQB^{+}94, CDB^{+}97, GR94b, GSR94, MdCP94, PBHS94b, PBHS94a, PBHK94, PHG^{+}94, PM94, Por94]. London [Ano92c, Bea95, Bla92, Bro97b, Fra96, Gol95, Hob91, How96b, Luc92, Luc96, Luc97b, Mac93, McA93, McC90a, McI91c, Nak93, Paw90, Paw96, Plu91, Pri95, Pri96b, Wil95]. Long [HQ91, Ken96, Kly98, CM96, Dah94, McI94, Mor93, SÍW91, WBCG98]. long-line [SÍW91]. long-range [Mor93]. Long-term [HQ91, Ken96, Kly98, CM96, Dah94]. longfin [CS98a]. longianalis [Iko96]. longimanus [LSP99]. longirostris [LAG95, Lev96]. longline [HBB97, HR92, HF90, KT98, Ken96, Løk91, LB92, LB95, LP97b, Pol91, SIE99]. longlines [He96, HGN99, NT96]. looking [JLMP96]. Lophius [MBH99]. Loricariidae [AMON95]. Loss [SFM91, BKB97, He96]. losses [CW98]. lost [LV98]. lotic [JAG^{+}99]. Louisiana [RRdS^{+}97]. low [Bha94, BMB97, He91, MDB92, Tro97]. Lowe [GH95]. lower [WN92]. Lowery [McC90a]. lowland [GW99, KD98b, LMA^{+}99]. Loxechinus [GM95]. Ltd [Bla90, Bro97c, McI97a]. Lubuk [HU94]. Lucas [ACGMGS^{+}99, LAG95]. lucioperca [HAN97a, HAN97b, MKLvD94]. lucius [HM99, SD94]. lumpsucker [Hol93]. lumpus [Hol93]. lunar [Gri99, Otu90]. lutjanid [MR97].

M [All96, Bro97d, Ear91, Ear93, Knu91, Loc93, McC90a, McI91a, McI97a, Pur93, Ste95, War95, Whi91]. M. [GM91, VA96]. mackerel [Bla90, CNW92, CMGB99, CWG^{+}91, DAB92, EL97, Gov94, Gov95, JHB96, KRS97, Mis93, PPS93, TF91, YAM98a]. Maclean [McI95a, Whi91]. macleayi [AJTP95]. MacLennan [Nak93]. macracanthus [LC99]. Macrobrachium [ES98, FF96, FF98, Yam98b]. macrochirus [KS93]. macrophthalma [SP93]. Macroscopic [KK97, Hol94]. macrozooplankton [Kyl96]. maculatus [EL97]. made [Ama90, KK95a]. Madeira [MNSC96]. maderensis [GH95]. magister [SB90]. Magnuson [Pri96a, MH90]. Mahé [MR97]. mahsena [Was91]. mail [Hop92]. main [Kly98]. Maine [SMT95, RD95]. major [Bas99, Bas00, LF93, YDM95, CAI96]. Majorca [MMN97]. Makaira [ACGMGS^{+}99, Fri95]. making [Mci95c]. Malalison [ACC96]. Malangen [NGNH91]. Malaysia [AL95]. Malaysian [IKA90]. Malcolm [Pri95]. Mallorca [MNM97, QAC^{+}98]. Mallotus [GB98b, TGdB98]. Malmberg [BM93, JB93a, JB93b]. Maluku [EGR^{+}97]. mammals [GLGB94, MBT^{+}99]. man [Ama90, GL91, KK95a]. man-made [Ama90, KK95a]. manage [She98]. Management [All96, Ano95r, BB97, DAB92, Hop92, McI95a, SHH98, All91, ACC96, BBC^{+}96, BRCK90, BFP99, Bla90, Bon97, BSA92, BPvD^{+}92, Cad97, CP98a, Cha98, Chu98, Cor96, Cor98, DPB91, DP97, Egg92, EL97, EGR^{+}97, FSSU98, GHR99, Gul98, Hoe98, Kes96b, LMNP98, LAG93, MG97, McC90a, MH90, McI95e, McI97b, Mci96b, Meg91, MM98, Nak98, Paw90, PB92, Pun95, Pun97, SBR98, San93, SB96, SB97a, SB98b, SAV92, Ste96, Ste99, SBM^{+}98, Sum98, Sym97, Tow98, Urb98, Whi91, Wil90, WH90a, You98, All90, All93, Bro97c, Pur93]. Managing [SBSW99, BAG98, Col98, DN98, MR97, PS98, McI91b]. Manila [Hop92, McI95a]. manipulative [AH93]. Mann [Pea93]. MANOVA [ME99]. mantis [AM93, AM93]. manual [Kai94]. many [MM98]. Manzalah [Bis92]. Mar [Hop92]. Marasco [All96]. March [Ano93m, Ano95l, Ano96t, Ano97u, Ano98v, Ano99k]. Mare [HG98]. Margaret [Bro97e]. mariae [SJ97]. mariculture [Sal95]. Marine [CK94, Acq92, BSA92, Cad97, CR99, GHR99, GR91, GLGB94, GP99, Hob91, JS98, Kes97a, KMNP97, Loc93, Loc97b, MA98, MK97, McI91b, McI96d, Mci95c, MBT^{+}99, Nak98, PJ96, PB92, PG95, RR93, Sum98, VS97, Wil90]. marine-mammals [MBT^{+}99]. Marini [RF97]. marinus [MMV98, DVY95, PLW99]. mark [BLQ91, Fra95, HG97]. mark-recapture [BLQ91, Fra95, HG97]. marker [DJ93]. market [SS93, Tro98]. marketing [MG92]. markets [Paw96]. Marking [PC91, VSG99]. Markov [DPB91, FS94]. marks [VM97]. marlin [ACGMGS^{+}99, Fri95]. Marshall [Pri96b]. Marta [MM96, RSM99]. Maryland [Pea93, SRSR90, TC92]. Masfjorden [SU92]. masou [YNEK93]. mass [TSS97]. materials [CWG^{+}91]. matrix [PLL97]. Matthew [Cri93]. Mattson [McI90]. Maturation [GHHEE98, SMEL98, CM99b, CTCN93, CQB^{+}94, KK97, PM94]. maturity [CM99a, PHG^{+}94, SÍW91]. maxima [LBBB95, AW90]. maximum [HH90, SPC93]. maximus [BF97]. May [Ano90m, Ano93q, Ano95j, Ano96s, Ano97n, Ano98r, Ano98x, Ano99h].

Mayden [How93]. mazara [ACGMGS^{+}99]. McCoy [NW98]. McGoodwin [Mci92]. McLoughlin [Kes96c]. MD [McI93, Pea93]. means [PB93, SD91]. measure [WH93]. measured [Ped97]. measurement [FX96, JP90, WTM92]. measurements [MDB92, TJE^{+}99b, TJ99, WF93]. Measures [Pol91, RN98]. Measuring [KW98a, PR96, PG95, TJE^{+}99b, TJ99]. meat [KD90]. Mechanical [MN92]. mechanisms [Azu95, DN98]. media [Bha94]. Mediterranean [MNM97, PLL97, QAC^{+}98, AMIA98, AM93, CCD^{+}99, GAA99, GG98, GM97, KBKO99, LAG95, Lev96, MSR95, MMN97, MNA97, PP94, RB96, SZ98, Ste90, Ste91, TF91, Tud99, WMB92]. medium [Eck98]. medium-sized [Eck98]. Megrey [Bro97b]. Megrim [LPP96, CDE93, LP97a, Lan99a]. melampygus [HLW96]. melanura [PCD^{+}98]. membrane [LH90]. Memorial [Mci97c]. mercury [TJA92]. Merlangius [Ism95a, IB99, PG96, SGK98, SG98]. merlangus [Ism95a, IB99, PG96, SGK98, SG98]. Merluccius [CAB99, DCPM93, Du 96, GM91, HC99, MNA97, NLT99, RF97, ST96, VO98]. Mesh [NW95, ABR98, Bri92, BAR99, BK94, BK96, CA99, Car93, CNW92, Dah94, FBB98, FvMH96, FX96, FM92, GWGR95, HCF^{+}99a, HCF^{+}99b, HH90, LLW96, LH90, OK96, PCEP98, RHO99, SLB96, SPC93, SPP97, Tho92c, WMCH92]. meshed [KW91, MDB92]. meshes [BKB97, LTS98]. Mesolongi [LS98, LS99]. meta [MM98]. meta-analysis [MM98]. metacercariae [LHS95]. metal [AJ90, TJA94]. metals [ATJ91]. Metapenaeus [AJTP95, CTCN93, SD90]. method [ADNT94, BKH99, BSVH99, FK95, HLK^{+}95, IKA90, KMNP97, LTS98, LS93, LLS97, LP95a, LM98, MK97, NW98, Okp91, PK99, TOSM96, TGdB98, TMW98, WS96]. methodological [Cav93]. methodologies [Kes96b]. methodology [EC98]. methods [AP96, AJ94, BBBF91, BI97, Cha98, FX96, HGC97, HG97, HMP96, MBH99, MA98, NLVF98, Orł98, PJ96, PPSvD95, PG94, Sch98, SP93, SC96, SCP97, TJE99a, TCA^{+}97, dSD97, McI93]. Mexican [PdP93]. Mexico [Kes97a, ACGMGS^{+}99, Bon97, EL97, GR94a, HJSNB98, MFCG98, NMCCMLB99, VCMCA97, VS97]. Michael [Bro97c, McI91a, Ste95, Sto90]. Michigan [FAC97]. Michoacán [VS97]. micro [TF94]. micro-simplex [TF94]. Microbiology [Hob91]. Microchirus [And98]. microconstituent [SZ98]. Micropogonias [LMC98, Nor95]. Micropterus [CM97]. microscopic [Hol94]. Microstomus [JHF93]. microstructure [MNA97]. microtags [LPH^{+}98]. mid [WH97]. middle [DK95, ET91]. midwater [BD93, CW95, FAC97, SLW97]. Miers [SD90]. Migration [AW90, WT93, AN95, CB98, GB98a, GG98, Gui98, Hol96, LMA^{+}99, SU92, Wes98]. migrations [JHF93, Sec99]. migratory [Ark93, BBC^{+}96, SMT95]. Miles [Ano92c]. million [Pau96]. Mills [Paw90]. mimus [CGG99b, CGG99a, DC98]. minimise [SBK95]. minimum [AB92, LFT91]. Ministry [Whi91]. minke [SHH98]. Minor [Ama90, LR90]. minutus [PP94]. miodon [Mts96]. Mirna [KD97]. Mitochondrial [GR94a, GR91, MMV98]. mm [HCF^{+}99a, HCF^{+}99b]. Mnemiopsis [BSVC99]. mobile [Hug98, JRM99]. Mobility [SM99]. Mochokidae [San96]. Model [LPSS98, BLQ91, Bas99, Bas00, BSVC99, BPvD^{+}92, CY91, CW98, CG97, CP98b, CIA96, DCS99, DPB91, Dic93a, Dic93b, Eml91, Fle92, Fra95, GP99, HPM94, HG98, Ism95b, LCU93, LFT91, MP96a, Mil93a, PCS96, PLL97, PBN96, Pun95, S.94, SRG96, SB90]. model-based [MP96a]. Modeling [NMCCMLB99, PKH^{+}95, GAA99]. Modelling [CLGK97, KW98b, SC96, SCP97, YB96, dC93, CP98b, Cru97, FvMH96, O’N93, SB96, TJE^{+}99b, TJ99, TB98]. Models [CW93, BB97, Bux92, CD96, Das99, DLT92, DS99, GFM94, Gri97, HG93, Lon99a, PH91, Sch91, TF94, Xia97, Xia98a]. modified [BKI96, TLT^{+}98]. mojarra [RSM99]. Moksness [Bro97b]. Molecular [Bea95]. Molina [GM95]. Moller [Fra96]. Mollusca [CDB^{+}97, GFM99, MSR95, PM94, dSD97]. moluccensis [KBKO99]. monitored [KD98b]. monitoring [CSHW96, HGC97, Okp91, RJS98, SI98, War95]. monk [KK95b]. monkfish [Aze95, MBH99]. monodon [RB93]. monofilament [CWG^{+}91, Njo91, WCMG91]. Monogenea [MM93]. monogenean [JB93a, JB93b]. Monte [Che96]. monthly [SCP97]. months [Ste90]. Morbihan [BD98]. Moreton [RT94]. morhua [Asp95, CM99a, CM99b, DJ93, Dah96, Du 95, FJM91, FL94, He91, HLM91, HLES93, JM90, KK93, LHML97, MMPL98, NLVF98, OKS98, PC91, RJR91, RTC99, Sal91, SU92, Sal95, Sve95, YA99]. mormyrid [Iko96]. mormyrus [KDCP96]. Morone [MR93, TC92]. Morphological [Tud99, GNN^{+}97, MKKO95]. morphology [RP99, VN98]. Morphometric [CTCN93, CCHW95, PHG^{+}94, BMB98, CF99, DLT92, ME99]. morphometrical [MS97]. Morrison [AKB93, Ken90, RT94]. Mortality [Gov95, SEO96, VLAM93, BKH99, CD96, CAI96, CIA96, CAB99, DS99, DK95, Ell93, ES98, GH95, Ism95b, Jen96, Kai94, KT98, KDCP96, MP97, MG97, MYMP91, PI93, RB96, She98, Sid91, SP93, TKH94, dC93]. mossambica [Bha94]. mossambicus [dS91]. motorized [TSS97]. Mountain [Mil97]. mouth [Smi90]. Movement [HF91, BHF99, DPB91, PS99]. Movements [HLW96, Mon90, Mor93, SK98]. Moyle [McI93]. Muck [Hop92]. Mugil [OB97]. Mugilidae [NE96]. Mullen [Ste95]. Muller [Hea96, TGdB98]. mullet [ELCM92, LMTE97, NE96, OB97, VP93]. mullets [MKKO95]. Mullidae [Ste90]. mulloway [BK94]. Mullus [LMTE97, LAG93, PS95a, PS96, VP93]. Mulsow [Hol96]. multi [Eml91, LBK97]. multi-locus [Eml91]. multibeam [GHL98, GG98, Gui98]. multifilament [Njo91]. multifiliis [BBF94]. multifleet [LPSS98]. multigear [AL95, Gob94]. multilevel [TJE^{+}99b, TJ99]. multilocus [LK99]. multimesh [KDD^{+}94, KR96]. multinomial [PCS96]. multiple [ESW97, WS96]. Multiplicative [dFB90]. MultiSampler’ [ESW97]. multispecies [AL95, BPvD^{+}92, DCS99, Gob94, LPSS98, TB98]. multivariate [SC96, SCP97, VMT97, WN92]. MULTSPEC [TB98]. Murphy [McI90]. muscle [AJ90]. mussel [Dol98, DKH99]. mussels [DKH99]. Mustelus [WTHC98]. mykiss [Lan99b, LMF92a, Mor93]. Mytilus [Dol98, DKH99].

N [Bro97d, Nak93, Pri96b, TJE^{+}99b]. NAFO [CM99b]. Namibia [BKH99, GM91]. Namibian [MBH99]. Nardo [LS98, LS99]. NASCO [WH90a]. Natal [FVF94, CMGB99, Gov94, Gov95]. National [Kes97a, McI96d, Pri96a]. Nations [Hea96]. native [BM93, CMK97]. NATO [All91]. Natural [Pur93, SP93, BKH99, CAB99, Egg92, Ell93, Løk90, PI93, Sid91, SG97, SRSR90, TM93, WK93]. naturally [TMD97]. nauplii [TW93]. navigation [JP90, PO98]. nebulosus [HGC94]. necrosis [Rob93, YNEK93]. need [ACC96]. neglected [MMPS96]. neighbouring [SM98a]. Nelson [Luc92]. Nematode [IB99]. nematodes [HLES93]. Nematonurus [PM98]. Nematopalaemon [ELK96]. neogastropod [Kid94]. neon [YMSU97]. neopilchardus [Fle92]. Nephrops [ABR98, Bri92, Bri95, BAR99, Car93, FFGG94, MMPHW99, MS97, MS95, SPP97, SAB^{+}93, TCA^{+}97, TH95]. neritic [MMPS96]. nerka [Kyl96]. net [Aco94, AJTP95, BPvD^{+}92, BD93, CAI96, HH90, KW91, Ken90, Kub92, KR96, LLW96, LH90, LM98, LMF92a, MB95, Mts96, PAL97, PS95a, PS96, RP92, RDG96, SMEL98, Smi90]. nets [Aco94, AA95, Bri92, FTHM96, HW99, KDD^{+}94, KR96, LMMK92, LMF92b, MHWMP99, MMH93, O’N93, PB97, Sal91, SX91]. netting [Dah94]. Newfoundland [Mci97c, Bow90, BMB97, DBJ98, MCC90b, RTC99, She98]. Newfoundland-Labrador [Bow90]. newly [AS91, BMB98]. News [Bla90, Bro97c, Hea96, Lai95a, McI95e, McI97a, McI97b, Pur93, She93, All93]. Nexpa [VS97]. Nggela [FD97]. Niger [IO98]. Nigeria [IO98, ELK96, ELK99, EI98, Iko96, KO90, Njo91, Otu90, SU93]. Nigerian [NE96]. nightlight [RDG96]. nigricans [Fri95]. nilitocus [EGR^{+}97]. niloticus [FD97]. Niño [LLL98]. Niño/Southern [LLL98]. nipponense [Yam98b]. No [All96, Mci97c, Pri95, Tur97, McI94]. noise [Sch91]. non [CMK97, MA97, PH91]. non-linear [PH91]. non-native [CMK97]. non-target [MA97]. nonlinear [TF94]. Nonparametric [SPC93]. Nord [PG96]. Nordmøre [RD95]. Noroeste [Kes97a]. Norse [Mci95c]. North [Han94, How93, JH97, Tur97, CCV^{+}98, GM97, SÍW91, SG97, DE90, ES99, Fur94, GSSM99, GSM99, JAC^{+}99, Lon99b, LR96, MMPHW99, MHWMP99, MS98, MMV98, NGH92, PLW99, Pow96, She90a, TJ99, YMSU97, Zhe96]. north-central [SG97]. north-coast [SÍW91]. north-western [GM97]. northeast [BP94b, DBJ98, DVY95, ES99, GSSM99, GSM99, JAC^{+}99, KCS^{+}99, LPP96, MBT^{+}99, PBHS94a, RMR^{+}99, SMS99, Zhe96]. northeastern [BMB97, HF91, LC99, WC95]. Northern [CK94, CMK97, AW90, DC98, FP99, GRG94, HLM91, HM99, Hus92, JH97, JT92, JAG^{+}99, KD97, Lan99a, LHML97, Mci97c, NGNH91, SD94, SMS99, Sve95, VA96, WBCG98, WT93, YNEK93, Dul97, Dul98, KDCP96, RL90, SAB^{+}93]. Northwest [BMB98, Knu91, Bow90, Bri95, Ken96, CM99b, RSB^{+}98]. northwestern [AMIA98, AM93, GLGB94, MSR95, SRG96, KK95b]. norvegicus [Car93, FFGG94, MMPHW99, MS97, MS95, SPP97, TCA^{+}97, TH95]. Norway [Car93, FL94, Han90, HLM91, HH90, JB93a, LHML97, NGNH91, NGH92, SU92, Sve95, TCA^{+}97]. Norwegian [BM93, BF97, Bri93, HJH^{+}93, HLES93, Hol93, Sal95, SBR98, TG95]. Nostrand [McV93]. note [ME99, She90b, dSD97]. notes [ZM92, dS94]. notialis [DLB91]. Nototodarus [NW98]. Nova [HC99]. November [Ano91n, Ano95o, Ano97q, Ano98y, Ano99g]. NSW [WA97]. number [ME99]. numbers [PM98]. nurse [CP98a]. nursery [CP98a, RF97]. nutrient [Kyl94]. nutrition [How96b, Gat90]. nutritional [CCV^{+}98]. NW [FFGG94, GR94b, MNA97, PLL97, RTC99, TJE^{+}99b]. nylon [CWG^{+}91, WCMG91].

O [McI90]. Obituary [McI96a]. objects [GMHSR98]. Oblada [PCD^{+}98]. Observation [GG98, JHB96, SL98]. observation-error [SL98]. Observations [EI98, Hol93, BTBP98, EE95, EE96a, He96, KK97, SJ95]. Observed [Hol96]. observer [LBK97]. observer-based [LBK97]. obtained [RHLE99]. O’C [Lai93]. Occurrence [TG95, HLES93, KO90]. Ocean [IH95b, Ral96, GS93, Kes97a, DPB91, Hsu99, IH95a, MS98, PM98, San93, SZ98, YMSU97, Han94]. Oceanic [Kes97a, Azu95, BK96, KLB98, LSP99, SÍW91]. oceanography [McI91a]. oceans [Zhe96, Bro97d]. ocellatus [GR91]. October [Ano90l, Ano91p, Ano92i, Ano94n, Ano95m, Ano96u, Ano97l, Ano98u, Ano99j]. Octopodidae [CGG99b, CGG99a]. octopus [HGHLC98, Iri91, WMB92, CGG99b, CGG99a, DC98, HGHLC98, Iri91, QAC^{+}98]. off [AM93, ACGMGS^{+}99, Avs94, BMB97, CCV^{+}98, Du 91, ELK96, GH95, GFM99, GM91, Gov94, Gov95, HC99, HGHLC98, HR92, IB99, KO90, LC99, LP95b, MSR95, MMN97, MNSC96, NE96, OM98, PL96, PL98, PS92, RTC99, SH97, SM98a, She98, SRG96, SK98, TW93, VS97, WTHC98, WMCW93, vdWB97]. officials [PSMSvD99]. officinalis [BD98, Dun99, dSD97]. offshore [AKB93, CLGK97, Sve91, YB96]. oglinum [VCMCA97]. Oguta [Njo91]. oils [BN92]. Okhotsk [Mil91, JH97]. Okuda [Ear91]. Olhão [SM97]. Oman [DAB92]. Omani [SJ97]. Ommastrephes [YMSU97]. Ommastrephidae [GRG94, SH97]. OMV [YNEK93]. Onboard [TJE^{+}99b, TJE99a]. Oncorhynchus [Kyl94, Kyl96, Lan99b, LMF92a, Mor93, NT96, RJS98, YNEK93]. One [Pau96, HMP96, LCU93, Sal91, WF93]. one-tenth [WF93]. ontogenetic [LMTE97]. Oozeki [Loc97b]. open [JP90]. open-sea [JP90]. opening [ESW97, Med98]. operant [FJM91]. operating [RN98]. operational [FF98, Pol91]. Operations [All91]. Opisthonema [PS92, VCMCA97]. opportunities [SM96]. Optic [KS93]. Optimal [HAS98, MP96a, KGM^{+}93, MK97]. optimisation [LP95a]. Optimization [MDB92]. Optimum [Cri94, ABR98, HH90]. options [LMNP98]. orange [SB97b]. Oregon [WN92, McI91a]. Oreochromis [dS91]. organisms [IASS96, Pea93, Gol95]. Organization [Hea96]. organizational [Cad97]. oriental [Yam98b]. oriented [Pea99]. origin [RF99]. original [EMS^{+}95]. Oscillation [LLL98]. Osmerus [PRMH99]. Osteichthyes [EI98, IO98]. other [Otu90, PL92, PPSvD95, RL90, SÍW91]. Otolith [LAA^{+}94, SZ98, VSG99, Dul95, Gra99, LHML97, MBH99, MNA97, NLT99, PG96, VM97]. otoliths [ELCM92, GL90, LP97a, Ped97, RvLV90, Thr99, ZM92]. otter [GW95, HC99, HCF^{+}99a, HCF^{+}99b]. outbreeding [Eml91]. Outdoor [LMF92b]. output [Eck98, MPDW97]. overexploitation [GP99]. oversight [Bra93]. Overview [WK93, Ano94p, BFP99, BP94b, FSSU98, Mac92]. ownership [Sip98]. Oxford [All93, Bro97c, Hea96, Lai93, Luc95, McI95e, McI96d, McI97a, McI97b, Pur93, She93]. oxygen [LBBB95]. oxytetracycline [PC91]. oyster [MM96, PKH^{+}95].

P [Hop92, McI90, McI97a]. Pacifastacus [GW99]. Pacific [DPB91, Han94, IH95a, IH95b, MS98, Tur97, YMSU97, Coo93, DAB92, FL93, HQ91, HF91, JH97, KT98, Kly98, LLL98, MRB96, NT96, NMCCMLB99, OB97, PPS93, SZ98, SBSW99, Zhe96]. page [Pur93]. Pagellus [PL98, PS96, SME95, SIE99]. pages [Mci97c, Ano90n, Ano90k, Ano90i, Ano90m, Ano90l, Ano90j, Ano91q, Ano91o, Ano91l, Ano91m, Ano91n, Ano91p, Ano91k, Ano92n, Ano92l, Ano92o, Ano92k, Ano92j, Ano92i, Ano92m, Ano93o, Ano93p, Ano93n, Ano93l, Ano93r, Ano93m, Ano93q, Ano94m, Ano94o, Ano94k, Ano94l, Ano94n, Ano95i, Ano95p, Ano95n, Ano95k, Ano95l, Ano95j, Ano95o, Ano95m, Ano95q, Ano96o, Ano96m, Ano96v, Ano96p, Ano96l, Ano96n, Ano96r, Ano96t, Ano96s, Ano96u, Ano96q, Ano97r, Ano97t, Ano97o, Ano97p, Ano97k, Ano97s, Ano97m, Ano97u, Ano97n, Ano97q, Ano97l, Ano98z, Ano98s, Ano98q, Ano98w, Ano98p, Ano98t, Ano98-27, Ano98v, Ano98r, Ano98x, Ano98y, Ano98u, Ano98o, Ano99n, Ano99i, Ano99l, Ano99f, Ano99q, Ano99m, Ano99p, Ano99k, Ano99h, Ano99g, Ano99j, Ano99o]. Pagrus [CAI96, FW95, OCU96, PL96, QK96, WMCW93]. pagurus [AB92]. paid [McI90]. paintings [McI90]. pair [GFM94]. Pakistan [ATJ91, TJA92, TJA94]. Palaemonidae [ES98, FF98]. Paleolimnological [UYD^{+}94]. Pandalus [Tho92c]. pandora [PL98, SME95]. panel [Bri92, BK96, Ken90]. panels [ABR98, BK94]. panga [BB97, BP98]. Panulirus [EE95, EE96a, Med98]. Paper [Mci95c, Han94, Han94]. paperback [Bro97c, Bro97e, Mil97, Pri96b, McI96c, Mci97c]. paradigm [Cad96]. paradox [Lon99a]. paradoxus [GM91]. parahaemolyticus [SS93]. parameter [CP98b, MJ93, SJ97]. parameters [Avs95, BHS99, CL95, Che96, HG93, HLK^{+}95, LP95b, NGNH91, Otu90, PCD^{+}98, PR99, S.94, SU92, TF91, WS96, Xia98b]. parametric [PCS96]. Paraná [AMON95]. Parapenaeus [LAG95, Lev96]. parasita [CS97]. Parasite [ST96, LHML97, WMCW93]. Parasites [Bri93, HLM91, MA98]. Park [Pea93, SU93]. parr [JAG^{+}99]. Part [Mil91, JH97, PB97, YNEK93]. particular [She90a, Wes98]. pass [LMA^{+}99]. passages [Bro97e]. passive [CSHW96, LMA^{+}99]. Past [Nak98]. Pat [McI90]. Patagonian [DCPM93, Iri91, LF99, RF97, SRMdC95]. Patch [McS96]. pathogen [LF93].

Pathological [Bru93]. pathology [Bru90, Pic93, Plu91]. pathwidth [RX95]. pattern [Ell93, Iko96, MKKO95]. Patterns [VS97, AN95, CQB^{+}94, DC98, ES99, GM97, GSSM99, GSM99, HLW96, JAC^{+}99, KCS^{+}99, LPP96, MRB96, MMV98, RMR^{+}99, TMK92, Zhe96, Kes97a]. Paul [Luc97b]. Pauly [Hop92, McI96c]. Pautzke [All96]. Paz [Kes97a]. Pearcy [Han94]. Pecten [MP96b]. pectinid [MP96b]. pedigree [Dah96]. Pelagic [HJSNB98, ADNT94, AJ94, BSVC99, BACH99, BEMW96, CNW92, EPCPW96, FvMH96, FS94, GMHSR98, Han93, Hur98, JLMP96, MBT^{+}99, PS99, RMXP93, SIE99, SEO96, WF93, YA99]. pelagics [PT95]. pelagicus [Wen92]. penaeid [CCHW95, TW92, WT93]. Penaeus [AN95, FTHM96, Gri99, Mon90, Sid91]. penetration [GWGR95]. Peninsula [SRG96]. penned [EMS^{+}95]. Penobscot [SMT95]. pens [MB95]. Pentium [HG98]. Pentland [McI94]. People [Mci92]. Pepsin [LHS95]. per-recruit [BB97, Gov95, Gri97]. Perca [Kub92, KR96]. perception [PSMSvD99]. perceptions [SM98b]. Perch [IH95b, IH95a, Kub92, KR96, SRSR90]. percid [Luc93]. Perciformes [Mil91, dS94]. Perciformes/2 [Mil91]. Percoidei [AOH90]. performance [CMK97, Pun95, Pun97, RD95, RM99, San93, SB98c, YF92]. perhaps [Lon98]. period [Bas99, Bas00, CM99a, Kub92]. periodicity [Gri99]. periods [KK95b, Otu90]. Periophthalmidae [Mil91]. perpetual [Bra93]. perspective [CM96, GS93, Ste99, Ull98]. perspectives [Boo99, CD96, Tro98]. Peru [Urb98, Hop92, SQM95]. Peruvian [Hop92, MJ93, SQM95]. Peter [Fra96, McI93]. Peters [Bha94, San96, dS91]. pH [Bha94]. phalericus [Avs94, Avs95, Dul98]. pharaonis [GHHEE98]. phases [Kes96d]. phenetic [VMT97]. phenomenon [WTHC98]. phenotypic [SF99]. Philippine [FS94]. Philippines [Hop92, ACC96, BAG98, PT95, SAV92]. Phillip [Col98]. Phoca [NGH92]. photographing [How96a]. physical [Das99]. Physiology [Luc90, War91]. phytoplankton [PKH^{+}95]. picarel [Ism95b, TF91, TF94, VT96, VMT97]. pictus [AOH90]. pike [HM99, SD94, Luc96]. pikeperch [HAN97a, HAN97b, MKLvD94]. pilchards [Fle92]. pilchardus [CCV^{+}98, Dul95]. Pilica [PK99]. Pillay [Lai95a]. pilot [HLM91, MK97]. pintadilla [SQM95]. Pisces [And98, BB97, Caz96, CCA97, CS97, Fri95, LS98, LS99, LP95b, NGNH91, PL98, Pla93, San96, TMK92, VCMCA97, Was91, vdWB97, NE96]. Piscirickettsia [LF93]. PIT [CSHW96]. Pitcher [Bea95, Luc97b]. place [Cad96]. placement [WK90]. plaice [BMB97, BMB98, MW93, PL92, RvLV90, WMCH92, WZ99]. Planning [Mil95, Mil96, CY91, MG92]. plans [MNJ91]. plastic [FF96]. Plata [Nor95]. Plateau [MR97]. platessa [RvLV90]. platessoides [BMB97, BMB98, MW93, PL92]. Platform [EE96a, EE95]. Platyhelminthes [MM93]. playback [EMS^{+}95]. plebejus [Gri99, Mon90]. Plectorhynchus [AOH90]. Plehn [Hol96]. Plenum [Wil95]. Pleuroncodes [RB93]. Pleuronectes [RvLV90, WZ99]. Pleuronectidae [NGNH91]. plumieri [RSM99, TR94]. plurilineatus [CMGB99]. point [Cad97, Tho92a]. points [DC97, Gov95]. Poland [PK99]. polar [Ral96]. policies [IH95a, IH95b, Mci92]. policy [Acq92, PW97]. Polish [Hor98, PB97]. Political [Hoe98]. Pollachius [Du 91, Sal95]. Pollack [Sal95]. pollock [EPCPW96, HCF^{+}99a, HCF^{+}99b, JH97, Meg91, Par98]. pollutants [Hea96]. Pollution [She93]. Polymesoda [RU98]. Polynesia [Pla93]. polynomial [Bas99, Bas00]. Pomoxis [MMH93]. pond [Bro97a, Gat90]. poor [PP94]. Population [Avs95, CQB^{+}94, ELK96, LS99, MM96, MPVB98, MSR95, Meg91, MdCP94, NGNH91, OB97, PL98, RU98, RSM99, SU92, SD90, AM93, AMON95, All90, BMB97, CCA97, DE90, DKH99, ELK99, FF96, GR91, HGC94, HH90, Kes97a, LHML97, LC99, MP96b, MS97, MSWP93, PS92, PRMH99, PO98, PvDM^{+}97b, PBHS94b, PR99, PKH^{+}95, PM98, RB93, RTC99, SB98a, Sch91, Ste99, Sve99, Tud99, TH95, WZ99, WMCW93, dC93, dSD97]. populations [BP96, BSVH99, Bri95, CL95, CP98b, Cio91, CGG99a, DRE97, Ell93, FAC97, GS93, GM97, Gui98, KDD^{+}94, Lan99b, LMC98, McS96, MM98, PK99, PKH^{+}95, RTC99, USS93, VMT97, Wen92, All96, Luc93]. porgy [PL96]. Port [Col98]. Porter [Knu91]. portion [TC92]. Portugal [SME95, SM98a, And98, Car93, CQB^{+}94, CDB^{+}97, FP99, GFM99, GBL^{+}97, SM97, SMEL98, Sha94]. Portuguese [CDE93, MdCP94, dCM94]. Portunus [Wen92]. positioning [ABR98]. Possible [DLT92, MDMG^{+}93, MPMC97, SMS99]. post [CW98, SRMdC95, STM^{+}97]. post-harvest [CW98]. post-larvae [SRMdC95]. post-smolt [STM^{+}97]. postage [McI90]. Potential [CM97, Ama90, CCV^{+}98, CB98, LM98, Løk90, MR97, PBN96, Sal95, VSJ^{+}99, WH90b]. pots [FL94, TMW98, YF92, Yam98b, ZS97]. power [CP94, RHO99, Xia98b]. Poza [RU98]. pp [All90, All91, All93, All96, Ano92c, Bea95, Bla90, Bla92, Bro97a, Bro97b, Bro97c, Bro97d, Bro97e, Bru90, Cri93, Ear91, Ear93, Fra96, Gat90, Han94, Hea96, Hob91, Hop92, How93, How96b, Kes96a, Kes96b, Kes96c, Kes97a, Knu91, Lai93, Lai95a, Lai95b, Loc93, Loc97b, Luc90, Luc92, Luc93, Luc95, Luc96, Luc97a, Luc97b, Mac93, McA93, McC90a, McI90, McI91a, McI91b, McI91d, McI91c, McI93, McI94, McI95a, McI95e, McI95f, McI95d, McI96c, McI96d, McI97a, McI97b, Mci92, Mci95c, Mci96b, Mil91, Mil97, Nak93, Pat94, Paw90, Paw96, Pea93, Plu91, Pri91, Pri95, Pur93, Ral96, She93, Ste95, Sto90, Tur97, War91, War95, Whi91, Wil90]. practical [EE96b, Gul98, Xia97]. practices [CCJ^{+}99, EGR^{+}97, WK93]. Prager [Sto90]. prawn [AN95, AKB93, Bae91, BREB98, BK94, BKI96, BK96, ELK96, FBB98, FVF94, FTHM96, KLB98, LK96, MBB95, RT94, RTBD95, Sid91, SD90, TS92, WK90, WBCG98, YF92, Yam98b]. prawn-trawl [BK94, BKI96, BK96]. prawns [AJTP95, Mon90, WK90]. pre [Ark93, EC98]. pre-defined [EC98]. pre-spawning [Ark93]. precautionary [DS99, Gar97, Kes96e, McI97d]. Precision [Loc97a, CAB99, Hol94, RvLV90]. predation [MMPS96]. predator [Fla98]. Predators [BMR^{+}90, McI97a]. Predictability [MCC90b]. Predictable [Mis92]. Prediction [Ste90, Par98]. predictions [PI93, Ull98]. predictive [CP94, DLT92, TS92]. Predominance [DBJ98]. Preface [BP94a]. preference [FJM91]. Preliminary [AS91, GRG94, SJ95, ZM92, NW95, RJR91]. preparation [Gra99]. presbyter [LP99]. presence [GB98b, WS96]. present [LHS95, Nak98, Por94]. presoaked [LJ92]. Press [Bro97e, Bru90, Gat90, Han94, How93, Knu91, Loc93, McI91d, McI94, McI95d, McI96d, Mci92, Mci95c, Mil97, Pri96a, Ral96, Ste95, War91, Wil95]. pressure [SFM91]. prevalence [Hol96]. previous [LHS95]. prey [FAC97, Fla98, McI97a]. Priacanthus [LC99]. price [All96, Bea95, Ear91, McA93, McI97a, Mil91, Pri96a, Pri96b]. Priede [War95]. primary [PBN96]. Prince [Coo93]. Principles [Lai95b]. prior [GK91]. priorities [Cha98]. Pristipomoides [MR97]. Pritcher [Paw96]. probabilities [GK91]. probability [BSVH99, HW99]. problems [Kes96d, Mci92, Ros92, Ull98, WN92, Xia97]. procedure [BMB98, Hol94, MG92, Pun95, TF94, SHH98]. procedures [Hol94, MA98]. Proceedings [Hop92, Kes96b, All96, Tur97]. process [KK97, LMNP98]. processes [JB93a, JB93b, Kes97a]. processing [CF99, Ear93, LL95, MG92, PW94]. produced [Qui93, TMD97]. Production [TW93, Eml91, FP99, HJ92, Kly98, Kyl94, Lon99a, PT95, PW97, PBN96, Pun95, Rob92, SB90, SG97, SL98, Xia97, Xia98a]. production-model [Pun95]. Productivity [SB98a]. products [Løk91, Hob91]. Prof. [Gar97, McI97d]. Professional [Ear93, Sip98]. Program [All96, Han94, Egg92, Tur97]. programming [MNJ91]. programs [Bra93, Sto90]. Progress [Lai95a]. project [TB98]. Proliferative [FB93]. prometheus [LP95b]. Promethichthys [LP95b]. propagated [EE93]. propagation [NYK93]. proper [AW90]. properties [FS94, LR98, MN92]. Property [Beg98, Tow98]. proposed [KK97]. prospects [Lai95a]. Protected [Sum98]. protecting [DN98]. protein [Bha94, DVY95]. provide [Fle92]. provident [McI91d]. Province [EGR^{+}97]. provision [BRCK90]. Psammobiidae [Urb98]. Psenopsis [WC95]. Psetta [AW90, LBBB95]. Pterogymnus [BB97]. Publication [Ano94p, Pea93]. Publications [Luc95, Lip99]. Published [Hea96]. Publisher [Ano97v]. Publishers [All91, Bro97d, Ear91, Kes96b, McI95f]. Publishing [Mci96b, Sto90]. Puerto [AA95, RAT91, TR94]. Pullin [McI95a]. pulsed [BC93]. punctifer [MG99]. puntazzo [VM97]. Purdom [McA93]. purpose [MA97]. purpuratus [MJ93, SG97]. purse [Fle92, MN92, MDB92, Mis92, Mis93]. pyrodinium [Whi91].

Q [Pri96b]. Qatar [AAP96]. Qld [Ano95r]. quadrituberculatus [WZ99]. Quality [LVIG90, HLK^{+}95, RL90, RLR92, RLS92, RKK90]. Quantification [BD93]. Quantitative [Bub96, SS93]. quantity [CSM95, Fur94]. queen [CMGB99]. Queensland [RT94, Wen92]. Quinn [All96]. quota [San93, SB96, SB97a]. quotas [Mil93a].

R [All96, Bea95, Gol95, Hea96, Hob91, Lai95a, Lai95b, McC90a, McI91a, McI95a, McI97a, Mci92, Pat94, Plu91, LS93]. rabbitfish [WH97]. radio [CM97, HM99, SMT95]. radio-tagged [SMT95]. radio-telemetry [CM97]. radiographs [Ped97]. Rafinesque [MMH93]. Ragnar [McI95e]. rainbow [Lan99b, LMF92a, Mor93]. ramada [MKKO95]. ranched [CMK97]. ranching [GS93, OKS98]. range [IKA90, Mor93]. Ranina [HW99, SBK95]. Rao [Bas00, Bas99]. RAPD [DRE97]. Rastrineobola [PBN96]. Rate [Løk90, Aco94, Ark93, Cio91, GAA99, GM91, Kid94, LS93, WTHC98, WK90, WAB98, dS91]. rates [Aze95, Cru97, DLT92, ET91, FW95, LPH^{+}98, LP97b, NW98, SEJ97, WBCG98, dS94]. Ratio [Jen96, HG97, KBKO99, MKLvD94, SEJ97, VM97, dSD97]. Rational [JS98]. ratios [CCV^{+}98]. Ravi [TJA94]. Rawal [TJA92]. Rayleigh [Yam98b]. reaction [CWG^{+}91, GW95]. Reactions [EMS^{+}95]. Reading [Ano96w, Dul95, LAA^{+}94]. real [KD98a]. realistic [CP98b]. Rear [McI94]. reared [HJ92, SMT95]. rebuilding [BP98]. recapture [BLQ91, Fra95, HG97, Sid91]. reciprocal [CA99]. Recksiek [Sto90]. Reconstructing [CA99]. Records [Mor93, STM^{+}97, GHR99]. recreational [Fri95, GHR99, KMNP97, MK97, SM96]. recruit [BB97, Bux92, CG97, Cri94, Gov95, Gri97, Ism95b, LFT91, Per92, RB96, S.94]. Recruitment [MRB96, Das99, ES98, FK95, GB98b, MPDW97, Pow96, S.94, WS96, Zhe96]. rectangular [MN92]. red [Bae91, CCD^{+}99, DLB91, GR91, GR94a, LMTE97, MPVB98, PL96, PW94, SP93, VP93, Whi91, ZS97, AOH90, Was91]. redfish [CLGK97]. reduce [BKI96, CW98, RN98]. reduced [PAL97]. Reducing [BK94, MM98, BK96, MA97]. Reduction [BREB98, Dah94, IVLK92, HC99, PdP93, RD95, RRdS^{+}97]. Reef [NW95, WBCG98, Aco94, Bux92, Gob92, RAT91, RHO99, SM97, SM98a]. reefs [AA95, ACC96, LM98, PW97]. referees [Ano96a]. reference [Azu95, Cad97, DC97, GLGB94, Gov95, She90a, Wes98]. reflection [CM96]. Regime [Cad96, Ste96]. regimes [DP97]. region [AMON95]. regional [PHG^{+}94, Tho92a]. registered [GSSM99]. regression [SC96, SCP97]. regressions [Mac97]. Regulatory [JB93a, JB93b]. rehabilitation [SB98b, Luc95]. Reinhardtius [DBJ98, HGN99]. Reinhold [McV93]. related [FS94, SRMdC95, Ste99]. Relating [Das99]. relation [BBF94, CSM95, FvMH96, LL96, MS97, MSWP93, SD94, TW93]. Relationship [TJA92, GB98b, HM92, Hsu99, KK95a, LS93, LLL98, PM98, She90a, Smi90, Xia98b]. relationships [CTCN93, CP94, DK96, DK97, GBL^{+}97, KD98a, PS95b, SQM95, Zhe96]. Relative [HGN99, Hol94, KGM^{+}93]. release [KT98, Løk90]. released [VLAM93]. releases [EE93]. releasing [IKA90]. relevance [HH93]. reliable [DJ93]. remarks [Per92]. Remote [War95]. remotely [ESW97]. removal [BSVH99, Kai94, LCU93, TSS97]. removing [HM95]. Renibacterium [FL93]. replication [KGM^{+}93]. reply [Gar97, McI97d]. Report [All96, McI96d, Kes96a, Tur97]. reports [Kes96c]. representation [WBCG98]. Reproduction [OM98, GH95, GHHEE98, GW99, OB97, SQM95, WA97, dS94]. Reproductive [LS98, MMN97, Mil99, WC95, CTCN93, GFM99, Hus92, Iko96, LR98, MG99]. requirements [Cad97]. Res [Bas00, EE96a, HCF^{+}99a, IH95b]. Resazurin [RKK90]. Research [Ano95e, DK97, McI96e, Mci97c, All91, Ano90e, Ano90f, Ano91f, Ano91g, Ano93g, Ano93h, Ano97e, Gul98, Hor98, Kes99, Mat92, Pau96, PS98, She90b, Ano94g, Gar97, McI97d, Bro97b, Ral96, Ano91e, Ano92d, Ano92e, Ano93f, Ano95b, Ano95c, Ano95d, Ano96d, Ano96e, Ano96f, Ano96g, Ano97c, Ano97d]. reserves [CR99, GP99, Sum98]. Reservoir [AL95, Kub92, Piv92, TMD97, ES98, PPSvD95, PWMM95, dS91]. reservoirs [ATJ91, DLT92, KW98c, dS94]. residuals [Mac97]. resolution [JP90]. Resource [Kes96c, ACC96, Gob92, GW92, MMPS96]. Resources [Ano95r, Hop92, McI95a, AA94, Ama90, BAG98, Bon97, DN98, DP97, EE96b, Nak98, fN94, Loc97b, McI95f, Pat94]. response [CAI96, HW99]. responses [Kyl94, LLW96, ZS97]. restriction [MMV98]. Results [AP96, NLVF98, RJR91, VN98]. Retained [KLB98, LBK97]. retention [Ken90, MW93]. retuning [SHH98]. return [She90a]. returning [YNEK93]. Reupell [ELK99]. reveals [LHS95]. reversibility [Li99]. Review [All90, All91, All93, All96, Ano92c, Bea95, Bla90, Bla92, Bro97a, Bro97b, Bro97c, Bro97d, Bro97e, Bru90, Cri93, Ear91, Ear93, Fra96, Gat90, Gol95, Han94, Hea96, Hob91, Hop92, How93, How96b, Kes96b, Kes96c, Kes97a, Knu91, Lai93, Lai95a, Lai95b, Loc93, Loc97b, Luc90, Luc92, Luc93, Luc95, Luc96, Luc97a, Luc97b, Mac93, McA93, McC90a, McI90, McI91a, McI91b, McI91d, McI91c, McI93, McI94, McI95a, McI95e, McI95f, McI95d, McI96c, McI96d, McI97a, McI97b, McV93, Mci92, Mci95c, Mci96b, Mci97c, Mil91, Mil97, Nak93, Pat94, Paw90, Paw96, Pea93, Plu91, Pri91, Pri95, Pri96a, Pri96b, Pur93, Ral96, RJS98, She93, Ste95, Sto90, Tho98, Tur97, War91, War95, Whi91, Wil95, Wil90, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h]. review [Ano98i, Ano98j, Ano99a, Ano99b, Ano99c, Ano99d, Bri93, CA95, HJH^{+}93, Hur98, LB92, PS92, PW97, Qui93, Ros92, RTC99, TB98, WH90b]. revised [PLW99, SHH98]. revisited [Boo99, Hol95]. revolution [Kes96d]. Rhinelepis [AMON95]. rhizophorae [MM96]. Rhizoprionodon [MFCG98]. Richard [How93, McI91b, She93]. richardsonii [SS95]. Ricker [Cri94, S.94]. Rico [AA95, RAT91, TR94]. rig [CBB97]. rigged [SB98c]. rights [Beg98, SAV92, Sip98, Tow98, dCB95]. rigid [IVLK92]. Rimov [Kub92]. rings [GM95]. Rio [Nor95]. RISSO [Avs95, Avs94, CRLG95, CCD^{+}99, CDE93, LP97a, SME95, TS92]. Ritchie [McI94]. Rites [Bro97e]. River [ATJ91, ELK99, Iko96, PK99, SMT95, SRSR90, TJA94, VS97, VLAM93, WN92, dCB95, BSVH99, FF96, GW99, HML97, KD98b, LMA^{+}99, She90a, Smi90, Yam98b, AMON95, BK94, CK94, CMK97, FP99, GL91, Han90, HU94, Hug98, IO98, LK96, Lyo98, SI98]. riverine [KD98a]. rivers [BHF99, EE93, NYK93, PMA95, PAL97]. RNA [CCV^{+}98]. RNA/DNA [CCV^{+}98]. roach [BBF94, HMP96, KR96, PRMH99]. Robert [Lai95b]. Roberts [Plu91]. Robinson [McV93]. rock [Cru97, JB93c, LVIG90, MB96, TMW98]. rockfish [PCEP98, SIE99]. Rodrigues [All91]. Roger [Pea93]. rohita [AKKD92]. role [BHS99, BFP99, EL97, GMHSR98, Har95, WH90a, dS91]. rose [LAG95]. Rosenfield [Pea93]. Rosenthal [McI95a]. Rotterdam [Ear91, Loc97b, Mil91]. roudi [LP95b]. rough [LS93]. roughy [SB97b]. route [KMNP97, MK97]. ROV [NM99]. royal [Bae91]. rubra [HGC97, HG97, SB96, SB97a, SB98b, WA97, WAB98]. Rudolphi [Asp95]. rule [She98]. rules [O’N93]. Rutilus [BBF94, HMP96, KR96, PRMH99].

S [Bro97e, Ear91, How96b, Luc92, McC90a, McI94, McI95a, McI97a, McV93, Pri96b, Urb98, War95, LS93, LLS97]. S. [GHHEE98]. sablefish [HF91, KT98, LFT91]. sac [JDDK95]. saddlead [PCD^{+}98]. Safe [CCJ^{+}99]. sagax [Fle92]. Saharan [CDB^{+}97]. Saila [Sto90]. Sainte [GGBC92]. Sainte-Croix [GGBC92]. saithe [Du 91, HCF^{+}99a, HCF^{+}99b]. sajori [LLS97]. Sakagawa [Kes96b]. Salamanca [RU98]. salar [BM93, CK94, CMK97, EMJ90, Han90, JB93a, JB93b, JAG^{+}99, Loc97a, MPDW97, RF99, SÍW91, She90a, SJ95, STM^{+}97, Smi90, SMT95]. salaris [BM93, JB93a, JB93b]. saliens [MKKO95]. Salmo [BM93, CK94, CMK97, EMJ90, Ell93, Han90, HH90, HJ92, JB93a, JB93b, JAG^{+}99, LV98, Lan99b, LCU93, Loc97a, MPDW97, PMA95, RJS98, RF99, Rub94, SÍW91, She90a, SJ95, STM^{+}97, Smi90, SMT95, TW93, TKH94, VA96]. Salmon [Mun90, Pur93, Tre98b, AP96, Azu95, BM93, BHF99, BPD90, Bri93, Coo93, CK94, CMK97, Egg92, EMJ90, Eml91, EE93, FL93, GK91, GS93, Han90, HM92, HJH^{+}93, JB93a, JB93b, Jen90, JAG^{+}99, Kyl94, Kyl96, Loc97a, MPDW97, MCC90b, NT96, NYK93, PM93, Qui93, RF99, RR93, SÍW91, SBSW99, She90a, She90b, SJ95, STM^{+}97, Smi90, SMT95, TM93, TW93, USS93, VLAM93, WH90a, WH90b, Mil97, Paw90]. salmonid [LF93, Pic93, RJS98, SI98, Wil95, YNEK93]. salmonids [Azu95, Bri93, Bru93, Buc93, FB93, HH93, MM93, Rob93, WK93, dC93, Han94]. salmoninarum [FL93]. salmonis [LF93, TW93]. Saloum [Gui98]. salpa [vdWB97]. Salps [MPMC97]. saltwater [Rub94]. same [CA99, DP97]. Sample [Erz90, KK93, PR99, PM94]. sampled [CT92, dSD97]. samples [SS93]. Sampling [RDG96, TJ99, All93, BACH99, CL95, CCJ^{+}99, ESW97, FBAM99, FF98, GW92, GS92, Hol96, KGM^{+}93, KR96, LM98, MP96a, MK97, PAL97, PvDM^{+}97a, PR99, TJE99a, TJE^{+}99b, TW92, WBB^{+}97]. San [ACGMGS^{+}99, Loc93, Cio91]. sand [BKB97, LP99, Wen92]. sandeel [PLW99]. sandeels [DVY95, MMV98]. Santa [MM96, RSM99]. Sardina [CCV^{+}98, Dul95]. sardine [AMIA98, CCV^{+}98, Dul95, Mts96, NMCCMLB99, PvDM^{+}97a, PvDM^{+}97b, Ste91]. sardine-anchovy [Ste91]. Sardinella [GH95]. Sardinops [Fle92, NMCCMLB99]. sargus [GM97]. Saronikos [KRS97]. Sarpa [vdWB97]. SAS [Mil93b]. Satchell [War91]. saturation [CD90]. Saul [Sto90]. Saunders [Plu91]. saxatilis [MR93, TC92]. scabbard [MNSC96]. Scad [dFB90]. scale [AL95, ACC96, BKB97, FvMH96, GBL^{+}97, HGC97, HGHLC98, KK97, MR93, O’N93, SAV92, SRG96, SPP96, WF93]. scale-loss [BKB97]. scaled [NE96]. scales [Beg98, TH95, Ste95]. Scaling [McI95d]. scallop [Cio91, MYMP91, MJ93, SG97]. scallops [Col98]. scanning [MKS99, PB93, SI98]. scattering [Tre98a]. Scavenging [LF99]. Schaefer [DCS99]. Schilbe [ELK99]. Schizothorax [SS95]. Schnute [Fra95, TF94]. school [AJTP95, FGS92, Hol96, Tre98b]. schooling [SJ95]. schools [GF92, LLW96, Mis92, Tor91]. Schreck [McI93]. sciaenid [Gri97]. Sciaenidae [Caz96]. Sciaenops [GR91]. Science [Ano95r, Bro97c, Bro97d, Hea96, Kes96b, McI95e, McI95f, McI97a, Kes96d, Sto90, Ull98]. Sciences [Kes96c]. Scientific [All93, Lai93, Luc95, McI97b, Mil91, Pur93, She93, WBCG98]. scientists [McI96c]. Scomber [CNW92, CS97, CWG^{+}91, JHB96, Mis93, PPS93, YAM98a]. Scomberomorus [CMGB99, DAB92, EL97, Gov94, Gov95]. Scombridae [CS97]. scombrus [CNW92, CWG^{+}91, JHB96, Mis93]. scope [Gul98]. Scophthalmus [BF97]. Scotia [HC99]. Scotland [Du 91, GSSM99, MSWP93]. Scottish [BM93, KDD^{+}94, PBHK94]. scouting [Mil96]. screening [Hol94]. Scyllarid [SK98]. SE [Han90]. Sea [AOH90, All96, EW96, Han94, JH97, Kes97a, MMPHW99, Pea93, QAC^{+}98, SP93, WMB92, AJ90, CG97, CAI96, CCJ^{+}99, DLB91, GG98, JP90, KDCP96, KD97, LV98, Lan99b, MDMG^{+}93, MSWP93, Mil91, NT96, OKS98, Rub94, SMEL98, She90a, SJ95, SIE99, SG97, TW93, Tur97, ABR98, AGA98, Asp95, Avs94, Avs95, BSVC99, Bri92, Bri95, BAR99, CCA97, CS97, DE90, Das99, Du 95, Du 96, Dul97, ES99, GB98b, GNN^{+}97, GM97, GSSM99, GSM99, HAN97a, HAN97b, HAS98, Hol93, Hus92, Ism95a, Ism95b, IB99, IZJT93, JAC^{+}99, JT96, KD90, KBKO99, LTT^{+}97, LR96, MMPL98, MHWMP99, MS95, Nak98, OB97, Orł98, PLW99, PLL97, RB96, Rub94, SHH98, SZ98, SG98, SP94, SPP97, TJE^{+}99b, TJ99, TB98, TLT^{+}98, TGdB98, TH95, VO98, VMT97, Was91, WZ99]. Sea [YB96, YA99, McI91d]. sea-ranching [OKS98]. seabirds [Fur94]. seabream [SME95]. seafrozen [LVIG90]. seal [KK95b]. seals [BPD90, NGH92]. Seaman [Loc93]. search [Mil97, Sym97]. seas [Chu98, ES99, Fur94, GSSM99, GSM99, JAC^{+}99, KCS^{+}99, MS98, RMR^{+}99, CLZS97]. seashore [McI91c]. season [HMP96, KBKO99, LR98, Med98, Tro97]. Seasonal [Dol98, CLZS97, HG93, KO90, PS99, PHG^{+}94, VO98]. Seasonality [ET91, FVF94, GHHEE98]. seatrout [HGC94]. Seattle [Knu91]. Second [Plu91, Ano95r]. section [Pur93]. sections [MDB92, ZM92]. sediment [AF99, ATJ91, Bro97a, TJA94]. seine [Fle92, LT92, MN92, Mis92, Mis93, PAL97, VCMCA97]. seines [Jen91, MDB92]. seining [TSS97]. SELECT [TMW98]. selected [FBB98, GBL^{+}97]. Selection [WH93, BC93, GS92, HCF^{+}99a, HCF^{+}99b, NW95, OCU96, PCS96, PCEP98, PR99, Pow96, WMCH92]. selective [Ken90, LB95, MG97, TSS97]. selectivities [TLT^{+}98]. Selectivity [CNW92, FM92, FTHM96, KW91, MMPL98, MMPHW99, AA95, BACH99, CP94, EC98, GFM94, Gob98, HGC94, HR92, HGN99, Jen91, LTT^{+}97, LB92, LMMK92, LMF92a, LR96, Mac92, MKLvD94, MHWMP99, MW92, Mil93b, Mts96, OK96, PPSvD95, PS95a, PS96, PB97, RP92, RP99, Sal91, SME95, SMEL98, SPC93, SIE99, SPP97, TOSM96, TMW98, YF92, Yam98b]. Semenora [Bro97d]. Semi [SIE99, BEMW96, LMF92a, RMXP93]. Semi-pelagic [SIE99, BEMW96, RMXP93]. semi-trammel [LMF92a]. semisquamatus [DVY95, MMV98]. semisulcatus [AN95, Sid91]. Senegal [Gui98]. separation [SMS99]. separator [AS91, CBB97, HC99, IVLK92, TOSM96]. Sepia [BD98, Dun99, GHHEE98, dSD97]. September [Ano90j, Ano91k, Ano92m, Ano95q, Ano96q, Ano98o, Ano99o]. Seri [Whi91]. Series [All91, Ano92c, Fra96, Luc92, McA93, Pri95, AN97, FK95, KR96, Par98, SC96, SCP97]. serpulid [KCS^{+}99]. Sessile [McI91c]. set [He96, MKLvD94, SX91]. setline [Kai94]. setnet [CLZS97]. sets [HBB97]. setting [LP97b]. several [MB96]. sex [Bux92, KBKO99, LS99, McI96c, PHG^{+}94, TJA92, dSD97]. sex-ratio [KBKO99]. Seychelles [MR97]. shad [DRE97]. shallow [GHL98, MKS99, RAT91, Tho98, Tre98a, Tre98b]. shape [LH90, MR93]. Shark [MS98, BN92, Bon97, CP98a, Cor98, DN98, HJSNB98, KK95b, LSP99, MFCG98, SBM^{+}98, WTHC98]. sharks [Hur98]. sharpnose [MFCG98]. Shearer [Pur93]. shelf [DBJ98, ES99, FFGG94, GSSM99, GSM99, HB97, JAC^{+}99, KCS^{+}99, Lan99a, MMPS96, NLT99, RMR^{+}99, RF97, TMK92, VS97, VT96]. Shell [CS98b]. shellfish [SS93]. shelter [HML97]. Shetland [BTBP98, STM^{+}97]. Shield [UYD^{+}94]. shielding [CSM95]. shift [LMTE97]. Shifts [MS91, Cad96, HMP96, Ste96]. shoals [LL95]. shore [LT92]. Short [TC92, Ark93, BRCK90, MRB96, SH97, SRG96, WBCG98]. short-duration [WBCG98]. short-finned [Ark93, SH97]. short-lived [MRB96]. Short-term [TC92]. showing [GP99]. shown [VS97]. Shrimp [Ken90, AM93, CCD^{+}99, CTCN93, ES98, EW96, FBAM99, HAS98, IVLK92, JS98, LAG95, OM98, PR99, RB96, RD95, Rob92, RRdS^{+}97, SM98b, Tho92c, TW92, VWL91, WBB^{+}97]. shrimps [CCHW95, WT93]. Shuttleworth [Luc90]. sibogae [OM98]. Sicilian [LAG93]. Sicily [RB96]. side [MKS99]. side-scanning [MKS99]. sidescan [Tre98b]. Siganidae [WH97]. Siganus [WH97]. sight [LP97a]. signal [ESNA95, LL95, Tor91]. significance [BSVH99, Buc93, PW97]. Sillago [BKB97]. siluroid [ELK99]. Silva [How96b]. silver [HC99, Wes98]. similarity [LMC98]. Simmonds [Nak93]. Simple [Mac97, BSVH99, LCU93, WS96, Xia98a]. simplex [TF94, Asp95]. Simulated [BKB97, CS98b, CSR99, MB95]. simulation [BPvD^{+}92, LPSS98]. simultaneous [FAC97, JLMP96]. sine [CT92, Gui98]. Sine-Saloum [Gui98]. single [DCPM93, RHLE99, SB98c]. sinking [MDB92]. sirm [KW91]. situ [CM97, JHB96]. situations [Mis92, Mis93, S.94]. six [AJ90]. Size [Gob94, GS92, HCF^{+}99a, HCF^{+}99b, Jen91, JB93c, LB95, MHWMP99, SPP97, TMW98, BKH99, BBBF91, BAR99, BPvD^{+}92, CA99, Car93, CM99a, Che96, CfyCaL^{+}96, CP94, Dah94, Dol98, EPCPW96, Erz90, FX96, Fle92, Gob98, HB97, HGC94, HH90, HJ92, IZJT93, KD98a, LB92, LFT91, MG97, NW95, NW98, OK96, OCU96, PCS96, PCEP98, PR99, RHO99, SD94, Sal95, SPC93, WK90, WBCG98, WA97, Yam98b, dC93]. size- [BPvD^{+}92]. size-at-maturity [CM99a]. Size-based [JB93c]. Size-selective [LB95, MG97]. Size-selectivity [TMW98]. size-specific [BKH99]. sized [Eck98]. sizes [AB92, DKH99, MA97, SLB96]. Skagerrak [BF97, DE90, JM91, TG95]. skin [LHS95]. Sledges [TW92]. slope [DBJ98]. Small [ACC96, O’N93, SPP96, AL95, Eck98, FF98, GBL^{+}97, HGHLC98, How96a, KW91, MCC90b, PT95, RLR92, SMEL98, SAV92, SRG96, YF92, Yam98b]. small-meshed [KW91]. small-pelagics [PT95]. Small-scale [ACC96, O’N93, SPP96, AL95, GBL^{+}97, HGHLC98, SAV92, SRG96]. smallmouth [CM97]. smaris [Ism95b, TF91, TF94, VT96, VMT97]. smelt [CS98a, LP99, PRMH99]. Smith [McI95d]. smolt [JSE92, MPDW97, STM^{+}97]. smolts [SJ95, SMT95]. smoothed [EMS^{+}95]. snakehead [dS94]. snapper [Har95, MPVB98, OCU96, QK96, WMCW93]. snapper-grouper [Har95]. snood [MKLvD94]. so-iuy [OB97]. Soak [Aco94, MR96, LP97b]. soaking [LMF92b, YDM95]. Social [Mci97c]. Society [McI93]. sociological [Mci97c]. sockeye [Azu95, Egg92, GK91, Kyl96, RR93]. Soft [Ken90, AKB93, Kes96c, Kes97a, RT94, HM99]. soils [Bro97a]. solar [NW98]. sole [And98, FM92, JHF93, MHWMP99]. Soleidae [And98]. solida [RU98, Urb98]. Solomon [FD97, BMR^{+}90]. Some [AH93, Cav93, GF92, Per92, CD96, DPB91, FF98, Otu90, dS94]. something [Lon99a]. sonar [GHL98, GG98, Gui98, KDD^{+}94, PB93]. sonars [Tre98b]. Sons [Cri93, Lai95b]. Sound [Coo93, Dol98, DKH99, EMS^{+}95]. sounder [GG98]. sounders [RHLE99]. sounding [JLMP96]. sounds [VWL91]. source [Ama90]. Sources [BPD90, Ric91]. South [AKB93, BB97, BP98, BKI96, BK96, CMGB99, FVF94, Gov94, Gov95, KLB98, LK96, Pea99, SS93, WAB98, And98, Car93, CDB^{+}97, Col98, GBL^{+}97, HU94, KBKO99, LMC98, OM98, RAT91, SME95, SM97, SM98a, SMEL98, Bux92, Gri97, Har95, JB93c, LLL98, MS91, MRB96, SPP96, VWL91, vdWB97]. south-east [Col98]. south-eastern [LMC98]. south-west [GBL^{+}97]. south-western [OM98, RAT91]. southeast [ELK96, JB93a]. southeastern [MP97, MPVB98, NW98, GMHSR98]. Southern [LLL98, Ral96, AGA98, Avs94, BP96, Bon97, CQB^{+}94, GFM99, GM95, HH90, KK97, MFCG98, MH97, McS94, PS99, SH97, She98, SMS99, TMW98, WTHC98]. southwest [Car93, San93, BBC^{+}96, SD90]. southwestern [AA95, CfyCaL^{+}96, LSP99]. sp [Eck98]. Spain [Mil99, CCV^{+}98, FFGG94, GL91, GR94b, Sha94]. Spanish [DAB92, EL97, GLGB94, GRG94, GSR94, Lan99a]. spanner [HW99, SBK95]. sparid [Bux92, JDDK95, VM97]. Sparidae [BB97, FW95, PL96, PL98, QK96, WMCW93, vdWB97]. sparids [GM97]. Sparus [KD97]. Spatial [JAG^{+}99, CCA97, Dol98, Fle92, GB98a, JAC^{+}99, KPH^{+}96, SP94, TH95]. spatially [Pea99]. spatio [HG98]. spatio-temporal [HG98]. spawner [Cri94]. spawner-recruit [Cri94]. spawners [DE90]. Spawning [Bow90, DCPM93, LV98, Lan99b, YAM98a, Ark93, KBKO99, Kub92, LR98, MR93, RF97, Wes98]. special [Azu95, Bux92, GLGB94]. Species [LB92, LL95, MM93, AJ90, Bux92, CD90, DK96, DK97, ET91, FBB98, GBL^{+}97, GSM99, Har95, JT96, KGM^{+}93, KD98a, LBK97, Luc97b, MA97, MRB96, PPSvD95, PS95b, PS96, RL90, RMR^{+}99, RKCB90, SBSW99, SIE99, SPP96, TLT^{+}98, TF91, WBCG98, YDM95]. species-specific [RL90]. specific [BKH99, PLL97, RL90]. Specifying [Sec99]. speed [MDB92]. Spicara [Ism95b, PS96, TF91, TF94, VT96, VMT97]. spined [JLMP96]. spiny [EE95, EE96a, EE96b, Med98, RSB^{+}98]. Spirinchus [CS98a]. split [GG98, Gui98, RJS98, RHLE99, SI98]. split-beam [Gui98, RJS98, SI98]. sport [VLAM93]. spot [LP97a, Lan99a, SIE99]. spotted [HGC94]. Spp [CD90, DBJ98, LVIG90, McS94, NT96, PS92, RJS98]. Sprague [Loc93]. sprat [Avs94, Avs95, Dul98, SRMdC95, TG95]. sprattus [Avs94, Avs95, Dul98, Avs94, Avs95, Dul98, SRMdC95, TG95]. spreaders [SX91]. Spriner [Pri91]. Spring [DE90]. Springer [Luc90]. Springer-Verlag [Luc90]. sprinid [TJA92]. spurdog [Vin91]. Squalus [CRLG95, RSB^{+}98, Vin91]. square [ABR98, Bri92, BK94, BK96, BKB97, CNW92, FBB98, FM92, HCF^{+}99a, HCF^{+}99b, SPP97, Tho92c, WMCH92]. square-mesh [ABR98, BK94, BK96, FBB98]. square-meshes [BKB97]. squat [RB93]. squid [Ark93, BBC^{+}96, BRCK90, BP94b, CQB^{+}94, DBJ98, Fur94, LF99, MS98, MdCP94, NW98, PBHS94a, PM94, Por94, RH90, RKCB90, SH97, Sha94, SRG96, YDM95, YMSU97, dCM94, Ear91]. Squilla [AM93]. Sri [Ama90, KW91, MNJ91, PPSvD95, PWMM95, SEJ97, dS91, dS94]. St [IO98, Mci97c, RTC99]. stability [ESNA95, Ken96]. stage [MG92, PLL97]. stage-specific [PLL97]. stages [Loc97b]. standard [FBF90]. standardization [PT95]. standing [FP97]. Stanford [How93, Mci92]. State [Ano95r, Bru90, McI95f, JH97, LAG93, PSMSvD99, SBR98, PB92]. States [BSA92, HJSNB98, MP97, MPVB98]. Statistical [CBB97, Aze95, BPSW90, BBBF91, BMB98, LP95a, Sch98]. statistics [PSMSvD99, SRG96, Sve99]. statolith [Ark93, RH90]. Status [Bon97, EE93, FL93, Kes96c, Mil99, Okp91, Por94, SRSR90, Tho92b]. steal [NT96]. steenbras [BKH99]. Steenstrup [PM94]. stickleback [JLMP96]. Stickney [Lai95b]. still [Cad96]. stimulants [FJM91, JM90]. stimuli [FGS92, GWGR95, GW95, LJ92]. stimulus [KW98a, KW98b]. Stizostedion [HAN97a, HAN97b, MKLvD94]. Stock [BRCK90, BFP99, DLB91, FD97, IZJT93, JLMP96, PG94, Ste99, Vin91, ABD94, Ark93, AJ94, Avs94, Bae91, BW99, BHS99, Boo99, BP98, BMB97, CF99, CM99a, Che96, CfyCaL^{+}96, Cor98, Das99, DCPM93, DKH99, Dun99, ELCM92, EL97, Fle92, FK95, FP97, GB98b, GGBC92, Hol96, HMP96, JHF93, KW98c, LSP99, LK99, LAG93, LAA^{+}94, MA98, MSP97, MP96a, Mci97c, Mil93a, ME99, Mon90, PR96, PPS93, PJ96, PLW99, PvDM^{+}97b, PdP93, RAT91, Ric91, S.94, Sal95, Sec99, She98, SMS99, SRSR90, Sve99, SF99, Thr99, TCA^{+}97, Wal99, Xia97, Xia98a, Zhe96]. stock-recruitment [GB98b, S.94, Zhe96]. stocked [LPH^{+}98, Mor93, Wes98]. Stocking [Kyl96, DLT92, HAN97a, HML97, VA96]. Stocks [SF99, ABB^{+}98, BM93, BBC^{+}96, Cad97, Egg92, EE93, EGR^{+}97, FP99, Lyo98, MR93, MP96b, McS94, MKS99, NGH92, PRMH99, PS98, PSMSvD99, Pla93, PKH^{+}95, Pow96, RKCB90, SÍW91, TM93, WK93, WH90b]. stomach [HM95]. Stomatopoda [AM93]. stood [Lon98]. storage [AKKD92, Dah94, LVIG90, MB95]. stow [AJTP95]. strains [CMK97]. Strait [RB96, WT93]. stranded [GLGB94]. strategic [CY91]. Strategies [All96, Sve91, All93, BPvD^{+}92, CL95, HBB97, Loc97b, PvDM^{+}97a, SB98b]. strategy [LV98, LS98, Mci95c, MR96, PR99]. stratification [Cav93]. stratified [HAS98]. straying [Qui93]. stream [LCU93]. streamer [WK90]. streams [AF99, FBAM99]. strength [CS98b, Mac97, RHLE99, SI98]. stress [CAI96, LH90, TKH94]. stressed [Pic93]. striata [dS94]. striatus [Caz96]. striped [KDCP96, LMTE97, MR93, RSM99, TC92]. Stromateidae [WC95]. structural [Acq92, FK95]. structure [Ark93, BHS99, CP98b, CQB^{+}94, ES99, FDD99, FGS92, GR91, GSSM99, GSM99, HB97, JAC^{+}99, KCS^{+}99, LHML97, LS99, LSP99, MSR95, OB97, PR99, RMR^{+}99, RTC99, TH95]. structured [BPvD^{+}92, CG97, GP99, HPM94, PMA95, SD91]. structures [Gob94]. studied [Tor91]. Studies [FJM91, GWGR95, GW95, Iko96, IO98, Mci97c, PM98, RAT91, Dah96, GR91, HF90, JB93a, JB93b, JSE92, LMC98, MW92, Mil93b, ME99, Orł98, Ral96, SS93, SBSW99, Wal99, Was91]. Study [LH90, ABR98, Avs94, CL95, Che96, CDB^{+}97, EL97, GB98a, GAA99, GRG94, HLM91, JLMP96, KR96, LPSS98, LHML97, LPH^{+}98, LMA^{+}99, MP96b, Med98, NLVF98, NYK93, NM99, Piv92, PB93, RN98, San93, SB96, Tro97, VN98, VWL91]. studying [ADNT94, LTS98]. sub [BTBP98, Ste99]. sub-units [Ste99]. sub-zero [BTBP98]. subdivision [CM99b]. Subject [Ano94q, Ano95s, Ano96x, Ano97w]. Sublethal [Hea96]. subsample [Che96]. subsistence [BMR^{+}90]. substrate [MS97]. substrates [IASS96]. Subtleties [Xia97]. subtropical [MH97]. success [HML97]. Successes [BSA92, Tho92a, Ven92, WTM92]. successful [LV98]. Suez [GHHEE98]. sugarcane [KMT^{+}95]. sugarcane-fish [KMT^{+}95]. Sugiyama [Ear91]. Suitability [TSS97, VA96]. suitable [BC93]. Sumatra [MG99, HU94]. supplementary [Ama90]. surface [Hop92, NT96]. surfaces [BD98]. surimi [Tro97, Tro98]. surmuletus [LMTE97, PS95a]. Surplus [SB90, CD96, Løk91, Lon99a, SL98]. Surrey [Bla90]. Surveillance [YNEK93, Mil95]. survey [BACH99, BPV95, EBG96, GGBC92, HGC97, HG97, HPM94, Hug98, KMNP97, LAG93, MK97, PG95, SRG96, KMNP97]. surveying [MKS99]. Surveys [PS98, CA99, Cav93, CM97, GF92, HAS98, HR92, Kes97b, MK97, Pat94]. Survival [MB95, Rub94, BKB97, CCV^{+}98, HJ92, LTS98, QK96, SLB96, WK90, WH93, Loc97b]. survivorship [JB93c]. susceptibility [BM93]. Susitna [VLAM93]. Sustainable [DC97, EE95, EE96a, SG97]. Sustaining [Ano95r]. SW [Mil99]. Sweden [MM93]. Swedish [EE93, Sve99, VSJ^{+}99]. Swift [War95]. Swimming [He91, KMT^{+}95, LBBB95, Mis92]. Symes [McI97b]. sympatric [LPH^{+}98]. sympatry [Azu95]. Symposium [All96, Tur97]. Syndenham [EI98]. Synodontis [San96]. synthesis [MM98]. System [Kes97a, RD95, CSHW96, ESW97, EI98, JSE92, LL95, MPDW97, SM97, SM98a, SI98]. Systematics [Luc92, How93]. Systemic [PM93, Bru90]. systems [JP90, McI96d].

T [All96, Bea95, Bro97d, How96b, Kes96b, Lai95a, Paw96, LLS97]. TAC [SBR98]. tackle [VLAM93]. tag [CSHW96, MB96, Sid91, WK90]. tag-recapture [Sid91]. tagged [BF97, HF91, SMT95, WK90]. Tagging [MB96, AMON95, Kid94, NP99]. tags [HLM91, LHML97, MA98, MW93, OKS98, PW94, QK96, Sve95, WK90]. Tai [fN94]. tailbeat [LHW98]. Taiwan [CfyCaL^{+}96, LL96, LC99, WC95]. taken [YDM95]. tale [McI94]. Tanganyika [PS99]. tangle [HW99]. tank [LMF92b]. target [Hol96, MA97, RHLE99, SI98]. Targeted [LK99]. targets [Eck98]. Tasker [McI97a]. Taylor [McI91b]. technical [PR96]. technique [EE96b, FJM91, How96a]. techniques [BBC^{+}96, CF99, Gra99]. Technische [Hop92, McI95a]. techno [Njo91]. techno-economics [Njo91]. technologies [Sve91]. technology [PB92, dCB95, Ear93]. TED [AKB93, Ken90, MBB95, RTBD95]. tehuelcha [Cio91]. tehuelchus [Iri91]. telemetry [CM97, War95]. teleost [Pri91, SS95]. Teleostei [AOH90, AMON95, PL96]. Teleostomi [Mil91]. tell [Lon99a]. temperature [Bow90, HQ91, LBBB95]. temperature-induced [HQ91]. temperatures [BTBP98, He91]. temporal [FS94, HG98, JAC^{+}99]. tenth [WF93]. Tenualosa [DRE97]. term [CM96, Dah94, HQ91, Ken96, Kly98, TC92]. terms [YF92]. terraenovae [MFCG98]. territorial [SAV92, dCB95]. test [RKK90]. Testing [Dic93b, Pow96]. thaleichthys [CS98a]. Thames [Hug98]. Their [All93, BAG98, CCV^{+}98, GM97, HM92, LP97a, McI97a, Pea99, PVM^{+}96, TMD97, TMK92, Wil90]. Theme [Kes96b]. Theoretical [ABD94, Coo93, Dic93a, Dic93b, MDB92]. theories [MMM99]. theory [LH90, LR90, Med98, PT95]. Theragra [EPCPW96, Meg91, Par98]. there [Cad96, DC97]. thermal [VSG99]. thickness [LR96]. thinking [Cad96]. Thorgaard [Wil95]. thread [PS92, VCMCA97]. thread-herrings [PS92]. threat [Lan99b]. Three [Lip99, PAL97, ATJ91, BACH99, Col90, DRE97, DMA^{+}91, DLT92, FX96, JLMP96, Kes96d, KDD^{+}94, MP96b, TLT^{+}98, TF91, UYD^{+}94, dSD97]. three-spined [JLMP96]. thresholds [CWG^{+}91, KS93]. throughout [Ell93]. Thunberg [AOH90]. Thunnus [Hsu99, SZ98]. thynnus [SZ98]. tides [Whi91]. tiger [AN95, Sid91]. tilapia [PPSvD95, Bha94]. Tim [McI95d]. time [Aco94, AN97, EMS^{+}95, FK95, HB97, KK95b, LP97b, LMF92b, NW98, Par98, PR96, She90a, SC96, SCP97, YAM98a, YDM95]. time-smoothed [EMS^{+}95]. times [MR96]. Tindall [Plu91]. tobianus [MMV98, DVY95]. Todaropsis [GRG94]. tonnes [Pau96]. Tony [Luc97b]. too [McI94]. tool [Arg92, Bri92, LAA^{+}94, Mat92, PS98, Tow98]. tools [Sum98]. Torres [WT93]. torsk [FL94]. Toshiaki [Mac93]. total [BBBF91]. tow [Cav93, GPV90]. towed [KW98b, MKS99]. tows [KGM^{+}93, WBCG98]. toxic [MPMC97]. Trace [AJ90, ELCM92, TJA94, ATJ91]. Trachurus [dFB90, KRS97, TF91]. Tracking [Sve95, TR94, War95]. Traditional [EGR^{+}97, IKA90, MP96a]. traffic [PO98]. training [Tho92b]. trajectories [SI98]. trammel [AA95, CAI96, FTHM96, LM98, LMF92a, Sal91]. trammel-nets [Sal91]. trammelnets [Gob92]. transducer [Gui98]. transfer [SJ95]. transferable [SB96, SB97a]. transmission [EKP93]. transmitter [LHW98]. transmitters [HM99]. transponder [CSHW96, LMA^{+}99]. trap [ABD94, CD90, Col90, ET91, Har95, HGHLC98]. Trapping [FF96, KMT^{+}95]. traps [CSM95, Gob98, NW95, RAT91, RHO99]. trawl [AS91, BEMW96, BREB98, BK94, BKI96, BK96, Bub96, CA99, Car93, Cav93, CW95, CBB97, Dic93a, Dic93b, ESW97, EBG96, FBB98, FVF94, FvMH96, FM92, FW95, GFM94, GW95, GPV90, GW92, GS92, GAA99, HC99, HAS98, HR92, IVLK92, JP90, Ken90, LLW96, LTS98, LAG93, LBK97, LTT^{+}97, MMPHW99, MA97, MMM99, MW92, Mil93b, MBT^{+}99, MBB95, O’N93, RMXP93, RTBD95, RM99, RN98, SB98c, SD90, SEO96, TLT^{+}98, WF93]. trawl-survey [LAG93]. trawler [CBB97, CCJ^{+}99, WBCG98]. trawlers [QAC^{+}98, WH93]. Trawling [Tro97, TKH94, AAP96, AS91, BTBP98, FAC97, JAC^{+}99, KCS^{+}99, KGM^{+}93, KLB98, LK96, VWL91]. trawls [ABR98, BACH99, BD93, CNW92, EPCPW96, HCF^{+}99a, HCF^{+}99b, HGN99, LR96, MW92, PR99, RX95, RRdS^{+}97, SLW97, WBB^{+}97, YA99]. treatment [Kyl94]. trees [CM96]. trematode [WMCW93]. Trends [PSMSvD99, AA94, EE95, EE96a, GSR94, JAC^{+}99, dCM94, GSSM99]. Trent [Lyo98]. trevally [HLW96]. Trevor [Luc90]. trials [CBB97, FL94, He96, LJ92, MDB92, SB98c]. Trichiurus [MH97, YB96]. Triglidae [CCA97, TMK92]. triglids [TMK92]. triostegus [Pla93]. trips [TJ99]. Trisopterus [PP94]. Trochidae [FD97]. trochus [FD97, EGR^{+}97, FD97]. trophic [BMR^{+}90, CB98, Kyl94]. trophic-level [Kyl94]. Trophodynamic [BSVC99]. Tropical [McI90, AKKD92, BREB98, DC97, EE96b, GHL98, Gob92, Gob94, Gui98, MRB96, PAL97, RMXP93, RDG96, WBB^{+}97, WT93, dS91]. tropics [AL95]. trouser [MW92]. trout [AF99, Ell93, HH90, HJ92, KDD^{+}94, LV98, Lan99b, LCU93, LMF92a, MPDW97, MDMG^{+}93, MSWP93, Mor93, PMA95, Rub94, TW93, TKH94, VA96]. trunculus [GFM99]. truth [Mci97c]. trutta [VA96, Ell93, HH90, HJ92, LV98, Lan99b, LCU93, MPDW97, PMA95, Rub94, TW93, TKH94]. Tsukayama [Hop92]. tubeheads [KCS^{+}99]. Tucker [McV93]. Tugela [FVF94]. Tuna [CM96, McI90, Pri96a, BMR^{+}90, DPB91, Pol91, SEJ97, SZ98, Pri96a, McI90]. tunnel [GWGR95, GW95]. turbot [AW90, BF97, LBBB95]. Turkey [KBKO99]. Turkish [Avs95, Ism95a, Ism95b, IB99, LTT^{+}97]. turtle [RT94]. twenty [McI96e]. twenty-five [McI96e]. twin [CBB97, SB98c]. twine [LR96]. Two [Xia98a, BKI96, Bux92, CS97, DP97, FBAM99, GHHEE98, Hol94, KK95b, Lon99b, MG92, PMA95, PAL97, Sal91, SMEL98, SIE99, SI98, SK98, TJA92, TF94, VWL91, WT93, Wen92]. two-stage [MG92]. type [EPCPW96]. types [MB96]. Tyrrhenian [CCA97, WMB92].

U [Mil91]. U.S.$48.50/£ [Ear91]. U.S.A [Knu91, Wil90]. UDN [Rob93]. UK [Bro97b, Bro97e, Ear93, Hea96, Mci96b, Bro97c, GSSM99, PBHS94b, Sha94]. Ulcerative [Rob93]. uncertainties [Chu98, You98]. Uncertainty [SB97a, Cha98, EL97, FSSU98, Hoe98, Li99, MR97, MM98]. undatum [Kid94, VSJ^{+}99]. under-ice [TSS97]. undersized [TKH94]. Understanding [McI96d]. underwater [He96]. Union [Cor96, Mci96b]. unit [Har95, SAB^{+}93, YDM95]. United [Hea96, Hug98, BP96, BSA92, HJSNB98, MP97, MPVB98]. units [Ste99, Xia98b]. univariate [SC96, SCP97]. University [All96, Bro97e, Bru90, Gat90, Han94, How93, Knu91, Luc93, McI91d, McI95d, Mci92, Mci97c, Mil97, Ral96, Ste95, Tur97, War91]. unstable [Luc93]. unwanted [BK96]. up- [JLMP96]. Upeneus [KBKO99]. upland [dS91]. upper [Iko96]. upwelling [Hop92]. Uruguay [Nor95]. Uruguayan [DLB91, ST96, NLT99]. US$15 [Han94, McI95a, Ste95]. US$15.00 [Hop92]. US$15.75 [McI90]. US$17.95 [Bro97e]. US$25 [Han94, Tur97]. US$25.00 [Knu91]. US$35 [Mci92]. US$50 [McI93]. US$59.50 [War91]. USA [Bro97d, Kes96b, McI95f, Pea93, AA94, SMT95, TMD97]. Use [Fle92, LMA^{+}99, RMXP93, Ric91, TF91, Cru97, DC97, FBF90, GHR99, GWGR95, GW95, Gob92, GGBC92, NM99, SAV92, SF99, Tho92c, WA97, Xia97, Xia98a]. used [BI97, DRE97, HGC97, ME99, OKS98, PG94, RD95, RMR^{+}99, She98, TGdB98]. usefulness [KR96]. uses [USS93]. using [ADNT94, Bae91, BB97, BK94, CG97, CW95, CBB97, Cri94, Das99, ESNA95, FF98, FJM91, FL94, GR91, GG98, Gui98, HPM94, IVLK92, Jen96, KGM^{+}93, LHML97, LK99, LMNP98, LH90, LS93, LLS97, Løk91, LB95, LFT91, Mac97, MW93, MS95, PR96, PBN96, Pla93, RJS98, RP92, RH90, SBSW99, Sid91, SJ97, SI98, Ste91, Sve95, Tre98b, TF94, VMT97, VM97, VLAM93]. Utilisation [EMJ90, Ear91]. utility [CF99].

V [Lai95a, McI95a, Mil91, GSM99, PKH^{+}95, SQM95]. Valenciennes [NE96]. Validation [VM97, GM95, LP97a, WBCG98]. validity [RvLV90]. valuable [GHHEE98]. value [San93, Tro98]. values [SAB^{+}93]. Vanuatu [FBF90]. Variability [Bri95, CCV^{+}98, Dol98, NW95, Tud99, TH95]. variable [PvDM^{+}97a, She98]. Variance [KGM^{+}93, PG95]. variances [HGC94]. Variation [Han93, AH93, BP96, CLZS97, CB98, Cru97, DJ93, FW95, GR94a, GM91, HQ91, PHG^{+}94, RT94, RN98, SF99]. Variations [San96, CM99a, NLT99, PL92, SÍW91, Tro97, VO98]. variegatus [SQM95]. Vector [Ste91]. vectors [CA99, MPMC97]. veined [MdCP94]. velocity [FvMH96, PAL97]. vendace [AJ94]. Veneridae [AGA98]. Venus [AGA98]. Verde [RU98]. Verlag [Luc90]. Vernon [McI91c]. verreauxi [MB96]. verrucosa [AGA98]. versus [MP96a, PW97, RL90, WMCH92]. vertebrae [Ped97]. vertical [ES99, GF92]. Vesijärvi [MP96a]. vessel [AH93, EMS^{+}95, GF92]. vessels [GSSM99, HM92]. VHS [EKP93]. VI [ES99]. viability [NLVF98, VN98]. Vibrio [SS93]. Victoria [SB96, SB97a, SB98b, McS94, PBN96, TMW98]. view [Tho92a]. views [Lon99b]. VII [KCS^{+}99]. villosus [GB98b, TGdB98]. Vinogradov [Bro97d]. virens [Du 91]. Virgin [Fri95, FB97]. virginica [PKH^{+}95]. Virtual [LC99, PRMH99]. virus [EKP93, YNEK93]. vision [Lon99b]. Vistula [PB97]. visual [CWG^{+}91, GWGR95, GW95, KW98a, KW98b, Loc97a]. Vital [Sve99]. viviparus [KK97]. Voigtlander [McI95f]. Vol [All91]. Volkov [Bro97d]. völlenhovenii [ES98]. Volume [Pri91, Ano90e, Ano90f, Ano91e, Ano91f, Ano91g, Ano92d, Ano92e, Ano93f, Ano93g, Ano93h, Ano94g, Ano95e, Ano95b, Ano95c, Ano95d, Ano96d, Ano96e, Ano96f, Ano96g, Ano97c, Ano97d, Ano97e, CT92]. volumes [Ano95a, Ano95s, Ano96c, Ano96x, Ano97b, Ano97w, McI96e]. vomerinus [MBH99]. VPA [Pun97]. VPA-based [Pun97]. vs [KD98a]. vulgaris [GM97, CQB^{+}94, CDB^{+}97, GR94b, HGHLC98, MdCP94, PBHS94a, PHG^{+}94, QAC^{+}98, VM97]. vulnerability [TMD97].

W [Bru90, LS98, LS99, Loc93, McI95f, Plu91, Pur93, Sto90]. Wageningen [Luc93]. Walbaum [CCV^{+}98, Dul95, HGN99, Lan99b]. Wales [AKB93, BKI96, KLB98, LK96, Pea99, WAB98, BK96, TJE^{+}99b]. walleye [EPCPW96, JH97, Meg91, Par98]. Ward [Hob91]. Warri [Iko96]. Washington [Han94, Knu91, McI91a, Mci95c, Mil97, Pri96a, Ste95, CS98a, PB92]. Watanabe [Loc97b]. Water [RL90, RLS92, AKKD92, ATJ91, Bae91, Bow90, BTBP98, BD93, ES98, LM98, RB96, RLR92, RU98, SJ95, TJA94, Tho98, TS92, WCMG91]. watercourse [JAG^{+}99]. waters [AN95, Azu95, BRCK90, BC93, dFB90, CfyCaL^{+}96, DMA^{+}91, DCPM93, DK96, DK97, GHL98, GRG94, GR94b, HLES93, LV98, LL96, LC99, MBH99, MKS99, NW98, PLL97, PS95a, PS95b, PS96, PBHS94b, PBHK94, RRdS^{+}97, RDG96, RF97, SEJ97, She90a, Tre98a, Tre98b, VSJ^{+}99, WC95, YAM98a]. watershed [EMJ90]. webbing [Ken90]. Webs [Ste95]. Weight [DK96, DK97, GBL^{+}97, PS95b, SQM95, SD91, Hsu99, MDB92, TJA92, Xia98b]. Weiss [TMD97]. Wells [JSE92]. West [ES98, PCEP98, GBL^{+}97, HJSNB98, KW91, RF99, Sal95, Sve99, TW93, Cav93, FD97, HR92, MG99, MDMG^{+}93]. westcoast [BKH99]. Western [CCD^{+}99, QAC^{+}98, GM97, JHF93, MMN97, MPDW97, MCC90b, MKKO95, MNM97, OM98, RAT91, SU92, VP93, dFB90, Fle92, GAA99, Kes96a, SB96, SB97a, SB98b, WMCW93]. whaling [Bro97e, SHH98]. whelks [VSJ^{+}99]. where [Mat92]. whiffiagonis [LPP96]. white [LMC98, MMH93, SRSR90, TMD97, TR94]. whitefish [LPH^{+}98]. whitetip [LSP99]. whiting [Bri92, BKB97, Ism95a, IB99, PG96, SLB96, SGK98, SG98, SMS99, Tro97, Tro98]. whole [LVIG90]. wickertrap [FF98]. wickertraps [FF96]. Wickins [Lai93]. widening [Cor96]. width [RN98]. wild [Bru93, Buc93, CGG99a, EKP93, EE93, FB93, HJH^{+}93, HLES93, HH93, MSWP93, Mil97, PM93, Qui93, Rob93, TW93, WH90b, YNEK93, dC93]. wild-caught [HLES93]. Wildlife [War95]. Wiley [All90, Cri93, Lai95b, Pat94, Wil90]. Willamette [WN92]. William [Coo93, Han94]. wind [Nor95]. window [MMPL98]. windows [MMPHW99]. Winds [Knu91]. Winfield [Luc92]. wings [LLW96]. winter [JRM99]. winter-fishing [JRM99]. witch [Bow90, NGNH91]. within [HBB97, HMP96, LLW96, RF97, WT93]. without [HML97, McI90, RB93]. Witold [McI90]. Women [Knu91]. Wood [SJ97]. worker [PHG^{+}94]. World [Ano95r, Kes96b, Mci92, Bro97d, Kes96d, McI95f]. worms [KCS^{+}99]. Wound [WK90].

X [Pri96b, McI95d, Bea95, How96b, Kes96b, Luc97a, McI95f, McI93, War95]. xi [McI90]. xiv [Ral96]. xxiv [Pea93]. XXIX [Mil91].

Yamashita [Loc97b]. Year [PL92, CM99b, HAN97a, Mac97, NLT99, Piv92]. Year-class [PL92, Mac97]. years [EE93, RB93]. yellow [ELCM92, Mil91, CLZS97, IZJT93, JT96]. yellow-eye [ELCM92]. yellowfin [DPB91]. yellowtail [PL92]. Yield [DS99, Bux92, CD96, CG97, DLT92, DC97, ES98, HH90, Ism95b, KD90, LFT91, MR97, Per92, PWMM95, RB96, RL90, RLS92, S.94, SM97, SM98a, WZ99]. yield-per-recruit [Bux92, CG97, RB96, S.94]. yield-predictive [DLT92]. yields [DP97, EE95, EE96a, PBN96]. yolk [JDDK95]. yolk-sac [JDDK95]. York [McI91b, McV93, Pat94, War95, Wil95, Wil90, MH90]. young [HQ91, HAN97a, NLT99]. young-of-the-year [HAN97a, NLT99].

Z [Mil91]. zambezensis [San96]. Zealand [Gra99, McS94, SB97b]. zero [BTBP98]. Zhujiang [CCHW95]. Zimbabwe [Mts96]. Zoarces [KK97]. Zone [SB96, SB97a, SB98b, HLW96, PAL97, SQM95, TG95, DLB91, ST96]. zones [CS97]. Zoophysiology [Pri91]. Zusammenarbeit [Hop92, McI95a].

References

Ahmed:1994:FRC

[AA94] Farid E. Ahmed and Roger D. Anderson. Fishery resources, consumption and import trends, and biotechnology developments in the USA. Fisheries Research, 19(1–2):1–15, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900116.

Acosta:1995:CES

[AA95] Alejandro R. Acosta and Richard S. Appeldoorn. Catching efficiency and selectivity of gillnets and trammel nets in coral reefs from southwestern Puerto Rico. Fisheries Research, 22(3–4):175–196, March 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400328T.

Al-Ansi:1996:EFQ

[AAP96] M. Al-Ansi and I. G. Priede. Expansion of fisheries in Qatar (1980–1992): growth of an artisanal fleet and closure of a trawling company. Fisheries Research, 26(1–2):101–111, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695003967.

Addison:1992:AML

[AB92] J. T. Addison and D. B. Bennett. Assessment of minimum landing sizes of the edible crab, Cancer pagurus L., on the east coast of England. Fisheries Research, 13(1):67–88, January 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290034Q.

Agnew:1998:AAS

[ABB^{+}98] D. J. Agnew, R. Baranowski, J. R. Beddington, S. des Clers, and C. P. Nolan. Approaches to assessing stocks of Loligo gahi around the Falkland Islands. Fisheries Research, 35(3):155– 169, May 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698000836.

Arena:1994:TET

[ABD94] Guillermo Arena, Luis Barea, and Omar Defeo. Theoretical evaluation of trap capture for stock assessment. Fisheries Research, 19(3–4):349–362, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900493.

Armstrong:1998:SOP

[ABR98] M. J. Armstrong, R. P. Briggs, and D. Rihan. A study of optimum positioning of square-mesh escape panels in Irish Sea Nephrops trawls. Fisheries Research, 34(2):179–189, March 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000787.

Amar:1996:SSF

[ACC96] E. C. Amar, R. M. T. Cheong, and M. V. T. Cheong. Smallscale fisheries of coral reefs and the need for communitybased resource management in Malalison Island, Philippines. Fisheries Research, 25(3–4):265–277, March 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004394.

Abitia-Cardenas:1999:DBM

[ACGMGS^{+}99] Leonardo Andres Abitia-Cardenas, Felipe Galvan-Magaña, Francisco Javier Gutierrez-Sanchez, Jesus Rodriguez-Romero, Bernabe Aguilar-Palomino, and Almei Moehl-Hitz. Diet of blue marlin Makaira mazara off the coast of Cabo San Lucas, Baja California Sur, Mexico. Fisheries Research, 44(1):95–100, November 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000533.

Acosta:1994:STN

[Aco94] Alejandro R. Acosta. Soak time and net length effects on catch rate of entangling nets in coral reef areas. Fisheries Research, 19(1–2):105–119, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900175.

Acquay:1992:ISA

[Acq92] Herbert K. Acquay. Implications of structural adjustment for Ghana’s marine fisheries policy. Fisheries Research, 14(1):59– 70, June 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900733.

Allison:1994:MSP

[ADNT94] E. H. Allison, A. Davies, B. P. Ngatunga, and A. B. Thompson. A method for studying pelagic fish communities in deep lakes using drifting gillnets. Fisheries Research, 20(1):87– 91, June 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900086.

Argent:1999:FSE

[AF99] David G. Argent and Patricia A. Flebbe. Fine sediment effects on brook trout eggs in laboratory streams. Fisheries Research, 39(3):253–262, January 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369800188X.

Arneri:1998:ADG

[AGA98] Enrico Arneri, Gianfranco Giannetti, and Bruno Antolini. Age determination and growth of Venus verrucosa L. (Bivalvia: Veneridae) in the southern Adriatic and the Aegean Sea. Fisheries Research, 38(2):193–198, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001465.

Abrahams:1993:SCV

[AH93] M. V. Abrahams and M. C. Healey. Some consequences of variation in vessel density: a manipulative field experiment. Fisheries Research, 15(4):315–322, January 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390082I.

Ashraf:1990:TMC

[AJ90] M. Ashraf and M. Jaffar. Trace metal concentrations in edible muscle of six species of fish from the Arabian sea. Fisheries Research, 8(3):287–291, January 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090028T.

Auvinen:1994:CPV

[AJ94] Heikki Auvinen and Juha Jurvelius. Comparison of pelagic vendace (Coregonus albula) stock density estimation methods in a lake. Fisheries Research, 19(1–2):31–50, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900132.

Andrew:1995:CAS

[AJTP95] N. L. Andrew, T. Jones, C. Terry, and R. Pratt. By-catch of an Australian stow net fishery for school prawns (Metapenaeus macleayi). Fisheries Research, 22(1–2):119–136, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400306H.

Andrew:1993:AMS

[AKB93] N. L. Andrew, S. J. Kennelly, and M. K. Broadhurst. An application of the Morrison soft TED to the offshore prawn fishery in New South Wales, Australia. Fisheries Research, 16(2):101– 111, March 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390046A.

Ali:1992:BCD

[AKKD92] Ayub Ali, Indrani Karunasagar, Iddya Karunasagar, and J. Dardignac. Bacteriological changes during iced storage of the tropical fresh water carp Labeo rohita. Fisheries Research, 13(2):189–197, February 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290026P.

Ali:1995:CRM

[AL95] Ahyaudin B. Ali and K. Y. Lee. Chenderoh Reservoir, Malaysia: a characterization of a small-scale, multigear and multispecies artisanal fishery in the tropics. Fisheries Research, 23(3–4):267–281, June 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400351V.

Allen:1990:BRF

[All90] K. Radway Allen. Book review: Fish population dynamics: the implications for management: J. A. Gulland (Editor). Wiley Interscience, Chichester, 2nd edn., 1988, 432 pp., hardcover, US $52.15, ISBN: 0-471-91151-8. Fisheries Research, 9(1):93– 96, July 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090046X.

Allen:1991:BRO

[All91] K. Radway Allen. Book review: Operations research and management in fishing: NATO ASI Series E, Vol. 189, A. Guimarães Rodrigues (Editor). Kluwer Academic Publishers, Dordrecht, 1990, 340 pp., £68, ISBN 0-7923-1051-9. Fisheries Research, 12(4):391–393, December 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783691900217.

Allen:1993:BRC

[All93] K. Radway Allen. Book review: Catch effort sampling strategies — Their application in freshwater fisheries management: I. G. Cowx (editor), fishing news books, Blackwell Scientific, Oxford, 1991, ix + 420 pp., hard cover, £49.50, ISBN 0-85238-177-8. Fisheries Research, 15(4):388–390, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390090T.

Allen:1996:BRP

[All96] K. Radway Allen. Book review: Proceedings of the International Symposium on Management Strategies for Exploited Fish Populations: G. Kruse, D. M. Eggers, R. J. Marasco, C. Pautzke and T. J. Quinn II (Editors), Alaska Sea Grant College Program Report No. 93-02, University of Alaska, Fairbanks, 1993, 825 pp., price US $30.00, ISBN 1-56612-021-7. Fisheries Research, 28(4):423–428, December 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005036.

Abello:1993:FDM

[AM93] Pere Abelló and Paloma Martín. Fishery dynamics of the mantis shrimp Squilla mantis (Crustacea: Stomatopoda) population off the Ebro delta (northwestern Mediterranean). Fisheries Research, 16(2):131–145, March 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390048C.

Amarasinghe:1990:MCR

[Ama90] U. S. Amarasinghe. Minor cyprinid resources in a manmade lake in Sri Lanka: a potential supplementary source of income for fishermen. Fisheries Research, 9(1):81–89, July 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090043U.

Abad:1998:AEA

[AMIA98] R. Abad, J. Miquel, M. Iglesias, and F. Alvarez. Acoustic estimation of abundance and distribution of sardine in the northwestern Mediterranean. Fisheries Research, 34(3):239– 245, April 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369700101X.

Agostinho:1995:CRA

[AMON95] Angelo Antonio Agostinho, Yasunobu Matsuura, Edson Kiyoshi Okada, and Keshiyu Nakatani. The catfish, Rhinelepis aspera (Teleostei; Loricariidae), in the Guaíra region of the Paraná River: an example of population estimation from catch-effort and tagging data when emigration and immigration are high. Fisheries Research, 23(3–4):333–344, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400347Y.

Abdulqader:1995:BMP

[AN95] E. A. A. Abdulqader and E. Naylor. Bionomics and migration patterns of the green tiger prawn, Penaeus semisulcatus De Haan, in Bahrain waters. Fisheries Research, 21(3–4):395– 407, January 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694002969.

Aglen:1997:ITS

[AN97] Asgeir Aglen and Odd Nakken. Improving time series of abundance indices applying new knowledge. Fisheries Research, 30(1–2):17–26, May 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000040.

Andrade:1998:AGB

[And98] José Pedro Andrade. Age and growth of the bastard sole, Microchirus azevia (Capello, 1868) (Pisces, Soleidae) from the south coast of Portugal. Fisheries Research, 34(2):205–208, March 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000933.

Anonymous:1990:Aa

[Ano90a] Anonymous. Announcement. Fisheries Research, 8(3):301, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090034S.

Anonymous:1990:Ab

[Ano90b] Anonymous. Announcement. Fisheries Research, 8(4):397, May 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090007I.

Anonymous:1990:AIa

[Ano90c] Anonymous. Author index. Fisheries Research, 8(4):399– 401, May 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090008J.

Anonymous:1990:AIb

[Ano90d] Anonymous. Author index. Fisheries Research, 9(4):375–377, October 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783690900562.

Anonymous:1990:CFRa

[Ano90e] Anonymous. Contents of fisheries research, volume 8. Fisheries Research, 8(4):402–404, May 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090009K.

Anonymous:1990:CFRb

[Ano90f] Anonymous. Contents of fisheries research volume 9. Fisheries Research, 9(4):378–380, October 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783690900573.

Anonymous:1990:EBa

[Ano90g] Anonymous. Editorial Board. Fisheries Research, 9(1):iii, July 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090035T.

Anonymous:1990:EBb

[Ano90h] Anonymous. Editorial Board. Fisheries Research, 10(1–2):iii, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090010S.

Anonymous:1990:PJb

[Ano90i] Anonymous. Pages 1–96 (July 1990). Fisheries Research, 9 (1):??, July 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1990:PS

[Ano90j] Anonymous. Pages 189–286 (September 1990). Fisheries Research, 9(3):??, September 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1990:PJa

[Ano90k] Anonymous. Pages 205–301 (January 1990). Fisheries Research, 8(3):??, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1990:PO

[Ano90l] Anonymous. Pages 287–380 (October 1990). Fisheries Research, 9(4):??, October 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1990:PM

[Ano90m] Anonymous. Pages 303–404 (May 1990). Fisheries Research, 8(4):??, May 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1990:PA

[Ano90n] Anonymous. Pages 97–187 (August 1990). Fisheries Research, 9(2):??, August 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1991:A

[Ano91a] Anonymous. Announcement. Fisheries Research, 10(3–4):365– 366, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190092T.

Anonymous:1991:AIa

[Ano91b] Anonymous. Author index. Fisheries Research, 10(3–4):367– 368, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190093U.

Anonymous:1991:AIb

[Ano91c] Anonymous. Author index. Fisheries Research, 11(3–4):397– 398, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783691900114.

Anonymous:1991:AIc

[Ano91d] Anonymous. Author index. Fisheries Research, 12(4):395–396, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783691900228.

Anonymous:1991:CFRa

[Ano91e] Anonymous. Contents of Fisheries Research volume 10. Fisheries Research, 10(3–4):369–370, January 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190094V.

Anonymous:1991:CFRb

[Ano91f] Anonymous. Contents of fisheries research volume 11. Fisheries Research, 11(3–4):399–400, August 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783691900125.

Anonymous:1991:CFRc

[Ano91g] Anonymous. Contents of fisheries research, volume 12. Fisheries Research, 12(4):397–398, December 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783691900239.

Anonymous:1991:C

[Ano91h] Anonymous. Corrigendum. Fisheries Research, 12(1):89, September 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190052H.

Anonymous:1991:EBa

[Ano91i] Anonymous. Editorial Board. Fisheries Research, 11(1):iii, February 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190029F.

Anonymous:1991:EBb

[Ano91j] Anonymous. Editorial Board. Fisheries Research, 12(1):iii, September 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190044G.

Anonymous:1991:PS

[Ano91k] Anonymous. Pages 1–89 (September 1991). Fisheries Research, 12(1):??, September 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1991:PF

[Ano91l] Anonymous. Pages 1–92 (February 1991). Fisheries Research, 11(1):??, February 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1991:PJ

[Ano91m] Anonymous. Pages 177–370 (January 1991). Fisheries Research, 10(3–4):??, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1991:PN

[Ano91n] Anonymous. Pages 187–282 (November 1991). Fisheries Research, 12(3):??, November 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1991:PD

[Ano91o] Anonymous. Pages 283–398 (December 1991). Fisheries Research, 12(4):??, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1991:PO

[Ano91p] Anonymous. Pages 91–185 (October 1991). Fisheries Research, 12(2):??, October 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1991:PA

[Ano91q] Anonymous. Pages 99–193 (April 1991). Fisheries Research, 11 (2):??, April 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1992:AIa

[Ano92a] Anonymous. Author index. Fisheries Research, 13(4):435– 437, April 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900663.

Anonymous:1992:AIb

[Ano92b] Anonymous. Author index. Fisheries Research, 14(4):343–344, September 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290043S.

Anonymous:1992:BRC

[Ano92c] Anonymous. Book review: Cichlid Fishes: Behaviour, Ecology and Evolution (Fish and Fisheries Series: 2): Miles H. A. Keenleyside (editor), Chapman and Hall, London, 1991, 378 pp., £45.00, ISBN 0-412-32200-5. Fisheries Research, 13(4):

430, April 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290064Z.

Anonymous:1992:CFRa

[Ano92d] Anonymous. Contents of Fisheries Research, volume 13. Fisheries Research, 13(4):438–440, April 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783692900674.

Anonymous:1992:CFRb

[Ano92e] Anonymous. Contents of Fisheries Research, volume 14. Fisheries Research, 14(4):345–346, September 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290044T.

Anonymous:1992:EBa

[Ano92f] Anonymous. Editorial Board. Fisheries Research, 13(1):iii, January 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290028R.

Anonymous:1992:EBb

[Ano92g] Anonymous. Editorial Board. Fisheries Research, 14(1): iii, June 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900685.

Anonymous:1992:EBc

[Ano92h] Anonymous. Editorial Board. Fisheries Research, 15(1–2): iii, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290001A.

Anonymous:1992:PO

[Ano92i] Anonymous. Pages 1–196 (October 1992). Fisheries Research, 15(1–2):??, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1992:PJb

[Ano92j] Anonymous. Pages 1–90 (June 1992). Fisheries Research, 14 (1):??, June 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1992:PJa

[Ano92k] Anonymous. Pages 1–94 (January 1992). Fisheries Research, 13(1):??, January 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1992:PD

[Ano92l] Anonymous. Pages 197–299 (December 1992). Fisheries Research, 15(3):??, December 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1992:PS

[Ano92m] Anonymous. Pages 251–346 (September 1992). Fisheries Research, 14(4):??, September 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1992:PA

[Ano92n] Anonymous. Pages 353–440 (April 1992). Fisheries Research, 13(4):??, April 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1992:PF

[Ano92o] Anonymous. Pages 95–200 (February 1992). Fisheries Research, 13(2):??, February 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1993:A

[Ano93a] Anonymous. Announcement. Fisheries Research, 18(3–4):393, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783693901665.

Anonymous:1993:AIa

[Ano93b] Anonymous. Author index. Fisheries Research, 15(4):392– 394, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390092L.

Anonymous:1993:AIb

[Ano93c] Anonymous. Author index. Fisheries Research, 16(4):367– 368, May 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390148Z.

Anonymous:1993:AIc

[Ano93d] Anonymous. Author index. Fisheries Research, 17(3–4):384– 386, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783693901403.

Anonymous:1993:AId

[Ano93e] Anonymous. Author index. Fisheries Research, 18(3–4): 403–404, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783693901676.

Anonymous:1993:CFRa

[Ano93f] Anonymous. Contents of Fisheries Research, volume 15. Fisheries Research, 15(4):395–397, January 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390093M.

Anonymous:1993:CFRb

[Ano93g] Anonymous. Contents of fisheries research, volume 16. Fisheries Research, 16(4):369–371, May 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693901492.

Anonymous:1993:CFRc

[Ano93h] Anonymous. Contents of fisheries research, volume 17. Fisheries Research, 17(3–4):387–389, August 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390141S.

Anonymous:1993:EBa

[Ano93i] Anonymous. Editorial Board. Fisheries Research, 16(1):iii, February 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390104F.

Anonymous:1993:EBb

[Ano93j] Anonymous. Editorial Board. Fisheries Research, 17(1–2): iii, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390001N.

Anonymous:1993:EBc

[Ano93k] Anonymous. Editorial Board. Fisheries Research, 18(1–2): iii, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783693900356.

Anonymous:1993:PF

[Ano93l] Anonymous. Pages 1–99 (February 1993). Fisheries Research, 16(1):??, February 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1993:PMa

[Ano93m] Anonymous. Pages 101–201 (March 1993). Fisheries Research, 16(2):??, March 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1993:PD

[Ano93n] Anonymous. Pages 163–408 (December 1993). Fisheries Research, 18(3–4):??, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1993:PAa

[Ano93o] Anonymous. Pages 203–285 (April 1993). Fisheries Research, 16(3):??, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1993:PAb

[Ano93p] Anonymous. Pages 251–390 (August 1993). Fisheries Research, 17(3–4):??, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1993:PMb

[Ano93q] Anonymous. Pages 287–371 (May 1993). Fisheries Research, 16(4):??, May 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1993:PJ

[Ano93r] Anonymous. Pages 301–397 (January 1993). Fisheries Research, 15(4):??, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1994:Ab

[Ano94a] Anonymous. Acknowledgements. Fisheries Research, 21(1–2): x, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900922.

Anonymous:1994:Aa

[Ano94b] Anonymous. Announcement. Fisheries Research, 19(3–4): 399, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900531.

Anonymous:1994:Ac

[Ano94c] Anonymous. Appendix 1. Fisheries Research, 21(1–2):305–307, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694901104.

Anonymous:1994:Ad

[Ano94d] Anonymous. Appendix 2. Fisheries Research, 21(1–2):309–311, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694901112.

Anonymous:1994:AIa

[Ano94e] Anonymous. Author index. Fisheries Research, 19(3–4):401– 403, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369490054X.

Anonymous:1994:AIb

[Ano94f] Anonymous. Author index. Fisheries Research, 20(4):297– 320, October 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900752.

Anonymous:1994:CFR

[Ano94g] Anonymous. Contents of Fisheries Research, volume 19. Fisheries Research, 19(3–4):404–405, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900558.

Anonymous:1994:EBa

[Ano94h] Anonymous. Editorial Board. Fisheries Research, 19(1–2):iii, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900108.

Anonymous:1994:EBb

[Ano94i] Anonymous. Editorial Board. Fisheries Research, 20(1): iii, June 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900019.

Anonymous:1994:EBc

[Ano94j] Anonymous. Editorial Board. Fisheries Research, 21(1–2):vii, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900906.

Anonymous:1994:PF

[Ano94k] Anonymous. Pages 1–191 (February 1994). Fisheries Research, 19(1–2):??, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1994:PJ

[Ano94l] Anonymous. Pages 1–95 (June 1994). Fisheries Research, 20 (1):??, June 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1994:PAa

[Ano94m] Anonymous. Pages 193–406 (April 1994). Fisheries Research, 19(3–4):??, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1994:POa

[Ano94n] Anonymous. Pages 293–381 (October 1994). Fisheries Research, 20(4):??, October 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1994:PAb

[Ano94o] Anonymous. Pages 97–291 (August 1994). Fisheries Research, 20(2–3):??, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1994:POb

[Ano94p] Anonymous. Publication overview. Fisheries Research, 20(4): 293, October 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900744.

Anonymous:1994:SI

[Ano94q] Anonymous. Subject index. Fisheries Research, 20(4):323– 381, October 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900760.

Anonymous:1995:AIV

[Ano95a] Anonymous. Author index, volumes 21–24. Fisheries Research, 24(4):351–358, November 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695900284.

Anonymous:1995:CFRb

[Ano95b] Anonymous. Contents of Fisheries Research, volume 22. Fisheries Research, 22(3–4):325–326, March 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695900276.

Anonymous:1995:CFRc

[Ano95c] Anonymous. Contents of Fisheries Research, volume 23. Fisheries Research, 23(3–4):385–386, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695900322.

Anonymous:1995:CFRd

[Ano95d] Anonymous. Contents of Fisheries Research, volume 24. Fisheries Research, 24(4):373–375, November 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695900306.

Anonymous:1995:CFRa

[Ano95e] Anonymous. Contents of fisheries research volume 21of fisheries research. Fisheries Research, 21(3–4):483–485, January 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695900004.

Anonymous:1995:EBa

[Ano95f] Anonymous. Editorial Board. Fisheries Research, 22(1–2):iii, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695900047.

Anonymous:1995:EBb

[Ano95g] Anonymous. Editorial Board. Fisheries Research, 23(1–2): iii, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695900012.

Anonymous:1995:EBc

[Ano95h] Anonymous. Editorial Board. Fisheries Research, 24(1): iii, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369590008X.

Anonymous:1995:PF

[Ano95i] Anonymous. Pages 1–162 (February 1995). Fisheries Research, 22(1–2):??, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:PMb

[Ano95j] Anonymous. Pages 1–194 (May 1995). Fisheries Research, 23 (1–2):??, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:PJc

[Ano95k] Anonymous. Pages 1–89 (July 1995). Fisheries Research, 24 (1):??, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:PMa

[Ano95l] Anonymous. Pages 163–328 (March 1995). Fisheries Research, 22(3–4):??, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:PO

[Ano95m] Anonymous. Pages 181–272 (October 1995). Fisheries Research, 24(3):??, October 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:PJb

[Ano95n] Anonymous. Pages 195–388 (June 1995). Fisheries Research, 23(3–4):??, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:PN

[Ano95o] Anonymous. Pages 273–375 (November 1995). Fisheries Research, 24(4):??, November 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:PJa

[Ano95p] Anonymous. Pages 315–486 (January 1995). Fisheries Research, 21(3–4):??, January 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:PS

[Ano95q] Anonymous. Pages 91–179 (September 1995). Fisheries Research, 24(2):??, September 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1995:SWF

[Ano95r] Anonymous. Second World Fisheries Congress — Developing and Sustaining World Fisheries Resources, The State of the Science and Management: 28 July–2 August 1996 Brisbane, Qld., Australia. Fisheries Research, 24(1):89, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695900144.

Anonymous:1995:SIV

[Ano95s] Anonymous. Subject index, volumes 21–24. Fisheries Research, 24(4):359–372, November 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695900292.

Anonymous:1996:AR

[Ano96a] Anonymous. Acknowledgement to referees 1995. Fisheries Research, 25(1):97, January 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578369690020X.

Anonymous:1996:A

[Ano96b] Anonymous. Announcement. Fisheries Research, 25(3–4):387– 388, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696900028.

Anonymous:1996:AIV

[Ano96c] Anonymous. Author index, volumes 25–28. Fisheries Research, 28(4):429–436, December 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696900132.

Anonymous:1996:CFRa

[Ano96d] Anonymous. Contents of Fisheries Research, volume 25. Fisheries Research, 25(3–4):389–390, March 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369690003X.

Anonymous:1996:CFRb

[Ano96e] Anonymous. Contents of Fisheries Research, volume 26. Fisheries Research, 26(3–4):391–392, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696900077.

Anonymous:1996:CFRc

[Ano96f] Anonymous. Contents of Fisheries Research, volume 27. Fisheries Research, 27(4):279–280, July 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696900168.

Anonymous:1996:CFRd

[Ano96g] Anonymous. Contents of Fisheries Research, volume 28. Fisheries Research, 28(4):457–459, December 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696900156.

Anonymous:1996:EBa

[Ano96h] Anonymous. Editorial Board. Fisheries Research, 25(1):iii, January 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696900181.

Anonymous:1996:EBb

[Ano96i] Anonymous. Editorial Board. Fisheries Research, 26(1–2): iii, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369690017X.

Anonymous:1996:EBc

[Ano96j] Anonymous. Editorial Board. Fisheries Research, 27(1–3): iii, June 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696900235.

Anonymous:1996:EBd

[Ano96k] Anonymous. Editorial Board. Fisheries Research, 28(1):iii, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696900223.

Anonymous:1996:PJa

[Ano96l] Anonymous. Pages 1–104 (January 1996). Fisheries Research, 25(1):??, January 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PAb

[Ano96m] Anonymous. Pages 1–118 (August 1996). Fisheries Research, 28(1):??, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PJb

[Ano96n] Anonymous. Pages 1–139 (June 1996). Fisheries Research, 27(1–3):??, June 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PAa

[Ano96o] Anonymous. Pages 1–198 (April 1996). Fisheries Research, 26(1–2):??, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PF

[Ano96p] Anonymous. Pages 105–200 (February 1996). Fisheries Research, 25(2):??, February 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PS

[Ano96q] Anonymous. Pages 119–242 (September 1996). Fisheries Research, 28(2):??, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PJc

[Ano96r] Anonymous. Pages 141–282 (July 1996). Fisheries Research, 27(4):??, July 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PMb

[Ano96s] Anonymous. Pages 199–394 (May 1996). Fisheries Research, 26(3–4):??, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PMa

[Ano96t] Anonymous. Pages 203–392 (March 1996). Fisheries Research, 25(3–4):??, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PO

[Ano96u] Anonymous. Pages 243–359 (October 1996). Fisheries Research, 28(3):??, October 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:PD

[Ano96v] Anonymous. Pages 361–459 (December 1996). Fisheries Research, 28(4):??, December 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1996:RL

[Ano96w] Anonymous. Reading list. Fisheries Research, 25(2):200, February 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696900107.

Anonymous:1996:SIV

[Ano96x] Anonymous. Subject index, volumes 25–28. Fisheries Research, 28(4):437–455, December 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696900144.

Anonymous:1997:A

[Ano97a] Anonymous. Announcement. Fisheries Research, 32(1):i, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697900023.

Anonymous:1997:AIV

[Ano97b] Anonymous. Author index, volumes 29–33. Fisheries Research, 33(1–3):167–175, December 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697816402.

Anonymous:1997:CFRa

[Ano97c] Anonymous. Contents of Fisheries Research, volume 29. Fisheries Research, 29(3):299–300, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578369790000X.

Anonymous:1997:CFRb

[Ano97d] Anonymous. Contents of Fisheries Research, volume 30. Fisheries Research, 30(3):261–262, June 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697900084.

Anonymous:1997:CFRc

[Ano97e] Anonymous. Contents of fisheries research, volume 33. Fisheries Research, 33(1–3):203, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697816426.

Anonymous:1997:EBa

[Ano97f] Anonymous. Editorial Board. Fisheries Research, 29(1):iii, January 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369790014X.

Anonymous:1997:EBb

[Ano97g] Anonymous. Editorial Board. Fisheries Research, 30(1–2): iii, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697900035.

Anonymous:1997:EBc

[Ano97h] Anonymous. Editorial Board. Fisheries Research, 31(1–2): iii, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697900072.

Anonymous:1997:EBd

[Ano97i] Anonymous. Editorial Board. Fisheries Research, 32(1):iii, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697900011.

Anonymous:1997:EBe

[Ano97j] Anonymous. Editorial Board. Fisheries Research, 33(1–3):iii, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697900138.

Anonymous:1997:PJa

[Ano97k] Anonymous. Pages 1–100 (January 1997). Fisheries Research, 29(1):??, January 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PO

[Ano97l] Anonymous. Pages 1–106 (October 1997). Fisheries Research, 32(1):??, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PJc

[Ano97m] Anonymous. Pages 1–167 (July 1997). Fisheries Research, 31 (1–2):??, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PMb

[Ano97n] Anonymous. Pages 1–172 (May 1997). Fisheries Research, 30 (1–2):??, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PDb

[Ano97o] Anonymous. Pages 1–203 (December 1997). Fisheries Research, 33(1–3):??, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PF

[Ano97p] Anonymous. Pages 101–202 (February 1997). Fisheries Research, 29(2):??, February 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PN

[Ano97q] Anonymous. Pages 107–199 (1 November 1997). Fisheries Research, 32(2):??, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PAb

[Ano97r] Anonymous. Pages 169–277 (August 1997). Fisheries Research, 31(3):??, August 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PJb

[Ano97s] Anonymous. Pages 177–262 (June 1997). Fisheries Research, 30(3):??, June 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PDa

[Ano97t] Anonymous. Pages 201–272 (1 December 1997). Fisheries Research, 32(3):??, December 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PMa

[Ano97u] Anonymous. Pages 203–300 (March 1997). Fisheries Research, 29(3):??, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1997:PAa

[Ano97v] Anonymous. Publisher’s acknowledgement. Fisheries Research, 30(1–2):167, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783697900047.

Anonymous:1997:SIV

[Ano97w] Anonymous. Subject index, volumes 29–33. Fisheries Research, 33(1–3):177–201, December 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697816414.

Anonymous:1998:BRa

[Ano98a] Anonymous. Book review. Fisheries Research, 38(1):91–95, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001477.

Anonymous:1998:BRb

[Ano98b] Anonymous. Book review. Fisheries Research, 38(1):95–97, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001489.

Anonymous:1998:BRc

[Ano98c] Anonymous. Book review. Fisheries Research, 38(1):97–99, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001519.

Anonymous:1998:BRd

[Ano98d] Anonymous. Book review. Fisheries Research, 38(2):207–208, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001490.

Anonymous:1998:BRe

[Ano98e] Anonymous. Book review. Fisheries Research, 38(2):208–209, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001507.

Anonymous:1998:BRf

[Ano98f] Anonymous. Book review. Fisheries Research, 39(1):95–96, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001921.

Anonymous:1998:BRg

[Ano98g] Anonymous. Book review. Fisheries Research, 39(1):97, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001933.

Anonymous:1998:BRh

[Ano98h] Anonymous. Book review. Fisheries Research, 39(1):97, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001945.

Anonymous:1998:BRi

[Ano98i] Anonymous. Book review. Fisheries Research, 39(1):98, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001957.

Anonymous:1998:BRj

[Ano98j] Anonymous. Book review. Fisheries Research, 39(1):98, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001969.

Anonymous:1998:C

[Ano98k] Anonymous. Correspondence. Fisheries Research, 34(3):315– 316, April 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698000897.

Anonymous:1998:E

[Ano98l] Anonymous. Editorial. Fisheries Research, 39(2):105, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001763.

Anonymous:1998:Ia

[Ano98m] Anonymous. Index. Fisheries Research, 38(3):307–314, November 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698002045.

Anonymous:1998:Ib

[Ano98n] Anonymous. Index. Fisheries Research, 38(3):315–337, November 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698002057.

Anonymous:1998:PS

[Ano98o] Anonymous. Pages 1–100 (1 September 1998). Fisheries Research, 38(1):??, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PF

[Ano98p] Anonymous. Pages 1–102 (February 1998). Fisheries Research, 34(1):??, February 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PDa

[Ano98q] Anonymous. Pages 1–104 (1 December 1998). Fisheries Research, 39(1):??, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PMb

[Ano98r] Anonymous. Pages 1–154 (May 1998). Fisheries Research, 35 (1–2):??, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PAb

[Ano98s] Anonymous. Pages 1–312 (1 August 1998). Fisheries Research, 37(1–3):??, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PJa

[Ano98t] Anonymous. Pages 1–74 (1 June 1998). Fisheries Research, 36(1):??, June 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PO

[Ano98u] Anonymous. Pages 101–210 (1 October 1998). Fisheries Research, 38(2):??, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PMa

[Ano98v] Anonymous. Pages 103–214 (March 1998). Fisheries Research, 34(2):??, March 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PDb

[Ano98w] Anonymous. Pages 105–228 (15 December 1998). Fisheries Research, 39(2):??, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PMc

[Ano98x] Anonymous. Pages 155–252 (15 May 1998). Fisheries Research, 35(3):??, May 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PN

[Ano98y] Anonymous. Pages 211–340 (1 November 1998). Fisheries Research, 38(3):??, November 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PAa

[Ano98z] Anonymous. Pages 215–324 (April 1998). Fisheries Research, 34(3):??, April 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1998:PJb

[Ano98-27] Anonymous. Pages 75–288 (15 June 1998). Fisheries Research, 36(2–3):??, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:BRa

[Ano99a] Anonymous. Book review. Fisheries Research, 39(3):317–319, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001891.

Anonymous:1999:BRb

[Ano99b] Anonymous. Book review. Fisheries Research, 39(3):319–321, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001908.

Anonymous:1999:BRc

[Ano99c] Anonymous. Book review. Fisheries Research, 39(3):321–322, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369800191X.

Anonymous:1999:BRd

[Ano99d] Anonymous. Book review. Fisheries Research, 42(1–2):201– 205, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000387.

Anonymous:1999:E

[Ano99e] Anonymous. Editorial. Fisheries Research, 40(2):103–106, March 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698002069.

Anonymous:1999:PF

[Ano99f] Anonymous. Pages 1–102 (1 February 1999). Fisheries Research, 40(1):??, February 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PN

[Ano99g] Anonymous. Pages 1–104 (November 1999). Fisheries Research, 44(1):??, November 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PMb

[Ano99h] Anonymous. Pages 1–110 (1 May 1999). Fisheries Research, 41(1):??, May 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PAb

[Ano99i] Anonymous. Pages 1–206 (August 1999). Fisheries Research, 42(1–2):??, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PO

[Ano99j] Anonymous. Pages 1–252 (October 1999). Fisheries Research, 43(1–3):??, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PMa

[Ano99k] Anonymous. Pages 103–206 (1 March 1999). Fisheries Research, 40(2):??, March 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PD

[Ano99l] Anonymous. Pages 105–200 (December 1999). Fisheries Research, 44(2):??, December 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PJb

[Ano99m] Anonymous. Pages 111–220 (June 1999). Fisheries Research, 41(2):??, June 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PAa

[Ano99n] Anonymous. Pages 207–310 (1 April 1999). Fisheries Research, 40(3):??, April 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PS

[Ano99o] Anonymous. Pages 207–310 (September 1999). Fisheries Research, 42(3):??, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PJc

[Ano99p] Anonymous. Pages 221–330 (July 1999). Fisheries Research, 41(3):??, July 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:1999:PJa

[Ano99q] Anonymous. Pages 229–324 (1 January 1999). Fisheries Research, 39(3):??, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Al-Ogaily:1990:BGP

[AOH90] Soliman M. Al-Ogaily and Altaf Hussain. Biology of grunt Plectorhynchus pictus (Thunberg) 1972, (Haemulidae, Teleostei, Percoidei) from the Red Sea (Jizan area). Fisheries Research, 9(2):119–130, August 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783690900595.

Adkison:1996:RBM

[AP96] Milo D. Adkison and Randall M. Peterman. Results of Bayesian methods depend on details of implementation: an example of estimating salmon escapement goals. Fisheries Research, 25(2):155–170, February 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369500405X.

Argyle:1992:ATA

[Arg92] Ray L. Argyle. Acoustics as a tool for the assessment of Great Lakes forage fishes. Fisheries Research, 14(2–3):179– 196, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290052U.

Arkhipkin:1993:AGS

[Ark93] Alexander Arkhipkin. Age, growth, stock structure and migratory rate of pre-spawning short-finned squid Illex argentinus based on statolith ageing investigations. Fisheries Research, 16(4):313–338, May 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390144V.

Ashok:1991:PBT

[AS91] M. K. Ashok and D. S. Sheshappa. Preliminary bottom trawling experiments with a newly designed separator fish trawl. Fisheries Research, 12(2):171–180, October 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190040M.

Aspholm:1995:ASR

[Asp95] Paul E. Aspholm. Anisakis simplex Rudolphi, 1809, infection in fillets of Barents Sea cod Gadus morhua L. Fisheries Research, 23(3–4):375–379, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400348Z.

Ashraf:1991:CTM

[ATJ91] M. Ashraf, Jaleel Tariq, and M. Jaffar. Contents of trace metals in fish, sediment and water from three freshwater reservoirs on the Indus River, Pakistan. Fisheries Research, 12(4):355–364, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190018B.

Avsar:1994:SDS

[Avs94] Dursun Avsar. A stock differentiation study of the sprat (Sprattus sprattus phalericus Risso) off the southern coast of the Black Sea. Fisheries Research, 19(3–4):363–378, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900507.

Avsar:1995:PPS

[Avs95] Dursun Avsar. Population parameters of sprat (Sprattus sprattus phalericus RISSO) from the Turkish Black Sea coast. Fisheries Research, 21(3–4):437–453, January 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694002888.

Aneer:1990:MTP

[AW90] Gunnar Aneer and Lars Westin. Migration of turbot (Psetta maxima L.) in the northern Baltic proper. Fisheries Research, 9(4):307–315, October 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783690900492.

Azevedo:1995:SAB

[Aze95] Manuela Azevedo. A statistical analysis of black monkfish catch rates in ICES division IXa. Fisheries Research, 24(4):281–289, November 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695003821.

Azuma:1995:BME

[Azu95] Teruo Azuma. Biological mechanisms enabling sympatry between salmonids with special reference to sockeye and chum salmon in oceanic waters. Fisheries Research, 24(4):291–300, November 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695003833.

Bethke:1999:CST

[BACH99] Eckhard Bethke, Fredrik Arrhenius, Massimiliano Cardinale, and Nils Håkansson. Comparison of the selectivity of three pelagic sampling trawls in a hydroacoustic survey. Fisheries Research, 44(1):15–23, November 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000545.

Baelde:1991:AAD

[Bae91] P. Baelde. Assessment of the Australian deep-water royal red prawn stock using commercial catch and effort data. Fisheries Research, 12(3):243–258, November 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190097Y.

Baticados:1998:FCC

[BAG98] Didi B. Baticados, Renato F. Agbayani, and Francisco E. Gentoral. Fishing cooperatives in capiz, central Philippines: their importance in managing fishery resources. Fisheries Research, 34(2):137–149, March 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000908.

Briggs:1999:CIM

[BAR99] R. P. Briggs, M. J. Armstrong, and D. Rihan. The consequences of an increase in mesh size in the Irish Sea Nephrops fishery: an experimental approach. Fisheries Research, 40(1):43–53, February 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698002240.

Basu:1999:ARP

[Bas99] T. K. Basu. Application of Rao’s polynomial growth curve model to analyse the length growth data of a freshwater major carp of India, Catla catla (Ham) during the early developmental period. Fisheries Research, 42(3):303–307, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369900051X. See erra-

tum [Bas00].

Basu:2000:EAR

[Bas00] T. K. Basu. Erratum to “Application of Rao’s polynomial growth curve model to analyse the length growth data of a freshwater major carp of India, Catla catla (Ham) during the early developmental period”: [Fish. Res. 42 (1999)

303–307]. Fisheries Research, 45(2):201, March 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001752. See [Bas99].

Booth:1997:MPP

[BB97] Anthony J. Booth and Colin D. Buxton. Management of the panga Pterogymnus laniarius (Pisces: Sparidae) on the Agulhas Bank, South Africa using per-recruit models. Fisheries Research, 32(1):1–11, October 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000453.

Bjerkeng:1991:FSD

[BBBF91] B. Bjerkeng, R. Borgstrøm, Å. Brabrand, and B. Faafeng. Fish size distribution and total fish biomass estimated by hydroacoustical methods: a statistical approach. Fisheries Research, 11(1):41–73, February 1991. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190032B.

Basson:1996:AMT

[BBC^{+}96] M. Basson, J. R. Beddington, J. A. Crombie, S. J. Holden, L. V. Purchase, and G. A. Tingley. Assessment and management techniques for migratory annual squid stocks: the Illex argentinus fishery in the Southwest Atlantic as an example. Fisheries Research, 28(1):3–27, August 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369600481X.

Brabrand:1994:EIM

[BBF94] Åge Brabrand, Tor Andreas Bakke, and Bjørn Andreas Faafeng. The ectoparasite Ichthyophthirius multifiliis and the abundance of roach (Rutilus rutilus): larval fish epidemics in relation to host behaviour. Fisheries Research, 20(1):49– 61, June 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900051.

Bird:1993:SSP

[BC93] D. J. Bird and I. G. Cowx. The selection of suitable pulsed currents for electric fishing in waters. Fisheries Research, 18(3–4):363–376, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693901632.

Buxton:1993:QWF

[BD93] N. G. Buxton and J. T. DeAlteris. Quantification of water flow and net drag in midwater trawls. Fisheries Research, 16(2):165–178, March 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390050H.

Blanc:1998:ASC

[BD98] A. Blanc and J. Daguzan. Artificial surfaces for cuttlefish eggs (Sepia officinalis L.) in Morbihan Bay, France. Fisheries Research, 38(3):225–231, November 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001647.

Beaumont:1995:BRM

[Bea95] A. Beaumont. Book review: Molecular genetics in fisheries: G. R. Carvalho and T. J. Pitcher (Editors), Chapman and Hall, London, 1995, 141 pp., price £24.99, ISBN 0-412-62950-X. Fisheries Research, 24(3):270–272, October 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695900209.

Begossi:1998:PRF

[Beg98] Alpina Begossi. Property rights for fisheries at different scales: applications for conservation in Brazil. Fisheries Research, 34(3):269–278, April 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000854.

Brewer:1996:AEF

[BEMW96] David Brewer, Steve Eayrs, Richard Mounsey, and You-Gan Wang. Assessment of an environmentally friendly, semipelagic fish trawl. Fisheries Research, 26(3–4):225–237, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004335.

Bergstad:1997:DTJ

[BF97] Odd Aksel Bergstad and Arild Folkvord. Dispersal of tagged juvenile turbot Scophthalmus maximus on the Norwegian Skagerrak coast. Fisheries Research, 29(3):211–215, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005437.

Begg:1999:SIR

[BFP99] Gavin A. Begg, Kevin D. Friedland, and John B. Pearce. Stock identification and its role in stock assessment and fisheries management: an overview. Fisheries Research, 43(1–3):1–8, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000624.

Bhaskar:1994:CLP

[Bha94] M. Bhaskar. Changes in the liver protein fractions of Tilapia mossambica (Peters) during acclimation to low and high pH media. Fisheries Research, 19(1–2):179–186, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900221.

Bernard:1999:HID

[BHF99] David R. Bernard, James J. Hasbrouck, and Steven J. Fleischman. Handling-induced delay and downstream movement of adult Chinook salmon in rivers. Fisheries Research, 44(1):37– 46, November 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000569.

Begg:1999:RLH

[BHS99] Gavin A. Begg, Jonathan A. Hare, and Daniel D. Sheehan. The role of life history parameters as indicators of stock structure. Fisheries Research, 43(1–3):141–163, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000715.

Booth:1997:CMU

[BI97] Anthony J. Booth and Edwin D. Isted. Comparison of methods used to convert length-frequency data. Fisheries Research, 29(3):271–276, March 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696005383.

Bishara:1992:FLM

[Bis92] Nabila F. Bishara. Fisheries of Lake Manzalah. Fisheries Research, 14(1):71–82, June 1992. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783692900744.

Broadhurst:1994:RCJ

[BK94] M. K. Broadhurst and S. J. Kennelly. Reducing the bycatch of juvenile fish (mulloway Argyrosomus hololepidotus) using square-mesh panels in codends in the Hawkesbury River prawn-trawl fishery, Australia. Fisheries Research, 19(3–4):321–331, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783694900477.

Broadhurst:1996:ECC

[BK96] M. K. Broadhurst and S. J. Kennelly. Effects of the circumference of codends and a new design of square-mesh panel in reducing unwanted by-catch in the New South Wales oceanic prawn-trawl fishery, Australia. Fisheries Research, 27(4):203– 214, July 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004696.

Broadhurst:1997:SEJ

[BKB97] M. K. Broadhurst, S. J. Kennelly, and D. T. Barker. Simulated escape of juvenile sand whiting (Sillago ciliata, Cuvier) through square-meshes: Effects on scale-loss and survival. Fisheries Research, 32(1):51–60, October 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000374.

Beyer:1999:MDS

[BKH99] J. E. Beyer, C. H. Kirchner, and J. A. Holtzhausen. A method to determine size-specific natural mortality applied to westcoast steenbras (Lithognathus aureti) in Namibia. Fisheries Research, 41(2):133–153, June 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000119.

Broadhurst:1996:AMC

[BKI96] M. K. Broadhurst, S. J. Kennelly, and B. Isaksen. Assessments of modified codends that reduce the by-catch of fish in two estuarine prawn-trawl fisheries in New South Wales, Australia. Fisheries Research, 27(1–3):89–111, June 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004572.

Blaber:1990:BRM

[Bla90] Stephen J. M. Blaber. Book review: The mackerel, its biology, assessment and the management of a fishery: Stephen J. Lockwood, Fishing News Books Ltd., Farnham, Surrey, England, 1988, 181 pp., £12.50, ISBN 0-85238-156-5. Fisheries Research, 8(3):294–295, January 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090030Y.

Blaxter:1992:BRE

[Bla92] J. H. S. Blaxter. Book review: Early life history of fish — an energetics approach: E. Kamler, Chapman and Hall, London, 1992, 267 pp., £35, ISBN 0-442-31428-0. Fisheries Research, 15(1–2):187–189, October 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290014K.

Baker:1991:GGM

[BLQ91] Timothy T. Baker, Robert Lafferty, and Terrance J. Quinn. A general growth model for mark-recapture data. Fisheries Research, 11(3–4):257–281, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190005Z.

Bakke:1993:CSN

[BM93] T. A. Bakke and K. MacKenzie. Comparative susceptibility of native Scottish and Norwegian stocks of Atlantic salmon, Salmo salar L., to Gyrodactylus salaris Malmberg: Laboratory experiments. Fisheries Research, 17(1–2):69–85, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390008U.

Bowering:1997:CPA

[BMB97] W. R. Bowering, M. J. Morgan, and W. B. Brodie. Changes in the population of American plaice (Hippoglossoides platessoides) off Labrador and northeastern Newfoundland: a collapsing stock with low exploitation. Fisheries Research, 30(3):199–216, June 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696005681.

Bowering:1998:AND

[BMB98] W. R. Bowering, R. K. Misra, and W. B. Brodie. Application of a newly developed statistical procedure to morphometric data from American plaice (Hippoglossoides platessoides) in the Canadian Northwest Atlantic. Fisheries Research, 34(2):191– 203, March 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000945.

Blaber:1990:DLF

[BMR^{+}90] S. J. M. Blaber, D. A. Milton, N. J. F. Rawlinson, G. Tiroba, and P. V. Nichols. Diets of lagoon fishes of the Solomon Islands: Predators of tuna baitfish and trophic effects of baitfishing on the subsistence fishery. Fisheries Research, 8(3):263– 286, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090027S.

Batista:1992:CSL

[BN92] I. Batista and M. L. Nunes. Characterisation of shark liver oils. Fisheries Research, 14(4):329–334, September 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290041Q.

Bonfil:1997:SSR

[Bon97] Ramón Bonfil. Status of shark resources in the southern Gulf of Mexico and Caribbean: implications for management. Fisheries Research, 29(2):101–117, February 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369600536X.

Booke:1999:SCR

[Boo99] Henry E. Booke. The stock concept revisited: perspectives on its history in fisheries. Fisheries Research, 43(1–3):9–11, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000636.

Bowering:1990:SWF

[Bow90] W. R. Bowering. Spawning of witch flounder (Glyptocephalus cynoglossus L.) in the Newfoundland-Labrador area of the northwest Atlantic as a function of depth and water temperature. Fisheries Research, 9(1):23–39, July 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090038W.

Boyle:1990:CBF

[Boy90] P. R. Boyle. Cephalopod biology in the fisheries context. Fisheries Research, 8(4):303–321, May 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090001C.

Boyle:1994:P

[BP94a] P. R. Boyle and G. J. Pierce. Preface. Fisheries Research, 21(1–2):ix, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783694900914.

Boyle:1994:FBN

[BP94b] Peter R. Boyle and Graham J. Pierce. Fishery biology of northeast Atlantic squid: an overview. Fisheries Research, 21(1–2):1–15, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900930.

Beaumont:1996:AVG

[BP96] A. R. Beaumont and S. M. J. Pether. Allozyme variation and gene flow between cockle Cerastoderma edule populations in southern United Kingdom. Fisheries Research, 28(3):263–275, October 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783696005061.

Booth:1998:ERP

[BP98] Anthony J. Booth and André E. Punt. Evidence for rebuilding in the panga stock on the Agulhas Bank, South Africa. Fisheries Research, 34(2):103–121, March 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000921.

Boyle:1990:SES

[BPD90] P. R. Boyle, G. J. Pierce, and J. S. W. Diack. Sources of evidence for salmon in the diet of seals. Fisheries Research, 10(1–2):137–150, December 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090019R.

Bergh:1990:SDC

[BPSW90] M. O. Bergh, E. K. Pikitch, J. R. Skalski, and J. R. Wallace. The statistical design of comparative fishing experiments. Fisheries Research, 9(2):143–163, August 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090061Y.

Bogstad:1995:CES

[BPV95] Bjarte Bogstad, Michael Pennington, and Jon Helge Vølstad. Cost-efficient survey designs for estimating food consumption by fish. Fisheries Research, 23(1–2):37–46, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400341S.

Buijse:1992:SAS

[BPvD^{+}92] Anthonie D. Buijse, Jos S. Pet, Wim L. T. van Densen, Marcel A. M. Machiels, and Rudy Rabbinge. A sizeand age-structured simulation model for evaluating management strategies in a multispecies gill net fishery. Fisheries Research, 13(2):95–117, February 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290021K.

Brannon:1993:POH

[Bra93] E. L. Brannon. The perpetual oversight of hatchery programs. Fisheries Research, 18(1–2):19–27, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783693900378.

Beddington:1990:SAP

[BRCK90] J. R. Beddington, A. A. Rosenberg, J. A. Crombie, and G. P. Kirkwood. Stock assessment and the provision of management advice for the short fin squid fishery in Falkland Islands waters. Fisheries Research, 8(4):351–365, May 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090004F.

Brewer:1998:ABR

[BREB98] David Brewer, Nicholas Rawlinson, Steve Eayrs, and Charis Burridge. An assessment of bycatch reduction devices in a tropical Australian prawn trawl fishery. Fisheries Research, 36(2–3):195–215, June 15, 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698000964.

Briggs:1992:ANS

[Bri92] R. P. Briggs. An assessment of nets with a square mesh panel as a whiting conservation tool in the Irish Sea Nephrops fishery. Fisheries Research, 13(2):133–152, February 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290023M.

Bristow:1993:PNF

[Bri93] Glenn A. Bristow. Parasites of Norwegian freshwater salmonids and interactions with farmed salmon — a review. Fisheries Research, 17(1–2):219–227, June 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390021X.

Briggs:1995:VNI

[Bri95] R. P. Briggs. Variability in northwest Irish Sea Nephrops populations. Fisheries Research, 23(1–2):175–187, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400332Q.

Brown:1997:BRB

[Bro97a] R. W. Brown. Book review: Bottom soils, sediment, and pond aquaculture: By C. E. Boyd, Chapman and Hall, 1995, 348 pp., £49, Hardback, ISBN 0-412-06941-5. Fisheries Research, 32(2):197, November 1, 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697900126.

Brown:1997:BRC

[Bro97b] R. W. Brown. Book review: Computers in fisheries research: Edited by Bernard A. Megrey and Erlend Moksness, Chapman and Hall, London, UK, 1996, 254 pp., £59 (hardback), ISBN 0-412-59550-8. Fisheries Research, 32(2):198–199, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000271.

Brown:1997:BRF

[Bro97c] R. W. Brown. Book review: Fisheries biology, assessment and management: By Michael King, Fishing News Books, 1995, 352 pp., £24.95, paperback, ISBN 0-85238-223-5. Available from Fishing News Books, Blackwell Science Ltd., Oxford, UK. Fisheries Research, 32(2):197, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697900114.

Brown:1997:BRH

[Bro97d] R. W. Brown. Book review: Hyperiid amphipods (Amphipoda, Hyperiidea) of the world’s oceans: By M. E. Vinogradov, A. F. Volkov and T. N. Semenora, Science Publishers, USA, 1996, 632 pp., $178, hardback, ISBN 1-886106-48-7. Fisheries Research, 32(2):197–198, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000283.

Brown:1997:BRR

[Bro97e] R. W. Brown. Book review: Rites and passages: the experience of American whaling, 1830–1870: By Margaret S. Creighton, Cambridge University Press, UK, 1995, 233 pp., £14.95 (US$17.95) paperback, ISBN 0-521-48448-0. Fisheries Research, 32(2):197, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697900102.

Bruno:1990:BRS

[Bru90] D. W. Bruno. Book review: Systemic pathology of fish: H. W. Ferguson (Editor). Iowa State University Press, Ames, IA, 1989, 276 pp., $54.95, ISBN 0-8138-0147-8. Fisheries Research, 9(2):187, August 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783690900643.

Bruno:1993:PCW

[Bru93] David W. Bruno. Pathological conditions of wild salmonids. Fisheries Research, 17(1–2):1–2, June 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390002O.

Brown:1992:SFF

[BSA92] Bradford E. Brown, Wilber R. Seidel, and Thomas Azarovitz. Successes and failures of fisheries acoustics in the management of United States East Coast marine fisheries. Fisheries Research, 14(2–3):129–134, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290048X.

Berdnikov:1999:TMB

[BSVC99] S. V. Berdnikov, V. V. Selyutin, V. V. Vasilchenko, and J. F. Caddy. Trophodynamic model of the Black and Azov Sea pelagic ecosystem: consequences of the comb jelly, Mnemiopsis leydei, invasion. Fisheries Research, 42(3):261–289, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000491.

Bravo:1999:SME

[BSVH99] R. Bravo, M. C. Soriguer, N. Villar, and J. A. Hernando. A simple method to estimate the significance level of the catch probability in the catch removal method in river fish populations. Fisheries Research, 44(2):179–182, December 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000818.

Bullough:1998:CDW

[BTBP98] L. W. Bullough, W. R. Turrell, P. Buchan, and I. G. Priede. Commercial deep water trawling at sub-zero temperatures — observations from the Faroe–Shetland Channel. Fisheries Research, 39(1):33–41, December 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001702.

Bublitz:1996:QEF

[Bub96] Christopher G. Bublitz. Quantitative evaluation of flatfish behavior during capture by trawl gear. Fisheries Research, 25(3–4):293–304, March 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695004319.

Bucke:1993:SDA

[Buc93] David Bucke. The significance of diseases and anomalies in wild salmonids. Fisheries Research, 17(1–2):209–217, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783693900208.

Buxton:1992:AYP

[Bux92] Colin D. Buxton. The application of yield-per-recruit models to two South African sparid reef species, with special consideration to sex change. Fisheries Research, 15(1–2):1–16, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290002B.

Begg:1999:HAF

[BW99] Gavin A. Begg and John R. Waldman. An holistic approach to fish stock identification. Fisheries Research, 43(1–3):35– 44, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369900065X.

Chopin:1995:CFE

[CA95] F. S. Chopin and T. Arimoto. The condition of fish escaping from fishing gears — a review. Fisheries Research, 21(3–4):315–327, January 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400301C.

Caddy:1999:RRV

[CA99] John F. Caddy and Alvaro J. Abella. Reconstructing reciprocal M vectors from length cohort analysis (LCA) of commercial size frequencies of hake, and fine mesh trawl surveys over the same grounds. Fisheries Research, 41(2):169– 175, June 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000156.

Cubillos:1999:EEN

[CAB99] Luis A. Cubillos, Ruben Alarcón, and Antonio Brante. Empirical estimates of natural mortality for the Chilean hake (Merluccius gayi): evaluation of precision. Fisheries Research, 42(1–2):147–153, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000429.

Caddy:1996:RSP

[Cad96] J. F. Caddy. Regime shifts and paradigm changes: is there still a place for equilibrium thinking? Fisheries Research, 25(3–4):219–230, March 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695004432.

Caddy:1997:CBM

[Cad97] John Caddy. Checks and balances in the management of marine fish stocks: organizational requirements for a limit reference point approach. Fisheries Research, 30(1–2):1–15, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005656.

Chopin:1996:CSR

[CAI96] F. S. Chopin, T. Arimoto, and Y. Inoue. A comparison of the stress response and mortality of sea bream Pagrus major captured by hook and line and trammel net. Fisheries Research, 28(3):277–289, October 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696005127.

Cardador:1993:NLN

[Car93] Fátima Cardador. Norway lobster (Nephrops norvegicus) from the southwest and south of Portugal — estimation of the effects of changing trawl mesh size and fishing effort by length cohort analysis. Fisheries Research, 17(3–4):259–271, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390129U.

Caveriviere:1993:SMC

[Cav93] A. Caverivière. Some methodological considerations on delta distribution, stratification and tow duration, for trawl surveys carried out in West Africa. Fisheries Research, 16(3):223– 237, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390095O.

Cazorla:1996:FCS

[Caz96] A. Lopez Cazorla. The food of Cynoscion striatus (Cuvier) (Pisces: Sciaenidae) in the Bahía blanca area, Argentina. Fisheries Research, 28(4):371–379, December 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005152.

Comeau:1998:DDV

[CB98] S. Comeau and D. Boisclair. Day-to-day variation in fish horizontal migration and its potential consequence on estimates of trophic interactions in lakes. Fisheries Research, 35(1–2): 75–81, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698000617.

Cotter:1997:SAT

[CBB97] A. J. R. Cotter, T. W. Boon, and C. G. Brown. Statistical aspects of trials of a separator trawl using a twin rig trawler. Fisheries Research, 29(1):25–32, January 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005218.

Colloca:1997:BSD

[CCA97] F. Colloca, M. Cardinale, and G. D. Ardizzone. Biology, spatial distribution and population dynamics of Lepidotrigla cavillone (Pisces: Triglidae) in the Central Tyrrhenian Sea. Fisheries Research, 32(1):21–32, October 1997.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000416.

Carbonell:1999:RSA

[CCD^{+}99] Aina Carbonell, Maria Carbonell, Montserrat Demestre, Antoni Grau, and Sebastià Monserrat. The red shrimp Aristeus antennatus (Risso, 1816) fishery and biology in the Balearic Islands, Western Mediterranean. Fisheries Research, 44(1):1–13, November 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369900079X.

Chu:1995:MAC

[CCHW95] K. H. Chu, Q. C. Chen, L. M. Huang, and C. K. Wong. Morphometric analysis of commercially important penaeid shrimps from the zhujiang estuary, China. Fisheries Research, 23(1–2): 83–93, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400342T.

Course:1999:SPS

[CCJ^{+}99] G. P. Course, A. J. R. Cotter, R. O. Jolliffe, B. J. Kay, and M. S. Rolfe. Safe practices for sampling commercial trawler catches at sea. Fisheries Research, 41(1):99–104, May 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000041.

Chicharo:1998:DNC

[CCV^{+}98] M. Alexandra Chıćharo, Luis Chıćharo, Luis Valdés, Eduardo López-Jamar, and Pedro Ré. Does the nutritional condition limit survival potential of sardine Sardina pilchardus (Walbaum, 1792) larvae off the north coast of Spain? RNA/DNA ratios and their variability. Fisheries Research, 39(1):43–54, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001672.

Castro:1990:ETS

[CD90] Kathleen M. Castro and Joseph T. DeAlteris. Effects of trap saturation and species interaction on the capture of Callinectes spp. crabs in the Guayas Estuary. Fisheries Research, 8(3):223–232, January 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090023O.

Caddy:1996:FEL

[CD96] John Caddy and Omar Defeo. Fitting the exponential and logistic surplus yield models with mortality data: some explorations and new perspectives. Fisheries Research, 25(1):39– 62, January 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004297.

Coelho:1997:CSD

[CDB^{+}97] M. Coelho, P. Domingues, E. Balguerias, M. Fernandez, and J. P. Andrade. A comparative study of the diet of Loligo vulgaris (Lamarck, 1799) (Mollusca: Cephalopoda) from the south coast of Portugal and the Saharan Bank (Central-East Atlantic). Fisheries Research, 29(3):245–255, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005401.

Castilho:1993:AGM

[CDE93] R. Castilho, M. T. Dinis, and K. Erzini. Age and growth of megrim Lepidorhombus boscii, Risso of the Portuguese continental coast. Fisheries Research, 16(4):339–346, May 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390145W.

Cadrin:1999:UIP

[CF99] Steven X. Cadrin and Kevin D. Friedland. The utility of image processing techniques for morphometric analysis and stock identification. Fisheries Research, 43(1–3):129–139, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000703.

Chu-fa:1996:EFE

[CfyCaL^{+}96] Tsai Chu-fa, Pei yu Chen, Ming anne Lee, Kwang yaw Hsia, and Kuo tien Lee. Effects of fishing effort on stock size and catch of larval anchovy in coastal waters of southwestern Taiwan. Fisheries Research, 28(1):71–83, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783696004742.

Chen:1997:ADS

[CG97] Y. Chen and G. N. G. Gordon. Assessing discarding at sea using a length-structured yield-per-recruit model. Fisheries Research, 30(1–2):43–55, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696005590.

Cortez:1999:GOM

[CGG99a] Tito Cortez, Angel F. González, and Angel Guerra. Growth of Octopus mimus (Cephalopoda, Octopodidae) in wild populations. Fisheries Research, 42(1–2):31–39, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000405.

Cortez:1999:GCO

[CGG99b] Tito Cortez, Angel F. González, and Angel Guerra. Growth of cultured Octopus mimus (Cephalopoda, Octopodidae). Fisheries Research, 40(1):81–89, February 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002033.

Charles:1998:LUF

[Cha98] Anthony T. Charles. Living with uncertainty in fisheries: analytical methods, management priorities and the Canadian groundfishery experience. Fisheries Research, 37(1–3):37–50, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001258.

Chen:1996:MCS

[Che96] Y. Chen. A Monte Carlo study on impacts of the size of subsample catch on estimation of fish stock parameters. Fisheries Research, 26(3–4):207–223, May 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004475.

Churchill:1998:LUI

[Chu98] Robin Churchill. Legal uncertainties in international high seas fisheries management. Fisheries Research, 37(1–3):225–237, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001398.

Chopin:1996:DCM

[CIA96] Frank Chopin, Yoshihiro Inoue, and Takafumi Arimoto. Development of a catch mortality model. Fisheries Research, 25(3–4):377–382, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783695004173.

Ciocco:1991:DIG

[Cio91] Néstor Fernando Ciocco. Differences in individual growth rate among scallop (Chlamys tehuelcha (d’Orb.)) populations from San José Gulf (Argentina). Fisheries Research, 12(1):31–42, September 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190047J.

Crozier:1994:MEA

[CK94] W. W. Crozier and G. J. A. Kennedy. Marine exploitation of Atlantic salmon (Salmo salar L.) from the River Bush, Northern Ireland. Fisheries Research, 19(1–2):141–155, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900205.

Castro:1995:SSS

[CL95] Margarida Castro and William Lawing. A study of sampling strategies for estimating growth parameters in fish populations. Fisheries Research, 22(1–2):59–75, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400309K.

Chen:1997:MLD

[CLGK97] Y. Chen, G. W. Liggins, K. J. Graham, and S. J. Kennelly. Modelling the length-dependent offshore distribution of redfish, Centroberyx affinis. Fisheries Research, 29(1):39–54, January 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005176.

Chen:1997:CCS

[CLZS97] Dagang Chen, Qun Liu, Xiaoqi Zeng, and Zhenming Su. Catch composition and seasonal variation of setnet fisheries in the Yellow and Bohai Seas. Fisheries Research, 32(1):61–68, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369700043X.

Caddy:1996:TTR

[CM96] John F. Caddy and Jacek Majkowski. Tuna and trees: a reflection on a long-term perspective for tuna fishing around floating logs. Fisheries Research, 25(3–4):369–376, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004491.

Cole:1997:PES

[CM97] Michael B. Cole and John R. Moring. Potential error with in situ surveys of smallmouth bass, Micropterus dolomieu Lacépède, as determined by radio-telemetry. Fisheries Research, 31(3):269–273, August 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000477.

Cardinale:1999:CSM

[CM99a] M. Cardinale and J. Modin. Changes in size-at-maturity of Baltic cod (Gadus morhua) during a period of large variations in stock size and environmental conditions. Fisheries Research, 41(3):285–295, July 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000211.

Chen:1999:GMC

[CM99b] Y. Chen and L. G. S. Mello. Growth and maturation of cod (Gadus morhua) of different year classes in the Northwest Atlantic, NAFO subdivision 3Ps. Fisheries Research, 42(1–2):87– 101, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000363.

Chale-Matsau:1999:AGQ

[CMGB99] J. R. Chale-Matsau, A. Govender, and L. E. Beckley. Age and growth of the queen mackerel Scomberomorus plurilineatus from KwaZulu-Natal, South Africa. Fisheries Research, 44(2):121–127, December 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000600.

Crozier:1997:CPN

[CMK97] W. W. Crozier, I. J. J. Moffett, and G. J. A. Kennedy. Comparative performance of native and non-native strains of Atlantic salmon (Salmo salar L.) ranched from the River Bush, Northern Ireland. Fisheries Research, 32(1):81–88, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000313.

Casey:1992:SSM

[CNW92] J. Casey, M. D. Nicholson, and S. Warnes. Selectivity of square mesh codends of pelagic trawls for Atlantic mackerel (Scomber scombrus L.). Fisheries Research, 13(3):267–279, March 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900814.

Collins:1990:CTF

[Col90] Mark R. Collins. A comparison of three fish trap designs. Fisheries Research, 9(4):325–332, October 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090051V.

Coleman:1998:CSM

[Col98] Noel Coleman. Counting scallops and managing the fishery in Port Phillip Bay, south-east Australia. Fisheries Research, 38(2):145–157, October 1, 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698001544.

Cooney:1993:TEC

[Coo93] R. Ted Cooney. A theoretical evaluation of the carrying capacity of Prince William Sound, Alaska, for juvenile Pacific salmon. Fisheries Research, 18(1–2):77–87, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783693900415.

Corten:1996:WGB

[Cor96] Ad Corten. The widening gap between fisheries biology and fisheries management in the European union. Fisheries Research, 27(1–3):1–15, June 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695004580.

Cortes:1998:DAA

[Cor98] Enric Cortés. Demographic analysis as an aid in shark stock assessment and management. Fisheries Research, 39(2):199–208, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001830.

Comas:1994:PPE

[CP94] JoséA. Perez Comas and Ellen K. Pikitch. The predictive power of empirical relationships describing size selectivity, with application to gadoid fish. Fisheries Research, 20(2–3):151– 164, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900809.

Carrier:1998:HMC

[CP98a] Jeffrey C. Carrier and Harold L. Pratt. Habitat management and closure of a nurse shark breeding and nursery ground. Fisheries Research, 39(2):209–213, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001842.

Chen:1998:CMR

[CP98b] Y. Chen and J. E. Paloheimo. Can a more realistic model error structure improve the parameter estimation in modelling the dynamics of fish populations? Fisheries Research, 38(1):9–17, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001155.

Coelho:1994:PSM

[CQB^{+}94] M. L. Coelho, J. Quintela, V. Bettencourt, G. Olavo, and H. Villa. Population structure, maturation patterns and fecundity of the squid Loligo vulgaris from southern Portugal. Fisheries Research, 21(1–2):87–102, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900973.

Charton:1999:EHE

[CR99] Jose Antonio Garcıá Charton and Ángel Pérez Ruzafa. Ecological heterogeneity and the evaluation of the effects of marine reserves. Fisheries Research, 42(1–2):1–20, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000430.

Cripps:1993:BRI

[Cri93] Simon J. Cripps. Book review: Introduction to aquaculture: Matthew Landau, John Wiley & Sons, 1992, 440 pp., £15.95 (hardback £45.95), ISBN 0-471-57020-6. Fisheries Research, 16(1):87–89, February 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390112K.

Crittenden:1994:OEC

[Cri94] Robert N. Crittenden. Optimum escapement computed using Ricker’s spawner-recruit curve. Fisheries Research, 20 (2–3):215–227, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783694900841.

Cannizzaro:1995:ADG

[CRLG95] L. Cannizzaro, P. Rizzo, D. Levi, and S. Gancitano. Age determination and growth of Squalus blainvillei (Risso, 1826). Fisheries Research, 23(1–2):113–125, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400333R.

Cruywagen:1997:UGL

[Cru97] Gerhard C. Cruywagen. The use of generalized linear modelling to determine inter-annual and inter-area variation of growth rates: the Cape rock lobster as example. Fisheries Research, 29(2):119–131, February 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005358.

Cremonte:1997:PFS

[CS97] Florencia Cremonte and Norma Haydee Sardella. The parasita fauna of Scomber japonicus Houttuyn, 1782 (Pisces: Scombridae) in two zones of the Argentine Sea. Fisheries Research, 31(1–2):1–9, July 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000246.

Chigbu:1998:FEL

[CS98a] Paulinus Chigbu and Thomas H. Sibley. Feeding ecology of longfin smelt (Spirinchus thaleichthys Ayres) in Lake Washington. Fisheries Research, 38(2):109–119, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001568.

Coffen-Smout:1998:SSC

[CS98b] S. S. Coffen-Smout. Shell strength in the cockle Cerastoderma edule L. under simulated fishing impacts. Fisheries Research, 38(2):187–191, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698001593.

Castro-Santos:1996:PIT

[CSHW96] Theodore Castro-Santos, Alex Haro, and Stephen Walk. A passive integrated transponder (PIT) tag system for monitoring fishways. Fisheries Research, 28(3):253–261, October 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783696005140.

Cyr:1995:CJC

[CSM95] Charley Cyr and Bernard Sainte-Marie. Catch of Japanese crab traps in relation to bait quantity and shielding. Fisheries Research, 24(2):129–139, September 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694003665.

Coffen-Smout:1999:BBD

[CSR99] S. S. Coffen-Smout and E. I. S. Rees. Burrowing behaviour and dispersion of cockles Cerastoderma edule L. following simulated fishing disturbance. Fisheries Research, 40(1):65–72, February 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698002148.

Crittenden:1992:CFD

[CT92] Robert N. Crittenden and G. L. Thomas. A comparison of four duration-in-beam estimators for the sine of the half-angle of the effective conical volume sampled. Fisheries Research, 14(2–3):197–208, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290053V.

Chu:1993:MRR

[CTCN93] K. H. Chu, Y. K. Tam, C. K. Chung, and W. L. Ng. Morphometric relationships and reproductive maturation of the shrimp, Metapenaeus ensis, from commercial catches in Hong Kong. Fisheries Research, 18(3–4):187–197, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390152W.

Collie:1993:MLD

[CW93] Jeremy S. Collie and Carl J. Walters. Models that ‘learn’ to distinguish among alternative hypotheses. Fisheries Research, 18(3–4):259–275, December 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693901573.

Chun-Woo:1995:DCM

[CW95] Lee Chun-Woo. Depth control of a midwater trawl gear using fuzzy logic. Fisheries Research, 24(4):311–320, November 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695003882.

Cheke:1998:MEI

[CW98] Robert A. Cheke and A. R. Ward. A model for evaluating interventions designed to reduce post-harvest fish losses. Fisheries Research, 35(3):219–227, May 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000745.

Cui:1991:LLT

[CWG^{+}91] G. Cui, C. S. Wardle, C. W. Glass, A. D. F. Johnstone, and W. R. Mojsiewicz. Light level thresholds for visual reaction of mackerel, Scomber scombrus L., to coloured monofilament nylon gillnet materials. Fisheries Research, 10(3–4):255–263, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190079U.

Charles:1991:SPM

[CY91] Anthony T. Charles and Chi-Wen Yang. A strategic planning model for fisheries development. Fisheries Research, 10(3–4):287–307, January 1991. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190081P.

Dudley:1992:MIP

[DAB92] Richard G. Dudley, Arundhati Prabhakar Aghanashinikar, and Edward B. Brothers. Management of the Indo–Pacific Spanish mackerel (Scomberomorus commerson) in Oman. Fisheries Research, 15(1–2):17–43, October 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290003C.

Dahm:1994:RMS

[Dah94] Erdmann Dahm. Reduction of mesh size of netting after long-term storage under atmospheric conditions. Fisheries Research, 20(1):81–85, June 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900078.

Dahle:1996:DFP

[Dah96] Geir Dahle. DNA fingerprint pedigree studies of Atlantic cod (Gadus morhua L.). Fisheries Research, 26(1–2):93–99, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004025.

Daskalov:1999:RFR

[Das99] Georgi Daskalov. Relating fish recruitment to stock biomass and physical environment in the Black Sea using generalized additive models. Fisheries Research, 41(1):1–23, May 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000065.

Dawe:1998:PSG

[DBJ98] E. G. Dawe, W. R. Bowering, and J. B. Joy. Predominance of squid (Gonatus spp.) in the diet of Greenland halibut (Reinhardtius hippoglossoides) on the deep slope of the northeast Newfoundland continental shelf. Fisheries Research, 36(2–3):267–273, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698000927.

desClers:1993:MID

[dC93] Sophie des Clers. Modelling the impact of disease-induced mortality on the population size of wild salmonids. Fisheries Research, 17(1–2):237–248, June 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390023Z.

Die:1997:SYI

[DC97] David J. Die and John F. Caddy. Sustainable yield indicators from biomass: are there appropriate reference points for use in tropical fisheries? Fisheries Research, 32(1):69–79, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000295.

Defeo:1998:HEP

[DC98] Omar Defeo and Juan Carlos Castilla. Harvesting and economic patterns in the artisanal Octopus mimus (Cephalopoda) fishery in a northern Chile cove. Fisheries Research, 38 (2):121–130, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783698001556.

deCastro:1995:EFG

[dCB95] Fábio de Castro and Alpina Begossi. Ecology of fishing on the Grande River (Brazil): technology and territorial rights. Fisheries Research, 23(3–4):361–373, June 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400343U.

daCunha:1994:RTP

[dCM94] Manuela Morais da Cunha and Ana Moreno. Recent trends in the Portuguese squid fishery. Fisheries Research, 21(1–2):231–241, December 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694901066.

DiGiacomo:1993:SAM

[DCPM93] Edgardo E. Di Giácomo, Jorge Calvo, María R. Perier, and Elba Morriconi. Spawning aggregations of Merluccius hubbsi, in Patagonian waters: evidence for a single stock? Fisheries Research, 16(1):9–16, February 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390106H.

Dedah:1999:CFS

[DCS99] S. Ould Dedah, R. E. Condrey, and R. F. Shaw. Conservative fit of the Schaefer model for a multispecies fishery. Fisheries Research, 41(2):155–168, June 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369900003X.

Dahle:1990:SAS

[DE90] G. Dahle and A. G. Eriksen. Spring and autumn spawners of herring (Clupea harengus) in the North Sea, Skagerrak and Kattegat; population genetic analysis. Fisheries Research, 9(2):131–141, August 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783690900609.

Borges:1990:MCA

[dFB90] Maria de Fatima Borges. Multiplicative catch-at-age analysis of Scad (Trachurus trachurus L.) from Western Iberian waters. Fisheries Research, 9(4):333–353, October 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090052W.

Dickson:1993:ECEa

[Dic93a] William Dickson. Estimation of the capture efficiency of trawl gear. i: Development of a theoretical model. Fisheries Research, 16(3):239–253, April 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390096P.

Dickson:1993:ECEb

[Dic93b] William Dickson. Estimation of the capture efficiency of trawl gear. II: Testing a theoretical model. Fisheries Research, 16(3):255–272, April 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390097Q.

Dahle:1993:HVC

[DJ93] G. Dahle and K. E. Jørstad. Haemoglobin variation in cod — a reliable marker for Arctic cod (Gadus morhua L.). Fisheries Research, 16(4):301–311, May 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390143U.

Dulcic:1995:AGM

[DK95] Jakov Dulčić and Miro Kraljević. Age, growth and mortality of damselfish (Chromis chromis L.) in the eastern middle Adriatic. Fisheries Research, 22(3–4):255–264, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400318Q.

Dulcic:1996:WLR

[DK96] J. Dulčić and M. Kraljević. Weight–length relationships for 40 fish species in the eastern Adriatic (Croatian waters). Fisheries Research, 28(3):243–251, October 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696005139. See corrigendum [DK97].

Dulcic:1997:CWL

[DK97] J. Dulčić and M. Kraljević. Corrigendum to “Weight– length relationships for 40 fish species in the eastern Adriatic (Croatian waters)” [Fisheries Research 28 (1996) 243– 251]. Fisheries Research, 30(1–2):169–171, May 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697005559. See [DK96].

Dolmer:1999:DBM

[DKH99] Per Dolmer, Per Sand Kristensen, and Erik Hoffmann. Dredging of blue mussels (Mytilus edulis L.) in a Danish sound: stock sizes and fishery-effects on mussel population dynamic. Fisheries Research, 40(1):73–80, February 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002185.

Defeo:1991:SAD

[DLB91] Omar Defeo, Virginia Little, and Luis Barea. Stock assessment of the deep-sea red crab Chaceon notialis in the Argentinian–Uruguayan Common Fishing Zone. Fisheries Research, 11(1):25–39, February 1991. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190031A.

DeSilva:1992:PYP

[DLT92] Sena S. De Silva, Yiongtai Lin, and Gangshen Tang. Possible yield-predictive models based on morphometric characteristics and stocking rates for three groups of Chinese reservoirs. Fisheries Research, 13(4):369–380, April 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783692900582.

DeSilva:1991:CAF

[DMA^{+}91] Sena S. De Silva, Jacques Moreau, U. S. Amarasinghe, Tawan Chookajorn, and R. D. Guerrero. A comparative assessment of the fisheries in lacustrine inland waters in three Asian countries based on catch and effort data. Fisheries Research, 11(2):177– 189, April 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190107Q.

Daves:1998:UIM

[DN98] Nancy K. Daves and Marta F. Nammack. US and international mechanisms for protecting and managing shark resources. Fisheries Research, 39(2):223–228, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001878.

Dolmer:1998:SSV

[Dol98] P. Dolmer. Seasonal and spatial variability in growth of Mytilus edulis L. in a brackish sound: comparisons of individual mussel growth and growth of size classes. Fisheries Research, 34(1):17–26, February 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000817.

Djama:1997:DEC

[DP97] Theodore Djama and T. J. Pitcher. The differential effects of changing management regimes on yields from two fisheries exploiting the same resources. Fisheries Research, 29(1):33– 37, January 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005206.

Deriso:1991:MMM

[DPB91] Richard B. Deriso, Richard G. Punsly, and William H. Bayliff. A Markov movement model of yellowfin tuna in the Eastern Pacific Ocean and some analyses for international management. Fisheries Research, 11(3–4):375–394, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190010D.

Dahle:1997:RFU

[DRE97] G. Dahle, M. Rahman, and A. G. Eriksen. RAPD fingerprinting used for discriminating among three populations of Hilsa shad (Tenualosa ilisha). Fisheries Research, 32(3):263–269, December 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000581.

deSilva:1991:GRR

[dS91] K. H. G. M. de Silva. Growth rate and the role of Oreochromis mossambicus (Peters) in the fishery of a tropical, upland, deep reservoir in Sri Lanka. Fisheries Research, 12(2):125–138, October 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190037G.

deSilva:1994:FGR

[dS94] K. H. G. M. de Silva. The fishery, growth rates and notes on reproduction of the snakehead Channa striata (Bloch) (Perciformes: Channidae) in some irrigation reservoirs of Sri Lanka. Fisheries Research, 19(3–4):257–268, April 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900434.

Defeo:1999:YMM

[DS99] Omar Defeo and Juan Carlos Seijo. Yield–mortality models: a precautionary bioeconomic approach. Fisheries Research, 40(1):7–16, February 1, 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698002203.

duSel:1997:NSR

[dSD97] Guy Pinczon du Sel and Jacques Daguzan. A note on sex ratio, length and diet of a population of cuttlefish Sepia officinalis L. (Mollusca: Cephalopoda) sampled by three fishing methods. Fisheries Research, 32(2):191–195, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000428.

DuBuit:1991:FFS

[Du 91] Marie Henriette Du Buit. Food and feeding of saithe (Pollachius virens L.) off Scotland. Fisheries Research, 12(4): 307–323, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783691900158.

DuBuit:1995:FFC

[Du 95] Marie Henriette Du Buit. Food and feeding of cod (Gadus morhua L.) in the Celtic Sea. Fisheries Research, 22(3–4):227– 241, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400322N.

DuBuit:1996:DHM

[Du 96] Marie Henriette Du Buit. Diet of hake (Merluccius merluccius) in the Celtic Sea. Fisheries Research, 28(4):381–394, December 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005164.

Dulcic:1995:EAG

[Dul95] Jakov Dulčić. Estimation of age and growth of sardine, Sardina pilchardus (Walbaum, 1792), larvae by reading daily otolith increments. Fisheries Research, 22(3–4):265–277, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400316O.

Dulcic:1997:GAE

[Dul97] Jakov Dulčić. Growth of anchovy, Engraulis encrasicolus (L.), larvae in the Northern Adriatic Sea. Fisheries Research, 31(3):189–195, August 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000301.

Dulcic:1998:LGS

[Dul98] Jakov Dulčić. Larval growth of sprat, Sprattus sprattus phalericus, larvae in the Northern Adriatic. Fisheries Research, 36(2–3):117–126, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698001088.

Dunn:1999:ASD

[Dun99] M. R. Dunn. Aspects of the stock dynamics and exploitation of cuttlefish, Sepia officinalis (Linnaeus, 1758), in the English Channel. Fisheries Research, 40(3):277–293, April 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698002239.

Donaghy:1995:DNA

[DVY95] M. J. Donaghy, E. Verspoor, and A. F. Youngson. Discrimination of the northeast Atlantic lesser sandeels, Ammodytes marinus, Ammodytes tobianus and Gymnammodytes semisquamatus by protein electrophoresis. Fisheries Research, 23(1–2): 1–9, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694003403.

Early:1991:BRU

[Ear91] J. Early. Book review: Utilisation of squid: M. Sugiyama, S. Kousa, M. Hanabe and Y. Okuda. A. A. Balkema Publishers, Rotterdam, 1989, 266 pp., price Dfl. 95.00/ U.S.$48.50/ £27.00, ISBN 90-6191-479-5. Fisheries Research, 10(3–4):353– 354, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190086U.

Early:1993:BRF

[Ear93] J. C. Early. Book review: Fish processing technology: George M. Hall (Editor), Blackie Academic and Professional, Glasgow, UK, 1992, 309 pp., £65, ISBN 0-216-93324-2. Fisheries Research, 16(4):363–365, May 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390147Y.

Everson:1996:CAT

[EBG96] Inigo Everson, Mark Bravington, and Catherine Goss. A combined acoustic and trawl survey for efficiently estimating fish abundance. Fisheries Research, 26(1–2):75–91, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004041.

Erzini:1998:AMF

[EC98] K. Erzini and M. Castro. An alternative methodology for fitting selectivity curves to pre-defined distributions. Fisheries Research, 34(3):307–313, April 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697001033.

Eckmann:1998:AEI

[Eck98] Reiner Eckmann. Allocation of echo integrator output to small larval insect (Chaoborus sp.) and medium-sized (juvenile fish) targets. Fisheries Research, 35(1–2):107–113, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698000654.

Eriksson:1993:SWH

[EE93] T. Eriksson and L.-O. Eriksson. The status of wild and hatchery propagated Swedish salmon stocks after 40 years of hatchery releases in the Baltic rivers. Fisheries Research, 18(1–2):147–159, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693900459.

Evans:1995:FES

[EE95] Christopher R. Evans and Arthur J. Evans. Fisheries ecology of spiny lobsters Panulirus argus (Latreille) and Panulirus guttatus (Latreille) on the Bermuda Platform: estimates of sustainable yields and observations on trends in abundance. Fisheries Research, 24(2):113–128, September 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694003676. See erratum [EE96a].

Evans:1996:EFE

[EE96a] C. R. Evans and A. J. Evans. Erratum to “Fisheries ecology of spiny lobsters Panulirus argus (Latreille) and Panulirus guttatus (Latreille) on the Bermuda Platform: estimates of sustainable yields and observations on trends in abundance”: [Fish. Res. 24 (1995)113–128]. Fisheries Research, 28(1):115, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783696894458. See [EE95].

Evans:1996:PFT

[EE96b] Christopher R. Evans and Arthur J. Evans. A practical field technique for the assessment of spiny lobster resources of tropical islands. Fisheries Research, 26(1–2):149–169, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695003916.

Eggers:1992:BCM

[Egg92] Douglas M. Eggers. The benefits and costs of the management program for natural sockeye salmon stocks in Bristol Bay, Alaska. Fisheries Research, 14(2–3):159–177, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290051T.

Evans:1997:TMP

[EGR^{+}97] S. M. Evans, M. E. Gill, A. S. W. Retraubun, J. Abrahamz, and J. Dangeubun. Traditional management practices and the conservation of the gastropod (Trochus nilitocus) and fish stocks in the Maluku Province (eastern Indonesia). Fisheries Research, 31(1–2):83–91, July 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000118.

Ezenwaji:1998:OBC

[EI98] H. M. G. Ezenwaji and N. M. Inyang. Observations on the biology of Clarias agboyiensis Syndenham, 1980 (Osteichthyes: Clariidae) in the Anambra floodriver system, Nigeria. Fisheries Research, 36(1):47–60, June 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000782.

Enzmann:1993:VWL

[EKP93] P.-J. Enzmann, M. Konrad, and K. Parey. VHS in wild living fish and experimental transmission of the virus. Fisheries Research, 17(1–2):153–161, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390015Y.

Ehrhardt:1997:RUF

[EL97] Nelson M. Ehrhardt and Christopher M. Legault. The role of uncertainty in fish stock assessment and management: a case study of the Spanish mackerel, Scomberomorus maculatus, in the US Gulf of Mexico. Fisheries Research, 29(2):145–158, February 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005280.

Edmonds:1992:TEO

[ELCM92] J. S. Edmonds, R. C. J. Lenanton, N. Caputi, and M. Morita. Trace elements in the otoliths of yellow-eye mullet (Aldrichetta forsteri) as an aid to stock identification. Fisheries Research, 13(1):39–51, January 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290032O.

Enin:1996:PDE

[ELK96] Udeme I. Enin, U. Lowenberg, and T. Kunzel. Population dynamics of the estuarine prawn (Nematopalaemon hastatus Aurivillius 1898) off the southeast coast of Nigeria. Fisheries Research, 26(1–2):17–35, April 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004114.

Etim:1999:DEP

[ELK99] Lawrence Etim, Peter E. Lebo, and Richard P. King. The dynamics of an exploited population of a siluroid catfish (Schilbe intermidius Reupell 1832) in the Cross River, Nigeria. Fisheries Research, 40(3):295–307, April 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002173.

Elliott:1993:PNM

[Ell93] J. M. Elliott. The pattern of natural mortality throughout the life cycle in contrasting populations of brown trout, Salmo trutta L. Fisheries Research, 17(1–2):123–136, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390012V.

Einarsson:1990:UFL

[EMJ90] S. M. Einarsson, D. H. Mills, and V. Johannsson. Utilisation of fluvial and lacustrine habitat by anadromous Atlantic salmon, Salmo salar L., in an Icelandic watershed. Fisheries Research, 10(1–2):53–71, December 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090015N.

Emlen:1991:HOD

[Eml91] John M. Emlen. Heterosis and outbreeding depression: a multi-locus model and an application to salmon production. Fisheries Research, 12(3):187–212, November 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190095W.

Engaas:1995:RPH

[EMS^{+}95] Arill Engås, Ole Arve Misund, Aud Vold Soldal, Berit Horvei, and Arne Solstad. Reactions of penned herring and cod to playback of original, frequency-filtered and time-smoothed vessel sound. Fisheries Research, 22(3–4):243–254, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400317P.

Erickson:1996:ECS

[EPCPW96] Daniel L. Erickson, Jose A. Perez-Comas, Ellen K. Pikitch, and John R. Wallace. Effects of catch size and codend type on the escapement of walleye pollock (Theragra chalcogramma) from pelagic trawls. Fisheries Research, 28(2):179–196, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783696004973.

Erzini:1990:SSG

[Erz90] Karim Erzini. Sample size and grouping of data for lengthfrequency analysis. Fisheries Research, 9(4):355–366, October 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090053X.

Etim:1998:GMR

[ES98] Lawrence Etim and Yacouba Sankare. Growth and mortality, recruitment and yield of the fresh-water shrimp, Macrobrachium völlenhovenii, Herklots 1851 (Crustacea, Palaemonidae) in the Fahe reservoir, Côte d’Ivoire, West Africa. Fisheries Research, 38(3):211–223, November 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001611.

Ehrich:1999:FEN

[ES99] S. Ehrich and C. Stransky. Fishing effects in northeast Atlantic shelf seas: patterns in fishing effort, diversity and community structure. VI. Gale effects on vertical distribution and structure of a fish assemblage in the North Sea. Fisheries Research, 40(2):185–193, March 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002112.

Empis:1995:DIF

[ESNA95] Jose M. A. Empis, Helena A. Silva, Maria L. Nunes, and Eduarda M. Andrade. Detection of irradiated fish using EPR of fish bone; signal intensity and stability. Fisheries Research, 21(3–4):471–475, January 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400300L.

Engaas:1997:MSR

[ESW97] Arill Engås, Roar Skeide, and Charles W. West. The ‘Multi-Sampler’: a system for remotely opening and closing multiple codends on a sampling trawl. Fisheries Research, 29(3):295– 298, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005450.

Eklund:1991:SFC

[ET91] Anne-Marie Eklund and Timothy E. Targett. Seasonality of fish catch rates and species composition from the hard bottom trap fishery in the Middle Atlantic Bight (US east coast). Fisheries Research, 12(1):1–22, September 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190045H.

Evans:1996:SFA

[EW96] S. M. Evans and R. I. Wahju. The shrimp fishery of the Arafura Sea (eastern Indonesia). Fisheries Research, 26(3–4):365– 371, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004483.

Fabrizio:1997:APF

[FAC97] Mary C. Fabrizio, Jean V. Adams, and Gary L. Curtis. Assessing prey fish populations in Lake Michigan: comparison of simultaneous acoustic-midwater trawling with bottom trawling. Fisheries Research, 33(1–3):37–54, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000611.

Feist:1993:PKD

[FB93] S. W. Feist and D. Bucke. Proliferative kidney disease in wild salmonids. Fisheries Research, 17(1–2):51–58, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390006S.

Friedlander:1997:FLH

[FB97] Alan M. Friedlander and James P. Beets. Fisheries and life history characteristics of dwarf herring (Jenkinsia lamprotaenia) in the US Virgin Islands. Fisheries Research, 31(1–2): 61–72, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000155.

Fievet:1999:ESB

[FBAM99] É Fièvet, P. Bonnet-Arnaud, and J.-P. Mallet. Efficiency and sampling bias of electrofishing for freshwater shrimp and fish in two Caribbean streams, Guadeloupe Island. Fisheries Research, 44(2):149–166, December 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000582.

Farmer:1998:DSF

[FBB98] M. J. Farmer, D. T. Brewer, and S. J. M. Blaber. Damage to selected fish species escaping from prawn trawl codends: a comparison between square-mesh and diamondmesh. Fisheries Research, 38(1):73–81, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001179.

Fletcher:1990:USI

[FBF90] W. J. Fletcher, I. W. Brown, and D. R. Fielder. The use of standard and inverse Leslie experiments to estimate the abundance of the coconut crab (Birgus latro L.) in Vanuatu. Fisheries Research, 9(4):317–324, October 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783690900506.

Foale:1997:SAT

[FD97] Simon Foale and Rob Day. Stock assessment of trochus (Trochus niloticus) (Gastropoda: Trochidae) fisheries at West Nggela, Solomon Islands. Fisheries Research, 33(1–3):1–16, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000623.

Fousekis:1999:ISE

[FDD99] Panos Fousekis, Anna Daouli, and Michael Demoussis. The internal structure of effort in Greek fisheries: an empirical analysis. Fisheries Research, 41(2):213–219, June 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000181.

Fouilland:1996:TEP

[FF96] Eric Fouilland and Odile Fossati. Trapping efficiency of plastic bottle “wickertraps” for population assessment of river Macrobrachium (Crustacea: Decapoda). Fisheries Research, 28(3):343–351, October 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696005073.

Fouilland:1998:ESO

[FF98] Eric Fouilland and Odile Fossati. Effects of some operational factors on Macrobrachium (Decapoda, Palaemonidae) sampling using small ‘wickertraps’. Fisheries Research, 34(1):87– 92, February 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000866.

Farina:1994:NNG

[FFGG94] A. Celso Fariña, Juan Freire, and Eduardo González-Gurriarán. Nephrops norvegicus in the Galician continental shelf (NW Spain): abundance and distribution. Fisheries Research, 19(3–4):333–347, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900485.

Freon:1992:CSS

[FGS92] Pierre Fréon, François Gerlotto, and Marc Soria. Changes in school structure according to external stimuli: description and influence on acoustic assessment. Fisheries Research, 15(1–2):45–66, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290004D.

Franco:1991:SBP

[FJM91] J. Franco, A. D. F. Johnstone, and A. M. Mackie. Studies of bait preference in the cod, Gadus morhua L.: characterisation of feeding stimulants using an operant conditioning technique. Fisheries Research, 10(3–4):229–242, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190077S.

Freeman:1995:STS

[FK95] S. N. Freeman and G. P. Kirkwood. On a structural time series method for estimating stock biomass and recruitment from catch and effort data. Fisheries Research, 22(1–2):77–98, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400304F.

Fryer:1993:HCS

[FL93] J. L. Fryer and C. N. Lannan. The history and current status of Renibacterium salmoninarum, the causative agent of bacterial kidney disease in Pacific salmon. Fisheries Research, 17(1–2): 15–33, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390004Q.

Furevik:1994:FTN

[FL94] Dag M. Furevik and Svein Løkkeborg. Fishing trials in Norway for torsk (Brosme brosme) and cod (Gadus morhua) using baited commercial pots. Fisheries Research, 19(3–4):219– 229, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369490040X.

Flaaten:1998:BPP

[Fla98] Ola Flaaten. On the bioeconomics of predator and prey fishing. Fisheries Research, 37(1–3):179–191, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001350.

Fletcher:1992:USM

[Fle92] W. J. Fletcher. Use of a spatial model to provide initial estimates of stock size for a purse seine fishery on pilchards (Sardinops sagax neopilchardus) in Western Australia. Fisheries Research, 14(1):41–57, June 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783692900722.

Fonteyne:1992:SES

[FM92] R. Fonteyne and R. M’Rabet. Selectivity experiments on sole with diamond and square mesh codends in the Belgian coastal beam trawl fishery. Fisheries Research, 13(3):221– 233, March 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900788.

Ni:1994:ADF

[fN94] Ru feng Ni. Analysis of dominant factors in enhancing the fishery resources in Lake Tai, China. Fisheries Research, 19(1–2):121–130, February 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900183.

Friedlander:1997:FHS

[FP97] Alan M. Friedlander and James D. Parrish. Fisheries harvest and standing stock in a Hawaiian Bay. Fisheries Research, 32(1):33–50, October 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000386.

Formigo:1999:FSA

[FP99] N. Formigo and T. Penczak. Fish stocks of the Âncora River, northern Portugal: production. Fisheries Research, 40(1):91–96, February 1, 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698002161.

Francis:1995:AMR

[Fra95] R. I. C. C. Francis. An alternative mark-recapture analogue of Schnute’s growth model. Fisheries Research, 23(1–2):95– 111, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400336U.

Fraser:1996:BRE

[Fra96] P. J. Fraser. Book review: Electric fishes — History and behaviour: Peter Moller. Chapman and Hall. Fish and Fisheries Series 17. London, 1995, 584 pp., £69, ISBN 0-412-37380-7. Fisheries Research, 28(2):239–240, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696004936.

Friedlander:1995:RFB

[Fri95] Alan Friedlander. The recreational fishery for blue marlin, Makaira nigricans (Pisces: Istiophoridae), in the US Virgin Islands. Fisheries Research, 22(3–4):163–173, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400329U.

Formacion:1994:MCP

[FS94] Sonia P. Formacion and Saul B. Saila. Markov chain properties related to temporal dominance changes in a Philippine pelagic fishery. Fisheries Research, 19(3–4):241–256, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900426.

Flaaten:1998:FMU

[FSSU98] Ola Flaaten, Anne Gro Vea Salvanes, Tore Schweder, and Øyvind Ulltang. Fisheries management under uncertainty — an overview. Fisheries Research, 37(1–3):1–6, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001222.

Fujimori:1996:SGE

[FTHM96] Yasuzumi Fujimori, Tadashi Tokai, Setsuhisa Hiyama, and Ko Matuda. Selectivity and gear efficiency of trammel nets for kuruma prawn (Penaeus japonicus). Fisheries Research, 26(1–2):113–124, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695003924.

Furness:1994:EQS

[Fur94] Robert W. Furness. An estimate of the quantity of squid consumed by seabirds in the eastern North Atlantic and adjoining seas. Fisheries Research, 21(1–2):165–177, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694901023.

Fennessy:1994:DSI

[FVF94] S. T. Fennessy, C. Villacastin, and J. G. Field. Distribution and seasonality of ichthyofauna associated with commercial prawn trawl catches on the Tugela Bank of Natal, South Africa. Fisheries Research, 20(2–3):263–282, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900876.

Ferro:1996:EVS

[FvMH96] R. S. T. Ferro, B. van Marlen, and K. E. Hansen. An empirical velocity scale relation for modelling a design of large mesh pelagic trawl. Fisheries Research, 28(2):197–230, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369600495X.

Francis:1995:DVT

[FW95] M. P. Francis and M. W. Williams. Diel variation in trawl catch rates of Pagrus auratus (Sparidae). Fisheries Research, 24(4):301–310, November 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695003845.

Ferro:1996:ITM

[FX96] R. S. T. Ferro and L. Xu. An investigation of three methods of mesh size measurement. Fisheries Research, 25(2):171–190, February 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004033.

Goni:1999:AGL

[GAA99] R. Goñi, F. Alvarez, and S. Adlerstein. Application of generalized linear modeling to catch rate analysis of Western Mediterranean fisheries: the Castellón trawl fleet as a case study. Fisheries Research, 42(3):291–302, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000399.

Garcia:1997:PAF

[Gar97] S. Garcia. The precautionary approach to fisheries: a reply to Prof. Kesteven [Fisheries Research 28 (1996) 109– 110]. Fisheries Research, 31(3):275–277, August 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000490. See [Kes96e].

Gatlin:1990:BRN

[Gat90] Delbert Gatlin. Book review: Nutrition of pond fishes: Balfour Hepher, Cambridge University Press, Cambridge, 1988, 388 pp. $69.00, ISBN 0-521-34150-7. Fisheries Research, 9(4):373– 374, October 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090055Z.

Gaudreau:1998:ISH

[GB98a] N. Gaudreau and D. Boisclair. The influence of spatial heterogeneity on the study of fish horizontal daily migration. Fisheries Research, 35(1–2):65–73, May 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000605.

Gjosaeter:1998:EPH

[GB98b] H. Gjøsæter and B. Bogstad. Effects of the presence of herring (Clupea harengus) on the stock-recruitment relationship of Barents Sea capelin (Mallotus villosus). Fisheries Research, 38(1):57–71, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698001143.

Goncalves:1997:WLR

[GBL^{+}97] J. M. S. Gonçalves, L. Bentes, P. G. Lino, J. Ribeiro, A. V. M. Canário, and K. Erzini. Weight–length relationships for selected fish species of the small-scale demersal fisheries of the south and south-west coast of Portugal. Fisheries Research, 30(3):253–256, June 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005693.

Gerlotto:1992:SEV

[GF92] François Gerlotto and Pierre Fréon. Some elements on vertical avoidance of fish schools to a vessel during acoustic surveys. Fisheries Research, 14(4):251–259, September 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290035R.

Galbraith:1994:DPT

[GFM94] R. D. Galbraith, R. J. Fryer, and K. M. S. Maitland. Demersal pair trawl cod-end selectivity models. Fisheries Research, 20(1):13–27, June 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900035.

Gaspar:1999:GRC

[GFM99] M. B. Gaspar, R. Ferreira, and C. C. Monteiro. Growth and reproductive cycle of Donax trunculus L., (Mollusca: Bivalvia) off Faro, southern Portugal. Fisheries Research, 41(3):309– 316, July 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369900017X.

Gonzalez:1998:OFM

[GG98] Luis Gonzalez and François Gerlotto. Observation of fish migration between the sea and a Mediterranean lagoon (Étang de l’Or, France) using multibeam sonar and split beam echo sounder. Fisheries Research, 35(1–2):15–22, May 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000551.

Guillard:1992:UGA

[GGBC92] Jean Guillard, Daniel Gerdeaux, Guy Brun, and Rémi Chappaz. The use of geostatistics to analyse data from an echo-integration survey of fish stock in Lake Sainte-Croix. Fisheries Research, 13(4):395–406, April 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783692900607.

Gabche:1995:GMR

[GH95] C. E. Gabche and H.-U. P. Hockey. Growth, mortality and reproduction of Sardinella maderensis (Lowe, 1841) in the artisanal fisheries off Kribi, Cameroon. Fisheries Research, 24(4):331–344, November 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695003717.

Gabr:1998:MFS

[GHHEE98] Howaida R. Gabr, Roger T. Hanlon, Mahmoud H. Hanafy, and Salah G. El-Etreby. Maturation, fecundity and seasonality of reproduction of two commercially valuable cuttlefish, Sepia pharaonis and S. dollfusi, in the Suez Canal. Fisheries Research, 36(2–3):99–115, June 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001076.

Gerlotto:1998:EMS

[GHL98] François Gerlotto, Carlos Hernandez, and Esteban Linares. Experiences with multibeam sonar in shallow tropical waters. Fisheries Research, 35(1–2):143–147, May 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000691.

Gartside:1999:EUF

[GHR99] Donald F. Gartside, Bradley Harrison, and Bret L. Ryan. An evaluation of the use of fishing club records in the management of marine recreational fisheries. Fisheries Research, 41(1):47– 61, May 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000077.

Geiger:1991:EGS

[GK91] Harold J. Geiger and Jeffery P. Koenings. Escapement goals for sockeye salmon with informative prior probabilities based on habitat considerations. Fisheries Research, 11(3–4):239– 256, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190004Y.

Gjosaeter:1990:BCG

[GL90] Jakob Gjøsæter and Kate Lønnhaug. Back-calculation of growth from broken cod otoliths. Fisheries Research, 9(3): 283–286, September 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605800074.

Granado-Lorencio:1991:EMF

[GL91] Carlos Granado-Lorencio. The effect of man on the fish fauna of the River Guadalquivir, Spain. Fisheries Research, 12(2):91– 100, October 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190034D.

Gonzalez:1994:DMM

[GLGB94] A. F. González, A. López, A. Guerra, and A. Barreiro. Diets of marine mammals stranded on the northwestern Spanish Atlantic coast with special reference to cephalopoda. Fisheries Research, 21(1–2):179–191, December 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694901031.

Gordoa:1991:DVF

[GM91] A. Gordoa and E. Macpherson. Diurnal variation in the feeding activity and catch rate of Cape hake (Merluccius capensis and M. paradoxus) off Namibia. Fisheries Research, 12(4):299–305, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783691900147.

Gebauer:1995:EVG

[GM95] Paulina Gebauer and Carlos A. Moreno. Experimental validation of the growth rings of Loxechinus albus (Molina, 1782) in southern Chile (Echinodermata: Echinoidea). Fisheries Research, 21(3–4):423–435, January 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694002925.

Gordoa:1997:AGS

[GM97] Ana Gordoa and Balbina Molí. Age and growth of the sparids Diplodus vulgaris, D. sargus and D. annularis in adult populations and the differences in their juvenile growth patterns in the north-western Mediterranean Sea. Fisheries Research, 33(1–3):123–129, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578369700074X.

Gomes:1998:RDO

[GMHSR98] C. Gomes, R. Mahon, W. Hunte, and S. Singh-Renton. The role of drifting objects in pelagic fisheries in the Southeastern Caribbean. Fisheries Research, 34(1):47–58, February 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000799.

Gordina:1997:NDM

[GNN^{+}97] A. D. Gordina, V. N. Nikolskiy, U. Niermann, F. Bingel, and A. A. Subbotin. New data on the morphological differences of anchovy eggs (Engraulis encrasicolus L.) in the Black Sea. Fisheries Research, 31(1–2):139–145, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005632.

Gobert:1992:IUT

[Gob92] Bertrand Gobert. Impact of the use of trammelnets on a tropical reef resource. Fisheries Research, 13(4):353–367, April 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290057Z.

Gobert:1994:SSD

[Gob94] Bertrand Gobert. Size structures of demersal catches in a multispecies multigear tropical fishery. Fisheries Research, 19(1–2):87–104, February 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900167.

Gobert:1998:DDS

[Gob98] Bertrand Gobert. Density-dependent size selectivity in Antillean fish traps. Fisheries Research, 38(2):159–167, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001192.

Gold:1995:BRG

[Gol95] John R. Gold. Book review: Genetics and evolution of aquatic organisms: A. R. Beaumont (Editor), 1994. Chapman & Hall, London, 65, ISBN 0-412-49370-5. Fisheries Research, 24(1): 82–86, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695900128.

Govender:1994:GKM

[Gov94] Anesh Govender. Growth of the king mackerel (Scomberomorus commerson) off the coast of Natal, South Africa — from length and age data. Fisheries Research, 20(1):63–79, June 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369490006X.

Govender:1995:MBR

[Gov95] Anesh Govender. Mortality and biological reference points for the king mackerel (Scomberomorus commerson) fishery off Natal, South Africa (based on a per-recruit assessment). Fisheries Research, 23(3–4):195–208, June 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400362Z.

Guenette:1999:ASM

[GP99] Sylvie Guénette and Tony J. Pitcher. An age-structured model showing the benefits of marine reserves in controlling overexploitation. Fisheries Research, 39(3):295–303, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001738.

Godo:1990:ETD

[GPV90] Olav Rune Godø, Michael Pennington, and Jon Helge Vølstad. Effect of tow duration on length composition of trawl catches. Fisheries Research, 9(2):165–179, August 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090062Z.

Gold:1991:GSM

[GR91] John R. Gold and Linda R. Richardson. Genetic studies in marine fishes. IV. An analysis of population structure in the red drum (Sciaenops ocellatus) using mitochondrial DNA. Fisheries Research, 12(3):213–241, November 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190096X.

Gold:1994:MDV

[GR94a] John R. Gold and Linda R. Richardson. Mitochondrial DNA variation among ‘red’ fishes from the Gulf of Mexico. Fisheries Research, 20(2–3):137–150, August 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900795.

Guerra:1994:LHL

[GR94b] Angel Guerra and Francisco Rocha. The life history of Loligo vulgaris and Loligo forbesi (Cephalopoda: Loliginidae) in Galician waters (NW Spain). Fisheries Research, 21(1–2):43–69, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900957.

Graynoth:1999:IOP

[Gra99] Eric Graynoth. Improved otolith preparation, ageing and back-calculation techniques for New Zealand freshwater eels. Fisheries Research, 42(1–2):137–146, August 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000296.

Gonzalez:1994:PSI

[GRG94] Angel F. González, Mario Rasero, and Angel Guerra. Preliminary study of Illex coindetii and Todaropsis eblanae (Cephalopoda: Ommastrephidae) in northern Spanish Atlantic waters. Fisheries Research, 21(1–2):115–126, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369490099X.

Griffiths:1997:APR

[Gri97] Marc H. Griffiths. The application of per-recruit models to Argyrosomus inodorus, an important South African sciaenid fish. Fisheries Research, 30(1–2):103–115, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005528.

Griffiths:1999:ELP

[Gri99] Shane P. Griffiths. Effects of lunar periodicity on catches of Penaeus plebejus (Hess) in an Australian coastal lagoon. Fisheries Research, 42(1–2):195–199, August 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000284.

Godo:1992:SSD

[GS92] Olav Rune Godø and Knut Sunnanå. Size selection during trawl sampling of cod and haddock and its effect on abundance indices at age. Fisheries Research, 13(3):293–310, March 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900836.

Gharrett:1993:PAI

[GS93] A. J. Gharrett and W. W. Smoker. A perspective on the adaptive importance of genetic infrastructure in salmon populations to ocean ranching in Alaska. Fisheries Research, 18(1–2):45– 58, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390039A.

Greenstreet:1999:FENb

[GSM99] Simon P. R. Greenstreet, Fiona E. Spence, and Julie A. McMillan. Fishing effects in northeast Atlantic shelf seas: patterns in fishing effort, diversity and community structure. V. Changes in structure of the North Sea groundfish species assemblage between 1925 and 1996. Fisheries Research, 40(2):153–183, March 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698002100.

Guerra:1994:SFL

[GSR94] A. Guerra, P. Sánchez, and F. Rocha. The Spanish fishery for Loligo: recent trends. Fisheries Research, 21(1–2):217–230, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694901058.

Greenstreet:1999:FENa

[GSSM99] Simon P. R. Greenstreet, Fiona B. Spence, Aileen M. Shanks, and Julie A. McMillan. Fishing effects in northeast Atlantic shelf seas: patterns in fishing effort, diversity and community structure. II. Trends in fishing effort in the North Sea by UK registered vessels landing in Scotland.

Fisheries Research, 40(2):107–124, March 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002070.

Guillard:1998:DMC

[Gui98] Jean Guillard. Daily migration cycles of fish populations in a tropical estuary (Sine-Saloum, Senegal) using a horizontal-directed split-beam transducer and multibeam sonar. Fisheries Research, 35(1–2):23–31, May 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000563.

Gullestad:1998:SRP

[Gul98] Peter Gullestad. The scope for research in practical fishery management. Fisheries Research, 37(1–3):251–258, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001416.

Godo:1992:EFD

[GW92] O. R. Godø and S. J. Walsh. Escapement of fish during bottom trawl sampling — implications for resource assessment. Fisheries Research, 13(3):281–292, March 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783692900825.

Glass:1995:SUVb

[GW95] C. W. Glass and C. S. Wardle. Studies on the use of visual stimuli to control fish escape from codends. II. The effect of a black tunnel on the reaction behaviour of fish in otter trawl codends. Fisheries Research, 23(1–2):165–174, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400331P.

Guan:1999:GRI

[GW99] Rui-Zhang Guan and Peter Roy Wiles. Growth and reproduction of the introduced crayfish Pacifastacus leniusculus in a British lowland river. Fisheries Research, 42(3):245–259, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000442.

Glass:1995:SUVa

[GWGR95] C. W. Glass, C. S. Wardle, S. J. Gosden, and D. N. Racey. Studies on the use of visual stimuli to control fish escape from codends. I. Laboratory studies on the effect of a black tunnel on mesh penetration. Fisheries Research, 23(1–2):157– 164, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400330Y.

Hansen:1990:EAS

[Han90] Lars P. Hansen. Exploitation of Atlantic salmon (Salmo salar L.) from the River Drammenselv, SE Norway. Fisheries Research, 10(1–2):125–135, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090018Q.

Hansson:1993:VHA

[Han93] Sture Hansson. Variation in hydroacoustic abundance of pelagic fish. Fisheries Research, 16(3):203–222, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390094N.

Hansen:1994:BRO

[Han94] Lars P. Hansen. Book review: Ocean ecology of North Pacific salmonids: William G. Pearcy. Washington Sea Grant Program, University of Washington Press, 1992, 179 pp., US$25 cloth, US$15 paper, ISBN 0-295-97192-4 cloth, and 0-295-97193-2 paper. Fisheries Research, 20(2–3):291, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900892.

Hansson:1997:BFS

[HAN97a] Sture Hansson, Fredrik Arrhenius, and Sture Nellbring. Benefits from fish stocking — experiences from stocking youngof-the-year pikeperch, Stizostedion lucioperca L. to a bay in the Baltic Sea. Fisheries Research, 32(2):123–132, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000507.

Hansson:1997:DGP

[HAN97b] Sture Hansson, Fredrik Arrhenius, and Sture Nellbring. Diet and growth of pikeperch (Stizostedion lucioperca L.) in a Baltic Sea area. Fisheries Research, 31(1–2):163–167, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000222.

Harris:1995:RTC

[Har95] Patrick J. Harris. The role of trap cameras in catch per unit effort calculations for species of the South Atlantic Bight snapper-grouper complex. Fisheries Research, 22(1–2):1–9, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400313L.

Harbitz:1998:OEA

[HAS98] Alf Harbitz, Michaela Aschan, and Knut Sunnanå. Optimal effort allocation in stratified, large area trawl surveys, with application to shrimp surveys in the Barents Sea. Fisheries Research, 37(1–3):107–113, August 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001301.

Haedrich:1997:CTS

[HB97] R. L. Haedrich and S. M. Barnes. Changes over time of the size structure in an exploited shelf fish community. Fisheries Research, 31(3):229–239, August 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000234.

He:1997:CAL

[HBB97] Xi He, Keith A. Bigelow, and Christofer H. Boggs. Cluster analysis of longline sets and fishing strategies within the Hawaii-based fishery. Fisheries Research, 31(1–2):147–158, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005644.

Halliday:1999:ESG

[HC99] R. G. Halliday and C. G. Cooper. Evaluation of separator grates for reduction of bycatch in the silver hake (Merluccius bilinearis) otter trawl fishery off Nova Scotia, Canada. Fisheries Research, 40(3):237–249, April 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369800229X.

Halliday:1999:ESS

[HCF^{+}99a] R. G. Halliday, C. G. Cooper, P. Fanning, W. M. Hickey, and P. Gagnon. Erratum to “Size selection of Atlantic cod, haddock and pollock (saithe) by otter trawls with square and diamond mesh codends of 130–155 mm mesh size”: [Fish. Res. 41 (1999) 255–271]. Fisheries Research, 44(2):197, December 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000880. See [HCF^{+}99b].

Halliday:1999:SSA

[HCF^{+}99b] R. G. Halliday, C. G. Cooper, P. Fanning, W. M. Hickey, and P. Gagnon. Size selection of Atlantic cod, haddock and pollock (saithe) by otter trawls with square and diamond mesh codends of 130–155 mm mesh size. Fisheries Research, 41(3): 255–271, July 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369900020X. See erra-

tum [HCF^{+}99a].

He:1991:SEA

[He91] Pingguo He. Swimming endurance of the Atlantic cod, Gadus morhua L., at low temperatures. Fisheries Research, 12 (1):65–73, September 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578369190050P.

He:1996:BLB

[He96] Pingguo He. Bait loss from bottom-set longlines as determined by underwater observations and comparative fishing trials. Fisheries Research, 27(1–3):29–36, June 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696004778.

Heath:1996:BRS

[Hea96] Alan G. Heath. Book review: Sublethal and chronic effects of pollutants on freshwater fish: R. Muller and R. Lloyd (Editors), Published by arrangement with the Food and Agriculture Organization of the United Nations by Fishing News Books, Blackwell Science, Oxford, UK, 1994, 371 pp., £55, ISBN 0-85238-207-3. Fisheries Research, 25(3–4):384–386, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696900016.

Huse:1990:FBS

[HF90] Ingvar Huse and Anders Fernö. Fish behaviour studies as an aid to improved longline hook design. Fisheries Research, 9(4):287–297, October 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090047Y.

Heifetz:1991:MDT

[HF91] Jonathan Heifetz and Jeffrey T. Fujioka. Movement dynamics of tagged sablefish in the northeastern Pacific. Fisheries Research, 11(3–4):355–374, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783691900095.

Hanumara:1993:IEP

[HG93] R. Choudary Hanumara and Sholeh Ghedari. Interval estimation of parameters in models for seasonal growth in fishes. Fisheries Research, 18(3–4):349–361, December 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390162Z.

Hart:1997:AEBb

[HG97] Anthony M. Hart and Harry K. Gorfine. Abundance estimation of blacklip abalone (Haliotis rubra) II. A comparative evaluation of catch-effort, change-in-ratio, mark-recapture and diver-survey methods. Fisheries Research, 29(2):171–183, February 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005292.

Huse:1998:EMP

[HG98] Geir Huse and Jarl Giske. Ecology in Mare Pentium: an individual-based spatio-temporal model for fish with adapted behaviour. Fisheries Research, 37(1–3):163–178, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001349.

Helser:1994:ESC

[HGC94] Thomas E. Helser, James P. Geaghan, and Richard E. Condrey. Estimating size composition and associated variances of a fish population from gillnet selectivity, with an example for spotted seatrout (Cynoscion nebulosus). Fisheries Research, 19(1–2):65–86, February 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900159.

Hart:1997:AEBa

[HGC97] Anthony M. Hart, Harry K. Gorfine, and Michael P. Callan. Abundance estimation of blacklip abalone (Haliotis rubra) I. An analysis of diver-survey methods used for large-scale monitoring. Fisheries Research, 29(2):159–169, February 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005279.

Hernandez-Garcia:1998:OOV

[HGHLC98] V. Hernández-Garcıá, J. L. Hernández-López, and J. J. Castro. The octopus (Octopus vulgaris) in the small-scale trap fishery off the Canary Islands (Central-East Atlantic). Fisheries Research, 35(3):183–189, May 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000800.

Huse:1999:RSG

[HGN99] Irene Huse, Agnes C. Gundersen, and Kjell H. Nedreaas. Relative selectivity of Greenland halibut (Reinhardtius hippoglossoides, Walbaum) by trawls, longlines and gillnets. Fisheries Research, 44(1):75–93, November 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000417.

Hesthagen:1990:EOG

[HH90] Trygve Hesthagen and Halgeir Holthe. Estimation of optimum gill net mesh size for maximum yield in an exploited population of brown trout (Salmo trutta L.) in southern central Norway. Fisheries Research, 9(1):57–66, July 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783690900403.

Hudson:1993:EDE

[HH93] Eric B. Hudson and Barry J. Hill. EC directive 91/67/EEC and its relevance to disease in wild salmonids. Fisheries Research, 17(1–2):229–236, June 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390022Y.

Hesthagen:1992:EFD

[HJ92] Trygve Hesthagen and Bjørn Ove Johnsen. Effects of fish density and size on survival, growth and production of hatchery-reared brown trout (Salmo trutta L.) in lakes. Fisheries Research, 15(1–2):147–156, October 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290010Q.

Heggberget:1993:IBW

[HJH^{+}93] Tor G. Heggberget, Bjørn O. Johnsen, Kjetil Hindar, Bror Jonsson, Lars P. Hansen, Nils A. Hvidsten, and Arne J. Jensen. Interactions between wild and cultured Atlantic salmon: a review of the Norwegian experience. Fisheries Research, 18(1–2):123– 146, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783693900448.

Holts:1998:PSF

[HJSNB98] David B. Holts, Alfredo Julian, Oscar Sosa-Nishizaki, and Norm W. Bartoo. Pelagic shark fisheries along the west coast of the United States and Baja California, Mexico. Fisheries Research, 39(2):115–125, December 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001787.

Hemmingsen:1993:OLA

[HLES93] W. Hemmingsen, D. A. Lysne, T. Eidnes, and A. Skorping. The occurrence of larval ascaridoid nematodes in wild-caught and in caged and artificially fed Atlantic cod, Gadus morhua L., in Norwegian waters. Fisheries Research, 15(4):379–386, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390088O.

Hattula:1995:ECM

[HLK^{+}95] Tapani Hattula, Tiina Luoma, Rauno Kostiainen, Jouko Poutanen, Marja Kallio, and Petri Suuronen. Effects of catching method on different quality parameters of Baltic herring (Clupea harengus L.). Fisheries Research, 23(3–4):209–221, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694003584.

Hemmingsen:1991:PBT

[HLM91] W. Hemmingsen, I. Lombardo, and K. MacKenzie. Parasites as biological tags for cod, Gadus morhua L., in northern Norway: a pilot study. Fisheries Research, 12(4):365–373, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190019C.

Holland:1996:MDP

[HLW96] Kim N. Holland, Christopher G. Lowe, and Bradley M. Wetherbee. Movements and dispersal patterns of blue trevally (Caranx melampygus) in a fisheries conservation zone. Fisheries Research, 25(3–4):279–292, March 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004424.

Healey:1992:RBD

[HM92] M. C. Healey and J. F. T. Morris. The relationship between the dispersion of salmon fishing vessels and their catch.

Fisheries Research, 15(1–2):135–145, October 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290009I.

Hartleb:1995:IGL

[HM95] Christopher F. Hartleb and John R. Moring. An improved gastric lavage device for removing stomach contents from live fish. Fisheries Research, 24(3):261–265, October 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369500390V.

Herke:1999:SHE

[HM99] Scott W. Herke and John R. Moring. “Soft” harness for external attachment of large radio transmitters to northern pike (Esox lucius). Fisheries Research, 39(3):305–312, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001970.

Huolila:1997:SCS

[HML97] Markus Huolila, Timo J. Marjomäki, and Eero Laukkanen. The success of crayfish stocking in a dredged river with and without artificial shelter increase. Fisheries Research, 32(2):185–189, November 1, 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000465.

Horppila:1996:DHS

[HMP96] Jukka Horppila, Tommi Malinen, and Heikki Peltonen. Density and habitat shifts of a roach (Rutilus rutilus) stock assessed within one season by cohort analysis, depletion methods and echosounding. Fisheries Research, 28(2):151–161, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783696004997.

Hobbs:1991:BRM

[Hob91] G. Hobbs. Book review: Microbiology of marine food products: D. R. Ward and C. R. Hackney (Editors). Chapman and Hall, London, 1990, 464 pp., £45.50, ISBN 0-442-23346-9. Fisheries Research, 12(2):183–184, October 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190042E.

Hoel:1998:PUI

[Hoe98] Alf Håkon Hoel. Political uncertainty in international fisheries management. Fisheries Research, 37(1–3):239–250, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001404.

Holst:1993:ODL

[Hol93] Jens Christian Holst. Observations on the distribution of lumpsucker (Cyclopterus lumpus, L.) in the Norwegian Sea. Fisheries Research, 17(3–4):369–372, August 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390136U.

Holst:1994:PTM

[Hol94] Jens Christian Holst. The precision of two macroscopic screening procedures relative to a microscopic procedure for screening of the fungus Ichthyophonus hoferi in herring (Clupea harengus L.). Fisheries Research, 19(1–2):187–190, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369490023X.

Holden:1995:BHR

[Hol95] Mike Holden. Beverton and Holt revisited. Fisheries Research, 24(1):3–8, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578369500377M.

Holst:1996:EPI

[Hol96] Jens Christian Holst. Estimating the prevalence of Ichthyophonus hoferi (Plehn and Mulsow) in a herring stock (Clupea harengus L.): Observed effects of sampling gear, target school density and migration. Fisheries Research, 28(1):85–97, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004653.

Hopkins:1992:BRP

[Hop92] Christopher C. E. Hopkins. Book review: The Peruvian upwelling ecosystem: Dynamics and interactions: D. Pauly P. Muck J. Dendo and I. Tsukayama (Editors), International Center for Living Aquatic Resources Management (ICLARM) Conference Proceedings 18, Instituto del Mar del Perú (IMARPE), Callao, Perú, Deutsche Gesellschaft für Technische Zusammenarbeit (GTZ), GmbH, Eschborn, Germany, and ICLARM, Manila, Philippines, 1989, 438 pp., US$15.00 (surface mail), ISBN 971-10-2247-8. Fisheries Research, 13(2):199– 200, February 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290027Q.

Horbowy:1998:CAL

[Hor98] Jan Horbowy. Comparison of age–length keys of Baltic cod derived from Polish commercial and research data. Fisheries Research, 36(2–3):257–266, June 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000952.

Howe:1993:BRS

[How93] Jeffrey C. Howe. Book review: Systematics, historical ecology, and North American freshwater fishes: Richard L. Mayden (editor), Stanford University Press, Stanford, CA, 1992, 969 pp., US $69.50, ISBN 0-8047 2162-9. Fisheries Research, 18(3–4):389–392, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693901654.

Howe:1996:TIP

[How96a] Jeffrey C. Howe. A technique for immobilizing and photographing small, live fishes. Fisheries Research, 27(4):261– 264, July 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369500470X.

Howe:1996:BRF

[How96b] Jeffry C. Howe. Book review: Fish nutrition in aquaculture: S. S. De Silva and T. A. Anderson, Chapman & Hall, London, 1995, 319 pp., £35, ISBN 0-412-55030-X. Fisheries Research, 26(3–4):387–390, May 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696900065.

Hilborn:1994:BED

[HPM94] Ray Hilborn, Ellen K. Pikitch, and Murdoch K. McAllister. A Bayesian estimation and decision analysis for an age-structured model using biomass survey data. Fisheries Research, 19(1–2):17–30, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900124.

Hagen:1991:LTG

[HQ91] Peter T. Hagen and Terrance J. Quinn. Long-term growth dynamics of young Pacific halibut: evidence of temperatureinduced variation. Fisheries Research, 11(3–4):283–306, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783691900062.

Hovgaard:1992:CLT

[HR92] Holger Hovgård and Frank F. Riget. Comparison of longline and trawl selectivity in cod surveys off West Greenland. Fisheries Research, 13(3):323–333, March 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783692900858.

Hsu:1999:LWR

[Hsu99] Chien-Chung Hsu. The length–weight relationship of albacore, Thunnus alalunga, from the Indian Ocean. Fisheries Research, 41(1):87–92, May 1, 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000028.

Hoggarth:1994:FEL

[HU94] Daniel D. Hoggarth and Agus D. Utomo. The fisheries ecology of the Lubuk Lampam River floodplain in south Sumatra, Indonesia. Fisheries Research, 20(2–3):191–213, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900833.

Hughes:1998:MHH

[Hug98] Simon Hughes. A mobile horizontal hydroacoustic fisheries survey of the River Thames, United Kingdom. Fisheries Research, 35(1–2):91–97, May 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698000630.

Hurley:1998:RFP

[Hur98] P. C. F. Hurley. A review of the fishery for pelagic sharks in Atlantic Canada. Fisheries Research, 39(2):107–113, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001775.

Hussain:1992:RBL

[Hus92] Syed Makhdoom Hussain. The reproductive biology of the lantern fish Benthosema fibulatum from the northern Arabian Sea. Fisheries Research, 13(4):381–393, April 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783692900593.

Hill:1999:RSC

[HW99] B. J. Hill and T. J. Wassenberg. The response of spanner crabs (Ranina ranina) to tangle nets — behaviour of the crabs on the nets, probability of capture and estimated distance of attraction to bait. Fisheries Research, 41(1):37–46, May 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000090.

Ibrahim:1996:IFA

[IASS96] Sakri Ibrahim, Mohd Azmi Ambak, Lokman Shamsudin, and Mohd Zaini Samsudin. Importance of fish aggregating devices (FADs) as substrates for food organisms of fish. Fisheries Research, 27(4):265–273, July 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696004730.

Ismen:1999:NIW

[IB99] A. Ismen and F. Bingel. Nematode infection in the whiting Merlangius merlangus euxinus off Turkish Coast of the Black Sea. Fisheries Research, 42(1–2):183–189, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000223.

Ianelli:1995:DAA

[IH95a] James N. Ianelli and Jonathan Heifetz. Decision analysis of alternative harvest policies for the Gulf of Alaska Pacific Ocean perch fishery. Fisheries Research, 24(1):35–63, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369594406G. See erratum

[IH95b].

Ianelli:1995:EDA

[IH95b] James N. Ianelli and Jonathan Heifetz. Erratum to “Decision analysis of alternative harvest policies for the Gulf of Alaska Pacific Ocean Perch Fishery”: [Fish. Res. 24 (1995) 35–63]. Fisheries Research, 24(4):345–349, November 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783696800099. See [IH95a].

Ibrahim:1990:ERT

[IKA90] Sakri Ibrahim, Gunzo Kawamura, and Mohd Azmi Ambak. Effective range of traditional Malaysian FAD as determined by fish-releasing method. Fisheries Research, 9(4):299–306, October 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090048Z.

Ikomi:1996:SGP

[Iko96] R. B. Ikomi. Studies on the growth pattern, feeding habits and reproductive characteristics of the mormyrid Brienomyrus longianalis (Boulenger 1901) in the upper Warri River, Nigeria. Fisheries Research, 26(1–2):187–198, April 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695003819.

Ikomi:1998:SAE

[IO98] R. B. Ikomi and O. Odum. Studies on aspects of the ecology of the catfish Chrysichthys auratus Geoffrey St. Hilaire (Osteichthyes; Bagridae) in the River Benin (Niger Delta, Nigeria). Fisheries Research, 35(3):209–218, May 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369800071X.

Iribarne:1991:IHP

[Iri91] Oscar Osvaldo Iribarne. Intertidal harvest of the Patagonian octopus, Octopus tehuelchus (d’Orbigny). Fisheries Research, 12(4):375–390, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190020G.

Ismen:1995:FWM

[Ism95a] Ali Ismen. Fecundity of whiting, Merlangius merlangus euxinus (L.) on the Turkish Black Sea coast. Fisheries Research, 22(3–4):309–318, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400327S.

Ismen:1995:GMY

[Ism95b] Ali Ismen. Growth, mortality and yield per recruit model of picarel (Spicara smaris L.) on the eastern Turkish Black Sea coast. Fisheries Research, 22(3–4):299–308, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369594955P.

Isaksen:1992:RFC

[IVLK92] B. Isaksen, J. W. Valdemarsen, R. B. Larsen, and L. Karlsen. Reduction of fish by-catch in shrimp trawl using a rigid separator grid in the aft belly. Fisheries Research, 13(3):335– 352, March 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900869.

Iversen:1993:SSD

[IZJT93] S. A. Iversen, D. Zhu, A. Johannessen, and R. Toresen. Stock size, distribution and biology of anchovy in the Yellow Sea and East China Sea. Fisheries Research, 16(2):147– 163, March 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390049D.

Jennings:1999:FEN

[JAC^{+}99] S. Jennings, J. Alvsvåg, A. J. R. Cotter, S. Ehrich, S. P. R. Greenstreet, A. Jarre-Teichmann, N. Mergardt, A. D. Rijnsdorp, and O. Smedstad. Fishing effects in northeast Atlantic shelf seas: patterns in fishing effort, diversity and community structure. III. International trawling effort in the North Sea: an analysis of spatial and temporal trends. Fisheries Research, 40(2):125–134, March 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002082.

Jorgensen:1999:SDA

[JAG^{+}99] L. Jørgensen, P.-A. Amundsen, H.-M. Gabler, M. Halvorsen, J. Erkinaro, and E. Niemelä. Spatial distribution of Atlantic salmon parr (Salmo salar L.) and bullhead (Cottus gobio L.) in lotic and lentic habitats of a diversified watercourse in northern Fennoscandia. Fisheries Research, 41(2):201–211, June 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000144.

Jansen:1993:RPMa

[JB93a] Peder A. Jansen and Tor A. Bakke. Regulatory processes in the monogenean Gyrodactylus salaris Malmberg — Atlantic salmon (Salmo salar L.) association. I. Field studies in southeast Norway. Fisheries Research, 17(1–2):87–101, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390009V.

Jansen:1993:RPMb

[JB93b] Peder A. Jansen and Tor A. Bakke. Regulatory processes in the monogenean Gyrodactylus salaris Malmberg — Atlantic salmon (Salmo salar L.) association. II. Experimental studies. Fisheries Research, 17(1–2):103–114, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783693900105.

Johnston:1993:SBE

[JB93c] S. J. Johnston and M. O. Bergh. Size-based estimates of survivorship for the South African rock lobster Jasus lalandii. Fisheries Research, 18(3–4):277–304, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783693901584.

Jug-Dujakovic:1995:EYS

[JDDK95] J. Jug-Dujaković, J. Dulčić, and I. Katavić. Embryonic and yolk-sac larval development of the sparid Dentex (Dentex) dentex (Linnaeus, 1758). Fisheries Research, 24(2):91–97, September 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369500372H.

Jensen:1990:ASG

[Jen90] Jens Møller Jensen. Atlantic salmon at Greenland. Fisheries Research, 10(1–2):29–52, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090014M.

Jennings:1991:SSD

[Jen91] S. Jennings. Size selectivity of drag seines employed for capture of juvenile bass, Dicentrarchus labrax (L.). Fisheries Research, 12(2):101–108, October 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190035E.

Jensen:1996:REM

[Jen96] A. L. Jensen. Ratio estimation of mortality using catch curves. Fisheries Research, 27(1–3):61–67, June 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004688.

Janusz:1997:SWP

[JH97] Jerzy Janusz and Jan Horbowy. The state of the walleye pollock in the northern part of the Okhotsk Sea, North Pacific. Fisheries Research, 30(1–2):87–102, May 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005619.

Juell:1996:SAO

[JHB96] Jon-Erik Juell, Jens Christian Holm, and Arvid K. Beltestad. In situ acoustic observation of food consumption in caged Atlantic mackerel (Scomber scombrus). Fisheries Research, 27(1–3):131–135, June 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696004791.

Jennings:1993:DMS

[JHF93] S. Jennings, G. J. Hewlett, and S. Flatman. The distribution, migrations and stock integrity of lemon sole Microstomus kitt in the western English Channel. Fisheries Research, 18(3–4):377–388, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693901643.

Jurvelius:1996:SAP

[JLMP96] Juha Jurvelius, Jouni Leinikki, Victor Mamylov, and Sergei Pushkin. Stock assessment of pelagic three-spined stickleback (Gasterosteus aculeatus): a simultaneous up- and downlooking echo-sounding study. Fisheries Research, 27(4):227– 241, July 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004645.

Johnstone:1990:LIB

[JM90] A. D. F. Johnstone and A. M. Mackie. Laboratory investigations of bait acceptance by the cod, Gadus morhua L.: identification of feeding stimulants. Fisheries Research, 9(3):219–230, September 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605800037.

Johannessen:1991:HLC

[JM91] A. Johannessen and E. Moksness. Herring larvae (Clupea harengus) in the Skagerrak area from December 1987 to April 1988. Fisheries Research, 11(2):155–170, April 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190105O.

Joll:1990:AHR

[JP90] L. M. Joll and J. W. Penn. The application of high-resolution navigation systems to Leslie–DeLury depletion experiments for the measurement of trawl efficiency under open-sea conditions. Fisheries Research, 9(1):41–55, July 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090039X.

Jurvelius:1999:EEW

[JRM99] Juha Jurvelius, Raimo Riikonen, and Timo J. Marjomäki. Estimating the effects of winter-fishing: mobile echosounding under ice. Fisheries Research, 39(3):313–316, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698002021.

Jakobsson:1998:RHC

[JS98] Jakob Jakobsson and Gunnar Stefánsson. Rational harvesting of the cod–capelin–shrimp complex in the Icelandic marine ecosystem. Fisheries Research, 37(1–3):7–21, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001234.

Johnson:1992:HSD

[JSE92] G. E. Johnson, C. M. Sullivan, and M. W. Erho. Hydroacoustic studies for developing a smolt bypass system at Wells Dam. Fisheries Research, 14(2–3):221–237, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290055X.

Johal:1992:AGC

[JT92] M. S. Johal and K. K. Tandon. Age and growth of the carp Catla catla (Hamilton, 1822) from northern India. Fisheries Research, 14(1):83–90, June 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783692900755.

Jin:1996:CFS

[JT96] Xianshi Jin and Qisheng Tang. Changes in fish species diversity and dominant species composition in the Yellow Sea. Fisheries Research, 26(3–4):337–352, May 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369500422X.

Kaimmer:1994:HIM

[Kai94] Stephen M. Kaimmer. Halibut injury and mortality associated with manual and automated removal from setline hooks. Fisheries Research, 20(2–3):165–179, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900817.

Kaya:1999:AGS

[KBKO99] Murat Kaya, H. Avni Benli, Tuncer Katagan, and Okan Ozaydin. Age, growth, sex-ratio, spawning season and food of golden banded goatfish, Upeneus moluccensis Bleeker (1855) from the Mediterranean and south Aegean Sea coasts of Turkey. Fisheries Research, 41(3):317–328, July 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000272.

Kaiser:1999:FEN

[KCS^{+}99] M. J. Kaiser, K. Cheney, F. E. Spence, D. B. Edwards, and K. Radford. Fishing effects in northeast Atlantic shelf seas: patterns in fishing effort, diversity and community structure. VII. The effects of trawling disturbance on the fauna associated with the tubeheads of serpulid worms.

Fisheries Research, 40(2):195–205, March 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002124.

Karacam:1990:AGM

[KD90] Hikmet Karacam and Ertug Düzgünes. Age, growth and meat yield of the European anchovy (Engraulis encrasicolus, L. 1758) in the Black Sea. Fisheries Research, 9(2):181–186, August 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783690900632.

Kraljevic:1997:AGG

[KD97] Miro Kraljević and Jakov Dulčić. Age and growth of gilthead sea bream (Sparus aurata L.) in the Mirna estuary, northern Adriatic. Fisheries Research, 31(3):249–255, August 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000167.

Kubecka:1998:ASV

[KD98a] J. Kubecka and A. Duncan. Acoustic size vs. real size relationships for common species of riverine fish. Fisheries Research, 35(1–2):115–125, May 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698000666.

Kubecka:1998:DCF

[KD98b] Jan Kubecka and Annie Duncan. Diurnal changes of fish behaviour in a lowland river monitored by a dual-beam echosounder. Fisheries Research, 35(1–2):55–63, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000599.

Kraljevic:1996:AGM

[KDCP96] M. Kraljević, J. Dulčić, P. Cetinić, and A. Pallaoro. Age, growth and mortality of the striped sea bream, Lithognathus mormyrus L., in the Northern Adriatic. Fisheries Research, 28(4):361–370, December 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696005188.

Kubecka:1994:BTP

[KDD^{+}94] J. Kubecka, A. Duncan, W. M. Duncan, D. Sinclair, and A. J. Butterworth. Brown trout populations of three Scottish lochs estimated by horizontal sonar and multimesh gill nets. Fisheries Research, 20(1):29–48, June 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900043.

Kendall:1990:SRC

[Ken90] Drew Kendall. Shrimp retention characteristics of the Morrison soft TED: a selective webbing exclusion panel inserted in a shrimp trawl net. Fisheries Research, 9(1):13–21, July 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090037V.

Kenchington:1996:LTS

[Ken96] T. J. Kenchington. Long-term stability and change in the commercial groundfish longline fishing grounds of the northwest Atlantic. Fisheries Research, 25(2):139–154, February 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004149.

Kesteven:1996:ARF

[Kes96a] G. L. Kesteven. Annual report 1994–1995: Fisheries department of Western Australia 58 pp. Fisheries Research, 28 (3):353–359, October 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783696900247.

Kesteven:1996:BRA

[Kes96b] G. L. Kesteven. Book review: Assessment methodologies and management: Proceedings of the World Fisheries Congress, Theme 5 Sakagawa, T. Gary (editor). Science Publishers, USA, 1995, 195 pp., £79, hard cover, ISBN 1-886106-10-X. Fisheries Research, 28(3):353–359, October 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369600505X.

Kesteven:1996:BRF

[Kes96c] G. L. Kesteven. Book review: Fishery status reports 1994: McLoughlin, Kevin et al. (editors), Bureau of Resource Sciences, Australia, 1995, 144 pp., A$20 soft cover, ISSN 1322-655X. Fisheries Research, 28(3):353–359, October 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696900259.

Kesteven:1996:FSA

[Kes96d] G. L. Kesteven. A fisheries science approach to problems of world fisheries or: three phases of an industrial revolution. Fisheries Research, 25(1):5–17, January 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004378.

Kesteven:1996:PA

[Kes96e] G. L. Kesteven. The precautionary approach. Fisheries Research, 28(1):109–110, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/0165783696005048. See replies [McI97d, Gar97].

Kesteven:1997:BRP

[Kes97a] G. L. Kesteven. Book review: Patterns in the ocean. Ocean processes and marine population dynamics: By Andrew Bakun. California Sea Grant College System, National Oceanic and Atmospheric Administration with Centro de Investigaciones Biologicas del Noroeste, La Paz, BCS Mexico. 1996, 323 pp., $24 soft cover, ISBN 1-888691-01-8. Fisheries Research, 32(1):102– 105, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369700009X.

Kesteven:1997:FIS

[Kes97b] G. L. Kesteven. Fishery independent surveys and data. Fisheries Research, 32(3):271–272, December 1, 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000325.

Kesteven:1999:FRF

[Kes99] G. L. Kesteven. Fisheries research and fisheries legislation. Fisheries Research, 40(3):207–211, April 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002197.

Kennelly:1993:VCB

[KGM^{+}93] S. J. Kennelly, K. J. Graham, S. S. Montgomery, N. L. Andrew, and P. A. Brett. Variance and cost-benefit analyses to determine optimal duration of tows and levels of replication for sampling relative abundances of species using demersal trawling. Fisheries Research, 16(1):51–67, February 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390109K.

Kideys:1994:ETG

[Kid94] Ahmet E. Kideys. Effect of tagging on the growth rate of the neogastropod, Buccinum undatum in the laboratory. Fisheries Research, 20(2–3):283–289, August 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900884.

Kenchington:1993:ECC

[KK93] T. J. Kenchington and E. L. R. Kenchington. An eighteenth century commercial length-frequency sample of Atlantic cod, Gadus morhua, based on archaeological data. Fisheries Research, 18(3–4):335–347, December 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390161Y.

Karenge:1995:RBH

[KK95a] Lawrence Karenge and Jeppe Kolding. On the relationship between hydrology and fisheries in man-made Lake Kariba, central Africa. Fisheries Research, 22(3–4):205–226, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400325Q.

Kobayashi:1995:ESD

[KK95b] Donald R. Kobayashi and Kurt E. Kawamoto. Evaluation of shark, dolphin, and monk seal interactions with Northwestern Hawaiian Island bottomfishing activity: a comparison of two time periods and an estimate of economic impacts. Fisheries Research, 23(1–2):11–22, May 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400338W.

Kosior:1997:MOM

[KK97] Maria Kosior and Jerzy Kuczyński. Macroscopic observations of a maturation process in eelpout (Zoarces viviparus L.) from the southern Baltic. a proposed gonadal maturation scale. Fisheries Research, 30(1–2):151–155, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000027.

Kennelly:1998:RDC

[KLB98] S. J. Kennelly, G. W. Liggins, and M. K. Broadhurst. Retained and discarded by-catch from oceanic prawn trawling in New South Wales, Australia. Fisheries Research, 36(2–3):217– 236, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698000915.

Klyashtorin:1998:LTC

[Kly98] Leonid B. Klyashtorin. Long-term climate change and main commercial fish production in the Atlantic and Pacific. Fisheries Research, 37(1–3):115–125, August 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001313.

Kawamura:1996:BGF

[KMNM96] Gunzo Kawamura, Tomohiro Matsushita, Mikio Nishitai, and Tatsuro Matsuoka. Blue and green fish aggregation devices are more attractive to fish. Fisheries Research, 28(1):99–108, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369600478X.

Kinloch:1997:EBR

[KMNP97] Martine A. Kinloch, David McGlennon, Gillian Nicoll, and Phillip G. Pike. Evaluation of the bus-route creel survey method in a large Australian marine recreational fishery: I. Survey design. Fisheries Research, 33(1–3):101–121, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000684.

Kawamura:1995:ESF

[KMT^{+}95] Gunzo Kawamura, Tatsuro Matsuoka, Takayuki Tajiri, Masataka Nishida, and Mitsuru Hayashi. Effectiveness of a sugarcane-fish combination as bait in trapping swimming crabs. Fisheries Research, 22(1–2):155–160, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400314M.

Knuth:1991:BRW

[Knu91] Barbara A. Knuth. Book review: Winds of change: Women in Northwest commercial fishing: C. J. Allison and M. A. Porter, University of Washington Press, Seattle, U.S.A., 1989, 232 pp., US$25.00, ISBN 0-295-96840-0, LC: 89-16422 (1990). Fisheries Research, 10(3–4):362–364, January 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190091S.

Kusemiju:1990:SOB

[KO90] K. Kusemiju and C. A. Onadeko. The seasonal occurrence and bionomics of the bonga, Ethmalosa fimbriata (Bowdich) off Aiyetoro Coast, Nigeria. Fisheries Research, 8(3):247–251, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090025Q.

Kulka:1996:ACS

[KPH^{+}96] D. W. Kulka, A. T. Pinhorn, R. G. Halliday, D. Pitcher, and D. Stansbury. Accounting for changes in spatial distribution of groundfish when estimating abundance from commercial fishing data. Fisheries Research, 28(3):321–342, October 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783696005085.

Kurkilahti:1996:CSU

[KR96] Mika Kurkilahti and Martti Rask. A comparative study of the usefulness and catchability of multimesh gill nets and gill net series in sampling of perch (Perca fluviatilis L.) and roach (Rutilus rutilus L.). Fisheries Research, 27(4):243– 260, July 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004637.

Karlou-Riga:1997:AGH

[KRS97] C. Karlou-Riga and A. Sinis. Age and growth of horse mackerel, Trachurus trachurus (L.), in the Gulf of Saronikos (Greece). Fisheries Research, 32(2):157–171, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000441.

Kawamura:1993:OCD

[KS93] Gunzo Kawamura and Takashi Shimowada. Optic critical duration and contrast thresholds in the freshwater fish, Lepomis macrochirus, as determined behaviourally. Fisheries Research, 17(3–4):251–258, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390128T.

Kaimmer:1998:ICM

[KT98] Stephen M. Kaimmer and Robert J. Trumble. Injury, condition, and mortality of Pacific halibut bycatch following careful release by Pacific cod and sablefish longline fisheries. Fisheries Research, 38(2):131–144, October 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001532.

Kubecka:1992:FFN

[Kub92] Jan Kubecka. Fluctuations in fyke-net catches during the spawning period of the Eurasian perch (Perca fluviatilis) in the Rimov Reservoir, Czechoslovakia. Fisheries Research, 15(1–2):157–167, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290011H.

Karunasinghe:1991:SEA

[KW91] W. P. N. Karunasinghe and M. J. S. Wijeyaratne. Selectivity estimates for Amblygaster sirm (Clupeidae) in the small-meshed gill net fishery on the west coast of Sri Lanka. Fisheries Research, 10(3–4):199–205, January 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190075Q.

Kim:1998:MBC

[KW98a] Y.-H. Kim and C. S. Wardle. Measuring the brightness contrast of fishing gear, the visual stimulus for fish capture. Fisheries Research, 34(2):151–164, March 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000878.

Kim:1998:MVS

[KW98b] Y.-H. Kim and C. S. Wardle. Modelling the visual stimulus of towed fishing gear. Fisheries Research, 34(2):165– 177, March 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000891.

Kubecka:1998:HBC

[KW98c] Jan Kubecka and Monika Wittingerova. Horizontal beaming as a crucial component of acoustic fish stock assessment in freshwater reservoirs. Fisheries Research, 35(1–2):99–106, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698000642.

Kyle:1994:ATL

[Kyl94] G. B. Kyle. Assessment of trophic-level responses and Coho salmon (Oncorhynchus kisutch) production following nutrient treatment (1981–1986) of Bear Lake, Alaska. Fisheries Research, 20(2–3):243–261, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900868.

Kyle:1996:SSS

[Kyl96] Gary B. Kyle. Stocking sockeye salmon (Oncorhynchus nerka) in barren lakes of Alaska: effects on the macrozooplankton community. Fisheries Research, 28(1):29–44, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783696004869.

Levi:1994:ORT

[LAA^{+}94] D. Levi, M. G. Andreoli, E. Arneri, G. Giannetti, and P. Rizzo. Otolith reading as a tool for stock identification. Fisheries Research, 20(2–3):97–107, August 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900779.

Levi:1993:ABT

[LAG93] Dino Levi, M. G. Andreoli, and G. B. Giusto. An analysis based on trawl-survey data of the state of the ‘Italian’ stock of Mullus barbatus in the Sicilian Channel, including management advice. Fisheries Research, 17(3–4):333–341, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390133R.

Levi:1995:FAR

[LAG95] D. Levi, M. G. Andreoli, and R. M. Giusto. First assessment of the rose shrimp, Parapenaeus longirostris (Lucas, 1846) in the central Mediterranean. Fisheries Research, 21(3–4):375– 393, January 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400298B.

Laird:1993:BRC

[Lai93] Lindsay Laird. Book review: Crustacean farming: Daniel O’C. Lee and John Wickins. Blackwell Scientific, Oxford, 1992, 392 pp., hard back £35, ISBN 0-632-02974-9. Fisheries Research, 15(4):390–391, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390091K.

Laird:1995:BRA

[Lai95a] Lindsay Laird. Book review: Aquaculture development: Progress and prospects: T. V. R. Pillay, Fishing News Books, 1994, 182 pp., £35, hardback. ISBN 0-85238-218-9. Fisheries Research, 24(2):179, September 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695900357.

Laird:1995:BRP

[Lai95b] Lindsay M. Laird. Book review: Principles of aquaculture: Robert R. Stickney, Wiley and Sons, 1994, 502 pp., £58, ISBN 0-471-57856-8. Fisheries Research, 24(1):86–87, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695900136.

Landa:1999:DDG

[Lan99a] J. Landa. Density-dependent growth of four spot megrim (L. boscii) in the northern Spanish shelf. Fisheries Research, 40(3):267–276, April 1, 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698002276.

Landergren:1999:SAR

[Lan99b] Peter Landergren. Spawning of anadromous rainbow trout, Oncorhynchus mykiss (Walbaum): a threat to sea trout, Salmo trutta L., populations? Fisheries Research, 40(1):55–63, February 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369800215X.

Lokkeborg:1992:SSS

[LB92] Svein Løkkeborg and Åsmund Bjordal. Species and size selectivity in longline fishing: a review. Fisheries Research, 13(3):311–322, March 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783692900847.

Lokkeborg:1995:SSE

[LB95] Svein Løkkeborg and Åsmund Bjordal. Size-selective effects of increasing bait size by using an inedible body on longline hooks. Fisheries Research, 24(4):273–279, November 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695003936.

Lagardere:1995:EIT

[LBBB95] Jean Paul Lagardère, Marie-Laure Bégout Anras, Hervé Breton, and Joan Baptista Company i Claret. The effects of illumination, temperature and oxygen concentration on swimming activity of turbot Psetta maxima (Linné 1758). Fisheries Research, 24(2):165–171, September 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400361Y.

Liggins:1997:DBO

[LBK97] G. W. Liggins, M. J. Bradley, and S. J. Kennelly. Detection of bias in observer-based estimates of retained and discarded catches from a multi species trawl fishery. Fisheries Research, 32(2):133–147, November 1, 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000532.

Liu:1999:VPA

[LC99] Kwang-Ming Liu and Yeu-Long Cheng. Virtual population analysis of the big eye Priacanthus macracanthus in the waters off northeastern Taiwan. Fisheries Research, 41(3):243– 254, July 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000259.

Lobon-Cervia:1993:SMD

[LCU93] Javier Lobón-Cerviá and Carmen G. Utrilla. A simple model to determine stream trout (Salmo trutta L.) densities based on one removal with electrofishing. Fisheries Research, 15(4):369– 378, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390087N.

Levi:1996:PLC

[Lev96] Dino Levi. Parapenaeus longirostris in the central Mediterranean. Fisheries Research, 28(1):111, August 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696004894.

Lannan:1993:PSM

[LF93] C. N. Lannan and J. L. Fryer. Piscirickettsia salmonis, a major pathogen of salmonid fish in Chile. Fisheries Research, 17(1–2):115–121, June 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390011U.

Laptikhovsky:1999:SFD

[LF99] Vladimir Laptikhovsky and Andrej Fetisov. Scavenging by fish of discards from the Patagonian squid fishery. Fisheries Research, 41(1):93–97, May 1, 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698002288.

Lowe:1991:BAM

[LFT91] Sandra A. Lowe, Jeffrey T. Fujioka, and Joseph M. Terry. Bioeconomic analysis of a minimum size limit for Gulf of Alaska sablefish using a yield per recruit model. Fisheries Research, 11(3–4):307–320, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783691900073.

Li:1990:SMS

[LH90] Xinyang Li and Enhuai Hou. Study of mesh stress and shape of an axisymmetrical net using the membrane theory. Fisheries Research, 9(3):269–282, September 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605800062.

Larsen:1997:PSC

[LHML97] G. Larsen, W. Hemmingsen, K. MacKenzie, and D. A. Lysne. A population study of cod, Gadus morhua L., in northern Norway using otolith structure and parasite tags. Fisheries Research, 32(1):13–20, October 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000404.

Lysne:1995:PDR

[LHS95] D. A. Lysne, W. Hemmingsen, and A. Skorping. Pepsin digestion reveals both previous and present infections of metacercariae in the skin of fish. Fisheries Research, 24(2):173–177, September 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369500387P.

Lowe:1998:NAT

[LHW98] Christopher G. Lowe, Kim N. Holland, and Thomas G. Wolcott. A new acoustic tailbeat transmitter for fishes. Fisheries Research, 36(2–3):275–283, June 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369800109X.

Li:1999:ERH

[Li99] Eric A. L. Li. Expandability, reversibility and harvesting under uncertainty. Fisheries Research, 44(2):167–174, December 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000776.

Lipinski:1999:TPC

[Lip99] M. R. Lipinski. Three publications on cephalopods. Fisheries Research, 44(2):193–196, December 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000892.

Lokkeborg:1992:ICS

[LJ92] Svein Løkkeborg and Tore Johannessen. The importance of chemical stimuli in bait fishing — fishing trials with presoaked bait. Fisheries Research, 14(1):21–29, June 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290070A.

Liggins:1996:CPT

[LK96] G. W. Liggins and S. J. Kennelly. By-catch from prawn trawling in the Clarence River estuary, New South Wales, Australia. Fisheries Research, 25(3–4):347–367, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004157.

Letcher:1999:TSI

[LK99] Benjamin H. Letcher and Tim L. King. Targeted stock identification using multilocus genotype ‘familyprinting’. Fisheries Research, 43(1–3):99–111, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000685.

Lu:1995:SIF

[LL95] H. J. Lu and K. T. Lee. Species identification of fish shoals from echograms by an echo-signal image processing system. Fisheries Research, 24(2):99–111, September 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694003687.

Lee:1996:LAE

[LL96] M. A. Lee and K. T. Lee. The larval anchovy (Engraulis japonicus) fishery in relation to the environmental factors in coastal waters of fangliao, Taiwan. Fisheries Research, 26(1–2):37– 48, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004122.

Lu:1998:RBN

[LLL98] Hsueh-Jung Lu, Kuo-Tien Lee, and Cheng Hsin Liao. On the relationship between El Niño/Southern Oscillation and South Pacific albacore. Fisheries Research, 39(1):1–7, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369800174X.

Li:1997:IIM

[LLS97] Hui Li, Qun Liu, and Zhengda Shen. Investigation of an iterative method for fitting von Bertalanffy growth equation using data on the halfbeak fish Hemirhamphus sajori (T et S). Fisheries Research, 29(3):289–294, March 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005371.

Lee:1996:BRL

[LLW96] Kuo-Tien Lee, Ming-Anne Lee, and Jeng-Pei Wang. Behavioural responses of larval anchovy schools herded within large-mesh wings of trawl net. Fisheries Research, 28(1):57– 69, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783696004857.

Lloyd:1998:PTN

[LM98] J. A. Lloyd and R. P. Mounsey. The potential of the trammel net as an alternative method for sampling fish on deep water reefs. Fisheries Research, 39(1):67–74, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001714.

Lucas:1999:UFB

[LMA^{+}99] Martyn C. Lucas, Tom Mercer, John D. Armstrong, Shaun McGinty, and Phillip Rycroft. Use of a flat-bed passive integrated transponder antenna array to study the migration and behaviour of lowland river fishes at a fish pass. Fisheries Research, 44(2):183–191, December 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000612.

Levy:1998:CGS

[LMC98] J. A. Levy, R. Maggioni, and M. B. Conceição. Close genetic similarity among populations of the white croaker (Micropogonias furnieri) in the south and south-eastern Brazilian coast. I. Allozyme studies. Fisheries Research, 39(1):87–94, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001660.

Losanes:1992:EEE

[LMF92a] Loel P. Losanes, Ko Matuda, and Yasuzumi Fujimori. Estimating the entangling effect of trammel and semi-trammel net selectivity on rainbow trout (Oncorhynchus mykiss). Fisheries Research, 15(3):229–242, December 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290017N.

Losanes:1992:OTE

[LMF92b] Loel P. Losanes, Ko Matuda, and Yasuzumi Fujimori. Outdoor tank experiments on the influence of soaking time on the catch efficiency of gillnets and entangling nets. Fisheries Research, 15(3):217–227, December 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290016M.

Losanes:1992:CES

[LMMK92] Loel P. Losanes, Ko Matuda, Tadanobu Machii, and Atsushi Koike. Catching efficiency and selectivity of entangling nets. Fisheries Research, 13(1):9–23, January 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290030W.

Leung:1998:EFM

[LMNP98] PingSun Leung, Jill Muraoka, Stuart T. Nakamoto, and Sam Pooley. Evaluating fisheries management options in Hawaii using analytic hierarchy process (AHP). Fisheries Research, 36(2–3):171–183, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698000976.

Labropoulou:1997:FHO

[LMTE97] M. Labropoulou, A. Machias, N. Tsimenides, and A. Eleftheriou. Feeding habits and ontogenetic diet shift of the striped red mullet, Mullus surmuletus Linnaeus, 1758. Fisheries Research, 31(3):257–267, August 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000179.

Lockwood:1993:BRA

[Loc93] Stephen J. Lockwood. Book review: Artificial habitats for marine and freshwater fisheries: W. Seaman, Jr., and L. M. Sprague (Editors), Academic Press, San Diego, 1991, 285 pp. US $39.95, hardback, ISBN 0-12-634345-4. Fisheries Research, 16(1):89–92, February 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390113L.

Locke:1997:PDC

[Loc97a] Andrea Locke. Precision of diving and canoe-based visual estimates of Atlantic salmon (Salmo salar) abundance. Fisheries Research, 29(3):283–287, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696005486.

Lockwood:1997:BRS

[Loc97b] Stephen J. Lockwood. Book review: Survival strategies in early life stages of marine resources: Y. Watanabe, Y. Yamashita and Y. Oozeki (Editors), A. A. Balkema, Rotterdam, 1996, 367 pp., £66, ISBN 90-5410-637-9. Fisheries Research, 30(1–2): 163–165, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005589.

Lokkeborg:1990:RRP

[Løk90] Svein Løkkeborg. Rate of release of potential feeding attractants from natural and artificial bait. Fisheries Research, 8(3):253–261, January 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090026R.

Lokkeborg:1991:FEA

[Løk91] Svein Løkkeborg. Fishing experiments with an alternative longline bait using surplus fish products. Fisheries Research, 12(1):43–56, September 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190048K.

Longhurst:1998:CPI

[Lon98] Alan Longhurst. Cod: perhaps if we all stood back a bit? Fisheries Research, 38(2):101–108, October 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001520.

Longhurst:1999:DBP

[Lon99a] Alan Longhurst. Does the benthic paradox tell us something about surplus production models? Fisheries Research, 41(2):111–117, June 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783699000168.

Longhurst:1999:DVT

[Lon99b] Alan Longhurst. Double vision: two views of the North Atlantic fishery. Fisheries Research, 40(1):1–5, February 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698002252.

Liu:1995:ACO

[LP95a] Qun Liu and Tony J. Pitcher. Application of a constrained optimisation method to the statistical analysis of fisheries catch-at-age data. Fisheries Research, 22(1–2):29–41, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400312K.

Lorenzo:1995:BPR

[LP95b] J. M. Lorenzo and J. G. Pajuelo. Biological parameters of the roudi escolar Promethichthys prometheus (Pisces: Gempylidae) off the Canary Islands. Fisheries Research, 24(1):65– 71, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694003654.

Landa:1997:ASF

[LP97a] J. Landa and N. Pérez. Ageing ‘on sight’ of four spot megrim, Lepidorhombus boscii (Risso) otoliths and their validation. Fisheries Research, 30(1–2):27–31, May 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000039.

Lokkeborg:1997:EST

[LP97b] Svein Løkkeborg and Teresa Pina. Effects of setting time, setting direction and soak time on longline catch rates. Fisheries Research, 32(3):213–222, December 1, 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000702.

Lorenzo:1999:AGS

[LP99] J. M. Lorenzo and J. G. Pajuelo. Age and growth of the sand smelt Atherina (Hepsetia) presbyter Cuvier, 1829 in the Canary Islands (Central-east Atlantic). Fisheries Research, 41(2):177–182, June 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000053.

Lehtonen:1998:ACM

[LPH^{+}98] Hannu Lehtonen, Heikki Peltonen, Outi Heikinheimo, Erkki Saarijärvi, Kari Saulamo, Mika Vinni, and Tauno Nurmio. Application of coded microtags to study growth rates of stocked sympatric whitefish (Coregonus lavaretus (L.)) forms. Fisheries Research, 39(1):9–15, December 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001751.

Landa:1996:MLW

[LPP96] J. Landa, C. Piñeiro, and N. Pérez. Megrim (Lepidorhombus whiffiagonis) growth patterns in the northeast Atlantic. Fisheries Research, 26(3–4):279–294, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369500419X.

Laloe:1998:MIF

[LPSS98] F. Laloë, N. Pech, R. Sabatier, and A. Samba. Model identification for flexible multifleet–multispecies fisheries: a simulation study. Fisheries Research, 37(1–3):193–202, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001362.

Lipinski:1990:MEB

[LR90] M. R. Lipinski and M. A. Roeleveld. Minor extension of the von Bertalanffy growth theory. Fisheries Research, 9(4):367– 371, October 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090054Y.

Lowry:1996:ETT

[LR96] N. Lowry and J. H. B. Robertson. The effect of twine thickness on cod-end selectivity of trawls for haddock in the North Sea. Fisheries Research, 26(3–4):353–363, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004181.

Laine:1998:CRP

[LR98] Päivi Laine and Marjut Rajasilta. Changes in the reproductive properties of Baltic herring females during the spawning season. Fisheries Research, 36(1):67–73, June 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000873.

Li:1993:MER

[LS93] Xinyang Li and Patrick Safran. A method to estimate the S–R relationship using rough information on the rate of exploitation. Fisheries Research, 17(3–4):293–309, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390131P.

Leonardos:1998:RSA

[LS98] Ioannis Leonardos and Apostolos Sinis. Reproductive strategy of Aphanius fasciatus Nardo, 1827 (Pisces: Cyprinodontidae) in the Mesolongi and Etolikon lagoons (W. Greece). Fisheries Research, 35(3):171–181, May 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000824.

Leonardos:1999:PAS

[LS99] Ioannis Leonardos and Apostolos Sinis. Population age and sex structure of Aphanius fasciatus Nardo, 1827 (Pisces: Cyprinodontidae) in the Mesolongi and Etolikon lagoons (W. Greece). Fisheries Research, 40(3):227–235, April 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002318.

Lessa:1999:AGS

[LSP99] Rosangela Lessa, Francisco Marcante Santana, and Renato Paglerani. Age, growth and stock structure of the oceanic whitetip shark, Carcharhinus longimanus, from the southwestern equatorial Atlantic. Fisheries Research, 42(1–2):21–30, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000454.

Leslie:1992:BSC

[LT92] J. K. Leslie and C. A. Timmins. Beach seine collections of freshwater larval fish at the shore of lakes. Fisheries Research, 15(3):243–251, December 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290018O.

Lehtonen:1998:IMS

[LTS98] Esa Lehtonen, Vesa Tschernij, and Petri Suuronen. An improved method for studying survival of fish that escape through meshes of trawl codends. Fisheries Research, 38(3):303–306, November 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001635.

Lok:1997:EDC

[LTT^{+}97] A. Lök, A. Tokaç, Z. Tosunoğlu, C. Metin, and R. S. T. Ferro. The effects of different cod-end design on bottom trawl selectivity in Turkish fisheries of the Aegean Sea. Fisheries Research, 32(2):149–156, November 1, 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000489.

Lucas:1990:BRP

[Luc90] M. C. Lucas. Book review: Physiology of elasmobranch fishes: Trevor J. Shuttleworth (Editor), Springer-Verlag, Berlin, 1988, 324 pp., DM 228.00, ISBN 0-387-18769-3. Fisheries Research, 8(3):297–299, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090033R.

Lucas:1992:BRC

[Luc92] Martyn C. Lucas. Book review: Cyprinid Fishes: Systematics, Biology and Exploitation (Fish and Fisheries Series: 3): Ian J. Winfield and Joseph S. Nelson (editors), Chapman and Hall, London, 1991, 667 pp., £60.00, ISBN 0-412-34920-5. Fisheries Research, 13(4):431–434, April 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783692900652.

Lucas:1993:BRD

[Luc93] Martyn C. Lucas. Book review: Dynamics and exploitation of unstable percid populations: A. D. Buijse, University of Wageningen, Wageningen, 1992, 167 pp., ISBN 90-5485-001-9. Fisheries Research, 16(3):284–285, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390103E.

Lucas:1995:BRR

[Luc95] Martyn C. Lucas. Book review: Rehabilitation of freshwater fisheries: I. G. Cowx (editor), Blackwell Scientific Publications, Oxford, 1994, 486 pp., £55, ISBN 0-85238-195-6. Fisheries Research, 22(3–4):321–322, March 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369590025X.

Lucas:1996:BRP

[Luc96] Martyn C. Lucas. Book review: Pike: Biology and exploitation: John F. Craig (editor). Chapman and Hall, London, 1996, 298 pp., £50, ISBN 0-412-42960-8. Fisheries Research, 28(2):240–242, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696824945.

Lucas:1997:BRC

[Luc97a] Martyn C. Lucas. Book review: Conservation of endangered freshwater fish in Europe: A. Kirchhofer and D. Hefti (Editors), Birkhäuser, Basel, 1996, 341 pp., DM 94, ISBN 3-7643-5321-X. Fisheries Research, 30(1–2):162–163, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005577.

Lucas:1997:BRI

[Luc97b] Martyn C. Lucas. Book review: The impact of species changes in African Lakes: Edited by Tony J. Pitcher and Paul J. B. Hart, Chapman and Hall, London, 1995, 601 pp., £85.00, ISBN 0-412-55050-4. Fisheries Research, 32(1):105–106, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369700026X.

Landergren:1998:SST

[LV98] Peter Landergren and Lars Vallin. Spawning of sea trout, Salmo trutta L., in brackish waters — lost effort or successful strategy? Fisheries Research, 35(3):229–236, May 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698000733.

Lakshmanan:1990:QCS

[LVIG90] P. T. Lakshmanan, P. R. G. Varma, T. S. G. Iyer, and K. Gopakumar. Quality changes in seafrozen whole and filleted rock cod (Epinephelus spp.) during storage. Fisheries Research, 9(1):1–12, July 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090036U.

Lyons:1998:HAF

[Lyo98] Jim Lyons. A hydroacoustic assessment of fish stocks in the River Trent, England. Fisheries Research, 35(1–2):83– 90, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698000629.

Matsushita:1997:ITL

[MA97] Yoshiki Matsushita and Rosidi Ali. Investigation of trawl landings for the purpose of reducing the capture of non-target species and sizes of fish. Fisheries Research, 29(2):133–143, February 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005346.

MacKenzie:1998:PBT

[MA98] K. MacKenzie and P. Abaunza. Parasites as biological tags for stock discrimination of marine fish: a guide to procedures and methods. Fisheries Research, 38(1):45–56, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001167.

MacLennan:1992:FGS

[Mac92] David N. MacLennan. Fishing gear selectivity: an overview. Fisheries Research, 13(3):201–204, March 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783692900766.

Mackie:1993:BRF

[Mac93] A. M. Mackie. Book review: Fish chemoreception: Toshiaki J. Hara, Chapman and Hall, London, 1992, 373 pp., £55.00, ISBN 0-412-35140-4. Fisheries Research, 16(3):278– 279, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390100L.

Maceina:1997:SAU

[Mac97] M. J. Maceina. Simple application of using residuals from catch-curve regressions to assess year-class strength in fish. Fisheries Research, 32(2):115–121, November 1, 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000519.

Mathisen:1992:HTF

[Mat92] Ole A. Mathisen. Hydroacoustics as a tool in fisheries research — “where are we and what are the future expectations?”. Fisheries Research, 14(2–3):91–93, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290045U.

Misund:1995:SHA

[MB95] Ole Arve Misund and Arvid K. Beltestad. Survival of herring after simulated net bursts and conventional storage in net pens. Fisheries Research, 22(3–4):293–297, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400326R.

Montgomery:1996:TER

[MB96] S. S. Montgomery and P. A. Brett. Tagging eastern rock lobsters Jasus verreauxi: effectiveness of several types of tag. Fisheries Research, 27(4):141–152, July 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696004833.

Mounsey:1995:DTE

[MBB95] Richard P. Mounsey, Graham A. Baulch, and Rik C. Buckworth. Development of a trawl efficiency device (TED) for Australian prawn fisheries. I. The AusTED design. Fisheries Research, 22(1–2):99–105, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400307I.

Maartens:1999:GML

[MBH99] Lima Maartens, Anthony J. Booth, and Thomas Hecht. The growth of monkfish Lophius vomerinus in Namibian waters, with a comparison of otolith and illicia methods of ageing. Fisheries Research, 44(2):139–148, December 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000788.

Morizur:1999:ICM

[MBT^{+}99] Y. Morizur, S. D. Berrow, N. J. C. Tregenza, A. S. Couperus, and S. Pouvreau. Incidental catches of marinemammals in pelagic trawl fisheries of the northeast Atlantic. Fisheries Research, 41(3):297–307, July 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000132.

McAndrew:1993:BRG

[McA93] Brendan McAndrew. Book review: Genetics and fish breeding: Colin E. Purdom, Chapman & Hall, Fish & Fisheries Series 8, London, 1992, 227 pp., price £39.95, ISBN 0-412-33040-7. Fisheries Research, 17(3–4):379–380, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390138W.

McClain:1990:BRF

[McC90a] W. R. McClain. Book review: Freshwater crayfish: Biology, management and exploitation: D. M. Holdich and R. S. Lowery (Editors), Croom Helm, London, 1988, 498 pp., £49.50, ISBN 0-88192-105. Fisheries Research, 8(3):296–297, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090032Q.

Michael:1990:PSC

[MCC90b] E. Michael, P. Chadwick, and Ross R. Claytor. Predictability in a small commercial Atlantic salmon fishery in western Newfoundland. Fisheries Research, 10(1–2):15–27, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090013L.

McIntyre:1990:BRT

[McI90] Alasdair D. McIntyre. Book review: Tuna and billfish — Fish without a country: James Joseph, Witold Klawe and Pat Murphy, paintings by George Mattson, Inter-American Tropical Tuna Commission, La Jolla, California, 4th edn., 1988, xi + 69 pp., US$15.75 postage paid from P. O. Box 271, La Jolla, CA 92038, ISBN 0-9603078-2-6. Fisheries Research, 8(3):295– 296, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090031P.

McIntyre:1991:BRC

[McI91a] Alasdair D. McIntyre. Book review: Coastal oceanography of Washington and Oregon: Michael R. Landry and Barbara M. Hickey (Editors). Elsevier, Amsterdam, 1989, 607 pp., US $163.25, ISBN 0-444-87308-2. Fisheries Research, 10(3–4):358– 360, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190089X.

McIntyre:1991:BRM

[McI91b] Alasdair D. McIntyre. Book review: Managing marine environments: Richard A. Kenchington, Taylor and Francis, New York, 1990, 248 pp., £35.00, ISBN 0-8448-1635-3. Fisheries Research, 12(2):181–182, October 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190041D.

McIntyre:1991:BRS

[McI91c] Alasdair D. McIntyre. Book review: Sessile animals of the seashore: Vernon A. Harris. Chapman and Hall, London, 1990, 379 pp., £37.50, ISBN 0-412-33760-6. Fisheries Research, 10(3–4):355–356, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190087V.

McIntyre:1991:BRP

[McI91d] Alasdair D. McIntyre. Book review: The provident sea: D. H. Cushing. Cambridge University Press, Cambridge, 1988, 329 pp, £37.50, ISBN 0-521-25727-1. Fisheries Research, 10(3–4):356–358, January 1991. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190088W.

Mcintyre:1992:BRC

[Mci92] Alasdair D. Mcintyre. Book review: Crisis in the World’s Fisheries: People, problems and policies: James R. McGoodwin, Stanford University Press, CA, 1991, 235 pp., US$35, ISBN 0-8047-1790-7. Fisheries Research, 13(4):429–430, April 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290063Y.

McIntyre:1993:BRM

[McI93] A. D. McIntyre. Book review: Methods for fish biology: Carl B. Schreck and Peter B. Moyle (editors), American Fisheries Society, Bethesda, MD, 1990, Cloth, 684 pp., US$50, ISBN 0-913235-58-X. Fisheries Research, 16(3):277–278, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390099S.

McIntyre:1994:BRN

[McI94] A. D. McIntyre. Book review: No day too long — An hydrographer’s tale: Rear Admiral G. S. Ritchie, Pentland Press, Durham, 1992, 250 pp., ISBN 1-872795-63-3. Fisheries Research, 20(1):93–95, June 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900094.

McIntyre:1995:BRE

[McI95a] A. D. McIntyre. Book review: Environment and aquaculture in developing countries: R. S. V. Pullin, H. Rosenthal and J. L. Maclean. International Centre for Living Aquatic Resources Management, Manila, and Deutsche Gesellschaft für Technische Zusammenarbeit, Eschborn, Germany, 1993, 360 pp., US$15, ISBN 971-8709-05-3. Fisheries Research, 23(1–2): 189–190, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695900020.

McIntyre:1995:E

[McI95b] A. D. McIntyre. Editorial. Fisheries Research, 24(1):1–2, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695900098.

Mcintyre:1995:BRG

[Mci95c] Alasdair D. Mcintyre. Book review: Global marine biological diversity — a strategy for building conservation into decision making: Elliott A. Norse (editor), Island Press, Washington, DC, 1993, 383 pp., Cloth US $50.00, ISBN 1-55963-255-0, Paper US $27.50, ISBN 1-55963-256-9. Fisheries Research, 22(3–4):319–321, March 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695900241.

McIntyre:1995:BRSb

[McI95d] Alasdair D. McIntyre. Book review: Scaling fisheries: Tim D. Smith, Cambridge University Press, 1994, 392 pp., £50, ISBN 0-521-39032-x, hardback. Fisheries Research, 24(1):80– 82, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369590011X.

McIntyre:1995:BRI

[McI95e] Alasdair D. McIntyre. Book review: The Icelandic fisheries: Evolution and management of a fishing industry: Ragnar Arnason. Fishing News Books, Blackwell Science, Oxford, 1995, 224 pp., £45 hard cover, ISBN 0 85238 210 3. Fisheries Research, 24(3):267–268, October 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695900187.

McIntyre:1995:BRSa

[McI95f] Alasdair D. McIntyre. Book review: The State of the world’s fishery resources: Clyde W. Voigtlander (Editor). International Science Publishers, USA. 1994, 204 pp., £85, hard cover, ISBN 1-881570-38-X. Fisheries Research, 24(1):79– 80, July 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695900101.

McIntyre:1996:O

[McI96a] A. D. McIntyre. Obituary. Fisheries Research, 28(1):1, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783696894434.

Mcintyre:1996:BRF

[Mci96b] Alasdair D. Mcintyre. Book review: Fisheries management in the European Union: Apostolos Karagiannakos, Avebury, Ashgate Publishing Company, UK, 1995, 286 pp., £39.95, ISBN 1 85972 070 6, hardback. Fisheries Research, 28(2):237–239, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783696004924.

McIntyre:1996:BRS

[McI96c] Alasdair D. McIntyre. Book review: On the sex of fish and the gender of scientists: Daniel Pauly. Chapman and Hall, 1994 250 pp., £35, ISBN 0-412-59540-0, paperback. Fisheries Research, 26(3–4):385–386, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696900041.

McIntyre:1996:BRU

[McI96d] Alasdair D. McIntyre. Book review: Understanding marine biodiversity: Report of the Committee on Biological Diversity in Marine Systems. National Academy Press, Oxford, 114 pp., hardback £24.95, ISBN 0-309-05225-4. Fisheries Research, 27(1–3):137–139, June 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696846741.

McIntyre:1996:FRA

[McI96e] Alasdair D. McIntyre. Fisheries research: after twenty-five volumes. Fisheries Research, 25(1):1–3, January 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696900193.

McIntyre:1997:BRA

[McI97a] Alasdair D. McIntyre. Book review: Aquatic predators and their prey: Edited by S. P. R. Greenstreet and M. L. Tasker, available from Fishing News Books, Blackwell Science Ltd., Oxford, 1996, 191 pp., price £69.50 (hardback), ISBN 0-85238-230-8. Fisheries Research, 32(1):101–102, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000258.

McIntyre:1997:BRFa

[McI97b] Alasdair D. McIntyre. Book review: Fisheries management in crisis: Keven Crean and David Symes (Editors), Fishing News Books/Blackwell Scientific, Oxford, 1996, 20 ill., 288 pp., £45 (hbk), ISBN 0-85238-231-6. Fisheries Research, 30(1–2):161– 162, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005565.

Mcintyre:1997:BRF

[Mci97c] Alasdaird D. Mcintyre. Book review: Fishing for truth: a sociological analyses of northern cod stock assessments from 1977 to 1990: Alan Christopher Finlayson, Institute of Social and Economic Research, Memorial University of Newfoundland, St. John’s, Social and Economic Studies No. 52 1994. 176 pages. Can $24.95., ISBN 0-919666-79-5, paperback. Fisheries Research, 29(2):199–200, February 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005334.

McIntyre:1997:PAF

[McI97d] Mr A. D. McIntyre. The precautionary approach to fisheries: a reply to Prof. Kesteven [Fisheries Research 28 (1996) 109–110]. Fisheries Research, 31(3):275, August 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697900096. See [Kes96e].

McShane:1994:EAA

[McS94] Paul E. McShane. Estimating the abundance of abalone (Haliotis spp.) stocks — examples from Victoria and southern New Zealand. Fisheries Research, 19(3–4):379–394, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900515.

McShane:1996:PDE

[McS96] Paul E. McShane. Patch dynamics and effects of exploitation on abalone (Haliotis iris) populations. Fisheries Research, 25(2):191–199, February 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695004165.

McVey:1993:BRC

[McV93] J. P. McVey. Book review: Channel catfish farming handbook: Craig S. Tucker and Edwin H. Robinson, Van Nostrand Reinhold, New York, 1991, ISBN 0-442-31836-7. Fisheries Research, 15(4):387–388, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390089P.

Misund:1992:OPS

[MDB92] O. A. Misund, W. Dickson, and A. K. Beltestad. Optimization of purse seines by large-meshed sections and low lead weight. Theoretical considerations, sinking speed measurements and fishing trials. Fisheries Research, 14(4):305–317, September 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290039V.

Moreno:1994:PBV

[MdCP94] Ana Moreno, Manuela Morais da Cunha, and João M. F. Pereira. Population biology of veined squid (Loligo forbesi) and European squid (Loligo vulgaris) from the Portuguese coast. Fisheries Research, 21(1–2):71–86, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900965.

McArdle:1993:FPF

[MDMG^{+}93] John F. McArdle, Caroline Dooley-Martyn, Fiona Geoghegan, Frank McKiernan, and Hamish Rodger. Furunculosis as a possible factor in the decline of sea trout in the West of Ireland. Fisheries Research, 17(1–2):201–207, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783693900194.

Misra:1999:NNM

[ME99] R. K. Misra and M. D. L. Easton. A note on the number of morphometric characters used in fish stock delineation studies employing a MANOVA. Fisheries Research, 42(1–2):191– 194, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000466.

Medley:1998:DTC

[Med98] P. A. H. Medley. A decision theory case study: Choosing a season opening for a spiny lobster (Panulirus argus L.) fishery. Fisheries Research, 36(2–3):159–170, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001015.

Megrey:1991:PDM

[Meg91] Bernard A. Megrey. Population dynamics and management of walleye pollock (Theragra chalcogramma) in the Gulf of Alaska, 1976–1986. Fisheries Research, 11(3–4):321–354, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783691900084.

Marquez-Farias:1998:FBD

[MFCG98] J. Fernando Márquez-Farias and J. Leonardo Castillo-Geniz. Fishery biology and demography of the Atlantic sharpnose shark, Rhizoprionodon terraenovae, in the southern Gulf of Mexico. Fisheries Research, 39(2):183–198, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001829.

Millar:1992:TSP

[MG92] H. H. Millar and E. A. Gunn. A two-stage procedure for planning marketing and fishing activities in fish-processing firms. Fisheries Research, 15(3):197–215, December 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290015L.

Martinez-Garmendia:1997:ELA

[MG97] Josué Martínez-Garmendia. Effects of length-at-age data on growth and management benchmark F_{0.1} estimates in the face of size-selective mortality. Fisheries Research, 32(3):233–247, December 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369700057X.

Maack:1999:CRB

[MG99] Gerd Maack and Michael R. George. Contributions to the reproductive biology of Encrasicholina punctifer Fowler, 1938 (Engraulidae) from West Sumatra, Indonesia. Fisheries Research, 44(2):113–120, December 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000843.

McHugh:1990:FMN

[MH90] J. L. McHugh and Emerson Hasbrouck. Fishery management in New York bight: Experience under the Magnuson Act. Fisheries Research, 8(3):205–221, January 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090022N.

Martins:1997:DAB

[MH97] Agnaldo Silva Martins and Manuel Haimovici. Distribution, abundance and biological interactions of the cutlassfish Trichiurus lepturus in the southern Brazil subtropical convergence ecosystem. Fisheries Research, 30(3):217–227, June 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005668.

Madsen:1999:SSS

[MHWMP99] N. Madsen, R. Holst, D. Wileman, and T. Moth-Poulsen. Size selectivity of sole gill nets fished in the North Sea. Fisheries Research, 44(1):59–73, November 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578369900048X.

Miller:1991:BRF

[Mil91] P. J. Miller. Book review: Fishes of the sea of Japan and the adjacent areas of the sea of Okhotsk and the yellow sea: Part 4, Teleostomi XXIX. Perciformes/2. Blennioidei-13. Gobioidei/ (CXLV. Fam. Anarhichadidae-CLXXV. Fam. Periophthalmidae), by G. U. Lindberg and Z. V. Krasyukova (Bruce B. Collette, Scientific Editor). A. A. Balkema, Rotterdam, 1989, 628 pp., price Dfl. 95.00/US $48.50/£26.00, ISBN 90-6191-415-9. Fisheries Research, 10(3–4):360–362, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783691900903.

Millar:1993:MAS

[Mil93a] Harvey H. Millar. A model for allocating stock quotas in the Canadian Atlantic groundfish fishery. Fisheries Research, 18(3–4):241–258, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693901562.

Millar:1993:ATS

[Mil93b] Russell B. Millar. Analysis of trawl selectivity studies (addendum): implementation in SAS. Fisheries Research, 17(3–4):373–377, August 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390137V.

Millar:1995:PAA

[Mil95] Harvey H. Millar. Planning annual allocation of fisheries surveillance effort. Fisheries Research, 23(3–4):345–360, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400344V.

Millar:1996:PFS

[Mil96] Harvey H. Millar. Planning fish scouting activity in industrial fishing. Fisheries Research, 25(1):63–75, January 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695003991.

Mills:1997:BRM

[Mil97] Derek Mills. Book review: Mountain in the clouds, a search for the wild salmon: Bruce Brown, University of Washington Press, 1995, 249 pp., paperback, ISBN 0-295-97475-3. Fisheries Research, 29(2):200–201, February 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005310.

Millan:1999:RCC

[Mil99] M. Millán. Reproductive characteristics and condition status of anchovy Engraulis encrasicolus L. from the Bay of Cadiz (SW Spain). Fisheries Research, 41(1):73–86, May 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000107.

Misund:1992:PSB

[Mis92] Ole Arve Misund. Predictable swimming behaviour of schools in purse seine capture situations. Fisheries Research, 14(4):319–328, September 1992. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290040Z.

Misund:1993:ABH

[Mis93] Ole Arve Misund. Avoidance behaviour of herring (Clupea harengus) and mackerel (Scomber scombrus) in purse seine capture situations. Fisheries Research, 16(2):179–194, March 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783693900518.

Mendo:1993:LBG

[MJ93] J. Mendo and E. Jurado. Length-based growth parameter estimates of the Peruvian scallop (Argopecten purpuratus). Fisheries Research, 15(4):357–367, January 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390086M.

McGlennon:1997:EBR

[MK97] David McGlennon and Martine A. Kinloch. Evaluation of the bus-route creel survey method in a large Australian marine recreational fishery II. Pilot surveys and optimal sampling allocation. Fisheries Research, 33(1–3):89–99, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000672.

Minos:1995:CCM

[MKKO95] G. Minos, G. Katselis, P. Kaspiris, and I. Ondrias. Comparison of the change in morphological pattern during the growth in length of the grey mullets Liza ramada and Liza saliens from western Greece. Fisheries Research, 23(1–2):143– 155, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400334S.

Machiels:1994:ESL

[MKLvD94] M. A. M. Machiels, M. Klinge, R. Lanters, and W. L. T. van Densen. Effect of snood length and hanging ratio on efficiency and selectivity of bottom-set gillnets for pikeperch, Stizostedion lucioperca L., and bream, Abramis brama. Fisheries Research, 19(3–4):231–239, April 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900418.

Mous:1999:TBD

[MKS99] P. J. Mous, J. Kemper, and A. Schelvis. A towed body designed for side-scanning hydroacoustic surveying of fish stocks in shallow waters. Fisheries Research, 40(1):97–98, February 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698002136.

Malmberg:1993:SGP

[MM93] G. Malmberg and M. Malmberg. Species of Gyrodactylus (Platyhelminthes, Monogenea) on salmonids in Sweden. Fisheries Research, 17(1–2):59–68, June 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390007T.

Mancera:1996:PDO

[MM96] Ernesto Mancera and Jaime Mendo. Population dynamics of the oyster Crassostrea rhizophorae from the Ciénaga Grande de Santa Marta, Colombia. Fisheries Research, 26(1–2):139– 148, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004017.

Myers:1998:RUB

[MM98] Ransom A. Myers and Gordon Mertz. Reducing uncertainty in the biological basis of fisheries management by meta-analysis of data from many populations: a synthesis. Fisheries Research, 37(1–3):51–60, August 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369800126X.

Muoneke:1993:CCE

[MMH93] Maurice I. Muoneke, O. Eugene Maughan, and Christopher C. Henry. Comparative capture efficiencies of frame and hoop nets for white crappie (Pomoxis annularis Rafinesque). Fisheries Research, 18(3–4):231–240, December 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390155Z.

McGilvray:1999:AII

[MMM99] J. G. McGilvray, R. P. Mounsey, and J. MacCartie. The AusTED II, an improved trawl efficiency device 1. Design theories. Fisheries Research, 40(1):17–27, February 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002215.

Massuti:1997:RBD

[MMN97] Enric Massutí and B. Morales-Nin. Reproductive biology of dolphin-fish (Coryphaena hippurus L.) off the island of Majorca (western Mediterranean). Fisheries Research, 30(1–2): 57–65, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005620.

Madsen:1999:SEE

[MMPHW99] N. Madsen, T. Moth-Poulsen, R. Holst, and D. Wileman. Selectivity experiments with escape windows in the North Sea Nephrops (Nephrops norvegicus) trawl fishery. Fisheries Research, 42(1–2):167–181, August 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000247.

Madsen:1998:SEW

[MMPL98] N. Madsen, T. Moth-Poulsen, and N. Lowry. Selectivity experiments with window codends fished in the Baltic Sea cod (Gadus morhua) fishery. Fisheries Research, 36(1):1–14, June 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369800085X.

Mianzan:1996:FPN

[MMPS96] H. W. Mianzan, N. Mari, B. Prenski, and F. Sanchez. Fish predation on neritic ctenophores from the Argentine continental shelf: a neglected food resource? Fisheries Research, 27(1–3): 69–79, June 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004599.

Mitchell:1998:DNA

[MMV98] A. Mitchell, E. McCarthy, and E. Verspoor. Discrimination of the North Atlantic lesser sandeels Ammodytes marinus, A. tobianus, A. dubius and Gymnammodytes semisquamatus by mitochondrial DNA restriction fragment patterns. Fisheries Research, 36(1):61–65, June 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698000812.

Machii:1992:MPR

[MN92] Tadanobu Machii and Yukio Nose. Mechanical properties of a rectangular purse seine in ideal conditions. Fisheries Research, 14(4):261–271, September 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290036S.

Morales-Nin:1997:GJM

[MNA97] Beatriz Morales-Nin and Yvonne Aldebert. Growth of juvenile Merluccius merluccius in the Gulf of Lions (NW Mediterranean) based on otolith microstructure and length-frequency analysis. Fisheries Research, 30(1–2):77–85, May 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369600553X.

Muthukude:1991:GPE

[MNJ91] Palitha Muthukude, James L. Novak, and Curtis Jolly. A goal programming evaluation of fisheries development plans for Sri Lanka’s coastal fishing fleet, 1988–1991. Fisheries Research, 12(4):325–339, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783691900169.

Morales-Nin:1997:LHF

[MNM97] Beatriz Morales-Nin and Joan Moranta. Life history and fishery of the common dentex (Dentex dentex) in Mallorca (Balearic Islands, western Mediterranean). Fisheries Research, 30(1–2):67–76, May 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696005607.

Morales-Nin:1996:AGB

[MNSC96] Beatriz Morales-Nin and Dalila Sena-Carvalho. Age and growth of the black scabbard fish (Aphanopus carbo) off Madeira. Fisheries Research, 25(3–4):239–251, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004327.

Montgomery:1990:MJE

[Mon90] S. S. Montgomery. Movements of juvenile eastern king prawns, Penaeus plebejus, and identification of stock along the east coast of Australia. Fisheries Research, 9(3):189–208, September 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605800013.

Moring:1993:RLR

[Mor93] John R. Moring. Records of long-range, downstream movements of stocked rainbow trout (Oncorhynchus mykiss). Fisheries Research, 16(2):195–199, March 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693900529.

Malinen:1996:OST

[MP96a]  Tommi Malinen and Heikki Peltonen.  Optimal sam-

pling and traditional versus model-based data analysis in acoustic fish stock assessment in Lake Vesijärvi. Fisheries Research, 26(3–4):295–308, May 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695004238.

Mattei:1996:PST

[MP96b] N. Mattei and M. Pellizzato. A population study on three stocks of a commercial Adriatic pectinid (Pecten jacobaeus). Fisheries Research, 26(1–2):49–65, April 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004130.

Manooch:1997:AGM

[MP97] Charles S. Manooch and Jennifer C. Potts. Age, growth and mortality of greater amberjack from the southeastern United States. Fisheries Research, 30(3):229–240, June 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005541.

Matthews:1997:JRS

[MPDW97] M. A. Matthews, W. R. Poole, M. G. Dillane, and K. F. Whelan. Juvenile recruitment and smolt output of brown trout (Salmo trutta L.) and Atlantic salmon (Salmo salar L.) from a lacustrine system in western Ireland. Fisheries Research, 31(1– 2):19–37, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000131.

Mianzan:1997:SPV

[MPMC97] H. W. Mianzan, M. Pájaro, L. Machinandiarena, and F. Cremonte. Salps: possible vectors of toxic dinoflagellates? Fisheries Research, 29(2):193–197, February 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005267.

Manooch:1998:PAR

[MPVB98] Charles S. Manooch III, Jennifer C. Potts, Douglas S. Vaughan, and Michael L. Burton. Population assessment of the red snapper from the southeastern United States. Fisheries Research, 38(1):19–32, September 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369800112X.

Margraf:1993:ESS

[MR93] F. J. Margraf and L. M. Riley. Evaluation of scale shape for identifying spawning stocks of coastal Atlantic striped bass (Morone saxatilis). Fisheries Research, 18(3–4):163–172, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783693901506.

Miller:1996:STF

[MR96] R. J. Miller and R. S. Rodger. Soak times and fishing strategy for American lobster. Fisheries Research, 26(3–4):199– 205, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004459.

Mees:1997:PYL

[MR97] C. C. Mees and J. A. Rousseau. The potential yield of the lutjanid fish Pristipomoides filamentosus from the Mahé Plateau, Seychelles: managing with uncertainty. Fisheries Research, 33(1–3):73–87, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000696.

Milton:1996:RPF

[MRB96] David A. Milton, Nicholas J. F. Rawlinson, and Stephen J. M. Blaber. Recruitment patterns and factors affecting recruitment of five species of short-lived clupeoids in the tropical South Pacific. Fisheries Research, 26(3–4):239–255, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004343.

McKenna:1991:SAD

[MS91] James E. McKenna and Saul B. Saila. Shifts in the Antarctic demersal fish community of South Georgia Island. Fisheries Research, 12(2):109–124, October 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190036F.

Mytilineou:1995:AGN

[MS95] Ch. Mytilineou and F. Sardá. Age and growth of Nephrops norvegicus in the Catalan Sea, using length-frequency analysis. Fisheries Research, 23(3–4):283–299, June 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694003506.

Maynou:1997:NNP

[MS97] Francesc Maynou and Francesc Sardà. Nephrops norvegicus population and morphometrical characteristics in relation to substrate heterogeneity. Fisheries Research, 30(1–2):139– 149, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005498.

McKinnell:1998:SBJ

[MS98] Skip McKinnell and Michael P. Seki. Shark bycatch in the Japanese high seas squid driftnet fishery in the North Pacific Ocean. Fisheries Research, 39(2):127–138, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001799.

Mackinson:1997:BCF

[MSP97] Steven Mackinson, Ussif Rashid Sumaila, and Tony J. Pitcher. Bioeconomics and catchability: fish and fishers behaviour during stock collapse. Fisheries Research, 31(1–2):11–17, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000209.

Martin:1995:PSE

[MSR95] Paloma Martín, Pilar Sánchez, and Montserrat Ramón. Population structure and exploitation of Bolinus brandaris (Mollusca: Gastropoda) off the Catalan coast (northwestern Mediterranean). Fisheries Research, 23(3–4):319–331, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400346X.

McVicar:1993:DWS

[MSWP93] A. H. McVicar, L. A. Sharp, A. F. Walker, and A. W. Pike. Diseases of wild sea trout in Scotland in relation to fish population decline. Fisheries Research, 17(1–2):175–185, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783693900172.

Mtsambiwa:1996:ELN

[Mts96] Morris Zororai Mtsambiwa. Evaluation of lift-net selectivity in the freshwater sardine Limnothrissa miodon fishery on Lake Kariba, Zimbabwe. Fisheries Research, 26(3–4):373– 379, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004556.

Munro:1990:SF

[Mun90] A. L. S. Munro. Salmon farming. Fisheries Research, 10 (1–2):151–161, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578369090020V.

Millar:1992:ATS

[MW92] R. B. Millar and S. J. Walsh. Analysis of trawl selectivity studies with an application to trouser trawls. Fisheries Research, 13(3):205–220, March 1992. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783692900777.

Morgan:1993:ERE

[MW93] M. J. Morgan and S. J. Walsh. Evaluation of the retention of external tags by juvenile American plaice (Hippoglossoides platessoides) using an aquarium experiment. Fisheries Research, 16(1):1–7, February 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390105G.

McLoughlin:1991:ASD

[MYMP91] R. J. McLoughlin, P. C. Young, R. B. Martin, and J. Parslow. The Australian scallop dredge: estimates of catching efficiency and associated indirect fishing mortality. Fisheries Research, 11(1):1–24, February 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190030J.

Nakken:1993:BRF

[Nak93] Odd Nakken. Book review: Fisheries acoustics: D. N. MacLennan and E. J. Simmonds, Chapman and Hall, London, 1992, 325 pp., £49.95, ISBN 0-412-33060-1. Fisheries Research, 16(1):85–86, February 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390111J.

Nakken:1998:PPF

[Nak98] Odd Nakken. Past, present and future exploitation and management of marine resources in the Barents Sea and adjacent areas. Fisheries Research, 37(1–3):23–35, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001246.

Njoku:1996:ACG

[NE96] D. C. Njoku and I. O. Ezeibekwe. Age composition and growth of the large-scaled mullet, Liza grandisquamis (pisces:

Mugilidae), Valenciennes, 1836 on the New Calabar estuary, off the Nigerian coast. Fisheries Research, 26(1–2):67– 73, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004203.

Nilssen:1992:EIH

[NGH92] Kjell T. Nilssen, Per E. Grotnes, and Tore Haug. The effect of invading harp seals (Phoca groenlandica) on local coastal fish stocks of North Norway. Fisheries Research, 13(1):25– 37, January 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290031N.

Nilsen:1991:PPW

[NGNH91] R. Nilsen, O. K. Gutvik, E. M. Nilssen, and C. C. E. Hopkins. Population parameters of the witch flounder, Glyptocephalus cynoglossus (L.) (Pisces: Pleuronectidae), from Malangen, northern Norway. Fisheries Research, 12(3):259–278, November 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190098Z.

Njoku:1991:CET

[Njo91] Daniel C. Njoku. Comparative efficiency and techno-economics of multifilament and monofilament gillnets on the Oguta Lake, Nigeria. Fisheries Research, 12(1):23–30, September 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190046I.

Norbis:1999:IAG

[NLT99] Walter Norbis, Marıá Inés Lorenzo, and Gabriel J. Torres. Intra-annual growth variations of young-of-the-year hake (Merluccius hubbsi) of the Uruguayan continental shelf based on otolith analysis. Fisheries Research, 44(2):129–137, December 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000594.

Nissling:1998:AEL

[NLVF98] Anders Nissling, Roger Larsson, Lars Vallin, and Karin Frohlund. Assessment of egg and larval viability in cod, Gadus morhua: methods and results from an experimental study. Fisheries Research, 38(2):169–186, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001210.

Norcross:1999:URS

[NM99] Brenda L. Norcross and Franz-Josef Mueter. The use of an ROV in the study of juvenile flatfish. Fisheries Research, 39(3):241–251, January 1, 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698002008.

Nevarez-Martinez:1999:MPS

[NMCCMLB99] Manuel O. Nevárez-Martínez, Ernesto A. Chávez, Miguel A. Cisneros-Mata, and Daniel Lluch-Belda. Modeling of the Pacific sardine Sardinops caeruleus fishery of the Gulf of California, Mexico. Fisheries Research, 41(3):273–283, July 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000235.

Norbis:1995:IWB

[Nor95] Walter Norbis. Influence of wind, behaviour and characteristics of the croaker (Micropogonias furnieri) artisanal fishery in the Rio de la Plata (Uruguay). Fisheries Research, 22(1–2):43–58, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400310S.

Nostvik:1999:CCT

[NP99] F. Nøstvik and T. Pedersen. Catching cod for tagging experiments. Fisheries Research, 42(1–2):57–66, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000375.

Nagasawa:1996:DPS

[NT96] Kazuya Nagasawa and Kazuaki Tadokoro. Do Pacific salmon (Oncorhynchus spp.) steal bait from surface longlines at sea? Fisheries Research, 26(3–4):381–384, May 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004548.

Newman:1995:MSS

[NW95] Stephen J. Newman and David McB. Williams. Mesh size selection and diel variability in catch of fish traps on the central Great Barrier Reef, Australia: a preliminary investigation. Fisheries Research, 23(3–4):237–253, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400353X.

Nowara:1998:ETS

[NW98] Gabrielle B. Nowara and Terence I. Walker. Effects of time of solar day, jigging method and jigging depth on catch rates and size of Gould’s squid, Nototodarus gouldi (McCoy), in southeastern Australian waters. Fisheries Research, 34(3):279– 288, April 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000957.

Nomura:1993:ESF

[NYK93] Tetsuichi Nomura, Mamoru Yoshimizu, and Takahisa Kimura. An epidemiological study of furunculosis in salmon propagation in Japanese rivers. Fisheries Research, 17(1–2):137– 146, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390013W.

Okumus:1997:PSG

[OB97] İbrahim Okumuş and Nadir Başçinar. Population structure, growth and reproduction of introduced Pacific mullet, Mugil so-iuy, in the Black Sea. Fisheries Research, 33(1–3):131–137, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000738.

Otway:1996:GDS

[OCU96] N. M. Otway, J. R. Craig, and J. M. Upston. Geardependent size selection of snapper, Pagrus auratus. Fisheries Research, 28(2):119–132, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696005000.

ONeill:1996:ECM

[OK96] F. G. O’Neill and R. J. Kynoch. The effect of cover mesh size and cod-end catch size on cod-end selectivity. Fisheries Research, 28(3):291–303, October 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696005012.

Okpanefe:1991:MMF

[Okp91] Moses O. Okpanefe. A method of monitoring fishery status. Fisheries Research, 11(2):171–175, April 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190106P.

Otteraa:1998:EAT

[OKS98] Håkon Otterå, Tore S. Kristiansen, and Terje Svåsand. Evaluation of anchor tags used in sea-ranching experiments with Atlantic cod (Gadus morhua L.). Fisheries Research, 35(3):237– 246, May 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000556.

Ohtomi:1998:RGJ

[OM98] Jun Ohtomi and Tatsuro Matsuoka. Reproduction and growth of jack-knife shrimp, Haliporoides sibogae, off south-western Kyushu, Japan. Fisheries Research, 38(3):271–281, November 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369800160X.

ONeill:1993:SSM

[O’N93] F. G. O’Neill. Small-scale modelling rules of trawl nets. Fisheries Research, 18(3–4):173–185, December 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390151V.

Orlowski:1998:AMA

[Orł98] Andrzej Orłowski. Acoustic methods applied to fish environmental studies in the Baltic Sea. Fisheries Research, 34(3):227– 237, April 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698000721.

Otubusin:1990:ELP

[Otu90] Samuel Olu. Otubusin. Effects of lunar periods and some other parameters on fish catch in Lake Kainji, Nigeria. Fisheries Research, 8(3):233–245, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090024P.

Penczak:1997:TSN

[PAL97] T. Penczak, A. A. Agostinho, and J. D. Latini. ‘three seine nets’ sampling applied to the littoral zone of two Brazilian tropical rivers with reduced velocity (Brazil). Fisheries Research, 31(1–2):93–106, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000015.

Park:1998:APW

[Par98] Hae-Hoon Park. Analysis and prediction of walleye pollock (Theragra chalcogramma) landings in Korea by time series analysis. Fisheries Research, 38(1):1–7, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001180.

Patterson:1994:BRS

[Pat94] Kenneth Patterson. Book review: Surveys of fisheries resources: Donald R. Gunderson, Wiley, New York, 1993, 248 pp., £36.95, ISBN 0-471-54735-2. Fisheries Research, 19 (3–4):395–397, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783694900523.

Pauly:1996:OHM

[Pau96] Daniel Pauly. One hundred million tonnes of fish, and fisheries research. Fisheries Research, 25(1):25–38, January 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369500436X.

Pawson:1990:BRE

[Paw90] M. G. Pawson. Book review: Ecology and management of Atlantic salmon: by Derek Mills. Chapman and Hall, London, 1989, 351 pp., hardcover, £45.00, ISBN: 0-412-32140-8. Fisheries Research, 9(1):91–92, July 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090044V.

Pawson:1996:BRH

[Paw96] Mike Pawson. Book review: Hake fisheries, ecology and markets: J. Alheit and T. J. Pritcher (Editors), Chapman and Hall, London, 1994, 496 pp., hardback £50, ISBN 0-412-57350-4. Fisheries Research, 25(3–4):383–384, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696900004.

Pedersen:1992:AHT

[PB92] M. G. Pedersen and J. F. Boettner. Application of hydroacoustic technology to marine fishery management in Washington State. Fisheries Research, 14(2–3):209–219, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290054W.

Presnyakov:1993:SFB

[PB93] V. V. Presnyakov and E. S. Borisenko. The study of fish behaviour under the ice of Lake Glubokoe by means of scanning sonar. Fisheries Research, 15(4):323–329, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390083J.

Psuty:1997:SGN

[PB97] I. Psuty and Wł. Borowski. The selectivity of gill nets to bream (Abramis brama L.) fished in the Polish part of the vistula lagoon. Fisheries Research, 32(3):249–261, December 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000568.

Pierce:1994:LHL

[PBHK94] Graham J. Pierce, Peter R. Boyle, Lee C. Hastie, and Linda Key. The life history of Loligo forbesi (Cephalopoda: Loliginidae) in Scottish waters. Fisheries Research, 21(1–2):17–41, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783694900949.

Pierce:1994:DSL

[PBHS94a] Graham J. Pierce, Peter R. Boyle, Lee C. Hastie, and M. Begoña Santos. Diets of squid Loligo forbesi and Loligo vulgaris in the northeast Atlantic. Fisheries Research, 21(1–2):149–163, December 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694901015.

Pierce:1994:DAF

[PBHS94b] Graham J. Pierce, Peter R. Boyle, Lee C. Hastie, and Aileen M. Shanks. Distribution and abundance of the fished population of Loligo forbesi in UK waters: analysis of fishery data. Fisheries Research, 21(1–2):193–216, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369490104X.

Pitcher:1996:FRA

[PBN96] Tony J. Pitcher, Alida Bundy, and William E. Neill. The fishery for Rastrineobola argentea in Lake Victoria: estimation of potential yields using a new approximate model based on primary production. Fisheries Research, 28(2):133–149, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783696004985.

Pedersen:1991:MCG

[PC91] Torstein Pedersen and Bjørn Carlsen. Marking cod (Gadus morhua L.) juveniles with oxytetracycline incorporated into the feed. Fisheries Research, 12(1):57–64, September 1991.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190049L.

Pallaoro:1998:BPS

[PCD^{+}98] A. Pallaoro, P. Cetinić, J. Dulčić, I. Jardas, and M. Kraljević. Biological parameters of the saddlead bream Oblada melanura in the eastern Adriatic. Fisheries Research, 38(2):199–205, October 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001209.

Perez-Comas:1998:CEM

[PCEP98] José A. Perez-Comas, Daniel L. Erickson, and Ellen K. Pikitch. Cod-end mesh size selection for rockfish and flatfish of the US West Coast. Fisheries Research, 34(3):247–268, April 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697001021.

Perez-Comas:1996:PMM

[PCS96] JoséA. Perez-Comas and John R. Skalski. A parametric multinomial model for size selection in alternate-haul experiments. Fisheries Research, 27(1–3):113–129, June 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004564.

Prince:1993:SRA

[PdP93] Jeremy D. Prince and Sergio A. Guzmán del Próo. A stock reduction analysis of the Mexican abalone (Haliotid) fishery. Fisheries Research, 16(1):25–49, February 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390108J.

Pearce:1993:BRD

[Pea93] Jack B. Pearce. Book review: Dispersal of living organisms into aquatic ecosystems: Aaron Rosenfield and Roger Mann (Editors), A Maryland Sea Grant Publication, College Park, MD, USA, 1992, xxiv + 471 pp., hard cover, US $35, ISBN 0-943676-5 6-8. Fisheries Research, 17(3–4):380–382, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390139X.

Pease:1999:SOA

[Pea99] B. C. Pease. A spatially oriented analysis of estuaries and their associated commercial fisheries in New South Wales, Australia. Fisheries Research, 42(1–2):67–86, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000351.

Pedersen:1997:CVO

[Ped97] Jens Pedersen. Comparison of vertebrae and otoliths measured directly and from radiographs. Fisheries Research, 29(3):277– 282, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005504.

Pereiro:1992:SCR

[Per92] J. A. Pereiro. Some conceptual remarks on yield per recruit. Fisheries Research, 13(4):423–428, April 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290062X.

Pierce:1994:SAM

[PG94] Graham J. Pierce and Angel Guerra. Stock assessment methods used for cephalopod fisheries. Fisheries Research, 21(1–2):255–285, December 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694901082.

Pennington:1995:MEC

[PG95] Michael Pennington and Olav Rune Godø. Measuring the effect of changes in catchability on the variance of marine survey abundance indices. Fisheries Research, 23(3–4):301– 310, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400345W.

Polat:1996:AWM

[PG96] Nazmi Polat and Aysun Gümücs. Ageing of whiting (Merlangius merlangus euxinus, nord., 1840) based on broken and burnt otolith. Fisheries Research, 28(2):231–236, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783696004754.

Pawlak:1991:CNL

[PH91] Carol Pawlak and R. Choudary Hanumara. A comparison of non-linear growth models for fisheries. Fisheries Research, 11(2):143–154, April 1991. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190104N.

Pierce:1994:MVL

[PHG^{+}94] Graham J. Pierce, Lee C. Hastie, Angel Guerra, Roger S. Thorpe, F. Godfrey Howard, and Peter R. Boyle. Morphometric variation in Loligo forbesi and Loligo vulgaris: regional, seasonal, sex, maturity and worker differences. Fisheries Research, 21(1–2):127–148, December 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694901007.

Pascual:1993:HGE

[PI93] Miguel A. Pascual and Oscar O. Iribarne. How good are empirical predictions of natural mortality? Fisheries Research, 16(1):17–24, February 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390107I.

Pickering:1993:EIP

[Pic93] A. D. Pickering. Endocrine-induced pathology in stressed salmonid fish. Fisheries Research, 17(1–2):35–50, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390005R.

Pivnicka:1992:KRC

[Piv92] Karel Pivnička. The Klíčava Reservoir, Czechoslovakia: a 30 year study of the fish community. Fisheries Research, 14(1):1– 20, June 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900696.

Pawson:1996:CMS

[PJ96] Michael G. Pawson and Simon Jennings. A critique of methods for stock identification in marine capture fisheries. Fisheries Research, 25(3–4):203–217, March 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004416.

Penczak:1999:AAB

[PK99] Tadeusz Penczak and Andrzej Kruk. Applicability of the abundance/biomass comparison method for detecting human impacts on fish populations in the Pilica River, Poland. Fisheries Research, 39(3):229–240, January 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369800201X.

Powell:1995:MOP

[PKH^{+}95] Eric N. Powell, John M. Klinck, Eileen E. Hofmann, Elizabeth A. Wilson-Ormond, and Matthew S. Ellis. Modeling oyster populations. V. Declining phytoplankton stocks and the population dynamics of American oyster (Crassostrea virginica) populations. Fisheries Research, 24(3):199–222, October 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369500370P.

Paz:1992:YCV

[PL92] J. Paz and M. G. Larrañeta. Year-class variations of American plaice (Hippoglossoides platessoides) and yellowtail flounder (Limanda ferruginea) in the Grand Banks and the abundance of other commercial fish. Fisheries Research, 15(1–2):121–133, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290008H.

Pajuelo:1996:LHR

[PL96] J. G. Pajuelo and J. M. Lorenzo. Life history of the red porgy Pagrus pagrus (Teleostei: Sparidae) off the Canary Islands, central east Atlantic. Fisheries Research, 28(2):163–177, September 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783696004961.

Pajuelo:1998:PBC

[PL98] J. G. Pajuelo and J. M. Lorenzo. Population biology of the common pandora Pagellus erythrinus (Pisces: Sparidae) off the Canary Islands. Fisheries Research, 36(2–3):75–86, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001106.

Planes:1993:EDS

[Pla93] S. Planes. Evidence of differentiated stocks of Acanthurus triostegus (Pisces, Acanthuridae) in French Polynesia, using electrophoretic data. Fisheries Research, 16(4):287–299, May 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390142T.

Pertierra:1997:ASS

[PLL97] Juan Pablo Pertierra, Jordi Lleonart, and Nancy C. H. Lo. Application of a stage-specific matrix model and length-cohort based analysis to assess the anchovy fishery in Catalan coastal waters (NW Mediterranean Sea). Fisheries Research, 30(1–2): 127–137, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005462.

Plumb:1991:BRF

[Plu91] John A. Plumb. Book review: Fish pathology: Second Edition, R. J. Roberts (Editor). Bailliere Tindall, W. B. Saunders, London, 1989, 467 pp., £75.00, ISBN 0-7020-1314-5. Fisheries Research, 10(3–4):351–353, January 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190085T.

Pedersen:1999:ALS

[PLW99] S. A. Pedersen, P. Lewy, and P. Wright. Assessments of the lesser sandeel (Ammodytes marinus) in the North Sea based on revised stock divisions. Fisheries Research, 41(3):221– 241, July 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000260.

Poppe:1993:SGH

[PM93] Trygve T. Poppe and Tor Atle Mo. Systemic, granulomatous hexamitosis of farmed Atlantic salmon: Interaction with wild fish. Fisheries Research, 17(1–2):147–152, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390014X.

Porteiro:1994:BLF

[PM94] F. M. Porteiro and H. R. Martins. Biology of Loligo forbesi Steenstrup, 1856 (Mollusca: Cephalopoda) in the Azores: sample composition and maturation of squid caught by jigging. Fisheries Research, 21(1–2):103–114, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900981.

Priede:1998:RBN

[PM98] I. G. Priede and N. R. Merrett. The relationship between numbers of fish attracted to baited cameras and population density: Studies on demersal grenadiers Coryphaenoides (Nematonurus) armatus in the abyssal NE Atlantic Ocean. Fisheries Research, 36(2–3):133–137, June 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001052.

Pauly:1995:CAS

[PMA95] D. Pauly, J. Moreau, and N. Abad. Comparison of agestructured and length-converted catch curves of brown trout Salmo trutta in two French rivers. Fisheries Research, 22 (3–4):197–204, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578369400323O.

Penczak:1998:AFP

[PO98] Tadeusz Penczak and Kenneth O’Hara. Accuracy of fish population estimates in a navigation canal with heavy traffic. Fisheries Research, 35(3):191–198, May 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000770.

Polacheck:1991:MET

[Pol91]  Tom Polacheck.  Measures of effort in tuna longline

fisheries: changes at the operational level. Fisheries Research, 12(1):75–87, September 1991. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190051G.

Porteiro:1994:PSS

[Por94] Filipe M. Porteiro. The present status of the squid fishery (Loligo forbesi) in the Azores archipelago. Fisheries Research, 21(1–2):243–253, December 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694901074.

Power:1996:TSR

[Pow96] M. Power. The testing and selection of recruitment distributions for North Atlantic fish stocks. Fisheries Research, 25(1):77–95, January 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783695003959.

Politou:1994:FEM

[PP94] C.-Y. Politou and C. Papaconstantinou. Feeding ecology of Mediterranean poor cod, Trisopterus minutus capelanus (Lacepede), from the eastern coast of Greece. Fisheries Research, 19(3–4):269–292, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900442.

Patterson:1993:SCF

[PPS93] K. R. Patterson, T. J. Pitcher, and T. K. Stokes. A stock collapse in a fluctuating environment: the chub mackerel Scomber japonicus (Houttuyn) in the eastern central Pacific. Fisheries Research, 18(3–4):199–218, December 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390153X.

Pet:1995:CME

[PPSvD95] J. S. Pet, C. Pet-Soede, and W. L. T. van Densen. Comparison of methods for the estimation of gillnet selectivity to tilapia, cyprinids and other fish species in a Sri Lankan reservoir. Fisheries Research, 24(2):141–164, September 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694003643.

Pascoe:1996:MCT

[PR96] Sean Pascoe and Catherine Robinson. Measuring changes in technical efficiency over time using catch and stock information. Fisheries Research, 28(3):305–319, October 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783696005024.

Polet:1999:EPS

[PR99] H. Polet and F. Redant. Effect of population structure, sampling strategy and sample size on the estimates of selection parameters for shrimp (Crangon crangon) trawls. Fisheries Research, 40(3):213–225, April 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002306.

Priede:1991:BRE

[Pri91] I. G. Priede. Book review: Electro-communication in teleost fishes: Behavior and experiments: B. Kramer. Zoophysiology Volume 29, Spriner, Berlin, 1990, 240 pp., Hard cover DM 198.00, ISBN 3-540-51927-0. Fisheries Research, 12(2):184– 185, October 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190043F.

Priede:1995:BRE

[Pri95] I. G. Priede. Book review: Environmental biology of fishes: Malcolm Jobling, Chapman & Hall, London, 1995, 455 pp., Fisheries & Fisheries Series No. 16, ISBN 0-412-58080-2. Fisheries Research, 24(3):268–270, October 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695900195.

Priede:1996:BRA

[Pri96a] I. G. Priede. Book review: An assessment Of Atlantic bluefin tuna: Committee to Review Atlantic Bluefin Tuna (Chairman John J. Magnuson). National Academy Press, Washington, DC, 1994, price £23.95, ISBN 0-309-05181-9. Fisheries Research, 27(4):275–277, July 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004629.

Priede:1996:BRB

[Pri96b] I. G. Priede. Book review: Biology of fishes: (2nd edition) Q. Bone, N. B. Marshall and J. H. S. Blaxter. Blackie, London, 1995, paperback price £19.99 ISBN 0-7514-0022-X, hardback ISBN 0-7514-0243-5. Fisheries Research, 28(1):113–114, August 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783696894446.

Peltonen:1999:ERR

[PRMH99] H. Peltonen, J. Ruuhijärvi, T. Malinen, and J. Horppila. Estimation of roach (Rutilus rutilus (L.)) and smelt (Osmerus eperlanus (L.)) stocks with virtual population analysis, hydroacoustics and gillnet CPUE. Fisheries Research, 44(1):25–36, November 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000570.

Patterson:1992:THO

[PS92] Kenneth R. Patterson and Miguel Santos. The threadherrings Opisthonema spp. off Ecuador: review and population dynamics. Fisheries Research, 14(4):273–294, September 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290037T.

Petrakis:1995:GNS

[PS95a] G. Petrakis and K. I. Stergiou. Gill net selectivity for Diplodus annularis and Mullus surmuletus in Greek waters. Fisheries Research, 21(3–4):455–464, January 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694002936.

Petrakis:1995:WLR

[PS95b] G. Petrakis and K. I. Stergiou. Weight–length relationships for 33 fish species in Greek waters. Fisheries Research, 21(3–4):465–469, January 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694002947.

Petrakis:1996:GNS

[PS96] G. Petrakis and K. I. Stergiou. Gill net selectivity for four fish species (Mullus barbatus, Pagellus erythrinus, Pagellus acarne and Spicara flexuosa) in Greek waters. Fisheries Research, 27(1–3):17–27, June 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696004766.

Pennington:1998:SRT

[PS98] Michael Pennington and Tore Strømme. Surveys as a research tool for managing dynamic stocks. Fisheries Research, 37(1–3):97–106, August 1, 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698001295.

Phiri:1999:DSM

[PS99] Harris Phiri and Kunio Shirakihara. Distribution and seasonal movement of pelagic fish in southern Lake Tanganyika. Fisheries Research, 41(1):63–71, May 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000089.

Pet-Soede:1999:TIC

[PSMSvD99] C. Pet-Soede, M. A. M. Machiels, M. A. Stam, and W. L. T. van Densen. Trends in an Indonesian coastal fishery based on catch and effort statistics and implications for the perception of the state of the stocks by fisheries officials. Fisheries Research, 42(1–2):41–56, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578369900034X.

Padilla:1995:APT

[PT95] Jose E. Padilla and Annabelle C. Trinidad. An application of production theory to fishing effort standardization in the small-pelagics fishery in central Philippines. Fisheries Research, 22(1–2):137–153, February 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400305G.

Punt:1995:PPM

[Pun95] A. E. Punt. The performance of a production-model management procedure. Fisheries Research, 21(3–4):349–374, January 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400302D.

Punt:1997:PVB

[Pun97] A. E. Punt. The performance of VPA-based management. Fisheries Research, 29(3):217–243, March 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005395.

Purdom:1993:BRA

[Pur93] C. E. Purdom. Book review: The Atlantic Salmon: Natural history, exploitation and future management: W. M. Shearer. Available from Fishing News Books, Blackwell Scientific, Oxford, 1992, hardback, 244 pp. with 105 illustrations and a central 4-page colour section, £29.50, ISBN 0-85238-188-3. Fisheries Research, 16(3):280–282, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390101C.

Pet:1997:CES

[PvDM^{+}97a] J. S. Pet, W. L. T. van Densen, M. A. M. Machiels, M. Sukkel, D. Setyohadi, and A. Tumuljadi. Catch, effort and sampling strategies in the highly variable sardine fisheries around East Java, Indonesia. Fisheries Research, 31(1–2):121–137, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000076.

Pet:1997:LBA

[PvDM^{+}97b] J. S. Pet, W. L. T. van Densen, M. A. M. Machiels, M. Sukkel, D. Setyohadi, and A. Tumuljadi. Length-based analysis of population dynamics and stock identification in the sardine fisheries around East Java, Indonesia. Fisheries Research, 31(1–2): 107–120, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000064.

Peja:1996:CAL

[PVM^{+}96] Nikolla Peja, Andrian Vaso, Aleko Miho, Ndoc Rakaj, and Alain J. Crivelli. Characteristics of Albanian lagoons and their fisheries. Fisheries Research, 27(4):215–225, July 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369500467X.

Pengilly:1994:ADI

[PW94] Douglas Pengilly and Leslie J. Watson. Automated detection of internally injected tags in red king crabs at crab processing facilities. Fisheries Research, 19(3–4):293–300, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900450.

Pickering:1997:ARF

[PW97] Helen Pickering and David Whitmarsh. Artificial reefs and fisheries exploitation: a review of the ‘attraction versus production’ debate, the influence of design and its significance for policy. Fisheries Research, 31(1–2):39–59, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000192.

Pet:1995:CSL

[PWMM95] J. S. Pet, J. W. M. Wijsman, P. J. Mous, and M. A. M. Machiels. Characteristics of a Sri Lankan reservoir fishery and consequences for the estimation of annual yield. Fisheries Research, 24(1):9–33, July 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694003609.

Quetglas:1998:BFO

[QAC^{+}98] Antoni Quetglas, Francesc Alemany, Aina Carbonell, Paolo Merella, and Pilar Sánchez. Biology and fishery of Octopus vulgaris Cuvier, 1797, caught by trawlers in Mallorca (Balearic Sea, Western Mediterranean). Fisheries Research, 36(2–3):237–249, June 15, 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698000939.

Quinn:1991:I

[QH91] Terrance J. Quinn and Jonathan Heifetz. Introduction. Fisheries Research, 11(3–4):iii–iv, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190001V.

Quartararo:1996:EDT

[QK96] N. Quartararo and R. E. Kearney. Effects of dart tags on the growth and survival of captive snapper, Pagrus auratus (Sparidae). Fisheries Research, 25(3–4):231–238, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004068.

Quinn:1993:RHS

[Qui93] Thomas P. Quinn. A review of homing and straying of wild and hatchery-produced salmon. Fisheries Research, 18(1–2):29–44, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783693900389.

Ralph:1996:BRB

[Ral96] Robert Ralph. Book review: The biology of the Southern Ocean studies in polar research: George I. Knox, Cambridge University Press, Cambridge, 1994, xiv + 444 pp., ISBN 0-521-32211-1. Fisheries Research, 26(3–4):386–387, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696900053.

Recksiek:1991:SFT

[RAT91] Conrad W. Recksiek, Richard S. Appeldoorn, and Ralph G. Turingan. Studies of fish traps as stock assessment devices on a shallow reef in south-western Puerto Rico. Fisheries Research, 10(3–4):177–197, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190074P.

Roa:1993:GEE

[RB93] Rubén Roa and Roberte Bahamonde. Growth and expansion of an exploited population of the squat lobster (Pleuroncodes monodon) after 3 years without harvesting. Fisheries Research, 18(3–4):305–319, December 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783693901595.

Ragonese:1996:GMY

[RB96] Sergio Ragonese and Marco L. Bianchini. Growth, mortality and yield-per-recruit of the deep-water shrimp Aristeus antennatus (Crustacea-Aristeidae) of the Strait of Sicily (Mediterranean Sea). Fisheries Research, 26(1–2):125–137, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695003940.

Riedel:1995:FAH

[RD95] R. Riedel and J. DeAlteris. Factors affecting hydrodynamic performance of the Nordmøre grate system: a bycatch reduction device used in the Gulf of Maine shrimp fishery. Fisheries Research, 24(3):181–198, October 1995.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369500375K.

Rooker:1996:SLF

[RDG96] Jay R. Rooker, George D. Dennis, and Denis Goulet. Sampling larval fishes with a nightlight lift-net in tropical inshore waters. Fisheries Research, 26(1–2):1–15, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004076.

Ruiz:1997:DHM

[RF97] Ana E. Ruiz and Ricardo R. Fondacaro. Diet of hake (Merluccius hubbsi Marini) in a spawning and nursery area within Patagonian shelf waters. Fisheries Research, 30(1–2):157– 160, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005449.

Reddin:1999:HIC

[RF99] David G. Reddin and Kevin D. Friedland. A history of identification to continent of origin of Atlantic salmon (Salmo salar L.) at west Greenland, 1969–1997. Fisheries Research, 43(1–3):221–235, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000740.

Rodhouse:1990:ADS

[RH90] P. G. Rodhouse and E. M. C. Hatfield. Age determination in squid using statolith growth increments. Fisheries Research, 8(4):323–334, May 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090002D.

Rudstam:1999:CTS

[RHLE99] Lars G. Rudstam, Sture Hansson, Torfinn Lindem, and Donald W. Einhouse. Comparison of target strength distributions and fish densities obtained with split and single beam echo sounders. Fisheries Research, 42(3):207–214, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000478.

Robichaud:1999:EIM

[RHO99] David Robichaud, Wayne Hunte, and Hazel A. Oxenford. Effects of increased mesh size on catch and fishing power of coral reef fish traps. Fisheries Research, 39(3):275–294, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001982.

Richards:1991:UCD

[Ric91] Laura J. Richards. Use of contradictory data sources in stock assessments. Fisheries Research, 11(3–4):225–238, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190003X.

Ringo:1991:FCG

[RJR91] Einar Ringø, Lilly Johansen, and Jan Raa. Feeding of cod, Gadus morhua (L.), larvae on an artificial diet. Preliminary results. Fisheries Research, 11(2):191–193, April 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190108R.

Ransom:1998:RMA

[RJS98] Bruce H. Ransom, Samuel V. Johnston, and Tracey W. Steig. Review on monitoring adult salmonid (Oncorhynchus and Salmo spp.) escapement using fixed-location split-beam hydroacoustics. Fisheries Research, 35(1–2):33–42, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000575.

Rosenberg:1990:ASA

[RKCB90] A. A. Rosenberg, G. P. Kirkwood, J. A. Crombie, and J. R. Beddington. The assessment of stocks of annual squid species. Fisheries Research, 8(4):335–350, May 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090003E.

Reddy:1990:RTE

[RKK90] Y. Mallikarjuna Reddy, Indrani Karunasagar, and I. Karunasagar. Resazurin test for estimating bacteriological quality of fishes. Fisheries Research, 9(1):75–79, July 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090042T.

Ranta:1990:WQV

[RL90] Esa Ranta and Kai Lindström. Water quality versus other determinants of species-specific yield of fish in Northern Finnish lakes. Fisheries Research, 8(4):367–379, May 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090005G.

Ranta:1992:FCW

[RLR92] Esa Ranta, Kai Lindström, and Martti Rask. Fish catch and water quality in small lakes. Fisheries Research, 13(1):1–7, January 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290029S.

Ranta:1992:WQF

[RLS92] Esa Ranta, Kai Lindström, and Kalervo Salojärvi. Water quality, fishing effort and fish yield in lakes. Fisheries Research, 15(1–2):105–119, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290007G.

Robins:1999:AII

[RM99] J. B. Robins and J. G. McGilvray. The AusTED II, an improved trawl efficiency device 2. Commercial performance. Fisheries Research, 40(1):29–41, February 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002227.

Rogers:1999:FEN

[RMR^{+}99] S. I. Rogers, D. Maxwell, A. D. Rijnsdorp, U. Damm, and W. Vanhee. Fishing effects in northeast Atlantic shelf seas: patterns in fishing effort, diversity and community structure. IV. Can comparisons of species diversity be used to assess human impacts on demersal fish faunas? Fisheries Research, 40(2):135–152, March 1, 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698002094.

Ramm:1993:USP

[RMXP93] David C. Ramm, Richard P. Mounsey, Yongshun Xiao, and Susan E. Poole. Use of a semi-pelagic trawl in a tropical demersal trawl fishery. Fisheries Research, 15(4):301–313, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390081H.

Rose:1998:SCG

[RN98] Craig S. Rose and Edmund P. Nunnallee. A study of changes in groundfish trawl catching efficiency due to differences in operating width, and measures to reduce width variation. Fisheries Research, 36(2–3):139–147, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369800099X.

Robin:1992:BSF

[Rob92] Jean-Paul Robin. The brown shrimp fishery of the Loire Estuary: production and by-catch of juvenile fish. Fisheries Research, 13(2):153–172, February 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290024N.

Roberts:1993:UDN

[Rob93] Ronald J. Roberts. Ulcerative dermal necrosis (UDN) in wild salmonids. Fisheries Research, 17(1–2):3–14, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390003P.

Rose:1992:RPN

[Ros92] G. A. Rose. A review of problems and new directions in the application of fisheries acoustics on the Canadian East Coast. Fisheries Research, 14(2–3):105–128, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290047W.

Reis:1992:DGN

[RP92] E. G. Reis and M. G. Pawson. Determination of gill-net selectivity for bass (Dicentrarchus labrax L.) using commercial catch data. Fisheries Research, 13(2):173–187, February 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290025O.

Reis:1999:FME

[RP99] E. G. Reis and M. G. Pawson. Fish morphology and estimating selectivity by gillnets. Fisheries Research, 39(3):263–273, January 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001994.

Rogers:1993:FAM

[RR93] Donald E. Rogers and Gregory T. Ruggerone. Factors affecting marine growth of Bristol Bay sockeye salmon. Fisheries Research, 18(1–2):89–103, October 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783693900426.

Rogers:1997:EST

[RRdS^{+}97] Donna R. Rogers, Barton D. Rogers, Janaka A. de Silva, Vernon L. Wright, and John W. Watson. Evaluation of shrimp trawls equipped with bycatch reduction devices in inshore waters of Louisiana. Fisheries Research, 33(1–3):55–72, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369700060X.

Rago:1998:IRI

[RSB^{+}98] P. J. Rago, K. A. Sosebee, J. K. T. Brodziak, S. A. Murawski, and E. D. Anderson. Implications of recent increases in catches on the dynamics of Northwest Atlantic spiny dogfish (Squalus acanthias). Fisheries Research, 39(2):165–181, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001817.

Rueda:1999:PDS

[RSM99] Mario Rueda and Adriana Santos-Martínez. Population dynamics of the striped mojarra Eugerres plumieri from the Ciénaga Grande de Santa Marta, Colombia. Fisheries Research, 42(1–2):155–166, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000338.

Robins-Troeger:1994:EMS

[RT94] J. B. Robins-Troeger. Evaluation of the Morrison soft turtle excluder device: prawn and bycatch variation in Moreton Bay, Queensland. Fisheries Research, 19(3–4):205–217, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900396.

Robins-Troeger:1995:DTE

[RTBD95] J. B. Robins-Troeger, R. C. Buckworth, and M. C. L. Dredge. Development of a trawl efficiency device (TED) for Australian prawn fisheries. II. Field evaluations of the AusTED. Fisheries Research, 22(1–2):107–117, February 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400308J.

Ruzzante:1999:REG

[RTC99] D. E. Ruzzante, C. T. Taggart, and D. Cook. A review of the evidence for genetic structure of cod (Gadus morhua) populations in the NW Atlantic and population affinities of larval cod off Newfoundland and the Gulf of St. Lawrence. Fisheries Research, 43(1–3):79–97, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000673.

Rueda:1998:PDF

[RU98] Mario Rueda and H.-Jörg Urban. Population dynamics and fishery of the fresh-water clam Polymesoda solida (Corbiculidae) in Cienaga Poza Verde, Salamanca Island, Colombian Caribbean. Fisheries Research, 39(1):75–86, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001684.

Rubin:1994:SDS

[Rub94] Jean-François Rubin. Survival and development of sea trout, Salmo trutta (L.), eggs in Baltic Sea saltwater. Fisheries Research, 20(1):1–12, June 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900027.

Rijnsdorp:1990:VPB

[RvLV90] A. D. Rijnsdorp, P. I. van Leeuwen, and T. A. M. Visser. On the validity and precision of back-calculation of growth from otoliths of the plaice, Pleuronectes platessa L. Fisheries Research, 9(2):97–117, August 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783690900584.

Ramm:1995:HGE

[RX95] David C. Ramm and Yongshun Xiao. Herding in groundfish and effective pathwidth of trawls. Fisheries Research, 24 (3):243–259, October 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016578369500373I.

S:1994:EPS

[S.94] Luis Cubillos S. Estimating the parameters of the stockrecruitment model of Ricker from a yield-per-recruit model in data-limited situations. Fisheries Research, 20(2–3):229– 242, August 1994. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369490085X.

Stewart:1993:DCN

[SAB^{+}93] D. A. Stewart, D. Angew, R. Boyd, R. Briggs, and P. Toland. The derivation of and changes in Nephrops per unit effort values for the Northern Ireland fishing fleet. Fisheries Research, 17(3–4):273–292, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390130Y.

Salvanes:1991:SCG

[Sal91] Anne Gro Vea Salvanes. The selectivity for cod (Gadus morhua L.) in two experimental trammel-nets and one gillnet. Fisheries Research, 10(3–4):265–285, January 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190080Y.

Salvanes:1995:PPP

[Sal95] Anne Gro Vea Salvanes. Pollack (Pollachius pollachius) stock size development and potential influence on cod (Gadus morhua) mariculture in a west Norwegian fjord. Fisheries Research, 24(3):223–242, October 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369500369L.

Sanders:1993:FPV

[San93] M. J. Sanders. Fishery performance and the value of future entitlements under quota management: a case study of a handline fishery in the southwest Indian Ocean. Fisheries Research, 18(3–4):219–229, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390154Y.

Sanyanga:1996:VAS

[San96] Rudo Angela Sanyanga. Variations in abundance of Synodontis zambezensis (Pisces: Mochokidae) Peters 1852, in the inshore fishery of Lake Kariba. Fisheries Research, 26(1–2):171– 186, April 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369500386X.

Siar:1992:ATU

[SAV92] Susana V. Siar, Renato F. Agbayani, and Jaime B. Valera. Acceptability of territorial use rights in fisheries: towards community-based management of small-scale fisheries in the Philippines. Fisheries Research, 14(4):295–304, September 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290038U.

Stocker:1990:SPM

[SB90] M. Stocker and T. H. Butler. Surplus production model analysis of crab (Cancer magister Dana) fisheries of British Columbia, Canada. Fisheries Research, 9(3):231–254, September 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605800049.

Sanders:1996:BEM

[SB96] M. J. Sanders and K. H. H. Beinssen. Bio-economic modelling of a fishery under individual transferable quota management: a case study of the fishery for blacklip abalone Haliotis rubra in the Western Zone of Victoria (Australia). Fisheries Research, 27(4):179–201, July 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695004661.

Sanders:1997:UAF

[SB97a] M. J. Sanders and K. H. H. Beinssen. Uncertainty analysis of a fishery under individual transferable quota management: Applied to the fishery for blacklip abalone Haliotis rubra in the Western Zone of Victoria (Australia). Fisheries Research, 31(3):215–228, August 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000337.

Smith:1997:GDO

[SB97b] P. J. Smith and P. G. Benson. Genetic diversity in orange roughy from the east of New Zealand. Fisheries Research, 31(3):197–213, August 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000349.

Salvanes:1998:PFF

[SB98a] Anne Gro Vea Salvanes and Beatriz M. Baliño. Productivity and fitness in a fjord cod population: an ecological and evolutionary approach. Fisheries Research, 37(1–3):143–161, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001337.

Sanders:1998:CMS

[SB98b] M. J. Sanders and K. H. H. Beinssen. A comparison of management strategies for the rehabilitation of a fishery: Applied to the fishery for blacklip abalone Haliotis rubra in the Western Zone of Victoria (Australia). Fisheries Research, 38(3):283–301, November 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001581.

Sangster:1998:GPC

[SB98c] G. I. Sangster and M. Breen. Gear performance and catch comparison trials between a single trawl and a twin rigged gear. Fisheries Research, 36(1):15–26, June 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000861.

Sumpton:1995:FGM

[SBK95] W. D. Sumpton, I. W. Brown, and S. J. Kennelly. Fishing gears that minimise the damage incurred by discarded spanner crabs (Ranina ranina): Laboratory and field experiments. Fisheries Research, 22(1–2):11–27, February 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400311J.

Stone:1998:FMU

[SBM^{+}98] Richard B. Stone, C. Michael Bailey, Sarah A. McLaughlin, Pamela M. Mace, and Margo B. Schulze. Federal management of US Atlantic shark fisheries. Fisheries Research, 39(2):215–221, December 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001854.

Sandberg:1998:BAT

[SBR98] Per Sandberg, Bjarte Bogstad, and Ingolf Røttingen. Bioeconomic advice on TAC — the state of the art in the Norwegian fishery management. Fisheries Research, 37(1–3):259–274, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001428.

Shaklee:1999:MFU

[SBSW99] James B. Shaklee, Terry D. Beacham, Lisa Seeb, and Bruce A. White. Managing fisheries using genetic data: case studies from four species of Pacific salmon. Fisheries Research, 43(1–3):45–78, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000661.

Stergiou:1996:MFA

[SC96] K. I. Stergiou and E. D. Christou. Modelling and forecasting annual fisheries catches: comparison of regression, univariate and multivariate time series methods. Fisheries Research, 25(2):105–138, February 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695003894.

Schnute:1991:INF

[Sch91] Jon T. Schnute. The importance of noise in fish population models. Fisheries Research, 11(3–4):197–223, August 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190002W.

Schweder:1998:FBM

[Sch98] Tore Schweder. Fisherian or Bayesian methods of integrating diverse statistical information? Fisheries Research, 37(1–3):61–75, August 1, 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698001271.

Stergiou:1997:MFM

[SCP97] K. I. Stergiou, E. D. Christou, and G. Petrakis. Modelling and forecasting monthly fisheries catches: comparison of regression, univariate and multivariate time series methods. Fisheries Research, 29(1):55–95, January 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696004821.

Smita:1990:PDP

[SD90] Paralkar Smita and M. Devaraj. Population dynamics of the prawn Metapenaeus dobsoni (Miers) exploited in the trawl fishery along the Southwest Coast of India. Fisheries Research, 8(4):381–395, May 1990. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090006H.

Storbeck:1991:WEF

[SD91] Frank Storbeck and Berent Daan. Weight estimation of flatfish by means of structured light and image analysis. Fisheries Research, 11(2):99–108, April 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190101K.

Salam:1994:BCN

[SD94] A. Salam and P. M. C. Davies. Body composition of northern pike (Esox lucius L.) in relation to body size and condition factor. Fisheries Research, 19(3–4):193–204, April 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900388.

Secor:1999:SDM

[Sec99] D. H. Secor. Specifying divergent migrations in the concept of stock: the contingent hypothesis. Fisheries Research, 43(1–3):13–34, October 1999. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000648.

Samaranayaka:1997:EHR

[SEJ97] Ariyapala Samaranayaka, Arill Engås, and Terje Jørgensen. Effects of hanging ratio and fishing depth on the catch rates of drifting tuna gillnets in Sri Lankan waters. Fisheries Research, 29(1):1–12, January 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696005243.

Suuronen:1996:MHE

[SEO96] Petri Suuronen, Daniel L. Erickson, and Ari Orrensalo. Mortality of herring escaping from pelagic trawl codends. Fisheries Research, 25(3–4):305–321, March 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004467.

Swain:1999:SCU

[SF99] Douglas P. Swain and Chris J. Foote. Stocks and chameleons: the use of phenotypic variation in stock identification. Fisheries Research, 43(1–3):113–128, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000697.

Smith:1991:LGD

[SFM91] P. J. Smith, R. I. C. C. Francis, and M. McVeagh. Loss of genetic diversity due to fishing pressure. Fisheries Research, 10(3–4):309–316, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190082Q.

Stotz:1997:AGP

[SG97] Wolfgang B. Stotz and Sergio A. González. Abundance, growth, and production of the sea scallop Argopecten purpuratus (Lamarck 1819): bases for sustainable exploitation of natural scallop beds in north-central Chile. Fisheries Research, 32(2):173–183, November 1, 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000106.

Seyhan:1998:FCW

[SG98] K. Seyhan and D. J. Grove. Food consumption of whiting, Merlangius merlangus, in the Eastern Irish Sea. Fisheries Research, 38(3):233–245, November 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001659.

Seyhan:1998:FBW

[SGK98] K. Seyhan, D. J. Grove, and J. King. Feeding behaviour of whiting, Merlangius merlangus, L. in captivity. Fisheries Research, 34(1):39–45, February 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000805.

Santos:1997:FFS

[SH97] Roberta Aguiar Santos and Manuel Haimovici. Food and feeding of the short-finned squid Illex argentinus (Cephalopoda: Ommastrephidae) off southern Brazil. Fisheries Research, 33(1–3):139–147, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000714.

Shaw:1994:EDA

[Sha94] Robert Shaw. An economic description and assessment of the squid catching industry in the UK, Portugal and Spain. Fisheries Research, 21(1–2):287–303, December 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694901090.

Shearer:1990:ASS

[She90a] W. M. Shearer. The Atlantic salmon (Salmo salar L.) of the North Esk with particular reference to the relationship between both river and sea age and time of return to home waters. Fisheries Research, 10(1–2):93–123, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090017P.

Shelton:1990:FRA

[She90b] R. G. J. Shelton. Fishery research and the Atlantic salmon, a note on the context. Fisheries Research, 10(1–2):1–4, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090011J.

Shelton:1993:BRP

[She93] R. G. J. Shelton. Book review: Pollution and freshwater fish: Richard Lloyd, The Buckland Foundation, Fishing News Books, Blackwell Scientific, Oxford, 1992, 176 pp., £25, ISBN 0-85238-187-5. Fisheries Research, 16(3):282–284, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390102D.

Shelton:1998:CBF

[She98] Peter A. Shelton. A comparison between a fixed and a variable fishing mortality control rule used to manage the cod stock off southern Labrador and the east coast of Newfoundland. Fisheries Research, 37(1–3):275–286, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369800143X.

Schweder:1998:ECH

[SHH98] Tore Schweder, Gro S. Hagen, and Einar Hatlebakk. On the effect on cod and herring fisheries of retuning the Revised Management Procedure for minke whaling in the greater Barents Sea. Fisheries Research, 37(1–3):77–95, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001283.

Steig:1998:AMS

[SI98] Tracey W. Steig and Thomas K. Iverson. Acoustic monitoring of salmonid density, target strength, and trajectories at two dams on the Columbia River, using a split-beam scanning system. Fisheries Research, 35(1–2):43–53, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698000587.

Siddeek:1991:ENM

[Sid91] M. S. M. Siddeek. Estimation of natural mortality of Kuwait’s grooved tiger prawn Penaeus semisulcatus (De Haan) using tag-recapture and commercial fisheries data. Fisheries Research, 11(2):109–125, April 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190102L.

Sousa:1999:SPL

[SIE99] Fernando Sousa, Eduardo Isidro, and Karim Erzini. Semipelagic longline selectivity for two demersal species from the Azores: the black spot sea bream (Pagellus bogaraveo) and the bluemouth rockfish (Helicolenus dactylopterus dactylopterus). Fisheries Research, 41(1):25–35, May 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369800232X.

Sipponen:1998:IOF

[Sip98] M. Sipponen. The impact of ownership of fishing rights on professional fishing in Finnish lakes. Fisheries Research, 34 (2):123–136, March 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578369700091X.

Scarnecchia:1991:EFL

[SÍW91] Dennis L. Scarnecchia, Árni Ísaksson, and S. E. White. Effects of the Faroese long-line fishery, other oceanic fisheries and oceanic variations on age at maturity of Icelandic north-coast stocks of Atlantic salmon (Salmo salar). Fisheries Research, 10(3–4):207–228, January 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190076R.

Shelton:1995:PAO

[SJ95] R. G. J. Shelton and A. D. F. Johnstone. Preliminary aquarium observations of the schooling behaviour of Atlantic salmon, Salmo salar L., smolts on transfer to sea water. Fisheries Research, 24(1):73–78, July 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694003632.

Siddeek:1997:GPE

[SJ97] M. S. M. Siddeek and D. W. Johnson. Growth parameter estimates for Omani abalone (Haliotis mariae, Wood 1828) using length-frequency data. Fisheries Research, 31(3):169– 188, August 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000362.

Stewart:1998:CMT

[SK98] John Stewart and Steven J. Kennelly. Contrasting movements of two exploited scyllarid lobsters of the genus Ibacus off the east coast of Australia. Fisheries Research, 36(2–3):127–132, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001040.

Su:1998:CFF

[SL98] Zhenming Su and Qun Liu. A continuous Fox-form of the surplus production observation-error estimator. Fisheries Research, 34(1):59–76, February 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000660.

Sangster:1996:CFE

[SLB96] G. I. Sangster, K. Lehmann, and M. Breen. Commercial fishing experiments to assess the survival of haddock and whiting after escape from four sizes of diamond mesh codends. Fisheries Research, 25(3–4):323–345, March 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004300.

Suuronen:1997:AEB

[SLW97] Petri Suuronen, Esa Lehtonen, and John Wallace. Avoidance and escape behaviour by herring encountering midwater trawls. Fisheries Research, 29(1):13–24, January 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696005231.

Sipponen:1996:FED

[SM96] M. Sipponen and M. Muotka. Factors effecting the demand for recreational fishing opportunities in Finnish lakes during the 1980s. Fisheries Research, 26(3–4):309–323, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783695004211.

Santos:1997:OAA

[SM97] Miguel Neves Santos and Carlos Costa Monteiro. The Olhão artificial reef system (south Portugal): Fish assemblages and fishing yield. Fisheries Research, 30(1–2):33–41, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369600567X.

Santos:1998:CCF

[SM98a] Miguel Neves Santos and Carlos Costa Monteiro. Comparison of the catch and fishing yield from an artificial reef system and neighbouring areas off Faro (Algarve, south Portugal). Fisheries Research, 39(1):55–65, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001696.

Sarda:1998:APD

[SM98b] F. Sardà and F. Maynou. Assessing perceptions: Do Catalan fishermen catch more shrimp on fridays? Fisheries Research, 36(2–3):149–157, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783698001027.

Schleusner:1999:MLB

[SM99] Clifford J. Schleusner and O. Eugene Maughan. Mobility of largemouth bass in a desert lake in Arizona. Fisheries Research, 44(2):175–178, December 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369900082X.

Santos:1995:ABG

[SME95] Miguel Neves Santos, Carlos Costa Monteiro, and Karim Erzini. Aspects of the biology and gillnet selectivity of the axillary seabream (Pagellus acarne, Risso) and common pandora (Pagellus erythrinus, Linnaeus) from the Algarve (south Portugal). Fisheries Research, 23(3–4):223–236, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400354Y.

Santos:1998:MGN

[SMEL98] Miguel Neves Santos, Carlos Costa Monteiro, Karim Erzini, and Gerard Lasserre. Maturation and gill-net selectivity of two small sea breams (genus Diplodus) from the Algarve coast (south Portugal). Fisheries Research, 36(2–3):185–194, June 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001003.

Smith:1990:RBR

[Smi90] G. W. Smith. The relationship between river flow and net catches of salmon (Salmo salar L.) in and around the mouth of the Aberdeenshire Dee between 1973 and 1986. Fisheries Research, 10(1–2):73–91, December 1990. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369090016O.

Skogen:1999:PSB

[SMS99] Morten D. Skogen, Terje Monstad, and Einar Svendsen. A possible separation between a northern and a southern stock of the northeast Atlantic blue whiting. Fisheries Research, 41(2):119– 131, June 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000193.

Spicer:1995:DMB

[SMT95] Agnes V. Spicer, John R. Moring, and Joan G. Trial. Downstream migratory behavior of hatchery-reared, radio-tagged Atlantic salmon (Salmo salar) smolts in the Penobscot River, Maine, USA. Fisheries Research, 23(3–4):255–266, June 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400352W.

Stergiou:1993:NMR

[SP93] K. I. Stergiou and C. Papaconstantinou. Natural mortality of red bandfish, Cepola macrophthalma (L.), in the Aegean Sea (Greece): comparison of direct and indirect estimation methods. Fisheries Research, 16(4):347–361, May 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390146X.

Stergiou:1994:SAC

[SP94] K. I. Stergiou and D. A. Pollard. A spatial analysis of the commercial fisheries catches from the Greek Aegean Sea. Fisheries Research, 20(2–3):109–135, August 1994. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900787.

Skalski:1993:NML

[SPC93] John R. Skalski and JoséAntonio Perez-Comas. Nonparametric maximum likelihood estimation of mesh size selectivity. Fisheries Research, 18(3–4):321–334, December 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783693901609.

Stergiou:1996:SSF

[SPP96] K. I. Stergiou, G. Petrakis, and C.-Y. Politou. Small-scale fisheries in the South Euboikos Gulf (Greece): species composition and gear competition. Fisheries Research, 26(3–4):325– 336, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004246.

Stergiou:1997:SSD

[SPP97] K. I. Stergiou, G. Petrakis, and Ch.-Y. Politou. Size selectivity of diamond and square mesh cod-ends for Nephrops norvegicus in the Aegean Sea. Fisheries Research, 29(3):203– 209, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005425.

Samame:1995:WLR

[SQM95] Manuel Samamé, B. Roger Quiroz, and Tadanobu Machii. Weight–length relationships and reproduction of the Peruvian pintadilla, Cheilodactylus variegatus V. (Cheilodactylidae), from the Callao fishing zone, Peru. Fisheries Research, 22(3–4):279–291, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400315N.

Simon:1996:SSS

[SRG96] Fernando Simón, Francisco Rocha, and Angel Guerra. The small-scale squid hand-jig fishery off the northwestern Iberian Peninsula: application of a model based on a short survey of fishery statistics. Fisheries Research, 25(3–4):253–263, March 1996. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695004254.

Sanchez:1995:DAP

[SRMdC95] Ramiro P. Sánchez, Alexander Remeslo, Adrián Madirolas, and Janina D. de Ciechomski. Distribution and abundance of post-larvae and juveniles of the Patagonian sprat, Sprattus fuegensis, and related hydrographic conditions. Fisheries Research, 23(1–2):47–81, May 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400339X.

Summers:1990:INA

[SRSR90] J. Kevin Summers, William A. Richkus, C. Foster Stroup, and Louis J. Rugolo. The influence of natural and anthropogenic environmental change on white perch stock status in the Choptank River, Maryland. Fisheries Research, 9(3):255–268, September 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605800050.

Sanjeev:1993:QSV

[SS93] S. Sanjeev and Jose Stephen. Quantitative studies on Vibrio parahaemolyticus in market samples of finfish and shellfish from Cochin (South India). Fisheries Research, 16(3):273– 276, April 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390098R.

Singh:1995:AGH

[SS95] Deepak Singh and Ramesh C. Sharma. Age and growth of a Himalayan teleost Schizothorax richardsonii (Gray) from the Garhwal Hills (India). Fisheries Research, 24(4):321–329, November 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695003766.

Sardella:1996:PCM

[ST96] Norma Haydée Sardella and Juan Tomás Timi. Parasite communities of Merluccius hubbsi from the Argentinian– Uruguayan Common Fishing Zone. Fisheries Research, 27 (1–3):81–88, June 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165783695004602.

Stergiou:1990:PMF

[Ste90] K. I. Stergiou. Prediction of the Mullidae fishery in the eastern Mediterranean 24 months in advance. Fisheries Research, 9(1): 67–74, July 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090041S.

Stergiou:1991:DFS

[Ste91] K. I. Stergiou. Describing and forecasting the sardineanchovy complex in the eastern Mediterranean using Vector AutoRegressions. Fisheries Research, 11(2):127–141, April 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190103M.

Steele:1995:BRW

[Ste95] J. H. Steele. Book review: Webs and scales: Michael M. Mullen. University of Washington Press, Washington, DC, 1993, 148 pp., US$15, ISBN 0-295-97245-9. Fisheries Research, 23(3–4):381–383, June 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695900314.

Steele:1996:RSF

[Ste96] John H. Steele. Regime shifts in fisheries management. Fisheries Research, 25(1):19–23, January 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783695004408.

Stephenson:1999:SCF

[Ste99] Robert L. Stephenson. Stock complexity in fisheries management: a perspective of emerging issues related to population sub-units. Fisheries Research, 43(1–3):247–249, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000764.

Shelton:1997:RPS

[STM^{+}97] R. G. J. Shelton, W. R. Turrell, A. Macdonald, I. S. McLaren, and N. T. Nicoll. Records of post-smolt Atlantic salmon, Salmo salar L., in the Faroe–Shetland Channel in June 1996. Fisheries Research, 31(1–2):159–162, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000143.

Stokes:1990:BRB

[Sto90] Kevin Stokes. Book review: Basic fishery science programs: Saul B. Saila, Conrad W. Recksiek and Michael H. Prager, Elsevier Science Publishing Co. Inc., Amsterdam, 1988, 230 pp., Dfl. 140.00, ISBN 0-444-43031-8. Fisheries Research, 8(3):293– 294, January 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369090029U.

Salvanes:1992:PPM

[SU92] Anne Gro Vea Salvanes and Øyvind Ulltang. Population parameters, migration and exploitation of the cod (Gadus morhua L.) in masfjorden, western Norway. Fisheries Research, 15(3):253–289, December 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290019P.

Solarin:1993:IBP

[SU93] B. B. Solarin and R. E. K. Udolisa. An investigation of brush park fishing in lagos lagoon, Nigeria. Fisheries Research, 15(4):331–337, January 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390084K.

Sumaila:1998:PMR

[Sum98] Ussif Rashid Sumaila. Protected marine reserves as fisheries management tools: a bioeconomic analysis. Fisheries Research, 37(1–3):287–296, August 1, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001441.

Svealv:1991:STO

[Sve91] Tore Sveälv. Strategies and technologies in offshore farming. Fisheries Research, 10(3–4):329–349, January 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369190084S.

Svendsen:1995:TJC

[Sve95] Y. S. Svendsen. Tracking juvenile cod (Gadus morhua L.) in northern Norway using acoustic tags. Fisheries Research, 23(3–4):311–318, June 1995. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694003492.

Svedang:1999:VPS

[Sve99] H. Svedäng. Vital population statistics of the exploited eel stock on the Swedish west coast. Fisheries Research, 40(3):251– 265, April 1, 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698002264.

Sun:1991:ACH

[SX91] Manchang Sun and Liuxiong Xu. Application of canvas horizontal spreaders to set nets. Fisheries Research, 12(2):157– 169, October 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190039I.

Symes:1997:FMS

[Sym97] David Symes. Fisheries management: in search of good governance. Fisheries Research, 32(2):107–114, November 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000520.

Secor:1998:OMA

[SZ98] David H. Secor and Vincent S. Zdanowicz. Otolith microconstituent analysis of juvenile bluefin tuna (Thunnus thynnus) from the Mediterranean Sea and Pacific Ocean. Fisheries Research, 36(2–3):251–256, June 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001039.

Tjelmeland:1998:MRM

[TB98] Sigurd Tjelmeland and Bjarte Bogstad. MULTSPEC — a review of a multispecies modelling project for the Barents Sea. Fisheries Research, 37(1–3):127–142, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001325.

Tsai:1992:STF

[TC92] Chu-Fa Tsai and Ai-Ling Chai. Short-term forecasting of the striped bass (Morone saxatilis) commercial harvest in the Maryland portion of Chesapeake Bay. Fisheries Research, 15(1–2):67–82, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290005E.

Tuck:1997:CMS

[TCA^{+}97] I. D. Tuck, C. J. Chapman, R. J. A. Atkinson, N. Bailey, and R. S. M. Smith. A comparison of methods for stock assessment of the Norway lobster, Nephrops norvegicus, in the Firth of Clyde. Fisheries Research, 32(1):89–100, October 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000350.

Tsangridis:1991:ULF

[TF91] A. Tsangridis and N. Filippousis. Use of length-frequency data in the estimation of growth parameters of three Mediterranean fish species: bogue (Boops boops L.), picarel (Spicara smaris L.) and horse mackerel (Trachurus trachurus L.). Fisheries Research, 12(4):283–297, December 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783691900136.

Tsangridis:1994:ATM

[TF94] A. Tsangridis and N. Filippousis. Analysis of two models for picarel (Spicara smaris L.) growth using Schnute’s micro-simplex nonlinear estimation procedure. Fisheries Research, 20(2–3):181–189, August 1994. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900825.

Torstensen:1995:OGS

[TG95] E. Torstensen and J. Gjøsæter. Occurrence of 0-group sprat (Sprattus sprattus) in the littoral zone along the Norwegian Skagerrak coast 1945–1992, compared with the occurrence of 0-group herring (Clupea harengus). Fisheries Research, 21(3–4):409–421, January 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369400297A.

Toresen:1998:AMU

[TGdB98] Reidar Toresen, Harald Gjøsæter, and Pedro de Barros. The acoustic method as used in the abundance estimation of capelin (Mallotus villosus Müller) and herring (Clupea harengus Linné) in the Barents Sea. Fisheries Research, 34(1):27–37, February 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000775.

Tully:1995:CSS

[TH95] O. Tully and J. P. Hillis. Causes and spatial scales of variability in population structure of Nephrops norvegicus (L.) in the Irish Sea. Fisheries Research, 21(3–4):329–347, January 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369400303E.

Thomas:1992:SFF

[Tho92a] G. L. Thomas. Successes and failures of fisheries acoustics — an international, federal, and regional point of view. Fisheries Research, 14(2–3):95–104, August 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290046V.

Thorne:1992:CST

[Tho92b] Richard E. Thorne. Current status of training and education in fisheries acoustics. Fisheries Research, 14(2–3):135– 141, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290049Y.

Thorsteinsson:1992:USM

[Tho92c] Gudni Thorsteinsson. The use of square mesh codends in the Icelandic shrimp (Pandalus borealis) fishery. Fisheries Research, 13(3):255–266, March 1992. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369290080D.

Thorne:1998:RES

[Tho98] Richard E. Thorne. Review: experiences with shallow water acoustics. Fisheries Research, 35(1–2):137–141, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369800068X.

Thresher:1999:ECO

[Thr99] Ronald E. Thresher. Elemental composition of otoliths as a stock delineator in fishes. Fisheries Research, 43(1–3):165– 204, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000727.

Tamsett:1999:STM

[TJ99] Duncan Tamsett and Gareth Janacek. Sampling trips for measuring discards in commercial fishing based on multilevel modelling of measurements in the North Sea from NE England. Fisheries Research, 42(1–2):103–115, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000326.

Tariq:1992:RBM

[TJA92] Jaleel Tariq, M. Jaffar, and M. Ashraf. Relationship between mercury concentration and length, weight and sex of two sprinid fish, Catla catla and Chela chanius, from Rawal Lake, Pakistan. Fisheries Research, 14(4):335–341, September 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290042R.

Tariq:1994:TMC

[TJA94] Jaleel Tariq, M. Jaffar, and M. Ashraf. Trace metal concentration, distribution and correlation in water, sediment and fish from the Ravi River, Pakistan. Fisheries Research, 19(1–2):131–139, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783694900191.

Tamsett:1999:CMO

[TJE99a] Duncan Tamsett, Gareth Janacek, and Mike Emberton. A comparison of methods for onboard sampling of discards in commercial fishing. Fisheries Research, 42(1–2):127–135, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000314.

Tamsett:1999:OSM

[TJE^{+}99b] Duncan Tamsett, Gareth Janacek, Mike Emberton, Bill Lart, and Grant Course. Onboard sampling for measuring discards in commercial fishing based on multilevel modelling of measurements in the Irish Sea from NW England and N Wales. Fisheries Research, 42(1–2):117–126, August 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000302.

Turunen:1994:TSM

[TKH94] Timo Turunen, Anne Käkelä, and Heikki Hyvärinen. Trawling stress and mortality in undersized (<40 cm) brown trout (Salmo trutta L.). Fisheries Research, 19(1–2):51–64, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783694900140.

Tokac:1998:CES

[TLT^{+}98] A. Tokaç, A. Lök, Z. Tosunoğlu, C. Metin, and R. S. T. Ferro. Cod-end selectivities of a modified bottom trawl for three fish species in the Aegean Sea. Fisheries Research, 39(1):17–31, December 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001726.

Thomas:1993:BIN

[TM93] G. L. Thomas and Ole A. Mathisen. Biological interactions of natural and enhanced stocks of salmon in Alaska. Fisheries Research, 18(1–2):1–17, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783693900367.

Travnichek:1997:AVB

[TMD97] Vincent H. Travnichek, Michael J. Maceina, and Rex A. Dunham. Angling vulnerability of black crappies, white crappies, and their naturally produced hybrid in Weiss Reservoir, Alabama, USA. Fisheries Research, 29(2):185–191, February 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369600522X.

Tsimenides:1992:DPT

[TMK92] N. Tsimenides, A. Machias, and A. Kallianiotis. Distribution patterns of triglids (Pisces: Triglidae) on the Cretan shelf (Greece), and their interspecific associations.

Fisheries Research, 15(1–2):83–103, October 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290006F.

Treble:1998:SSL

[TMW98] Rodney J. Treble, Russell B. Millar, and Terence I. Walker. Size-selectivity of lobster pots with escape-gaps: application of the SELECT method to the southern rock lobster (Jasus edwardsii) fishery in Victoria, Australia. Fisheries Research, 34(3):289–305, April 1998. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000726.

Toresen:1991:AAE

[Tor91] Reidar Toresen. Absorption of acoustic energy in dense herring schools studied by the attenuation in the bottom echo signal. Fisheries Research, 10(3–4):317–327, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190083R.

Tokai:1996:MDS

[TOSM96] Tadashi Tokai, Shigeyuki Omoto, Ryozo Sato, and Ko Matuda. A method of determining selectivity curve of separator grid. Fisheries Research, 27(1–3):51–60, June 1996. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004718.

Townsend:1998:BIP

[Tow98] Ralph E. Townsend. Beyond ITQs: property rights as a management tool. Fisheries Research, 37(1–3):203–210, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001374.

Tulevech:1994:ATA

[TR94] S. M. Tulevech and C. W. Recksiek. Acoustic tracking of adult white grunt, Haemulon plumieri, in Puerto Rico and Florida. Fisheries Research, 19(3–4):301–319, April 1994.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900469.

Trevorrow:1998:BSL

[Tre98a] Mark V. Trevorrow. Boundary scattering limitations to fish detection in shallow waters. Fisheries Research, 35(1–2):127– 135, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698000678.

Trevorrow:1998:SHS

[Tre98b] Mark V. Trevorrow. Salmon and herring school detection in shallow waters using sidescan sonars. Fisheries Research, 35(1– 2):5–14, May 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369800054X.

Trondsen:1997:TBW

[Tro97] Torbjørn Trondsen. Trawling blue whiting for surimi: a study of catch variations in low season. Fisheries Research, 32(3):201–212, December 1, 1997. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783697000635.

Trondsen:1998:BWS

[Tro98] Torbjørn Trondsen. Blue whiting surimi: new perspectives on the market value. Fisheries Research, 34(1):1–15, February 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369700088X.

Tobar:1992:ADL

[TS92] R. Tobar and F. Sardà. Annual and diel light cycle as a predictive factor in deep-water fisheries for the prawn Aristeus antennatus Risso, 1816. Fisheries Research, 15(1–2):169–179, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290012I.

Turunen:1997:SMU

[TSS97] Timo Turunen, Ilkka Sammalkorpi, and Petri Suuronen. Suitability of motorized under-ice seining in selective mass removal of coarse fish. Fisheries Research, 31(1–2):73–82, July 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000180.

Tudela:1999:MVM

[Tud99] Sergi Tudela. Morphological variability in a Mediterranean, genetically homogeneous population of the European anchovy, Engraulis encrasicolus. Fisheries Research, 42(3):229–243, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000521.

Turnock:1997:BRP

[Tur97] Benjamin J. Turnock. Book review: Proceedings of the International Symposium on North Pacific Flatfish (1994: Anchorage, Alaska): Alaska sea grant college program report no. 95-04, University of Alaska Fairbanks, 643 pp., hardback US$25, ISBN 1-56612-034-9. Fisheries Research, 29(2):201– 202, February 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783696005322.

Turnbull:1992:SDS

[TW92] C. T. Turnbull and R. A. Watson. Sledges for daytime sampling of juvenile penaeid shrimp. Fisheries Research, 14(1): 31–40, June 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290071Z.

Tully:1993:PNL

[TW93] O. Tully and K. F. Whelan. Production of nauplii of Lepeophtheirus salmonis (Krøyer) (Copepoda: Caligidae) from farmed and wild salmon and its relation to the infestation of wild sea trout (Salmo trutta L.) off the west coast of Ireland in 1991. Fisheries Research, 17(1–2):187–200, June 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165783693900183.

Ulltang:1998:EPF

[Ull98] Øyvind Ulltang. Explanations and predictions in fisheries science — problems and challenges in a historical and epistemological perspective. Fisheries Research, 37(1–3):297–310, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001453.

Urban:1998:DMC

[Urb98] H.-Jörg Urban. Description and management of a clam fishery (Gari solida, psammobiidae) from Bahía Independencia, Peru (14^{o}S). Fisheries Research, 35(3):199–207, May 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698000757.

Utter:1993:CUE

[USS93] Fred M. Utter, James E. Seeb, and Lisa W. Seeb. Complementary uses of ecological and biochemical genetic data in identifying and conserving salmon populations. Fisheries Research, 18(1–2):59–76, October 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390040E.

Uutala:1994:PAD

[UYD^{+}94] Allen J. Uutala, Norman D. Yan, Aruna S. Dixit, Sushil S. Dixit, and John P. Smol. Paleolimnological assessment of damage to fish communities in three acidic, Canadian Shield lakes. Fisheries Research, 19(1–2):157–177, February 1994. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783694900213.

Vehanen:1996:CNF

[VA96] Teppo Vehanen and Jouni Aspi. Classification of northern Finnish lakes and the suitability for the stocking for brown trout (Salmo trutta (L.) m. lacustris). Fisheries Research, 27(1–3):37–49, June 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696004729.

Vega-Cendejas:1997:BTH

[VCMCA97] M. E. Vega-Cendejas, G. Mexicano-Cíntora, and A. M. Arce. Biology of the thread herring Opisthonema oglinum (Pisces: Clupeidae) from a beach seine fishery of the Campeche Bank, Mexico. Fisheries Research, 30(1–2):117–126, May 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005474.

vanderWalt:1997:AGS

[vdWB97] B. A. van der Walt and L. E. Beckley. Age and growth of Sarpa salpa (Pisces: Sparidae) off the east coast of South Africa. Fisheries Research, 31(3):241–248, August 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000210.

Venema:1992:SFF

[Ven92] Siebren C. Venema. Successes and failures of fisheries acoustics in developing countries. Fisheries Research, 14(2–3):143– 158, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900504.

Vince:1991:SIS

[Vin91] M. R. Vince. Stock identity in spurdog (Squalus acanthias L.) around the British Isles. Fisheries Research, 12(4):341–354, December 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190017A.

Vincent-Lang:1993:MCS

[VLAM93] Doug Vincent-Lang, Marianna Alexandersdottir, and Doug McBride. Mortality of Coho salmon caught and released using sport tackle in the Little Susitna River, Alaska. Fisheries Research, 15(4):339–356, January 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390085L.

Villanueva:1997:VOI

[VM97] Roger Villanueva and Balbina Molí. Validation of the otolith increment deposition ratio using alizarin marks in juveniles of the sparid fishes, Diplodus vulgaris and D. puntazzo. Fisheries Research, 30(3):257–260, June 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000088.

Vidalis:1997:DBP

[VMT97] Kosmas Vidalis, George Markakis, and Nikolaos Tsimenides. Discrimination between populations of picarel (Spicara smaris L., 1758) in the Aegean Sea, using multivariate analysis of phenetic characters. Fisheries Research, 30(3):191–197, June 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005711.

Vallin:1998:CMI

[VN98] Lars Vallin and Anders Nissling. Cell morphology as an indicator of viability of cod eggs — results from an experimental study. Fisheries Research, 38(3):247–255, November 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578369800157X.

Velasco:1998:EHM

[VO98] F. Velasco and I. Olaso. European hake Merluccius merluccius (L., 1758) feeding in the Cantabrian Sea: seasonal, bathymetric and length variations. Fisheries Research, 38(1):33–44, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001118.

Vassilopoulou:1993:FHR

[VP93] V. Vassilopoulou and C. Papaconstantinou. Feeding habits of red mullet (Mullus barbatus) in a gulf in western Greece. Fisheries Research, 16(1):69–83, February 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369390110S.

Vera:1997:PMF

[VS97] Juan Madrid Vera and Pilar Sánchez. Patterns in marine fish communities as shown by artisanal fisheries data on the shelf off the Nexpa River, Michoacán, México.

Fisheries Research, 33(1–3):149–158, December 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000593.

Volk:1999:OTM

[VSG99] Eric C. Volk, Steven L. Schroder, and Jeffrey J. Grimm. Otolith thermal marking. Fisheries Research, 43(1–3):205–219, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000739.

Valentinsson:1999:APF

[VSJ^{+}99] Daniel Valentinsson, Frank Sjödin, Per R. Jonsson, Per Nilsson, and Colin Wheatley. Appraisal of the potential for a future fishery on whelks (Buccinum undatum) in Swedish waters: CPUE and biological aspects. Fisheries Research, 42(3):215–227, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783699000508.

Vidalis:1996:ADG

[VT96] K. Vidalis and N. Tsimenidis. Age determination and growth of picarel (Spicara smaris) from the Cretan continental shelf (Greece). Fisheries Research, 28(4):395–421, December 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783696004882.

VanDolah:1991:SES

[VWL91] Robert F. Van Dolah, Priscilla H. Wendt, and Martin V. Levisen. A study of the effects of shrimp trawling on benthic communities in two South Carolina sounds. Fisheries Research, 12(2):139–156, October 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190038H.

Worthington:1997:DCB

[WA97] Duncan G. Worthington and Neil L. Andrew. Does covariation between growth and reproduction compromise the use of an alternative size limit for the blacklip abalone, Haliotis rubra, in NSW, Australia? Fisheries Research, 32(3):223–231, December 1, 1997. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000544.

Worthington:1998:IIC

[WAB98] D. G. Worthington, N. L. Andrew, and N. Bentley. Improved indices of catch rate in the fishery for blacklip abalone, Haliotis rubra, in New South Wales, Australia. Fisheries Research, 36(2–3):87–97, June 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698001064.

Waldman:1999:ICS

[Wal99] John R. Waldman. The importance of comparative studies in stock analysis. Fisheries Research, 43(1–3):237–246, October 1999. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000752.

Wardle:1991:BRP

[War91] C. S. Wardle. Book review: Physiology and form of fish circulation: Geoffrey H. Satchell. Cambridge University Press, 1991, 235 pp., £32.50 (US$59.50), ISBN 0 521 39519 4. Fisheries Research, 12(3):279–280, November 1991. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783691900992.

Wardle:1995:BRW

[War95] C. S. Wardle. Book review: Wildlife telemetry: Remote monitoring and tracking of animals: I. G. Priede and S. M. Swift (Editors). Ellis Horwood, New York, 1992, 708 pp., US $162, ISBN 0-13-957994-X. Fisheries Research, 23(1–2):190– 191, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783695900039.

Wassef:1991:CGS

[Was91] Elham A. Wassef. Comparative growth studies on Lethrinus lentjan, Lacépède 1802 and Lethrinus mahsena, Forsskål 1775 (Pisces, Lethrinidae) in the Red Sea. Fisheries Research, 11(1):75–92, February 1991. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369190033C.

Wassenberg:1997:EFS

[WBB^{+}97] T. J. Wassenberg, S. J. M. Blaber, C. Y. Burridge, D. T. Brewer, J. P. Salini, and N. Gribble. The effectiveness of fish and shrimp trawls for sampling fish communities in tropical Australia. Fisheries Research, 30(3):241–251, June 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783696005516.

Wassenberg:1998:VSD

[WBCG98] T. J. Wassenberg, C. Y. Burridge, M. Connell, and N. Gribble. A validation of short-duration scientific tows as a representation of long commercial-length tows: comparing the catch rates, size composition and species composition of prawn trawler bycatch in the far northern Great Barrier Reef, Australia. Fisheries Research, 36(1):35–46, June 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698000769.

Wang:1995:RBJ

[WC95] S. B. Wang and C. T. Chen. Reproductive biology of Japanese butterfish Psenopsis anomala (Stromateidae) from coastal waters of northeastern Taiwan. Fisheries Research, 23(1–2):127– 141, May 1995. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369400335T.

Wardle:1991:ECA

[WCMG91] C. S. Wardle, G. Cui, W. R. Mojsiewicz, and C. W. Glass. The effect of colour on the appearance of monofilament nylon under water. Fisheries Research, 10(3–4):243–253, January 1991. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369190078T.

Weng:1992:SCP

[Wen92] H. T. Weng. The sand crab (Portunus pelagicus (Linnaeus)) populations of two different environments in Queensland. Fisheries Research, 13(4):407–422, April 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290061W.

Westin:1998:SME

[Wes98] Lars Westin. The spawning migration of European silver eel (Anguilla anguilla L.) with particular reference to stocked eel in the Baltic. Fisheries Research, 38(3):257–270, November 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001623.

Ward:1993:COT

[WF93] J. N. Ward and R. S. T. Ferro. A comparison of onetenth and full-scale measurements of the drag and geometry of a pelagic trawl. Fisheries Research, 17(3–4):311–331, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390132Q.

Windsor:1990:IMA

[WH90a] M. L. Windsor and P. Hutchinson. International management of Atlantic salmon — the role of NASCO. Fisheries Research, 10(1–2):5–14, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016578369090012K.

Windsor:1990:PIB

[WH90b] M. L. Windsor and P. Hutchinson. The potential interactions between salmon aquaculture and the wild stocks — a review. Fisheries Research, 10(1–2):163–176, December 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090021M.

Wassenberg:1993:SAD

[WH93] T. J. Wassenberg and B. J. Hill. Selection of the appropriate duration of experiments to measure the survival of animals discarded from trawlers. Fisheries Research, 17(3–4):343– 352, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369390134S.

Wassef:1997:BBR

[WH97] Elham A. Wassef and Hayam A. Abdul Hady. Breeding biology of rabbitfish Siganus canaliculatus (Siganidae) in mid Arabian Gulf. Fisheries Research, 33(1–3):159–166, December 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783697000751.

White:1991:BRB

[Whi91] Alan W. White. Book review: Biology, epidemiology and management of pyrodinium red tides: Gustaaf M. Hallegraeff and Jay L. Maclean (Editors), Fisheries Department, Ministry of Development, Bandar Seri Begawan 1921, Brunei Darussalam, 1989, 286 pp., US $40.00 (including Airmail), ISBN 971-10-2264-8, ISSN 0115-4435. Fisheries Research, 12(3):280–282, November 1991. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369190100T.

Wilson:1990:BRM

[Wil90] John H. Wilson. Book review: Marine invertebrate fisheries: their assessment and management: John F. Caddy (Editor). John Wiley, New York, U.S.A., 1989, 752 pp., £51.10, ISBN: 0-471-83237-5. Fisheries Research, 9(1):92–93, July 1990. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369090045W.

Wilkins:1995:BRG

[Wil95] N. P. Wilkins. Book review: Genetic conservation of salmonid fishes: J. G. Cloud and G. H. Thorgaard (editors), Plenum Press, New York and London, 1993, $89.50, ISBN 0-306-44532-8. Fisheries Research, 22(3–4):323–324, March 1995. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695900268.

Wassenberg:1990:WHP

[WK90] T. J. Wassenberg and J. Kerr. Wound healing in prawns tagged with streamer tags, and the effect of tag placement and prawn size on survival rate. Fisheries Research, 9(3):209–218, September 1990. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605800025.

Washington:1993:OIE

[WK93] Percy M. Washington and Ann M. Koziol. Overview of the interactions and environmental impacts of hatchery practices on natural and artificial stocks of salmonids. Fisheries Research, 18(1–2):105–122, October 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783693900437.

Wurtz:1992:DAO

[WMB92] M. Wurtz, G. Matricardi, and P. Belcari. Distribution and abundance of the octopus Eledone cirrhosa in the Tyrrhenian Sea, Central Mediterranean. Fisheries Research, 13(1):53–66, January 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290033P.

Walsh:1992:CSA

[WMCH92] Stephen J. Walsh, Russell B. Millar, Chris G. Cooper, and William M. Hickey. Codend selection in American plaice: diamond versus square mesh. Fisheries Research, 13(3):235– 254, March 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165783692900799.

Williams:1993:DTI

[WMCW93] A. Williams, M. Moran, N. Caputi, and C. Walters. Didymozoid trematode infection of snapper, Pagrus auratus (Sparidae), off Western Australia: parasite population biology and fishery implications. Fisheries Research, 16(2):113–129, March 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390047B.

Ward:1992:DFA

[WN92] David L. Ward and Anthony A. Nigro. Differences in fish assemblages among habitats found in the lower Willamette River, Oregon: application of and problems with multivariate analysis. Fisheries Research, 13(2):119–132, February 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290022L.

Wang:1996:SME

[WS96] You-Gan Wang and Ian F. Somers. A simple method for estimating growth parameters from multiple lengthfrequency data in presence of continuous recruitment. Fisheries Research, 28(1):45–56, August 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696004845.

Watson:1993:MGT

[WT93] R. A. Watson and C. T. Turnbull. Migration and growth of two tropical penaeid shrimps within Torres Strait, northern Australia. Fisheries Research, 17(3–4):353–368, August 1993. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369390135T.

Walker:1998:PAC

[WTHC98] Terence I. Walker, Bruce L. Taylor, Russell J. Hudson, and Jason P. Cottier. The phenomenon of apparent change of growth rate in gummy shark (Mustelus antarcticus) harvested off southern Australia. Fisheries Research, 39(2):139–163, December 15, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001805.

Washington:1992:SFA

[WTM92] Percy M. Washington, G. L. Thomas, and D. A. Marino. Successes and failures of acoustics in the measurement of environmental impacts. Fisheries Research, 14(2–3):239–250, August 1992. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016578369290056Y.

Wilderbuer:1999:EPD

[WZ99] Thomas K. Wilderbuer and Chang Ik Zhang. Evaluation of the population dynamics and yield characteristics of Alaska plaice, Pleuronectes quadrituberculatus, in the eastern Bering Sea. Fisheries Research, 41(2):183–200, June 1999. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000120.

Xiao:1997:SPP

[Xia97] Yongshun Xiao. Subtleties in, and practical problems with, the use of production models in fish stock assessment. Fisheries Research, 33(1–3):17–36, December 1997. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783697000659.

Xiao:1998:TSA

[Xia98a] Yongshun Xiao. Two simple approaches to use of production models in fish stock assessment. Fisheries Research, 34(1):77– 86, February 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000647.

Xiao:1998:WUP

[Xia98b] Yongshun Xiao. What are the units of the parameters in the power function for the length–weight relationship? Fisheries Research, 35(3):247–249, May 15, 1998. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783698000794.

Yousif:1999:APD

[YA99] Aly Yousif and Asgeir Aglen. Availability of pelagic distributed cod (Gadus morhua L.) to bottom trawls in the Barents Sea. Fisheries Research, 44(1):47–57, November 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699000557.

Yamada:1998:STS

[YAM98a] Tomohide Yamada, Ichiro Aoki, and Isamu Mitani. Spawning time, spawning frequency and fecundity of Japanese chub mackerel, Scomber japonicus in the waters around the Izu Islands, Japan. Fisheries Research, 38(1):83–89, September 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783698001131.

Yamane:1998:ARD

[Yam98b] Takeshi Yamane. Application of the Rayleigh distribution to size selectivity of small prawn pots for the oriental river prawn, Macrobrachium nipponense. Fisheries Research, 36(1):27–33, June 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698000848.

Ye:1996:MIB

[YB96] Yimin Ye and John Beddington. Modelling interactions between inshore and offshore fisheries: the case of the East China Sea hairtail (Trichiurus haumela) fishery. Fisheries Research, 27(4):153–177, July 1996. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165783696004808.

Yatsu:1995:EST

[YDM95] Akihiko Yatsu, Michael Dahlberg, and Skip McKinnell. Effect of soaking time on catch-per-unit-effort of major species taken in the Japanese squid driftnet fishery in 1990. Fisheries Research, 23(1–2):23–35, May 1995. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369400337V.

Yamane:1992:CPS

[YF92] Takeshi Yamane and Akio Fujiishi. Catch performance of small prawn pots in terms of selectivity and escape. Fisheries Research, 15(3):291–299, December 1992. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369290020T.

Yatsu:1997:AGN

[YMSU97] Akihiko Yatsu, Satoshi Midorikawa, Takahiro Shimada, and Yuji Uozumi. Age and growth of the neon flying squid, Ommastrephes bartrami, in the North Pacific Ocean. Fisheries Research, 29(3):257–270, March 1997. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783696005413.

Yoshimizu:1993:SCI

[YNEK93] M. Yoshimizu, T. Nomura, Y. Ezura, and T. Kimura. Surveillance and control of infectious hematopoietic necrosis virus (IHNV) and Oncorhynchus masou virus (OMV) of wild salmonid fish returning to the northern part of Japan 1976– 1991. Fisheries Research, 17(1–2):163–173, June 1993. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016578369390016Z.

Young:1998:IUI

[You98] Oran R. Young. Institutional uncertainties in international fisheries management. Fisheries Research, 37(1–3):211–224, August 1, 1998. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783698001386.

Zheng:1996:HSR

[Zhe96] Jie Zheng. Herring stock-recruitment relationships and recruitment patterns in the North Atlantic and Northeast Pacific oceans. Fisheries Research, 26(3–4):257–277, May 1996. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165783695004351.

Zhang:1992:PNA

[ZM92] Ziyang Zhang and Erlend Moksness. Preliminary notes on age information from sections of decalcified otoliths of Atlantic herring. Fisheries Research, 15(1–2):181–186, October 1992. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016578369290013J.

Zhou:1997:BRR

[ZS97] Shijie Zhou and Thomas C. Shirley. Behavioural responses of red king crab to crab pots. Fisheries Research, 30(3):177– 189, June 1997. CODEN FISRDJ. ISSN 0165-7836 (print), REFERENCES 270

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783697000052.